A Bibliography of Publications in the Hewlett-Packard Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

28 November 2022 Version 1.25

Title word cross-reference

2 [WJ79]. _{2} [Can93, Kri81b, Mer81]. f(x) = 0 [Kah79]. μ [Ano94-31, BBF^{+}97, Raj95]. π/4 [Bir91].

-law [Ano94-31].

0.01-to-20-GHz [BBI^{+}93]. 0.01-to-40-GHz [Zel91]. 0.1-Micrometer [RNR^{+}98]. 0.35- [BBF^{+}97, Raj95].

1 [Bol75, DSB94, FMN84, Mae74, MS82, RO86, REB86, Sch86, Toe86]. 1-GHz [RO86, REB86, Sch86, Toe86]. 1-Hz [Bol75]. 1-MHz [AI84, Mae74]. 1-MIPS [FMN84, MS82]. 1.0625 [CDLN96].

1.0625-Gbit [CDLN96]. 1.0625-Gbit/s [CDLN96]. 1.55-mm [Ano95-47]. 1.8-GHz [KN93]. 1/4 [DSB94]. 1/4-Inch [DSB94]. 10 [DA75, Hil93a, San72]. 10-Hz-to-13-MHz [DA75]. 10-Hz-to-150-MHz [CCH^{+}91]. 10-Megasample-per-Second [Hil93a]. 10-MHz [Hil93b]. 10-MHz-to-18-GHz [San72]. 100 [Bes75, HT92, Jer74]. 100-dB [Jer74]. 100-MHz [Wit92b, Bes75, HT92]. 100000-sample-per-second [Goe89]. 100VG [AT95, Ano95-36, Bro95, CCC^{+}95, CJ95, GS95]. 100VG-AnyLAN [AT95, Ano95-36, Bro95, CCC^{+}95, CJ95, GS95]. 11.8-in [HFKS95]. 110-dB [Hil93b]. 1100 [Jek75]. 1100-MHz [Jek75]. 12-bit [GY88, SIBH88]. 12-megabyte [Smi79]. 120 [Rya95]. 120-MHz [Rya95]. 1200C [BCK^{+}94, BTE^{+}94, CAJ^{+}94, DMW94, PPMW94]. 125 [SIBH88]. 125-MHz [SIBH88]. 128K [WSBK83]. 128K-Bit [WSBK83]. 1300-nm [RWT97]. 13MHz [NTA72]. 15-port [Bro95]. 150 [LGV^{+}98]. 150-MHz-Bandwidth [LGV^{+}98]. 16-Bit [RR83, Goe89]. 16-volt [KE74]. 1611A [Ogd78]. 18 [BFHT83, LE83]. 18-MHz [BFHT83, LE83]. 18C [LS87]. 1980A [Kar82]. 1980A/B [Kar82]. 1980s [Cla79]. 1990s [HPF93]. 1996 [Ano96-28]. 1a [MRH94d].

2 [DSB94, HE75, Ujv94]. 2-to-18-GHz [HE75]. 2-to-26.5-GHz [HZ83]. 2.488-Gbit [GLP^{+}97, WSYD97]. 2.488-Gbit/s [GLP^{+}97, WSYD97]. 2.5-GHz [US93]. 2/ [PS81, Sol81]. 20 [KE74]. 20-GHz [DVl91, HH91, HCHR91]. 20-MHz [KE74]. 200 [MP81]. 200-kHz [MP81]. 2100A [Cou71]. 21MX [Lan77, RH74, Rig77]. 2392A [SBB85]. 25-Line [Adl86]. 250 [GZC74]. 250-Megasample-Per-Second [DWD88]. 250-MHz [GZC74]. 26.5 [Bol80, NML90]. 26.5-GHz [Bol80]. 26.5-to-75-GHz [NML90]. 280 [MT81]. 280-MHz [MT81]. 28C [LS87]. 2D [MMNR89].

3-GHz [SPD^{+}93, Wag93]. 3-MIPS [WST84]. 3.0 [PT90]. 30-bit [WYSE97]. 30-lb [Smi69]. 300-lpm [LT85]. 300-MHz [BKE^{+}90, SES90]. 3000/IBM [Ish80]. 3054A [Wol81]. 32-Bit [ACK^{+}84, BZS83, BBD^{+}84, GJW83, BFHT83, Fia83, GOS84, HKC84b, Jon84, LMM84, LWBB84]. 33 [Hol79, MK79]. 3326A [Spa85]. 33T [Ano91e]. 3421A [Leo83]. 35 [Coc72]. 3820A [Was80]. 38G [Don96]. 3D [Ano94w, CG98, LB98, Tha89]. 3V [KM97].

4 [Bol74, Zel79]. 4-GHz [Bol74]. 4-Inch [DSB94]. 4-MHz [GY88]. 4-to-1300-MHz [Zel79]. 4.5 [Bol75]. 4.5-GHz [Bol75]. 40 [Say80]. 40-GHz [Say80]. 400 [Jek80, Ter78]. 400-Ipm [Ter78]. 400-to-1600-MHz [Jek80]. 400-to-1600-MHz/ [Jek80]. 41 [Wol83]. 41C [CR81, JM80, Ste80a]. 45C [Bec80, JF80]. 486/33T [Ano91e]. 48SX [BBE^{+}91, HW91, Rip91, SMB^{+}91, WP91]. 4940A [Lee74].

5 [DSB94, WD84, Yan92]. 5-Hz-to-200-MHz [WD84]. 5-Hz-to-500-MHz [Yan92]. 5-volt [DiF75]. 5.9 [Kuh76]. 5.9-12.4-GHz [Kuh76]. 5/10 [Har87]. 50 [Sti77]. 50-mbit [YPS76]. 50-mbit/s [YPS76]. 50-Megabyte [Sti77]. 50-to-110-GHz [SA91]. 500 [Ano97a, BWW84, Vif76, WLHC84]. 500-kHz-to-1.3-GHz [Vif76]. 500-MHz [BKE^{+}90, Hei88, KW90, Ano97a]. 500C [MH92, MJKS92, RL92]. 500C/DeskWriter [MH92, MJKS92, RL92]. 50Hz [NTA72]. 50Hz-to-13MHz [NTA72]. 5335A [LaF80]. 5345A [Fel74]. 53C710 [Ano92b]. 54601A [Ano92-36].

600-pixel-per-inch [WYSE97]. 64000 [AVB80, McD80]. 65 [Pet74, Sto74].

7090A [DF86]. 7100LC [Ano95-39, BBJ^{+}95, BKQW95, LBLS95a, LBLS95b]. 71400C [Ano91f]. 71B [BD84, Don86, Lin84, Mey84, Mil84, WAZH84]. 7300LC [BT97, JU97, WT97]. 75 [AT83, Pat83, Yor83]. 7550A [TRH85].

8 [Koh72, MB93, Sch93a]. 8-Bit [MB93]. 8-Gigasample-per-Second [MB93, Sch93a]. 8-MHz [Koh72]. 8.5-digit [GSVS89, Ste89b]. 8.5-digital [Goe89]. 80 [EDF82, Har82]. 8000 [BBR^{+}97, DMJCB97, MGBB97]. 8146A [Bel93, JS93, Mai93]. 85 [Leo83, NMS80]. 85/3054A [Wol81]. 850-nm [SFHE85]. 8505A [DH76]. 8509A [Ano95u]. 8509A/B [Ano95u]. 8662A [Kov81]. 8703A [Ano91f]. 8901A [Nae79]. 8903A [Boy80].

9000 [Gup95]. 9100A [Coc68, Nea68, Osb68]. 9490 [BS98]. 9660/HSG [KGTM90]. 9826A [DU82, KHI82, Spr82]. 9826A/9836A [KHI82]. 9835A [Chu79, Hal79, US79]. 9835A/B [Chu79, Hal79, US79]. 9836A [DU82, KHI82, Spr82]. 9915A [Swe81].

A-law [Ano94-31, BFS94]. A-series [WS84b]. a.c [Han70]. Abbreviations [Ano94p]. Absolute [Ano93a]. absorbance [WTB^{+}90]. AC [Fol72, Rob82b, Swe89, WB83a]. accelerates [SK80]. Accelerator [Mar95b, SOG98]. Acceptable [Ano92a]. accepts [Jer74]. Access [Ano94v, Man90, BCH^{+}89, CKP89, HG79, SWP90, WJ80]. accessibility [Inh75]. Accessible [Ano95x]. Accessing [EY96]. accessories [DVl91, KCP77, ST75]. AccuPage [WHBP95]. Accuracy [Bar82, Bir91, Duf87, HH91, LeV88, MR93, SC83b, Yan92, Bru79a, Coc72, Edw75, FR74, Geb78, Han70, KS72, Mar77, MW73, Ric81b, RW76, Rus93, SKB78, Sch70]. Accurate [NK84, Van98, AP76, BBT92, Kri81b, Ras87, Sch87, She71, Sol81]. accurately [Cha77]. achievement [Lea89]. Achieving [TD92]. Acoustic [LGV^{+}98, Shr82, Wai93, CE81, IKS81, Lar83b, Mil83]. acoustical [KB69b]. Acquisition [Cam94, HT92, KYKM85, MB93, Toe97, da 94, Ham93, HRH81, Hey81, Leo83, Res83, RO86, Wol75]. acquisition/control [HRH81, Res83]. across [Bac91]. Action [Ano92r, Bar96]. Active [ES87, OST75, Gue74a, Spa85, US93]. activity [Die75]. adaptable [FMN84]. adapted [FRAC85]. Adapting [Ano92b, Ogd78, Rob79, Wat88]. adaptive [gJH81, PN88]. adapts [GJ74, Ogd80].

ADC [Ano93b, Bak84, MR93]. Add [REOB78]. Add-on [REOB78]. added [EDF82]. Adding [BHW70, NMS80]. Adds [BD84, Don86, Mil84]. Adhesive [RL92]. adjustments [Sei91]. Administering [EY96]. Administration [Ano93-38, Not93]. Adopting [LH95]. Advanced [Ama86, BP82, BH81, BMS87, BPA^{+}94a, Coz78, Cun98, DJS83, EAW^{+}86, FHMP97, FCK97, GW77, SKK^{+}86, SAS98, Yan92, Bee82, BFMP84, GTM95, HDH^{+}87, Hun92, WD84]. AdvanceNet [CGS^{+}86]. Advances [Mat90, Mat92]. advantages [Kei71]. Advisor [Moo92, Ano92x, Bha92, BKD94, Dou92, Kre92, Pru92]. Against [Har84]. Agent [Spe96, Sto96, PW89]. Agents [Ste89a]. agile [RF72]. Ahead [Ano96-31, Ano96-32]. AI [Cag86]. Aid [BYA86, WS84b]. Aida [SMH86]. Aided [BP82, DV85, FR95, Fan72, PSA83, PGW73]. Aids [PS83, TE86, Ano67, Por80]. air [MM90]. airport [KB69a]. alarm [Haj96]. AlGaAs [SDC^{+}88]. algebraic [JY72]. Algorithm [Ano93-42, Can93, Kri81a, LHL83, Smi84a]. Algorithms [Coc72, FO95, HDH96, AM95, Egb77a, Egb77b, Egb77c, Egb78]. Alignment [Has92, KW93]. AlInAs [RNR^{+}98]. AlInAs/GaInAs/GaAs [RNR^{+}98]. all-haul [BP93]. All-semiconductor [Fin72, TW76]. Allows [Mal87]. almost [Mic71, Ogd80]. alone [DF86, YK77]. Alpha [Har87]. alphanumeric [UAH77]. also [BB78]. Alternative [ARL^{+}94, Ric76]. ALU [ALT82, Mur82]. Aluminium [FKO^{+}93]. American [Hoo91, Whi93]. AMIGO [Car79, WM79]. AMIGO/ [Car79, WM79]. Amplifier [Ano95-47, Ano96-29, KWR^{+}87, KM97, LSM^{+}95, Cho90, DeV70, Hei88, Jer74, MLJP93, Mil75, Pec74, Rie94]. Amplifiers [Sti97, Kih82, Oli71]. amplitude [Gue74b]. amplitude/delay [Gue74b]. amplitudes [BB78]. Analog [Cro79, DWD88, Gar92, Gro91, Hig88, Hil93a, KV94, Lym87, McD84, Mur92, MPR82, NVBH87, RK94a, San86, SWI97, TOB97, WP94a, Bes75, CPR75, CPK88, Dam74, GY88, Goe89, Hil93b, LS76, MC85, Rie94, SIBH88]. Analog-Digital [Gro91]. Analog-To-Digital [DWD88, MPR82, Hil93a, KV94, RK94a, WP94a, CPK88, GY88, Goe89, Hil93b, Rie94]. analyser [ALRS89, Boy80, Gue74b]. analyses [SWM78]. Analysis [Adc87b, ABC^{+}88, Ano92d, Ano93-30, Ano95k, AWH^{+}87b, AWH^{+}87a, BW92, BDS85, Bra96, CP87, Cor96, DG88, DV85, Edw87, EH84, Eps84, Gra96, HNL87, HKM96, HR82, KP88, LY94, Mil94, MM95, MKW^{+}96, Moo92, PK84, RB82, Sch96, SD95, WR85, AKL69, Ano67, Bet97, Bha92, BNB93, BF67, BBG91, CM72, Cra92, Edw79b, Fro77, GRS93, HH91, IOS80, Kei71, Kha81, Kil94, Ks69, Kin83, KHW89, Lin79b, LM93, MM90, McK75, Mil90, MB96, Nad77, NSA89, Nic88, NPG^{+}83, PCW77, Pru92, QW83, RR69, Rob82a, Rot70, RG69, RS70, SK69, Wha68, WB83a, WHHC89, WHH91]. analysis-concepts [Nad77]. Analytical [Ano93v, SD95, PB76]. Analyzer [Adc87a, Ano91f, Ano92-33, Ano95u, Ano96j, Ano96k, Ano97f, AWH^{+}87b, AWH^{+}87a, BW92, BKD94, DWB^{+}97, EEH^{+}84, HDH96, JRM^{+}86, KYKM85, KRDS83, LW82, MD96, MFW87, NG87, Neu84, PH96, Roe96, Uph88, War97a, Yan92, Yon94, ABM73, AP69, Ano67, ARGA85, Bak84, BHW70, BBDK93, BDOT85, BVAP87, CCH^{+}91, Cha77, CM75, CW72, CGH^{+}93, DH76, DA75, DW92b, DO77, Dou92, Due74, ES77, Edw79a, EAW^{+}86, Elo93, Erd79, GHUW68, GW77, Haa78, HW73, Haw71, Hei88, HCHR91, HH78, HMLN78, JM91, KN93, Kei71, Kis70, LD79, LS76, Lin79a, MW78, MB80, Mor76, MH80, MP83, NYS81, NS88, NPG^{+}83, Ogd78, Ogd80, PF78, RW70, Sap77, SSMK92, Sha72, Sha78a, SM75, SD78, Smi77, Sol81, Ste88, Tho80, Urq72, Vif76, Wai93, WN79b, WA93, WD84]. analyzer [Wri96, YN93, Zel79, ZHN83, Bru79b, Far74, Nae79]. Analyzers [Ano95v, Mat86, BS93, BCO^{+}93, CDS71, NTA72, NML90, REOB78]. Analyzing [Ano93c, Chu88, HBHC84]. angle [MS80]. anisotropic [HH90]. Annotation [Ano96q]. answer [CR73]. Anticipating [WR86]. any [Kah79, Ogd80]. AnyLAN [AT95, Ano95-36, Bro95, CCC^{+}95, CJ95, GS95]. anywhere [MW73]. API [LB98]. APIs [BT98]. APL [Joh77b, Leo77, Mun77, Van77b]. Aplets [Don96]. APLGOL [Joh77b]. Appendix [Ano96a, Bel98, MRH94c, MRH94d, MRH94b, MRH94a, WWS^{+}98]. Application [Ano96b, Cox96, Dei94, GLM95, Gro91, HD98, Nak96, Rus92, SDDC96, Smi94, Swe96, War86, WJ91, WCH94, BF89, Cow70, Dra77, GC90, Kru89, Kur81, MOT90, Ste89a, Vog91, Win81]. Application-Specific [Gro91]. Applications [Ake85, Ano93d, Ano94m, Ano95j, BKAWBCTK98, BP82, Can93, Cha96a, Col88b, CG98, EY96, FR87c, Gup95, Hal75, Kal93, KF95, KH88b, LY94, LB98, LMNQ96, LH95, MD86, May88, NIWM98, Not93, QH83, RNR^{+}98, SPA^{+}84a, Spe96, THHW95, Wai93, WN79a, Wat88, YN98, ZFE^{+}86, AC93, AK79, Bec93, BBE^{+}91, BC74, BP70, Cro89, DO77, Edw79a, FGH^{+}89, FD78, GS95, HMZ82, HRH81, ICNT93, Jon89, Lie88, MRM83, Nic90, Pra80, RW81, Smi80, SB80, Tho88, YK77]. Applied [DW87]. Applying [Gil96, Hig95]. Approach [Ano94-36, Ano97h, DGHP85, MSS87, MB85, BP83, Han80b, Joh80, Kru97, Lyn80a, NSA89, Raf93, SW80, SF95, Sel97, Wat81, Win81]. Approaches [Han94, RC98, SD95]. Arbitrary [Ano93-46, HSK87, KH88b, HK88, KMT88]. architecting [Sel97]. Architectural [CGS^{+}86, JM93]. Architecture [ALT82, Ame85, Ano93h, Ano94-33, Ano95m, Ano97a, BW92, Bee92, BKMM87, BHK^{+}86, CHKS86, Des92, Dic84, FSB^{+}87, GSG^{+}87, HSK87, HT92, KHLM90, LCPG98, MBC95, Nel82, PT92, Roe96, Wat87b, WR85, Wor92, YLJ^{+}87, BW85, BTE^{+}94, BLY92, Cag90, Chi90, Cou71, Fis88, Mar89, O’B93, Rei91, Rue80, SMT^{+}97, Tal88a, TSv92, WH91, BGK^{+}87, MBF^{+}87, MBH^{+}89, PB87, RJH^{+}87, WCT87]. Archiving [Loo97]. Area [AT95, Gd86, JRM^{+}86, LAB97, Met86, BWW84]. Arithmetic [Ano95-41]. Arrangements [HPL96]. Array [GFK83, MW94, SR95, TLCB98, Kal95, Lar83b, McK83, Mil83, WM78]. Arrays [BBF^{+}97, BP70]. Arsenide [Gib86]. Art [Rip87, TRH85]. arterial [MH76]. artificial [GUW92]. artwork [BH81, TH81]. ASAP [BH81]. ASCII [Leo77, Nad75, Tre78]. ASIC [Bur95, MBL95, MMPT97]. ASICs [MY95]. Aspect [SSMM81]. assembler [AS90, Mil84, Yac80]. assembler/disassembler [AS90]. Assemblies [Dul85, KWR^{+}87, Ste98, HRH81]. Assembly [Ano93-28, Ano96k, Buf96, CAJ^{+}94, Hal79, Pal96, Gas88, MH92]. Assembly-Language [Buf96]. Assessment [SD95, VCMH96, Gil89]. Assignment [Kru94a, Kru94b]. Assistant [SMH86]. associated [Kei71]. Assurance [AY98, Mar88]. assure [BLKH77]. assures [DJS83, Ter78]. Assuring [Bru79a]. Astronomers [She71]. Asynchronous [BYA86]. ATLAS [FG75]. ATM [DNC96, DGH^{+}97, MM97a]. attained [Sta77]. attenuation [Inh75]. Attenuator [MSS87, Ada71, SF95, Zam73]. attenuators [Ada71, KV74, Vet91]. Audio [BFS94, Ton85, Ujv94, Boy80, Foo80, KB69b]. audio-frequency [KB69b]. Audit [Loo97]. August [Ano96-28]. Author [Ano94b]. Authorization [Cas95b]. Authors [Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano96c]. autochanger [DSC^{+}90, OB90, SH90a]. Autoloader [Ano94i, Ano94j, Ano94t, DSB94]. autoloading [DPA^{+}88]. Automated [BF79, Bob73, CAJ^{+}94, Duf87, Dul85, Gri69, MH92, McD84, MB96, RSW^{+}86, Whi85, Gor91, Sch97, WCT87]. Automatic [Bal84, Ban69, BBT92, BF67, Bol75, Gar83, GDKK94, Kar94, KH84, Rya95, SJ80, Sim94a, APA92, Bol74, Bol80, Bre74, CWF^{+}79, CW72, Ewy69, ES69, FG75, FPS72, HT76, Hue73, Koe80, Mae74, NTA72, PN88, Pee73, RW70, RS70, Sch79, Sch73, Sha72, Tho80, Urq76, YK77]. automatically [Goo77, NSM70]. Automating [AM73, FS86, Jac88, Kur81, San72]. Automation [BP82, Dia94, MS95, PT90, Wat88, AGI^{+}91, ACCM83, Jon81]. Automotive [BKAWBCTK98, Ste98]. Autopeaking [Zan83]. AUTOPLOT [BP83]. AUTOPLOT/ [BP83]. autoranging [GBHS81, Ris82, Tho80]. autosampler [KP90]. Availability [Reh88, WA97]. Average [Ano92c, Ano92p, Cas95a, Ket74, Sch87]. averager [AP69]. Averaging [Ano93e, Ano93y, Chu74, RR69, Sch70]. axis [HPH81, KT81, KW93].

B [Kar82, Ano95u, Chu79, Hal79, LHIR98, NZ97, Oli71, US79]. B-Class [LHIR98]. B-ISDN [NZ97]. Backbones [BB97b]. backplane [Tal90]. Backscatter [Ano95f]. Backup [Ano94t, BDM85]. Balancing [Kru94a]. Band [McK75]. Band-selectable [McK75]. Bandpass [Wai93]. Bandwidth [Ano95p, ANSW95, Arm95, LGV^{+}98, Per86]. Bang [Ano92d]. Bang-Bang [Ano92d]. Bank [Wai93]. bantam [CJ71]. bar [CR81, ULW81]. Barium [NML90]. barrier [LM73]. Base [BF82, BK95, BHK^{+}86, EKW93, FGH^{+}89, Het79, McI74, Mor70, PN88]. Baseband [BBDK93, FST92, Lau87, LW82, Rob82a]. Based [Ano92m, Ano92-31, Bru97, CS88, CBL96, GV86, HR82, IGHG^{+}94, Mur97, NML90, TD92, BGG88, Boy80, DD97, EH81, Kis70, LFWEJ89, MBW79, Mat79a, Mat84, Moy78, Mye96, SSV90, SMT^{+}97, Sho89, Wag75, Wol75, GSG^{+}87, GLM95, KV94]. bases [TPG79]. Basic [NTJ^{+}91, AD75, MR91, FK84, Kov79, LTC^{+}84, LWBB84, MS74, MRH80, Pee79a, Sch76, Spa72a]. BASIC-language [Spa72a]. batch [AG75]. batch/spool [AG75]. Battery [BLSM83, GG77]. Battery-Operated [BLSM83]. battery-powered [GG77]. BCD [Nad75]. be [Spa72a, HH90]. Beam [Can93, Kam82, AN90, BOOS81, CLL81, EHS81, Gar83, Ham81, HHS76, gJH81, HMO73, KGK81, LM81, MW80, NE81]. Beam-Recrystallized [Kam82]. Behavior [DF90, MRH94e, HR91, Wec88]. Behavioral [SWI97]. bench [Bre74, MRM83]. Benchmark [MBL95]. Benchtop [Kis97]. Benefits [BCH81, MW80]. Best [Pet96]. Best-Fit [Pet96]. Better [AM95, Inh75, LM73, QM80]. between [LM93, Sha98]. Beyond [BW85, Ano93-45, SM82]. bias [But95]. bidirectional [And70]. Big [Mac75]. Binary [Ano94-27, Don71]. Bipolar [WSYD97, HLT^{+}83, Pec74, PW82, WJ79]. Bit [ACK^{+}84, BZS83, BBD^{+}84, GJW83, MBC95, MPR82, RR83, BFHT83, Fia83, GY88, GOS84, Goe89, HKC84b, Jon84, LMM84, LWBB84, SIBH88, WYSE97, MB93, WSBK83]. Bits [Ano92f, Ano93b, Ano93p, Gro92]. Black [Ben95]. Bleeding [TLCB98]. Blending [FO95]. Block [Ano97c, Rya95, CP85]. Block-Level [Ano97c]. block-mode [CP85]. blocking [VBK89]. blood [HT86, MH76, Rom91]. Board [Ano92e, Bal84, Hol84, Jac88, Kru94b, Reg88, ST84, SM86, Bas79, CWF^{+}79, FM84, GB87, HDH^{+}87, Mat84, Par79, Pet74, Sch79, SM79, TWCD93]. Board-Level [Ano92e]. Boards [Gas88, BL87]. Bond [HPL96]. Bonding [PHL96]. Boolean [Ano95o, Ern95, Fos98]. Boot [SMQ88]. both [ABW78, Emm72, Leo77, Nad75]. Branch [HNL87]. Bridge [Ano95-46, FP90]. Bridges [FP90]. Bridging [Ano95g, LM93, Fro90]. brief [AK81]. brighter [Har76]. brings [Har82]. Broadband [AKL69, Ano96j, Ano96k, BS95, DY75, LD79, SGC93, DF82, Han72, Kih82, KCP77, MC91, Spa84b, Wri96]. Broadcast [YN98]. broadens [BC74]. Broker [Ano95x, Ano93w, Ano93x, Ano93-41, GAL^{+}93, YGR95]. budget [TRH85]. Bugs [Ben95]. Build [BT98, Hig95]. builder [WB90]. Building [Cha96a, Ewy69, MK97, Ric91, Ive88]. Built [EMT83, Pea95, RK94a, Sei91, HA92, MT76]. Built-In [EMT83, Pea95, RK94a, Sei91, HA92, MT76]. Bulk [JM80]. Bus [BCF96, Fid96, Kar82, Rya95, BB73, Die75, JS74, RHM79, Lou75, Mat79a]. bus-oriented [RHM79]. Business [ARL^{+}94, BEM95, Jen98, SDDC96, BM79, Cla79, CR73, Ham79, Kov79, Mat79b, MD79, Tem83, WN79a, Wec87]. Buy [Hig95]. Buy-versus-Build [Hig95].

C [ACM92, Ano93f, Atc92, Bur86, DD97, LHIR98, MH92, MJKS92, MB96, Neu91, O’K93, QAKP90, RL92, Wil97b]. C-Class [LHIR98]. C-terminal [MB96]. C1553A [Ano94t]. Cable [Van98, GH73, Low69]. Cables [Ano95l]. Cabling [MCD97]. Cache [CP87, KNTV96]. Caching [BK85b]. CAD [Ano94k, BK95, WYO^{+}87]. CAI [AE71]. CaLan [Van98]. Calculating [Ano92f, War91]. Calculation [Reh88, WJ91]. Calculator [BBD^{+}96, BMS87, BPA^{+}94a, EMT83, MC82, BC74, CS76, CHM76, Coc68, CAV69, CR81, Egb77a, Egb77b, Egb77c, Egb78, Eve83, HW91, Her73, Kah79, Kah80, Lil72, Mar77, McC87, Mon68, MCCB76, MT76, MWC80, Nea68, Osb68, Rip91, SMB^{+}91, Spa72a, Tun74, WRT72, Wic87, WP91, Wol75, Zel70]. calculator-based [Wol75]. Calculators [LS87, Rip87, CFW75, DE76, HWS87, HDH^{+}87, NT75, Pat87, Spa72b]. Calculus [Ano95-28]. Calibrated [CK76, AP69, Fos74, Kei71, Kis70, Unt68]. calibrates [Goo77]. Calibration [BOOS81, Ger98, GSVS89, Hen91, Kaf88, MCM93, RK94a, RK94b, AM73, BBT92, Han70, MRRS93, PS81, Ste89b, Tho80, WB83b]. Calibrations [BTW83]. Call [Ano95h]. Caller [Ano95i]. Caller-ID [Ano95i]. camera [Bre74, Pax70]. can [Sch70, SFW78, Spa72a, Tho72]. Capabilities [Ano91f, BTBW83, BD84, Don86, Neu84, Pea95, Tuc95, ABD^{+}97, AG75, CM72, DF79, GDV73, Lai81, Mon68, MCCB76, NK78, SMR^{+}86, WN79b]. Capability [AWH^{+}87b, AWH^{+}87a, Ban84, BDS85, BLSM83, DSW83, BW72, Coz78, Emm72, GT71, Hal79, Hef73, HW82, LB80, Mac75, Mar77, MLS79, MFPE72, NMS80, Nie90, NSW76, SD78, LC96]. capability-both [Emm72]. capability/cost [Hef73]. capable [Ste89b]. Capacitance [AI84, Das96, FK79, NSM70, OS75]. Capacitance/Conductance [AI84]. Capacitive [Car81]. capacitors [Kam72]. Capacity [DSB94, Edw84, Mic72, VBK89, Wil84]. Capillary [Ano95j, Hol95, LRS83, RWW95, SD95, Zim95, GTM95, Kal95, Sch95, SMD95]. Capture [Bur95, Bia90, RBB78]. capturing [Chu88]. Card [AT83, HSYJ83, Ing76, LB80, Pug90, ST90, BE90, Joh90, Mat90, Pat83, SSV90, SH90b, Tag74, Tal90, TBTM92, Vog91, Zel70, AW90, Lon90]. card-based [SSV90]. Card-programmable [Ing76].

Cardiac [BJ82, Fan72, O’C86]. Cardiograph [BDS85]. Cards [Ano92e, Ada71, CLL^{+}92]. Carlo [NE81]. carrier [ZC69]. carrying [CM75]. Cartridge [Ano92g, Ano92h, BJJN88, Gen85, Gil89, Top89, BCK^{+}94, DMW94, MH92, MJKS92, Mon92, RL92, Ric76]. Cartridges [Ano92u, Ano94x, BCRS88, CAJ^{+}94]. Case [BKQW95, BB91, LM86, LHIR98, BF89, Cho90, Car95]. Cassette [BGS83]. catheterization [Fan72]. cathode [Cha70]. cathode-ray [Cha70]. CATV [Low69]. Caused [Har84]. Cavity [CWFS82, NMY95, RLB^{+}95, MRRS93, SL93, TZ93]. Cavity-Tuned [CWFS82]. CD [CKP89, KGTM90, Mey90b, Nie90, SS90]. CD-ROM [CKP89, KGTM90, Mey90b, Nie90, SS90]. CDE [Ano96b, Bar96, Cha96b, EY96, Joy96, OR96]. CDMA [Whi93]. CE [Ano95w, Ano95z, Ano95-44, BBW95, Jeg95, Sch95]. Cell [MY95, GTM95]. Cells [Rya95]. Cellular [Ano93g, Bir91, Hoo91, Whi93]. centered [BF89]. centers [ARGA85]. Central [BB73, BN80, Gol80, GJ74]. Centralized [But94]. Ceramic [TLCB98]. cesium [HHS76, HMO73]. Chain [HLH86, KWR^{+}87, BBI^{+}93, Kru97]. Challenge [Lon92]. Challenges [Bir97, LHIR98, Cra92, UAH^{+}92, WRT72]. change [Ano71b, HH90]. Channel [Ake85, AS85, Ban91, Kal93, San86, Gue74b, Mil75, Mut74, Sch87, CDLN96, Pri96, SP96]. channels [CM75, CPR75, Dam74]. Character [BP76]. Characteristics [SS83b]. Characterization [ADG^{+}96, BS95, Dia94, LGV^{+}98, Mil94, Wec88, Kha81, Kir71]. characterize [PHK77]. characterizing [BP93]. Chart [TE86, BP83, MW73, Mic71]. chartmaker [MS83, SDS83]. Chassis [Lon92, MHKS94, JKP^{+}88]. Checklist [Ano93l]. checkout [Gor69]. Checks [Ano95-41]. Chemical [Ili82, GRS93]. ChemStation [Ano95m]. Chip [BB97a, KGH^{+}87, MC84, Mur97, NSMO98, SLM^{+}95, TBB^{+}97, YLJ^{+}87, BT97, Fia83, For77, SP96, WJ79]. Chips [Ano90a, Fin72]. chipset [CDLN96, YWP^{+}92]. Choosing [ES69]. Chromatograph [KH84, DJS83, Ste74b, Wie90]. Chromatography [JR84, LRS83, NH94, ACCM83, BFP90, GV84, KP90, Row83, Sch75, SW90a, Smi73, SW90b, WMZ84]. chronicling [Lea89]. Chronological [Ano94a]. CIM [HCF^{+}88]. Circuit [Col88b, DMJCB97, Dul85, Gas88, Har84, Hen84, Jac88, KP88, LBKY86, Mar88, McD84, Pal96, Reg88, SM79, Tan96, WSYD97, BL87, Bas79, Bec76, BCH81, Cho90, Cro79, CWF^{+}79, EZ81, FL80, Har83, HDH^{+}87, HBB88, Hei88, Hol82a, Mat92, PW82, Ryt77, Sch79, Ste81]. circuit-board [Sch79]. Circuitry [HPL96]. Circuits [Ama86, BHW86, CS88, Kam82, Lam95, Lau87, WR97, ZFE^{+}86, BM81, CDE^{+}87, Cha77, DL77, FL73, FEG77, GJM^{+}92, HMPF74, Ham93, HLT^{+}83, NTH^{+}84, SDA81, WJ79]. Class [BK95, LHIR98, NK96, Pug90, Oli71, Wil97b]. class-B [Oli71]. Classes [BBD^{+}96]. cleaner [Bru73]. Cleanroom [Hea94]. Client [Ano93h, Ano94-33, HKH93, Kon95, LH95]. Client/Server [Ano93h, Ano94-33, HKH93, Kon95, LH95]. Clients [Ano96b]. Clip [BP72, Pri76]. Clip-and-read [BP72]. Clock [Toe97, WSYD97, WP94b, WP94c, LE83]. Clocks [Hoe97, Let92]. closed [MM79]. closed-loop [MM79]. Clusters [Ano93w, GL88]. CMOS [Ano94-36, AT83, Dia94, Dic84, EM98, GBO94, Har83, JM80, KM97, Nis87]. CMOSC [HJL^{+}83]. CO [Can93, Kri81b, Mer81, PS81, Sol81]. CO/ [PS81, Sol81]. Coatings [SD95]. Coaxial [KCP77]. COBOL [Car95]. Code [Bir96, BB91, Gil96]. Code-Domain [Bir96]. CodeAdvisor [DD97]. codes [CR81, ULW81]. Coding [CJ95, Sim94b]. coherence [BCH^{+}93]. Coherent [KNTV96, Mag83]. Coinjection [Ano92k]. Collaboration [GWY94]. Collaborative [Yod94]. Collection [Whi85, Ric81a]. Color [Ano92g, Ano92h, Ano94x, Ano95-48, BJJN88, Bar95, BCRS88, EGH^{+}88, FH80, HSW88, HASP94, MTMP83, Moa98, SB84, STMM88, TCWN88, WMBB94, ACK79, BB80, BLKH77, Bru77, GS93, HMW83, KA92, MH92, Mon92, PHK77, Pra80, RL92, TBBM83, WM94, WYSE97]. Column [Sch84, SD95, Kal95, KGK81]. Columns [LRS83]. combination [Kin81, Wol81]. Combinational [Lam95, GB87]. combined [PB82, YN93]. Combines [AWH^{+}87b, AWH^{+}87a, NVBH87, Mil75, SKB78]. come [Res83]. Command [CHP85, Kel92, Wol86, Voi84]. Commands [RK94b]. Commercial [BKO87, HMM96, AC93]. Common [Ano95m, CBL96, Lou72, KA95]. Commonality [MRH94c]. Communication [NDS86, Wat87b, Jon84, YWP^{+}92]. Communications [Bir97, BYA86, Dan96, DN86, Det72, DU82, HGWS97, HDH96, HKM96, McF98, MD96, MBD^{+}98, MKW^{+}96, PH96, Pit97, RWT97, RNR^{+}98, TBB^{+}97, WSYD97, AS90, BP77, Bra78, Bur85, CS76, CM75, Due74, Fis72, Han79, ICNT93, Low69, Mai90, NK78, Rue80, SSV90, Sta81a, Urq76, YPS76]. Communications-oriented [Det72]. Compact [ACK^{+}84, BLSM83, BGS83, Hef73, HL85, da 94, BHW77, BM79, BFMP84, CP85, CS80, GDV73, HMZ82, KP90, LK80, MS80, SFW78]. Compactness [Goo72]. companies [FS81a]. comparable [BCC^{+}94]. comparator [BP72]. Comparing [FJ89, Mur92]. Comparison [DR98]. Compatibility [Pal86, BSY89, BW72, Edw75, LB80]. compatible [DE76, Nad75]. compensation [Gar83]. Competitive [BKQW95]. compiled [CKP89]. Compiler [Ano94l, GJ98, AVB80, LTC^{+}84, PB87, Van77a, Van77c]. Compilers [CHK86, ST83, Han92, Ram92, San92]. Complementary [SNN88, Oli71]. complementary-pair [Oli71]. complete [Foo80]. completely [Bol74]. Complex [MBC95, MRH94e, SHB84, BCO^{+}93, Cha77, Far74, Haa78, Ive88, NTH^{+}84, Pir97]. Complexity [Moa98, War89]. Component [Ano91a, Ano91b, Ano91f, Ano94-38, BW92, DF91, DR95, Kai91, Rei91, SW91, Tiv91, WJ91, WH91, Wes91, ALRS89, BGG88, Cro79, HH91, HCHR91, JM91, WHHC89, WHH91]. Components [BTW83, Cox96, AKL69, KCP77, MN79, PK75, Sei81]. Compound [Col86]. Comprehensive [WS84b]. Compression [Moa98, BKV89]. computation [Lyn80b, Pat87]. Computational [Ano93w, GAL^{+}93]. Computer [AGO^{+}85, Ame85, BK85a, Bar84, BZS83, BP82, BLSM83, BBD^{+}84, BKO87, CGS^{+}86, DF90, Dic84, DU82, Don86, DV85, ELC^{+}86, FKL86, Fro86, GV86, GJW83, Hal82, HSYJ83, HL85, KE85, Lin84, MC84, Nea68, Oh82a, PDW85, Pea85, PGW73, Reh88, SJL91, SW86, Sny75, SM86, SHB^{+}82, Tuc95, WR86, WST84, WW84, AK79, Bak81, Bau80, BM79, BBG91, Bow72, BFMP84, Car69, Cat78, CP85, Chu79, Cla79, CPR75, Cou71, Dic74, DR81, DP84, EW78, EDF82, Edw79c, Eve83, FR95, FRAC85, Fan72, FG73, FK79, Fos73, FMN84, FH80, GOS84, GP84, Hal79, Ham79, HB72, HKC84b, Hey81, HM84, Joh77a, Jon81, Jon84, Jun78, Kis70, LMM84, Leo83, LK80, LLS88, LFWEJ89, Lyn80a, MM90, Mat79a, MS82, Mat79b, MK79, MD79, MBH^{+}89, Mey84, Mil84, MRH80, MCCB76]. computer [MRM83, Mye96, Nel79, PSA83, Per88, Pou68, Rig77, RJH^{+}87, Sha76, Spa72a, SS90, SK69, Sto74, SMK84, TBL^{+}83, Tun74, US79, Van77b, Van74, WN79a, WLHC84, WAZH84, Wil78, WM85, Win81, Won79, SPA^{+}84a, Suk84, WBSC84, Yor83]. Computer-Aided [BP82, DV85, PGW73, FR95, Fan72, PSA83]. computer-based [Kis70, LFWEJ89]. computer-controlled [Car69]. computer-like [MCCB76]. Computer-testing [Nea68]. computerlike [Mon68]. Computers [BKMM87, BHK^{+}86, BGK^{+}87, BK85b, CHKS86, CHK86, DN86, FSB^{+}87, FT86, GSG^{+}87, Gd86, HJL^{+}83, KM84, MD86, Met86, Pal86, RSW^{+}86, Row86, Spr82, Wol83, ASCC84, BWW84, CKP89, CS78, Geb78, GJM^{+}92, Har82, HL86, LP92, Mar89, McI74, MG92, TL89, Wal70, WS84b, WCT87]. Computing [ARL^{+}94, ALT82, GR69, HMM96, Kon95, LEV96, Mar70, Cat78, Chu79, CAV69, HB88, Ing70, Mar71]. Computing-counter [Mar70, Ing70]. Concept [MHKS94, MLSW83, DL73]. concepts [HS68, Nad77]. Concurrent [BKS93, CL98, Lon92]. Conditioning [BMD^{+}88, da 94, GY88]. conditions [MN79, Van74]. Conductance [AI84, FK79]. CONE [DKW90]. Configurable [Ano93-37]. Configuration [Ano91c, Lub91, Sch93c, Mey84, Wag88]. Configuring [WM79]. Conformance [Dub94]. Connection [Ano93k]. Connectionless [Ano90c]. Connectivity [BB97b, PPMW94]. Connector [Cou94, Rad87, RTD91]. Conservative [DGHP85]. Considerations [Cle93, Fid96, Bec80, DW92b, HH91, Mai93, MH89]. Consistency [Gem85]. Consistent [DF90, Bac91]. console [KM79]. Constant [KM97]. Constant-Transconductance [KM97]. Constellation [McK87]. constraints [Ker80]. Constructing [Gup95, Swe96]. Construction [WYO^{+}87].

Consultant [Wec87]. consultants [YGR95]. Consumers [Ano92u]. contact [Joh80]. contained [DL73, Foo80, Lo81]. Contemporary [BLMJ77]. Contents [Ano96-28]. Context [Ano93i]. Continuous [Bur93, MH76, MLS79, Wec88]. continuous-sweep [MLS79]. Contrast [Ben95]. Control [Ano94i, BP85, Can93, CDG^{+}88, Hig85a, JAD83, LW82, Nak96, Nel82, Sch84, Sha85, SW86, Van86, Wol83, ACCM83, BBW95, Bre74, Bru78, Chi81, Cla81a, Cou72, CH82, Dac76, FS81a, FS81b, GHK^{+}93, Ham81, HRH81, Hol82a, HL86, Kas72, Kla78, Kur81, Lai81, Leo83, Lin79a, MW78, Moy78, Res83, Sch75, SL93, TTML81, WMZ84, Woo78]. controlled [BB78, Bra78, Car69, Gor80, MH89, MW80, Pee73, QW83, Ste74b, Wai75, Wal70]. Controller [Adl86, AP90, LK85, MG85, AS90, Bee82, Boy80, LL91, Mai90, MMPT97, Roy81, Van77a, BGK^{+}87]. controllers [SS83a]. Controlling [AMSW85, MC82, Won79]. Controls [DH77, PL84]. Convenience [SC83b, FR74, Inh75, KS72]. Convenient [DSW83, PS83, Edw79b]. Conversion [Kel94, GY88, LHL83]. Converter [Ano96-29, DWD88, Gar92, MPR82, RK94a, WP94a, Bol74, CPK88, Cra72, DH76, Goe89, Hil93a, Hil93b, MC91, O’N80, Rie94, SIBH88, Say80]. Converters [KV94]. convincing [Ric76]. Convolution [Ano92o]. Cooling [Ano93j]. cooperation [BC81b]. coordinate [Joh80]. Coordinated [Ano96-30]. Copper [MCD97]. coprocessor [MG92]. Core [Cha96b, Lin97]. Corporate [ARL^{+}94, DH98]. Corrected [Ano93o]. Correction [Gal84, Ano90b, HH91, Mey90b, TW76]. corrections [RW81]. correctness [WT97]. correlating [RP88]. Correlation [Ano96r, RR69, She96, BGG88, SNN88]. correlation-based [BGG88]. correlator [AP69, MFPE72]. correlator/averager/probability [AP69]. Cost [Ano92-31, Ano97a, BDM85, BM79, BEM95, BCF96, BGS83, Cau85, Cla81b, Dou75, DGN98, Gib86, GLW85, Hal82, HM85, HL85, LAB97, Mal87, Mar95b, MD79, Pea95, QH83, RSW^{+}86, Spr82, WST84, WMBB94, Wol83, ABD^{+}97, Azm82, BD73, BDOT85, CP85, Cle79, CHM76, CS80, CK76, CF68, Dic78, Edw79c, FCC79, FL80, FHDS92, HVSM96, HWS87, Hay74, Hef73, Het79, HMM73, HC75, HL86, KA92, KPC^{+}82, LT85, MC85, MEM82, MO74, MBW79, Res83, RS75, Ric76, RC79, SJ75, SMT^{+}97, Smi79, SM87, Sta81b, SM79, TD83, TC79, War91, WDDH84, Wit92b]. Cost-Effective [Ano97a, GLW85, Hal82, BM79, Cla81b, Dou75, MD79, SM79]. costly [NT75]. Costs [Ano94y, BJM71, KSMM97]. Count [Ano96s]. Counter [Ano97a, GR69, SHB84, Ban69, Bol75, BB78, Bol80, CAF78, CK76, HC75, Ing70, Jek75, Koe80, LM86, LJ75, Mac73, Mac75, Mar70, Mar71, MBW79, McO80, Mut74, Per86, SJ80, Sch73, Sch80, WJ79]. Counters [Gib86, LM86, MO74]. Coupled [Kis97]. coupler [And70]. course [MB74]. Cover [Ano96d, Ano96e]. Covers [Ano92k]. CP [Har82]. CP/M [Har82]. CPU [CHK^{+}96, YLJ^{+}87]. CPUs [DR98]. Crash [Ran88]. Creating [AC90, Don96, Fro86, RC98]. Creation [PS83]. Creative [PDW85]. Creativity [Gem85]. creator [ADRS83]. Critical [War86, FD78, Nik97, Pir97]. Cross [Ano95k, Ano95l, NSMO98]. crowded [FEG77]. CRT [Bau80, Dou75, EH78, Lan75, Leo77, NSW76, Wai75]. Crystal [Ano95-29, CK76, Maz80, Ste81]. Cu [HH90]. Cube [Lam95]. Current [Bec76, DiF75, Emm72, IK69, KGK81, Lou75, Pee73, PP69]. Curve [Adc87b, Ano93-42, BD84]. Curve-Fit [Ano93-42]. Curve-Fitting [BD84]. Custom [Ano90a, BHW86, Bur95, Dic84, FHMP97, Hea85, PK84, PL84, Tha89, BP83, Lyn80a]. Customer [Cam94, Cot96, MB85, Reh97]. Customer-Driven [Cam94]. Customer-Oriented [Cot96]. Customers [BT98]. Customizations [Joy96]. Customizing [Cou94]. Cut [Ano92j, AK81, BW81]. cutter [APA92]. CW [SFO80]. Cycle [Cot96, Saf96]. Cyclic [Ano95-41].

D [EAW^{+}86, Mac93]. d.c [Pet70]. d.c.-to-v.h.f [Pet70]. DAC [Ano92-36, HMM73]. Data [Ano93-33, Ano94-31, Ano96f, Ano96t, Bar96, BKV89, Bia90, BFS94, BHK^{+}86, BYA86, Cam94, CS88, DN86, DU82, DSB94, Don86, Gen85, HGWS97, HSK87, Hur90, Jon84, KP88, Kal93, KYKM85, KRDS83, Lan94, Leo83, MLPN88, MB93, MCD97, MP83, NIWM98, NK78, Pie87, RP88, Sch96, Sha85, ST90, THHW95, WSYD97, Whi85, Wol81, da 94, AS90, And70, CLL81, CDE^{+}87, CS76, CPR75, CDS71, Dam74, Due74, EH78, Erd79, FGH^{+}89, Gue74b, Hal93, Ham93, Han79, HRH81, Het79, Hey81, HBNB83, Mai90, McI74, Mor76, Mun77, NTH^{+}84, Res83, Ric81a, Ric76, RBB78, RO86, SKWR95, SFW78, Sha89, SM75, Sta81a, Ste74b, Ste80b, SK69, TPG79, Wol75, YWP^{+}92]. Data-Driven [Lan94]. data-transmission [Gue74b]. Database [Ano93h, HKH93, AC93, Bet97, DD97]. Databases [Loo97]. Datacom [AP90]. Dataflow [Bee92]. dB [Hil93b, Jer74]. DBMS [Loo97]. dc [PP69, HK88, MiITT82, Vet91, Mer73]. dc-to-20-GHz [Mer73]. DC-to-50-GHz [Vet91]. DCE [Ano93-41, Cas95b, GH95, GLM95, KT95, Kon95, LH95]. DCE-Based [GLM95]. DDS [Ano94t, DSB94, Ujv94]. DDS-2 [DSB94, Ujv94]. Debug [Ano93k, GNR93, PDW85, Tan89]. Debugger [Ano94z, IGHG^{+}94]. Debugging [ADG^{+}96, Ano92i, Ano93-35, Ano94l, GNR93, McD80]. decentralized [Erd79]. Decision [Tan96, MS83]. Decisions [BKQW95, Gra96, Hig95]. Decoder [LBLS95b, LBLS95a]. decodes [Pru92]. DECT [Elo93]. Dedicated [HPM87, Ogd80]. Deep [NSMO98, FK79]. deep-level [FK79]. Deep-Submicrometer [NSMO98]. Defect [Ano92-32, HKH93, Bla86, DD97, NSA89, War89]. defects [BR91, War91]. Defibrillator [BJ82]. Defibrillator-Monitor [BJ82]. Definition [Ano93r, Ano96a, BF89]. Definitions [Ano92j]. deflection [BL87]. Delay [GBO94, Lam95, FH74, Gue74b, IKS81]. delays [FD78]. Delivering [LHIR98]. Delivers [HL85, CHM76]. delivery [GV84, SW90b]. Demand [ASTW95]. Demodulation [Ano93d, BVAP87]. Demodulator [Kaf88, Ril79]. demountable [Mat92]. Dense [TD92]. density [FL80]. Departments [Ano96g]. dependable [Gol80]. Deposition [DLJ85, Ili82]. Derivation [Sha98]. Deriving [Ste74b]. descriptions [Gil96]. Design [ABG^{+}92, ALRS89, AGS85, Ano92g, Ano92l, Ano92z, Ano93l, Ano93s, Ano94j, Ano94y, Ano94w, Ano95p, Ano95z, Ano95-43, Ano96j, Ano96-29, Ano97e, BJJN88, BBJ^{+}95, BKQW95, BBF^{+}97, BCH^{+}93, Bra96, BMS87, CLG87, Cau85, CR89, CHK^{+}96, Col88a, CSWR85, Cow70, Cra92, Cra85, DWB^{+}97, DD86, DR95, DW92b, DDT83, DVE85, DMJCB97, DS82, ELC^{+}86, Edw84, EEH^{+}84, EM98, FHM^{+}85, Fid96, FST92, FHMP97, FSB^{+}87, FPUB88, Ful96, GHUW68, Har86, HCHR91, HHT^{+}97, HMM96, Jac88, JJ85, KP88, KGH^{+}87, LS87, Lon92, Mai93, MSS87, Mar88, May88, MTWG85, MH89, MFW87, MKW^{+}96, Mur97, NG87, Oh82a, OR96, Pea85, Pet96, PBD88, Pon80a, Reg88, RSM91, RMVT95, Rya95, SDA81, SPD^{+}93, Sim94b, SBB85, SGD83, Smi84b, ST90, SC89, SM86, Ste83, SHB^{+}82, TCWN88, TM91, WP94a]. Design [WYSE97, WMZ84, WJ80, Wor92, YF92, BLMJ77, BBDK93, BGG88, Bor90, BM81, BSY89, Bro95, Cam82, CJHE93, Car81, Cho90, CWF^{+}79, CGH^{+}93, CBH91, DSC^{+}90, DeV81, DR81, FR95, Fel74, FFS89, Gil81, GL88, HJH^{+}97, HKZ74, HG81, Hig88, Hor79, JKP^{+}88, JS93, JM93, JU97, KT81, KSV78, KPC^{+}82, Ker80, Kil94, Kil97, Kre92, KSMM97, LH82, LM93, MC85, MEM82, MSCK92, MP83, MRRS93, MOT90, Neg81, NS88, OB90, Osb68, OGB84, PCW77, PGW73, PPJ91, RH92, Roe95, Ryt77, SK80, SCI^{+}81, SF95, SMB^{+}91, SSMM81, Sti77, Tal88b, TBBM83, Top89, TC79, TH81, Vob88, Wag75, WBSC84, WP94b, WJ82, Win81, Wya92, YN93, RM88]. Design-for-Testability [DR95]. designed [BB80, CS80, Dic78, FCC79, Gra81, HVSM96, LK80, MO74, Sha76]. designer [Kas72, NH92]. Designing [Bec79, BL83, Cos76, GS93, HCW^{+}82, HLT^{+}83, Mei82, MPR82, PB76, Sch78a, Sha78a, Swe81, Tag74, TOB97, Wal76, WST84, WS84a, Wat81]. DesignJet [Ano92k, Ano92l, Lon92, MSCK92, PT92]. designs [FHDS92]. desk [GH73]. DeskJet [BCK^{+}94, BTE^{+}94, CAJ^{+}94, PPMW94, Rho88, DMW94, Gas88, HW88, JKP^{+}88, MH92, MLPN88, MJKS92, RL92]. Desktop [ABW78, BBD^{+}84, CBL96, Jon81, Joy96, Pea85, Bak81, BFMP84, Chu79, EW78, FK79, FHDS92, Hal79, US79]. DeskWriter [ACM92, MH92, MJKS92, RL92]. detaching [SB84]. Detailed [Ano97b]. Details [MRH94c, MRH94d]. Detection [Ano92p, Ano95h, Ano95-42, Gal84, Ano90b, DD97, Kal95, KW80, Ran88, TW76, WDDH84]. detector [But95, CRSY73, GTM95, Kal95, LNH^{+}84, Smi73, YPS76]. Detectors [Ano93z, Bec87, Uph88, KBHS93, MCM93, WTB^{+}90]. determines [HFOS81]. Determining [Pet96]. Developing [Ano93m, Cox96, CHNJ97, DL97, Ful96, GC90, Hun92, KF95, LH95, MW78, Reg88, ST83, Smi94, Spe96]. Development [AS85, Ano91b, Ano92h, Ano94u, Ano95m, Ano95n, Azm82, BJJN88, BLTY81, BCK^{+}94, BLNP86, BA85, BEM95, Bur95, BHH^{+}88, Cam94, CL98, Cou94, CT85, DSW83, Dra77, FG87, GSFF90, HBB88, Jan96, JM91, LC96, MZ96, MM97b, Mil96, OSN^{+}85, OR96, PSS^{+}88, PL81, PT92, ROKH94, Rya95, STMM88, TE86, War92, ACM92, BH92, Ber88, Bri94, BFP90, Cle79, CG87, CBH91, DeV80, DGM80, Fel80, FJ89, Ger90, Goo91, GCDW91, Jes89b, KSW89, Kra89, LF91, Lie97, LD93, LFWEJ89, MJKS92, MMPT97, Mur82, Nie85, Rho88, SK80, Wil97a]. Device [GSFF90, Oh82a, Tem83, Gue74a, Kir71, RF72, Rom91, SMK84, TST82]. Device-independent [Tem83]. Devices [Col88b, Dia94, ES87, Har84, IC85, Dia81, ES89, Pli89]. Diagnose [GV86]. Diagnosing [Pet94]. Diagnosis [Ano92-32, AM95, O’C86]. Diagnostic [AD82, Lon90, SS83b, Hol82a]. diagnostics [Hol79, LaF80, WS84b]. Diagram [Mil94]. Diagrams [Ano96h]. dictionary [Rob92]. Dielectric [Wak97]. Difference [DR98]. Differential [MD96]. Difficult [OH82b, BCO^{+}93]. Diffraction [Ano93n, WK93]. diffused [Bec93]. digit [GSVS89, Ste89b, Tho72]. Digital [Ano93a, Ano93q, Bal84, Bir91, Bir97, BGS83, Byr96, Chi81, DWD88, DO77, Edw75, EL82, Gar92, Gro91, HDH96, HPM87, Hil93a, HKM96, Hol84, gJH81, HLH86, Kah92, Kjo92, KV94, Lau87, McK87, MB93, MD96, Mil94, MM95, MKW^{+}96, Mon88, Mur92, MPR82, NVBH87, OS75, PK84, PH96, Ped86, Pet96, PBD88, QH83, RK94a, Rot70, Sha85, ST84, Sor88, Tal88a, TOB97, Toe97, Ujv94, Van86, Wai93, War92, War97b, WP94a, AH72, BLMJ77, Bes75, BVAP87, Car69, CCH^{+}91, CR89, CM72, CPK88, CRSY73, Cze89, Don69, Don71, DL73, ES77, EH78, Erd79, FL73, Far74, FCC79, Fro77, GDV73, GY88, Goe89, Goo72, GW77, Gro92, Gro79, GT71, Gun94, HMZ82, HT76, HBB88, Hil93b, Hoo91, HBNB83, Ing76, KH69, LS76, Lou72, Mae74, MB74, MW78]. digital [NSV92, PK75, Par79, PCW77, PB76, RHM79, Rie94, REOB78, SIBH88, SWM78, Sch78a, Ste89b, Swe89, TD83, Wat87a, WB83a, YPS76]. Digital-to-Analog [Gar92, SIBH88]. Digitally [Ano93o, Pee73, Lau80]. Digitally-controlled [Pee73]. digitizer [Car78]. Digitizing [BMD^{+}88, HT92, Mon88, PF88, REB86, Sch93a, Sch86, Toe86, Wit92a, Chu88, Cze89, EKW93, Mil88, Nic88, RO86, Rus93, Wit92b]. dimension [MHCH77]. Dimensional [WB83b, FR95, HH90, QW83]. Dimensioning [WYO^{+}87]. dimensions [BHW77]. Diode [EJG86, ZFE^{+}86, Kal95]. Diodes [BKAWBCTK98, DBB^{+}95, FKO^{+}93, SDC^{+}88, Bec93, But95, Cow70, LM73, ZC69]. Direct [IK69, YN98, SWP90, Vif76, WK93]. direct-drive [WK93]. direct-reading [Vif76]. Direction [Nis87]. DirectModel [CG98]. director [MW80]. Directory [CFK^{+}90, KT95]. DIS [GSFF90]. disassembler [AS90]. Disc [BDM85, BJWW84, BK85b, DGHP85, DLJ85, Edw84, HL72, MG85, OSN^{+}85, Smi84b, BB84, Bow72, GO72, HB72, Smi79, Sti77, Voi84, Wil84]. Discless [HM88, SMQ88, Wag88, Wan88, BT88, GL88, Ran88]. discrete [Mil75]. discriminator [AP76]. Discs [AGS85, HRE85, MSB85]. Disk [Gal84, SR95, DSC^{+}90, OB90, SH90a, SWP90]. Dispersion [GLP^{+}97, JR84, HH95]. Display [Ano95-29, Ano95-43, Ano96l, BB80, Dul85, ES77, Gra81, HPM87, Hig85b, HFKS95, LW82, OMD83, SAS98, SD82, Wit92a, ACL89, Bau80, BCH^{+}89, HG79, HMZ82, Kin81, LS76, MC85, Maz80, MFPE72, Nic88, Por80, Pra80, RF72, Roy75, SB84, Sta81a, SA80, SW81, UAH77]. displaying [Far74, LHL83]. displays [BP69, EH78, How79]. Dissecting [Gra89]. distance [MS80, Smi80, SB80]. Distortion [Ano93b, Oli71, CF68, Dre74, FH74, Gue74b, Pon80a]. Distributed [Ano95-34, CHK^{+}95, Dic74, KF95, LMNQ96, LEV96, Sak78, Sha78b, Ada71, Cra92, GC90, Gue79, HB88, Sch78b, BGK^{+}87]. distributed-thin-film [Ada71]. Distributing [GAL^{+}93]. Distribution [Ano94y, WP94c, LE83]. disturbances [DR81]. Dither [Ano93p]. Dithering [Ano93-47, Toe97]. Divider [Ano93j]. DMA [Ano90a]. Dmi [NDS86]. Dmi/3000 [NDS86]. DMM [HMM73]. Do [Ano92-34, Sch70]. Document [Gor82, WHBP95]. Documents [Moa98, RA85, SJ94]. does [ABW78, BJM71]. Domain [Bir96, Cor96, MS95, MD96, BGG88, Bel93, BP93, FRS88, KH88a, KBHS93, Lin69, RP88, SNN88, SM75, Smi69, Unt68, Vob88, War72, Cha91]. Doped [LSM^{+}95, Sti97]. Doppler [Che86, HT86, HLH86, Kar86, Kee89, Mag86, O’C86]. Doppler-shifted [HT86]. Dot [Cra85, DiV84, PL84, ISLW78, Jep85]. Dot-Matrix [DiV84, PL84, ISLW78]. Dots [Ano92-34, MLPN88]. double [WA93]. double-pass [WA93]. doubler [Zel91]. doubles [Rig77]. down [Cra72, DH76, Due74, SWM78]. down-converter [DH76]. Downtime [AD82]. DQPSK [Bir91]. Drafting [Ano94x, Has92, APA92, BAUR92, BSME92, GAP92, MSCK92, PL81, SSMM81]. drawing [BAUR92]. Drive [BJWW84, BGS83, CDG^{+}88, CT85, DVE85, Edw84, Gal84, Gil89, Gre85, HL72, RRC^{+}85, Smi84b, Top89, Ujv94, BKV89, Bow72, DPA^{+}88, FS81b, HB72, Hol82c, KT81, Mey90b, Nie90, PHK77, Ric97, SH90a, Smi79, SS90, Sti77, WK93, WJ82]. drive/computer [Bow72]. Driven [Cam94, Lan94, Sta88]. driver [ACM92, CDLN96, HJH^{+}97, Kil94, Tan89]. Drivers [Ano93-37]. Drives [BDM85, Wil84]. Drop [Ano94r]. Dry [Mar82]. DSEE [Lub91]. DSP [Ano93q]. DTMF [Ano95h]. Dual [ALT82, JBS93, Mur82, RD75, Sta81a]. Dual-output [JBS93]. Dump [SMH86]. DVM [JRC81, WJT71]. Dynamic [Ano93b, EH84, EEH^{+}84, Eps84, FPUB88, HRE85, MRH94e, PK84, SA82, WSBK83, Yan92, BS93, Bol80, Cat82, FR95, FRS88, JRC81, Tal68, Van77a, Van77c, Wad68]. dynamically [McK83].

E-series [Lan77, Rig77, Sta77]. E5200A [Ano96j, Ano96k, Wri96]. Early [FS94]. ease [Dic78, GW77, Suk84]. eases [WB83b]. easier [Har76]. easier-to-use [Har76]. easily [Fos74, GJ74]. Easy [Bac91, Car78, Cha77, DH77, BBT92, CP79, HM71, MK79, RBB78, SKB78, SDS83, ULW81, Wol81]. easy-to-service [MK79]. Easy-to-Use [DH77, Bac91, Car78, Cha77, CP79, RBB78, SDS83]. ECG [BDS85, DV85, GG77, Gro91, LDNT72]. Echocardiographic [Pop83]. ECL [ALT82, HR82, Let92]. Economic [Hig95]. Economical [ELN88, KV74, FG73, MB80, Zel79]. Edge [CHNJ97, DBB^{+}95]. editor [Wil97b]. EDP [ARGA85]. Education [BPMF87, Reh97]. educational [HMPF74]. Effect [Ano94y, HH90, Mag86, DR81, RW81]. Effective [Ano92f, Ano97a, GLW85, Hal82, Kir71, NSP97, AC90, BM79, Cla81b, Dou75, Gro92, LDNT72, MD79, SM79]. effectiveness [Hor79, HC75]. Effects [Ano97e, DR98, Pli89]. Efficiency [FKO^{+}93, Car78, DP73, PB76]. efficient [BP77, Bow72, RC79]. EFL [WJ79]. EISA [TBTM92, Ano91c, Ano91d, Ano92e, CL92, CLL^{+}92, RTD91]. Electrical [AGS85, BBR^{+}97, Dia94, Fid96, Kha81]. electrocardiograph [DKMS81]. Electroluminescent [Hig85b]. Electromagnetic [Wak97, Bec79, Won79]. electromechanical [BMTW78]. Electromigration [Mer82, MMNR89]. electron [BOOS81, CLL81, EHS81, Ham81, gJH81, KGK81, LM81, NE81]. Electronic [ACCM83, BMS87, BW76, Bur95, Ful83b, GFK83, HSMA^{+}86, MHKS94, MSCK92, Smi80, GR91, HBHC84, Hor69, Mon68, NR72, Rob92, WRT72]. Electronics [Ano94i, AT83, BLSM83, CDG^{+}88, Gor82, GLW85, GDKK94, HSW88, KE85, OS76, Sha85, BB84, BBW95, FS81b, Lan83, Lo81, MB74, MEM82, MD79, Ric97, RC79]. Electrophoresis [Ano95j, Hol95, RWW95, SD95, Zim95, GTM95, Kal95, Sch95, SMD95]. Electrostatic [SS82]. Embedded [GNR93, KA95, BNB93, LD93]. Embedding [GUW92]. EMC [Ano96j, Wya92]. emergency [GG77]. EMI [Hoe97]. emissions [WB89]. Emitting [BKAWBCTK98, DBB^{+}95, FKO^{+}93, THHW95, SDC^{+}88]. Emulating [DGH^{+}97]. Emulation [Ano93-29, BKMM87]. emulator [Fis88]. Emulators [DGM80, RR83]. Enable [BT98]. Encapsulation [Cro89]. Encapsulator [Fro90]. Encina [Gup95]. Encina/9000 [Gup95]. enclosure [Inh75]. Encoder [Ano93a, ELU81, Her73, WJ82]. encoders [ELN88]. encroachment [Chi82]. End [Ano92k, BMD^{+}88, SM86, GJM^{+}92, Kri81a, PCW77]. end-tidal [Kri81a]. end-tidal/respiration-rate [Kri81a]. engineer [VDF72]. engineered [Ham79]. Engineering [BP82, CL98, Col86, Hea94, Kel94, Lon92, SKK^{+}86, SGD83, TLCB98, BR91, BH92, BKS93, Mat79b, SB84, SW90a]. Engineers [RA85]. Enhanced [BDFL95, DWB^{+}97, Ili82, LBLS95b, Lym87, MRH80, ABD^{+}97, ACK79, BN80, Bol80, Hef73, LTC^{+}84, LBLS95a, SH77]. Enhancement [Ano92x, BKAWBCTK98, HSB^{+}94]. Enhancements [Ano95-39, FO95, US79]. Enhances [Duf87, FH80, Geb78, Hor79, Lam74, Mey84, REOB78]. Ensured [Gen85]. Ensures [AGO^{+}85, Sch84]. Enterprise [Ano94m, Ano94n, Jen98, MRH94e, Sel97]. Entry [SD78, NBZ^{+}97, NK78]. entry-level [NBZ^{+}97]. Envelope [FH74]. Environment [Ano91b, Ano92w, Ano93f, Ano93k, Ano93-39, BKQW95, Buf96, CHK^{+}95, Car95, CBL96, DKW90, Fer94, GNR93, Kon95, LMNQ96, Man94, MS94, Pet94, Ric94, TOB97, WPS94, Wor92, Yod94, BH92, Bha92, Bla86, Cag90, CW92, CFK^{+}90, Cro89, Dra77, Ful89, Gas88, Gor91, Lie97, Mac93, Mat84, Ric91, Rob79, Sho89, SW89]. Environmental [Ano94y]. Environmentally [Ano94o]. Environments [LEV96, BDF84, FGH^{+}89, GG77, HB88, HHS76, Wil97a]. Epoxy [TLCB98]. equation [Kah79, McC87, Vog91]. Equipment [GDKK94, Kar94, MHKS94, RL92]. Equivalence [Fos98]. Erbium [LSM^{+}95, Sti97]. Erbium-Doped [LSM^{+}95, Sti97]. Error [Ano90b, Ano93r, Gal84, Kar94, Kat90, Mey90b, CRSY73, Due74, HH91, Raf93, YPS76]. Errors [Ano92p, McD84, Sha98]. ESCA [KT73]. Estimating [FS94]. etching [Mar82]. evaluates [Kah80]. Evaluating [McK87, Haa78, YPS76]. Evaluation [Ano94-29, Har87, MBL95, Ohm67, Rus92, Sch97]. Even [Van98]. Event [IGHG^{+}94, MRH94b, She96]. Event-Based [IGHG^{+}94]. every [DG70]. Evolution [Cas95b, Lie97, Lou72, WP91]. Evolutionary [Cot96, MZ96, QM80, Wic87]. evolvable [MK97]. Exacting [CS82, SAMS73]. Example [Ano95-46, Fel74]. Examples [Ano92m, Nad77]. excellent [BS93]. Exception [Ano95n]. exchange [SKWR95]. Execution [FS86, HM88]. executive [Ave75, SH77, HKC^{+}84a, Rob79]. exerciser [Ike83]. expand [HMM73, US79]. expandable [SMB^{+}91, WP91]. Expanded [LLS88, LB80, NSW76]. Expanding [Tho77, WN79b]. expands [BW72, JS74]. Expectations [Ano91g]. Experimental [GLP^{+}97]. Experimenting [Wil78]. Experiments [MRH94c, MRH94d]. Expert [But94, Ros88, SMH86]. Explanations [MRH94d]. exposure [Bre74]. exposures [NE81]. Express [AW90, Lon90, BE90, Bor90, Joh90, Pug90, SH90b, ST90, Tal90]. Extend [Uph88]. Extended [EDF82, SMR^{+}86]. extender [Gue79]. Extending [Chu79, EJG86]. Extends [EMT83, Kal93, Cat78, DiP80, GP84, MFPE72]. Extensible [Ano93m, PW89, TBL^{+}83]. Extension [KA95]. Extensive [Ban84, DSW83, Hol84]. extensively [DiF75]. External [MRRS93, RLB^{+}95, SL93, GSVS89, Roe95, TZ93]. External-Cavity [RLB^{+}95, MRRS93, SL93, TZ93]. Extraction [HT86]. Eye [Ano96h, Mil94]. Eyeline [Ano96l].

F [Cat78, Geb78]. F-Series [Cat78, Geb78]. F8 [Ogd78]. Fabrication [Bec93, DLJ85, BC81a, Slo89]. Facilities [WPS94, Joh77b]. facility [Dys89, LFWEJ89, Nel79, SW89]. Fact [Ano96t]. Factor [DSB94]. Factory [GSFF90, Jen98, Ric81a]. Factory-Floor [GSFF90]. Failure [Gra96]. Failures [Hen84, WR86, BP72, Gra89]. families [Que76]. Family [BDS85, BK85b, DR95, DiV84, DGN98, ELC^{+}86, Kru94a, Kru94b, MTWG85, Rub86, WR97, Bru77, Cat78, Dou75, DP84, EDF82, FM84, FR87a, FRN83, HPF93, HLT^{+}83, HKC84b, Jep85, KNF^{+}97, KB85, MJKS92, MP89, MBH^{+}89, ORB91, Ogd78, Ste74a, WB89]. farad [OS75]. Fast [FHMP97, OH82b, Ujv94, WCH94, Aue81, ABW78, Cha70, GO72, Goo77, HA92, HMZ82, Has81, MH89, Mor70, Pon80a, Yan93, Zel70]. fast-reading [Goo77]. fast-settling [Pon80a]. fast-switching [Has81, Yan93]. Fast-Writing [OH82b, Cha70]. Faster [NTA72, HB72]. FASTRACE [Bas79, Gro79]. Fault [Lam95, SR95, AM95, GH73, Gro79, Low69, TW76]. Faults [Ano95g, Bec76, Erd79]. Faxing [SJ94]. FDDI [BKD94, Bha94]. Feature [CHP85]. featured [Mye96]. Features [AFG^{+}86, DR95, Ste83, Wen85, FK84, Lan77, SH90b, Str77]. Federated [Mar95a]. Ferrite [NML90]. ferrites [Nic90]. FET [MP81, WWS^{+}98]. Fetal [BJPH95, CRSS77]. few [HJW75]. Fiber [Ano95-47, Ano95-36, CHNJ97, ES85, LSM^{+}95, RWT97, Sti97, WSYD97, BP93, FR87a, FR87b, Han79, HH90, Hen91, KW93, MR91]. Fiber-Optic [Ano95-47, CHNJ97, RWT97, WSYD97, Han79]. Fibre [CDLN96, KBHS93, Pri96, SP96]. fiducial [gJH81]. Field [LGV^{+}98, ARGA85, BDOT85, Fro77, MRM83]. Fields [Ano94-30]. Fighting [Ano95o]. Figure [SC83b, IK69]. File [Ano96a, Man90, BT88, KGTM90]. files [TPG79]. Filling [ST75]. Film [AGS85, DGHP85, DLJ85, HRE85, MSB85, OSN^{+}85, Ada71, BA85, Mer73, Mil75, Toe86]. film/semiconductor [Jac74]. Films [Ili82, Zam73]. Filter [Ano93s, Hil93a, gJH81, Lau80]. Filtering [Mon88]. filterphotometric [WDDH84]. Filters [NML90, PK84, Wai93, IKS81, Nae79, Sch78a]. Final [SW91]. finance [CR73]. financial [Mar77]. find [Bec76, She71]. finds [BP72]. Fine [MM72]. Fine-line [MM72]. Finite [DR98]. Finite-Difference [DR98]. Finstrate [MLSW83]. Firmware [Ano92q, Ano92-33, Ano93m, Ano94j, Ano95p, BP85, BFP90, DCN^{+}88, FG87, GLW85, Hal85, Han80a, HDH96, HFOS81, JM87, KSW89, Kla78, MFW87, Moy78, NS88, RRC^{+}85, Sap77, Tal88b, Wai75, WW84, Wor92, BRHS89, BBW95, FJ89, JF80, Kil97, Lei71, LD93, MSCK92, MP83, O’B93, PPJ91, SW81]. First [AGO^{+}85, FSB^{+}87, Dou75, Maz80, PN88, WCT87]. first-generation [PN88]. First-Pass [AGO^{+}85]. Fit [Ano93-42, Pet96]. Fitter [Adc87b]. Fitting [BD84]. Five [Gen85, Tho72]. five-digit [Tho72]. Fixed [Moa98, Smi79]. Fixed-Rate [Moa98]. fixture [Ham93]. fixturing [DMW94]. flame [Smi73]. flame-ionization [Smi73]. Flat [Ano95-43, HFKS95]. Flexibility [Gor82, BP76, FCC79, Mey84, Suk84]. Flexible [Hal82, MBC95, Ste81, SW90b, da 94, Fel80, HL86, SA80]. Flip [BB97a, TBB^{+}97]. Flip-Chip [BB97a, TBB^{+}97]. Floating [MC84, Pon80b]. Floating-Point [MC84]. Floor [GSFF90, Jen98]. Flow [Sch75, SDDC96, ACCM83, CLL81, Haa78, HT86, MM90]. Fluid [NH94]. Fluxes [VCMH96]. Fly [Kel94]. FM [CK76, McH87, Ril79]. Focus [Ful96, McK83]. fold [MM72]. foot [DG70]. Forecast [Sha98]. Form [DSB94, Far74]. Formal [CBH91, LF91]. Format [Ano94x, Has92, BSME92, Hal93]. Formation [SS82]. Formats [Ano94-31, Ano95y, BFS94]. Formatting [Sha85]. Formed [Ili82]. Forth [Ano92n, Mil84]. FORTRAN [GMMM92]. Four [ACK79, SBK^{+}97, BLKH77, PHK77, SW81]. Four-color [ACK79, BLKH77, PHK77]. Four-Way [SBK^{+}97]. Fourier [Kis70, McK75, Rot70]. Fractional [Ano93o]. Fractional-N [Ano93o]. Frame [Ano95y, Dub94]. Framework [GLM95, Raj95, CG87]. Free [BDFL95, Ter78]. Freeform [ME95]. Frequencies [Eps84, Ks69, Per86]. Frequency [Adc87a, AS85, Ano92o, Ano93j, Ast77, AWH^{+}87b, AWH^{+}87a, BS95, CSWR85, FHM^{+}85, Hoe97, Lin69, Mar71, PH96, San86, Ste88, Yon94, ALRS89, AH89, Ban69, BS90, Bau69, BVAP87, Bol74, Bol80, BP70, CDE^{+}87, Cha70, CM72, Cos76, GW77, Gro92, HW73, Haw71, HHS76, Hei88, HMO73, Jek75, KB69b, Kei71, KS72, Koe80, LM73, Mac73, MN79, MO74, Mor70, NYS81, NS88, PF78, Rus92, Say80, SFO80, Sei81, SGC93, Sor74, Spa85, SS89, Thr68, Unt68, Urq72, Vif76, Wai93, WKT83, Wec88, Wha68, Zel79, Zel91]. Frequency-domain [Lin69, Unt68]. friction [Ter78]. friction-free [Ter78]. Friendly [Ano94o, BHP85, Car79, MK79, Pee79a, TBBM83, WS84b]. Front [BMD^{+}88, PCW77]. Front-End [BMD^{+}88, PCW77]. Fuji [Saf96]. full [FG73, HS80, Mye96]. full-featured [Mye96]. full-scale [FG73]. full-spectrum [HS80]. Fully [Dul85, Lin97, Unt68, DF82, Gor80, Kei71, Tun74]. function [BF79, DF79, FRN83, GP84, Hef73, HLT^{+}83, Pie75, RD75, SMR^{+}86]. Functional [JU97, Kas72, MGBB97, RHM79]. functionality [BLY92]. functionally [Lan75]. Functions [Ano94-28, EMT83, Egb77b, Egb77c, Egb78, GL88, KH69, SBM85]. fundamental [SGC93]. Fundamentals [Ped86]. further [Kru89]. Fused [LRS83]. Fusion [Ano96u, Ano96-27, Cot96, Hol95, DL97]. Future [Van97]. Fuzzy [Kru94a]. fx [SOG98].

G [YWP^{+}92]. G-link [YWP^{+}92]. G1009A [MB96]. GaAs [FST92, RNR^{+}98]. gain [Han72, NYS81, NTA72]. gain-phase [Han72, NYS81, NTA72]. GaInAs [RNR^{+}98]. gains [O’K93]. Gallium [FKO^{+}93, Gib86]. gap [Fro90, LM93]. gaps [ST75]. gaps-modular [ST75]. Gas [LRS83, ACCM83, DJS83, Row83, Smi73]. Gate [BBF^{+}97, RNR^{+}98]. Gate-Length [RNR^{+}98]. Gbit/s [GLP^{+}97, WSYD97, CDLN96]. GDMO [Ano96a]. General [GT71, KM97, Ogd80, ADRS83, BLMJ77, CF68, GOS84, MC91, OGB84, Sha76]. General-Purpose [KM97, GT71, Ogd80, ADRS83, BLMJ77, CF68, GOS84, MC91, OGB84, Sha76]. Generalization [KH88a]. generated [Won79]. generates [FD78]. Generating [Don69, EH78, NMY95]. Generation [ABC^{+}94, Ano92y, Ano94r, Bal84, BGR^{+}88, BKO87, CHK86, DR98, Edw87, McD84, Sim94a, ACS88, BP83, BB84, BCZ93, Cag90, CP79, Fro90, Ger90, GHUW68, HHS^{+}73, HS68, HH78, KT73, KMT88, LS83, MCCB76, PN88, Par79, Ric97, Sha71, Sor74, SP92, Tho77, WTB^{+}90, DNC96]. Generator [Ano93-27, CSWR85, FHM^{+}85, GC87, HSK87, HZ83, JM87, JJ85, Kal93, KWR^{+}87, LK85, Lau87, NVBH87, Ton85, Wen85, Ain81, AH89, Ast77, Aue81, BBI^{+}93, BKS93, Bur85, Chi81, CK76, CRSY73, DF79, Don71, FL73, FFS89, GZC74, Hal75, HK75, Has81, Hay74, Hef73, HRZE77, HLT^{+}83, HBNB83, HKA79, KS72, KW90, KE74, Lar83a, PK75, PB81, PB82, Ras87, RHM79, RD75, SCI^{+}81, SES90, SPD^{+}93, SAMS73, Str77, SS89, Wag93, Win80, WB89, YPS76]. Generators [CS82, BKE^{+}90, FRN83, HPF93, Kih82, MP89, Pag71]. Geometry [RO97, WYO^{+}87]. Get [Ano92-34, Ban69, Sch70]. gets [GT71]. GHz [Ano93-45, BBI^{+}93, Bol74, Bol75, Bol80, DVl91, EJG86, HK75, HH91, HHS^{+}73, HZ83, HE75, HCHR91, KN93, KCP77, Koe80, Kuh76, MCM93, Mer73, NML90, RO86, REB86, San72, Say80, SA91, SPD^{+}93, Sch86, SFO80, SGC93, Toe86, US93, Vet91, Vif76, Wag93, WHH91, Zel91]. gigabit [SP96, YWP^{+}92]. gigabit-rate [YWP^{+}92]. Gigabyte [HGWS97]. Gigabyte-per-Second [HGWS97]. Gigasample [MB93, Sch93a, CPK88, Mil88]. Glitch [Ano97a]. Global [Ano97c, Bur89, Kus96]. Globalization [Tiv91]. Glossary [Ano93t, Ano94q, Ano94p, Ano95q, Ano95r, Ano95s, Ano96i]. go [MW73]. go-anywhere [MW73]. graduations [DG70]. grain [Woo78]. graphic [ABW78, RF72]. Graphical [CAV69, DD86, DF90, Fer94]. Graphics [Ano94q, BTBW83, BM86, BHW86, BL83, CG98, Cun98, DDT83, HASP94, KPC^{+}82, LB98, Mal87, Mar95b, MTMP83, MC82, Pea95, Sch86, SOG98, SW86, STMM88, TCWN88, Tha89, Van86, Azm82, BB80, Bar81, BR78, BM81, BSY89, Coz78, Dic78, FH80, Hea85, Kin81, LWBB84, Moy78, Pra80, SB84, Sta81b, TBL^{+}83, Tem83, TBBM83, TWCD93]. Graphing [BBD^{+}96, BPA^{+}94a]. Grating [Ano93n, WK93]. Gravity [Ano93u]. Gravity-Sensing [Ano93u]. greater [GDV73, Inh75]. Greedy [Kru94b]. Grid [TLCB98]. grip [KT81]. Group [Ano95-33, Ano95-34]. Growth [CGS^{+}86, Kru88, HW82, Kru89]. Growth-Oriented [CGS^{+}86]. GUEST [HR82, Mei82]. guided [KSMM97]. guidelines [Nad77]. guides [Jon89].

Half [CT85, Gre85, Sha85, BKV89, DPA^{+}88]. Half-Inch [CT85, Gre85, Sha85, BKV89, DPA^{+}88]. hand [DL73]. hand-held [DL73]. Handheld [BMS87, Dic84, Kah80, LS87, Lin84, MC82, Rip87, ULW81, Wol83, GP84, HDH^{+}87, McC87, Mey84, Mil84, Pat87, WAZH84, Wec87]. Handle [Pea85]. handler [Tal90]. Handling [Ano95n, NZ97, Zim95, HPH81, LM81, Raf93]. Handpulled [AT83, HSYJ83]. hands [Edw75]. hands-off [Edw75]. Hard [Ano92l]. Hardware [Ano93-29, Ano93-37, CP87, CWF^{+}79, DVE85, EEH^{+}84, FSB^{+}87, GE86, JAD83, KNTV96, LB98, NG87, Osb68, SOG98, SMB^{+}91, SHB^{+}82, SW81, WR86, WBSC84, BBDK93, BB73, BM79, CGH^{+}93, Gil96, KSV78, Rob79, Ste88, WH91, WB83a]. Harmonic [Ano92p, BB78]. harsh [BDF84]. Harvesting [DL98]. Hatching [WYO^{+}87]. Haul [GLP^{+}97, BP93]. hazards [Con93]. HDL [Lin97]. Head [BJWW84, Ste83, AMSW85]. head/tape [AMSW85]. Header [Ano94-30]. Heads [HMSFR87, Say80]. Healthcare [SSW96]. heartbeat [Hor69]. Heat [BB97a]. held [DL73]. Help [CHW96, Smi94, WCH94, SW89, Sta88]. Helps [Ped86, BLKH77, Kas72]. heterodyne [Bol74, BB78]. heterogeneous [Wil97a]. Heterojunctions [Ano93-40]. Heuristic [Kru94b]. Hewlett [Hal79, RJH^{+}87, Cag86, Chu79, CHK86, Fle83, Lan83, Lou75, MBH^{+}89, PB87, RL86, Sch95, Sch78b, Sta87, US79]. Hexagonal [Nic90]. Hierarchy [RK94b, TE86]. High [AI84, ALT82, Ano92w, Ano92-36, Ano93z, BDH91, BGR^{+}88, BCRS88, BMD^{+}88, BEM95, BLSM83, Bro82, Bru97, BCF96, BL83, Cam94, CSS^{+}91, CS93, CWFS82, CT85, Cze89, DWD88, DR98, DBB^{+}95, DiV84, DSB94, DP73, EM76, Edw84, FL80, FKO^{+}93, FHDS92, FSDC82, Gar92, GSG^{+}87, GE86, Gor82, Gra96, Hal82, Han70, HM85, Hig85b, HT92, HJW75, HASP94, KYKM85, KRDS83, KBHS93, KSMM97, Kuh76, LB98, Let92, LRFL83, Mal87, MY95, MTWG85, MBD^{+}98, Mil90, Mil94, MM95, Mon88, Pea95, PBD88, RO97, Rie94, RNR^{+}98, Rub86, REB86, Sch93b, Smi69, Smi73, ST83, SA82, SA80, SMGR85, SS89, Tal68, TBB^{+}97, TD92, WM94, WHBP95, WR85, WHHC89, WHH91, WA97, Yan92, Yon94, Zel69, ALRS89, ASCC84, Azm82, BS93, BD73, BLTY81, Bau69, BBI^{+}93, Bau80, Bel93, BW85, BSY89]. high [BP70, Bur85, BHH^{+}88, Car78, CR89, Cha70, Chi81, CF68, Coz78, CRSY73, DL77, DiF75, EHS81, EH78, Emm72, EH81, ELN88, FCC79, FRS88, FS81b, For77, Goo77, GTM95, Ham93, HMW83, HKZ74, HVSM96, HMZ82, Has81, HE75, HW82, HBNB83, HMO73, JRC81, JM91, Kal95, KB71, Kam72, KN93, KGK81, KW80, Koe91, LT85, LM73, LNH^{+}84, Mac73, MVG77, MP89, MG92, Mor70, O’B93, ORB91, OS75, Pax70, PW82, PF78, Pir97, PB81, Pli89, Rad87, Ras87, Ric81b, RSM91, Roy75, RW76, SJ80, SA91, Sch73, Sch75, Sei81, Slo89, SD78, SC89, Spa85, Sta81b, Sti77, SMD95, TTML81, TD83, TC79, Urq72, WJT71, WK93, Wie90, WB83a, Wil84, Win80, WJ79, Yan93]. high- [LT85]. high-accuracy [RW76]. High-Capability [BLSM83]. high-capacitance [OS75]. High-Capacity [DSB94, Edw84, Wil84]. high-current [DiF75, KGK81]. High-Dynamic-Performance [SA82]. High-Efficiency [FKO^{+}93]. High-Frequency [Yon94, ALRS89, Bau69, BP70, Cha70, LM73, Mor70, Sei81, Urq72]. High-Interference [Ano92w]. high-isolation [Yan93]. High-Level [ST83, Pir97, SD78, Spa85]. High-Performance [ALT82, BDH91, BEM95, BCF96, BL83, Cam94, CSS^{+}91, CWFS82, CT85, DWD88, FSDC82, GSG^{+}87, GE86, Let92, LRFL83, MY95, Pea95, RO97, Rub86, EM76, FHDS92, KSMM97, REB86, Sch93b, ASCC84, BLTY81, BBI^{+}93, Bel93, BSY89, Bur85, Chi81, DL77, EH81, ELN88, FS81b, Ham93, HKZ74, HE75, HW82, HMO73, KN93, KW80, MP89, MG92, O’B93, ORB91, PW82, Pli89, RSM91, SJ80, SA91, Sch73, SC89, SMD95, TD83, Wie90, WJ79]. High-Power [BGR^{+}88, DR98, DBB^{+}95, Kuh76]. high-precision [BW85, Rad87].

High-pressure [Bro82, Sch75]. high-purity [Has81, PB81]. High-Quality [DiV84, Hal82, Hig85b, HASP94, WHBP95, WM94, Azm82, Bau80, FCC79, HMZ82, Sta81b]. High-Resolution [Ano92-36, Gar92, Mal87, CS93, Cze89, Smi69, BHH^{+}88, Car78, Coz78, Goo77, HMW83, Mac73, PF78, Roy75, WK93]. High-Return [Gra96]. High-Sample-Rate [Bru97]. High-sensitivity [HJW75, CS93, GTM95, Kal95]. High-spectral-purity [SS89]. High-Speed [Ano93z, BMD^{+}88, Gor82, HM85, KYKM85, KRDS83, LB98, MTWG85, MBD^{+}98, Mil94, MM95, Mon88, PBD88, RNR^{+}98, TBB^{+}97, AI84, KBHS93, Mil90, SA80, WHHC89, WHH91, BD73, CRSY73, EHS81, EH78, FRS88, For77, HBNB83, JM91, KGK81, Koe91, LNH^{+}84, MVG77, Pax70, Pli89, Slo89, TTML81, WJT71, WB83a]. high-stability [Ras87]. high-technology [Sti77]. High-Throughput [HT92, SMGR85, Rie94]. high-value [Kam72]. High-Volume [BCRS88]. Higher [Ano95t, Fos98, Jun78, MCD97, Wes77, Inh75, WKT83]. Higher-Level [Ano95t]. Higher-Performance [Fos98, Jun78]. Higher-Speed [MCD97, WKT83]. Highly [BDM85, BT97, EW78, WJ80]. History [Loo97, Mag86, Nie85]. hold [Lar83a, Ups71]. holding [Gue74a]. Homojunctions [Ano93-40]. hopping [MH89]. hospital [Fan72]. Host [Ber88, Fis88]. hot [EZ81, ZC69]. Housing [MHKS94]. HP [DNC96, ABC^{+}94, AC93, ABG^{+}92, AW90, ADG^{+}96, AS90, ACM92, AP90, AGI^{+}91, ABJ78, AFG^{+}86, AC90, Ano91e, Ano91d, Ano91f, Ano92r, Ano92s, Ano92q, Ano92t, Ano92u, Ano92y, Ano92-36, Ano93i, Ano93k, Ano93v, Ano93w, Ano93x, Ano94n, Ano94t, Ano94-28, Ano94-32, Ano94-38, Ano95u, Ano95w, Ano95x, Ano95v, Ano95z, Ano95-43, Ano95-44, Ano96j, Ano96k, Ano96l, AE71, AT83, Bac91, BWW84, Bam83, BE90, Bar95, BT88, BB73, BB84, BCK^{+}94, Bec93, BBE^{+}91, Bee82, Bee92, BBW95, Bel93, BKMM87, BKD94, BD84, BBR^{+}97, BTE^{+}94, BLY92, BS98, Bor90, Boy80, BFS94, BHK^{+}86, BGK^{+}87, BOOS81, BK85b, BKO87, Cag90, CP87, CLL81, CGS^{+}86, Cat78, CL92, Cha91, CHK^{+}96, Chi90, CS78, CLL^{+}92, CHKS86, Coc72, CAJ^{+}94, CR81, CFK^{+}90, Cor96, CR92, CHP85, COO^{+}93]. HP [Cun98, DMW94, DH77, DF86, DF91, DOR97, DF90, Don86, Don96, DMJCB97, Dou92, DP84, ELC^{+}86, Edw79c, FR95, FKL86, FM84, Fle83, FK79, FSB^{+}87, FP90, Fro90, FT86, Ful83a, Ful89, GWY94, GC90, GSFF90, GSG^{+}87, Gas88, GR91, GMMM92, GAL^{+}93, Gd86, GJ98, Gre90, GNR93, Gue79, Gup95, HG79, Haj96, Ham81, HW88, Har82, Har83, HW91, HR91, HDH96, HJG93, Hol79, HL85, Hor79, HM84, HWK95, Hun92, Hur90, Ish80, JS93, Jeg95, JM80, Joh90, JU97, Joy96, Jun78, KGTM90, Kar82, KF95, KG92, KHLM90, KM84, KH82, Kov79, Kre92, KT90, Kus96, LCB90, LL91, LP92, LT90, LS87, LCPG98, Leo83, LHIR98, LG92, Lin84, Lon90, LLS88, Mag94, MHKS94, Mai93, Mai90, MB85, MBF^{+}87, MGBB97, Man90, MKR95, Mar89, Mar88, MH92, Mat79a]. HP [MK79, MD86, MLPN88, Maz80, MJKS92, ME90, MSCK92, Met86, Mey90a, Mey84, Mil84, MB96, Moo92, Mor93, MOT90, Nad75, NMS80, NTJ^{+}91, Nea68, Not93, OY88, PSS^{+}88, PT90, PPMW94, Pat83, Pee79a, Pee79b, Pet74, PA97, Pug90, QH83, Ran88, RL92, RM88, RA85, Rho88, Rid90, Rip91, RH92, ROKH94, Ros88, SKWR95, SSW96, Sch95, SW91, SSV90, SMQ88, SH90b, She96, SJL91, Sho89, SBB85, SR95, ST90, SKG90, SMB^{+}91, SS83a, SH77, Spa85, SAS98, SW89, SS90, Ste89a, Ste80a, Ste98, Sto74, Sto96, STE90, SMK84, SJ94, Tal90, TPG79, TBTM92, Tiv91, TW90, TD83, Tre78, TRH85, Tuc95, TWCD93, Van98, Vog91, WLHC84, Wan88, Was80, WBSC84, WHBP95, WAZH84, WJ91, Wes91, WP91, WN79b, WMZ84, Wil84, WS84b, WB90, Wol81, Wol83, Wol86]. HP [WW84, Wri96, Wya92, WA97, WCT87, YF92, YLJ^{+}87, Yod94, Yor83, Zim95]. HP-18C [LS87]. HP-28C [LS87]. HP-35 [Coc72]. HP-41 [Wol83]. HP-41C [CR81, JM80, Ste80a]. HP-65 [Pet74, Sto74]. HP-71B [BD84, Don86, Lin84, Mey84, Mil84, WAZH84]. HP-75 [AT83, Pat83, Yor83]. HP-85 [Leo83, NMS80, Wol81]. HP-85/3054A [Wol81]. HP-GL [Tre78]. HP-IB [DH77, FK79, Gue79, Kar82, Nad75, Was80, WN79b, Wol86]. HP-IL [Fle83, Har83, QH83, SS83a, TD83, Yor83]. HP-PAC [MHKS94]. HP-RL [Ros88]. HP-RT [Ano93i, Mor93]. HP-SL [GR91, HR91]. HP-UX [AFG^{+}86, Ano92y, BT88, BLY92, Cha91, CHKS86, CR92, KGTM90, KG92, LP92, Mar89, Not93, OY88, Ran88, SSV90, SMQ88, SAS98, STE90, WLHC84, Wan88]. HP/CAI [AE71]. HPWord [Smi84a]. HQMOS [FSDC82]. HSG [KGTM90]. HSTL [EM98]. hub [Bro95]. Human [Ano92l, Mat79b, Sta87, TM91, FPS72, Ham79]. human-engineered [Ham79]. Human-engineering [Mat79b]. hundred [DG70]. Hybrid [Adl86, Dan96, Toe86, ZC69, FL80, HDH^{+}87, Mil75]. Hydrodynamics [AMK85]. Hydrophone [LGV^{+}98]. hyphenation [Smi84a]. Hz [CCH^{+}91, DA75, WD84, Yan92, Bol75, Tut72].

I/O [ABD^{+}97, Bee82, DN86, EM98, FK84, GJW83, LMM84, Mic72, NMS80, NTJ^{+}91, Nel82, SLM^{+}95, SSMM81]. IB [DH77, FK79, Gue79, Kar82, Nad75, Was80, WN79b, Wol86]. IBM [Ish80]. IC [Ano93j, BP72, Bak81, BKQW95, Ben95, Bro82, DiF75, FHMP97, FCK97, Gor69, Gro91, Ing76, Lo81, Mat90, OLR82, PL84, RHK^{+}83, SM82, Tal88a, Whi85]. ICs [NK84]. ID [Ano95f, Ano95i, But95]. IEEE [AT95, Ano95y, MCD97]. IF [Ano93q, Ano95p, ANSW95, CCH^{+}91, Nae79, Per86]. Ignition [Mag94]. II [Egb77b, HKC84b, Mil83, MRH94b, Sha76]. III [ABH^{+}83, Egb77c, MFMS83, MRH94c, YLJ^{+}87]. illumination [Bur89]. Illustrated [RA85]. Image [Ano94-28, MW94, SS82, SC83a, HLM89, SP92]. Images [Moa98, OMD83, LHL83, Mag83, Rus92]. Imaging [Ben95, Che86, JAD83, Kar86, KD83, MW94, MS94, Pop83, Ban83, Kin83, Lar83b, Mil83]. immunity [Han72]. Impact [Ano92b, DiV84, MTWG85, PL84, Raj95, SD95, BP76]. Impairment [DNG84, CDE^{+}87, NTH^{+}84]. Impairments [DGH^{+}97]. IMPATT [Cow70]. Impedance [Yon94, Bec76, Bot71, IOS80, NYS81, WKT83, Zel69]. Implementation [Ano95-39, DF91, GE86, HPL96, Pet96, Pug90, Mey90b, Sch78a, WLHC84, Win81]. Implemented [DR95]. Implementing [HSMA^{+}86, LC96, Mac93]. Improve [Ano93e, Ano94s, LeV88, AM95, GHK^{+}93, OST75, SB83]. Improved [Ano95p, BAUR92, FR74, Hig95, MW94, BW81, Edw75, FRS88, Rob82b, Sha76, Voi84]. Improvement [Gra96, Saf96, Sny75]. Improves [Gib86, EAW^{+}86, Her73]. Improving [Ano93y, BB97a, GCDW91]. In-Circuit [Har84, McD84, Cro79]. In-Line [DLJ85]. In-service [CDE^{+}87, PS81]. Inaccuracies [Ano95o]. Inch [CT85, Gre85, Sha85, BKV89, DPA^{+}88, Gil89, Top89, WYSE97, DSB94]. includes [Fin72]. incoming [Ing76]. incorporates [Koh72]. Increased [MH80, Gra81]. Increases [Neu84, Car78, Mut74, Reh84]. Incremental [Cot96, ELU81, KS72, Van77c]. independent [Ber88, Fis88, Tem83]. Index [Ano92a, Ano94a, Ano94b, Ano94c]. Indium [FKO^{+}93]. individualized [HRZE77]. Inductance [Das96]. Induction [Wak97]. Inductively [Kis97]. Industrial [Ano95z, Cle93, SB80, BDF84, BC81b, Cla81b, RBB78, Smi80]. Industry [LHIR98, WSJ87]. inertia [BLTY81]. inexpensive [KB85, ULW81]. Infared [MBD^{+}98]. Information [BPMF87, Bir97, DH98, SSW96, HT86, MHCH77]. informative [Bot71]. infrared [HWS87]. InGaAs [Ano93z, Slo89]. Ingress [Van98]. Inhibit [TLCB98]. Initiative [Ano96y]. Injection [KH84, Kat90, Sch95]. injector [KP90]. Ink [AMK85, Ano92g, EGH^{+}88, MJKS92, PSS^{+}88, BLKH77, Bru73, KB85]. ink-jet [KB85]. Inkjet [Ano92g, Ano92h, Ano92-34, Ano94s, Ano94x, BJJN88, BCRS88, EGH^{+}88, HSB^{+}94, Has92, HSW88, HASP94, WMBB94, ACS88, Bar88, BSME92, BCC^{+}94, Bri94, BHH^{+}88, DCN^{+}88, HCF^{+}88, KA92, Ric97, SMT^{+}97, WM94, ABC^{+}94]. innovation [WP91]. Innovative [Hor79, KM79]. InP [Ano93z, Slo89]. InP/InGaAs/InP [Ano93z, Slo89]. Input [Hei88, KNTV96, KM97, Mal87, BS93, DiP80, Ega72, ES89, HW91, HJW75, LD79, LG92, Mat79a, MS82, Mut74, Nic88, Pee79b, SMK84]. Input/Output [KNTV96, Ega72, HW91, LG92, Mat79a, MS82, Nic88, Pee79b]. ins [HEW82]. Insight [Byr96, Raj95]. inspection [Gil96, Ing76, Kee89, Mac93]. Inspections [BB91, FS94]. Installable [Bar84]. Instant [Mag94]. Instruction [Fia83, Geb78]. Instrument [Ano95z, BW92, BP85, EH84, Eps84, HKM96, Jeg95, Kjo92, Lai81, MiITT82, Nel82, SW86, SA82, Wol83, Zim95, ABM73, Boy80, DL73, EH81, GUW92, Gue79, Hal93, Hoo91, HL86, Inh75, JS74, Jon81, MHCH77, NH92, PB76, ST75, Smi69, SMD95, Tal88b]. Instrumentation [SR88, SHB^{+}82, TBB^{+}97, Wal70, HMZ82, KSW89, Mer73, MC91, Rad87, RS75]. Instruments [Ano92m, Ger98, Kel92, RSW^{+}86, WR97, DL97, FR87a, FK79, Jes89a, Kei71, LM86, Lou72, NR72, SJ75, Wag75]. Insulator [Col88b]. Integral [Pea85, NYS81, Hea85]. integrals [Kah80]. integrate [Wil97a]. Integrated [ABC^{+}88, Car95, Cha96a, CS88, Col88b, DS82, GDKK94, Kam82, LBKY86, MKR95, Mar95b, MG85, NDS86, SB83, SGD83, SHB^{+}82, Tan96, ZFE^{+}86, BM79, BT97, BM81, BCH81, EW78, ELU81, Gor80, HG79, HMPF74, Ham93, Har83, HBB88, HLT^{+}83, Lie97, Mat92, PW82, Ric91, Rob82a, Roy81, WJ80]. Integrating [DH98, HW88, May88, Rob92, STE90, Tuc95, Cze89]. Integration [AW90, AGI^{+}91, AT83, Byr96, GNR93, Goe89, Gre90, LK80]. integrator [Fro90, Ste74b]. Integrity [Gen85]. intelligence [GUW92, Kla78]. Intelligent [DOR97, Fit82, MiITT82, SMGR85, Bru78, Bru77]. Intensity [BS95, Che86]. Interaction [Ano92l, HWK95, Por80]. Interactions [Cou94]. Interactive [Fit82, Ful83b, JY72, LB98, Lin79a, SWM78, SW86, BM81, FS81a, Ish80, Kin81, LTC^{+}84, LT90, NGW92, Pee79a, Ryt77, SMK84, WB90]. interchange [Hal93]. Interconnect [Eri92, FCK97, NSMO98, RO97, Raj95, Fle83]. Interconnection [HGWS97, RL86, Jes89a]. Interface [Ano92l, Ano92-31, Ano93c, BP85, Buf96, BPA^{+}94b, DH77, DF90, Fer94, Fle83, HSK87, JS74, LK85, Lou75, Mat79a, NZ97, Not93, QH83, RWW95, Rya95, Sta87, TM91, Tre78, AC90, AMSW85, Bac91, Bow72, BHP85, Cha91, Die75, Dou92, FPS72, Gil81, GAP92, Gra81, Har83, HL86, Hun92, JS93, JF80, Kai91, Lal93, Lan83, LT90, Lou72, Mai90, MOT90, NR72, NGW92, Per88, Sho89, SKG90, SS83a, Ste89a, Vob88, WB90]. Interfaces [GSFF90, BBE^{+}91]. Interfacing [LB98, LHIR98, Mor76, Was80]. Interference [Ano92w, Bec79, Inh75, Wat87a, Won79]. interferometer [BGH74]. interferometry [BGR71, CS93]. interleaved [MR93]. Internal [Ano93-27, Coc68, DBB^{+}95, LK85, Nak96]. Internally [HZ83, JJ85]. Internals [Ano94z]. International [WS84a, PH70]. Internationalized [RC98]. Interoperability [Ano93-41, Mey90a]. Interpolation [Ano93s]. Interpreter [Ano92n, Mat84]. interpreter-based [Mat84]. interpretive [Van77c]. Interprocess [Wat87b]. Interrupt [Ano92b]. interval [Chu74, CAF78, Hei88, NS88, OST75, Ste88]. intervals [Sch70, Wes77]. Intrabuilding [Pie87]. Introducing [CW72, GR69]. Introduction [AT95, AN90, McS92, Wat87b, Wes91, Cag86, Pri96]. Intuitive [WCH94]. Invalid [Ano95-27]. invasive [MH76]. Inverse [Egb77c]. Inverter [Ano94-36]. inverts [Ban69]. ionization [Smi73]. IP2 [Saf96]. Ipm [Ter78]. IR [Kri81b, Sol81]. IrDA [MBD^{+}98]. ISDN [NZ97]. islands [AGI^{+}91]. ISO [KGTM90]. isolation [Gro79, Yan93]. Isolator [CSS^{+}91]. Issue [Ano96m, Ano96n]. Issues [RR88, WP94b]. itself [Goo77, Tho72]. IV [Egb78, MRH94d, WM78]. IVI [MOT90, TW90, WB90]. IVIBuild [AC90, WB90]. iX [Lal93].

jams [Hol82b]. jet [KB85]. Jets [AMK85]. Jitter [Ano94-36, MM95, EKW93]. joint [Kee89]. Jones [Ano95-28]. Joy [Ano93u]. Just [BLNP86, Pea85, Gas88]. Just-In-Time [BLNP86, Gas88].

Kayak [Cun98]. kernel

[BLY92, GOS84, KG92, OY88, SSV90]. Kettering [Mer81]. key [Ing70, Kah79]. keyboard [Bla75, DP84, SB84]. kHz [Vif76, MP81, RG69]. kind [DKMS81, Sch70]. kit [ELU81]. Knowledge [DL98]. Kunz [KH88a].

L [QAKP90, WJ79]. lab [Bre74, BS79, Pec74]. lab-bench [Bre74]. Laboratories [Ano94n]. Laboratory [Ano93v, DGH^{+}97, GBHS81, Loo97, AM73, MW73]. Laboratory-performance [GBHS81]. LAN [Cra92, FP90, GUW92]. Language [Ano93-43, BK85a, Buf96, CHP85, DH77, KA95, KHI82, LBKY86, Ros88, ST83, TE86, CP79, FG75, JY72, KMT88, LTC^{+}84, LWBB84, Mil84, MRH80, PW89, Pee79a, Pet68, Pir97, Rob92, Van77c, Voi84, Spa72a]. Languages [Tre78]. LANs [MCD97]. Large [Ano92-27, Ano94x, BJWW84, Has92, Smi84b, Ave75, BSME92, HKC^{+}84a, Joh80, PL81, Van77b]. Large-Format [Ano94x, Has92, BSME92]. large-scale [Joh80, Van77b]. Laser [BTBW83, BTW83, BCC^{+}94, Can93, Gro83, Hal82, Her73, Hol82a, NMY95, RW76, SS82, Ste83, THHW95, BGR71, BGH74, CDLN96, DCN^{+}88, Gar83, JBS93, Kee89, Mai91, MLJP93, MRRS93, OS76, QW83, SL93, CH82, Hol82c, LH82]. Laser-comparable [BCC^{+}94]. laser-quality [DCN^{+}88]. Laser/calculator [Her73]. Lasers [RLB^{+}95, RWT97, BS90, TZ93]. latest [Har82]. launching [WM79]. law [Ano94-31, BFS94, Ano94-31, BFS94]. Layer [Pug90, RHK^{+}83, ST90, Chi90]. Layers [Ano95-31, BE90]. layout [Bak81]. lb [Smi69]. LC [AN90, LNH^{+}84, Sch84, WDDH84, WTB^{+}90]. LC/MS [AN90]. LCD [Adl86]. LCR [HT76, Mae74, MN79, YK77]. LCZ [WKT83]. leadership [DW92a]. Leading [CHNJ97]. Leading-Edge [CHNJ97]. leads [DKMS81]. Learning [Ano92l]. LED [FR87c, Ste98]. LEDs [Ano93-40]. Legal [Ano94-29]. Length [RNR^{+}98]. less [GZC74, NT75]. Level [Ano92a, Ano92e, Ano95t, Ano97c, DR95, ST83, ACCM83, Cos76, FK79, MM90, NBZ^{+}97, Pir97, PH70, Pra75, RU76, Ric81b, SD78, Spa85, Tho80, Urq76, Wal76, Win80]. Leveled [HZ83]. Leveling [BDH91, Cat82, Lar83a]. Levels [Ben95, Gen85, MH76]. Leverage [WP94a]. libraries [CR92, Mar89]. Library [Ano94-28, BK95, Hol84, Sch96, Wol86, BSY89, SWP90, STE90, Vog91]. Life [Cot96, Ter78]. lifecycle [Gre90]. Lift [HM85]. Light [BKAWBCTK98, DBB^{+}95, FKO^{+}93, KW80, NMY95, Por80, SDC^{+}88, CS93, Jon89]. Light-Emitting [BKAWBCTK98, DBB^{+}95, FKO^{+}93, SDC^{+}88]. Lighting [Ste98]. Lightwave [Ano91f, Ano93z, Ano95u, CSS^{+}91, PPJ91, ALRS89, Bec93, DVl91, HH91, ICNT93, JM91, KBHS93, MR91, MCM93, Mil90, MC91, RSM91, WHHC89, WHH91, HCHR91]. Lightweight [BJPH95]. Like [War97a, MCCB76]. Lilliputian [BJM71]. Line [Adl86, DLJ85, GLW85, Hur90, MTWG85, Pie87, BLKH77, BD73, CDS71, DR81, Jep85, KB69b, LT85, MM72, Pli89, Sta88, Ter78]. Linear [Adc87a, MW94, TC79, Ryt77]. linearity [Hil93b]. lines [AD75, IKS81]. Linewidth [BS90]. Link [CHNJ97, ST90, THHW95, CM75, Han79, HWS87, Ish80, Rob82a, Urq72, Sta87, YWP^{+}92]. linked [Wil78]. linker [Ste80c]. Linking [Jen98, AGI^{+}91]. Links [Ano95-36, BYA86, BB73, BP93]. Liquid [Ano95-29, JR84, KH84, Wak97, BFP90, GV84, KP90, Maz80, Sch75, SW90a, SW90b, WMZ84, Wie90]. Lisp [KA95]. listens [Nad75]. Lithography [GN82, OLR82, BOOS81, EHS81, Ham81, gJH81, KGK81, LM81]. little [BJM71]. load [OY88]. Local [AT95, Gd86, JRM^{+}86, LAB97, BWW84, Chi82, MLS79]. localization [GHK^{+}93]. localized [NSP97]. locates [EZ81]. locating [Erd79]. location [GH73]. locator [Low69]. Locked [TOB97].

Log [Cas95a]. Logarithmic [Egb78, Jer74]. Logging [Don86, SFW78, TW76, Wol81]. Logic [AH72, ALT82, Ano97f, Ben95, DSW83, DDT83, HR82, Neu84, SK80, War97a, ABM73, Bec76, BC76, Car69, Cha77, COO^{+}93, DeV80, DiF75, Gor69, Haa78, Lan75, Mor76, MH80, Ogd78, Ogd80, Pet74, Pri76, Que76, SWM78, SM75, SD78, Smi77, WN79b, ZHN83, Far74]. logic-circuit [Bec76]. Login [MKR95]. Logon [Wat88]. Long [GLP^{+}97, Ter78]. Long-Haul [GLP^{+}97]. Loop [Ano92d, TOB97, AH89, Gue74a, MM79, Fle83, Lan83]. loop-holding [Gue74a]. loss [Mai91, Sch93b]. lot [Fra71]. Loudness [Ano67, Ohm67, BF67]. Low [Ano92-31, AWH^{+}87b, AWH^{+}87a, BDM85, BEM95, BCF96, BGS83, CLG87, Cau85, DBB^{+}95, DGN98, Eps84, GLP^{+}97, HJL^{+}83, HM85, Het79, HL85, JR84, KA92, LAB97, Mal87, Mar95b, MEM82, MHV96b, MHV96a, Moa98, Pea95, QH83, Res83, RSW^{+}86, San86, SCI^{+}81, SM86, Spr82, VCMH96, WST84, Wit92b, WMBB94, Wol83, AP76, Ain81, ABD^{+}97, Azm82, BD73, BLTY81, Bec76, BDOT85, BVAP87, BCH^{+}93, CP85, CM72, CS80, CK76, CF68, Dic78, Edw79c, EKW93, FCC79, FHDS92, GJM^{+}92, GW77, HW73, HVSM96, HWS87, Haw71, Hay74, HMM73, HL86, KB71, KPC^{+}82, Ks69, LT85, Mac73, MC85, MBW79, McH87, NYS81, PF78, Pon80a, Pra75, Ric76, RC79, Smi79, SM87, Sta81b, TD83, TC79, Wha68, WDDH84]. low-coherence [BCH^{+}93]. Low-Complexity [Moa98]. Low-Cost [Ano92-31, BDM85, BEM95, BCF96, BGS83, Cau85, DGN98, HM85, HL85, LAB97, Mal87, Mar95b, Pea95, QH83, RSW^{+}86, Spr82, WST84, WMBB94, Wol83, Het79, KA92, MEM82, Res83, Wit92b, Azm82, BD73, BDOT85, CP85, CK76, CF68, Dic78, Edw79c, FHDS92, HWS87, HMM73, HL86, LT85, MC85, MBW79, Ric76, Smi79, SM87, Sta81b, TD83, WDDH84]. Low-Dispersion [GLP^{+}97, JR84]. low-distortion [Pon80a]. Low-End [SM86, GJM^{+}92]. Low-Frequency [AWH^{+}87b, AWH^{+}87a, San86, BVAP87, CM72, GW77, HW73, Mac73, NYS81, PF78]. low-impedance [Bec76]. low-inertia [BLTY81]. Low-Internal-Reflection [DBB^{+}95]. Low-Mass [HM85, BLTY81]. Low-Noise [CLG87, SCI^{+}81, AP76, Ain81, McH87]. Low-Power [HJL^{+}83]. Low-Temperature [MHV96b, VCMH96, MHV96a]. lower [CHM76, Kei71, SMT^{+}97]. lower-cost [SMT^{+}97]. Lowers [Gib86]. lpm [LT85]. LSI [BS98, DL77, EM76, Gun94, Lyn80a, Sch78a, WJ79].

m [BBF^{+}97, BFS94, Raj95, Har82]. m-law [BFS94]. Machine [BTW83, Kru94a, Mon92, CR73, CH82, Fel74, HCF^{+}88, RW76]. Machines [Sim94b, Anz93, Saf96]. Macros [Ano96o]. made [HM71]. Magnetic [AGS85, HRE85, HSYJ83, KH88b, Sei81, NSW76, Tag74]. Magnetically [NML90]. Magnetooptical [Spo90]. Magnetostatic [IC85]. Magnetostatic-Wave [IC85]. magnifies [Mic72]. Mail [Ful83b, HSMA^{+}86, WPS94, GR91]. main [Elw74]. Mainframe [Ano92-31, SGD83, Wor92, CJHE93, Ish80]. Mainframes [ARL^{+}94, Ano92t]. maintains [BW72]. Maintenance [MB85, DMW94]. major [MW80]. Make [Ano95o, Ups71]. makes [EH81, HW73, ULW81]. Making [Ano92u, Boy80, Dac76, DF90, Kri81b, AD75, Pee73]. Malfunctioning [GV86]. Managed [Nak96]. Management [AP90, Ano92r, Ano95-33, Ano95-34, Ano96v, Bra96, BK95, BHK^{+}86, Cha96a, FP90, HKH93, KHLM90, Kru88, LEV96, Lub91, MB85, Man90, May88, MM97b, NH94, Not93, Row86, SDDC96, AC93, Dys89, GC90, Gau81, Haj96, Han80b, Het79, HM88, Kru97, McI74, Rid90, Sch93c, SKG90]. Management/800 [Man90]. Manager [Ano92r, KT90, MD86, Spe96, BKD94, FP90, RM88]. Manager/Agent [Spe96]. manages [Ave75]. Managing [ACL89, Anz93, ISLW78, Mil96, RM88, Rho88, TPG79, Ric91]. Manifold [Ano92g]. manipulate [SFW78]. Manipulating [Don71]. Manipulation [PS83]. Manual [Ano93-38, SW86, Sta88]. Manufacturability [Ano93l, Ful96, Cho90, Pie75]. Manufacturing [BLNP86, Cra85, Gro83, Kjo92, KV94, MSB85, MRH94e, PT90, Rip87, Rip91, Whi85, BKS93, CW92, FS81a, GTM95, HA92, LCB90, UAH^{+}92]. Manufacturing-Oriented [Kjo92]. many [Que76, SBM85]. MAP [Chi90, CFK^{+}90, Gre90, LCB90, Man90, Mey90a]. Mapping [Bha94]. Marco [Neg81]. mark [gJH81]. Marker [Ano95-27]. Market [WS84a, But95]. Mass [BC74, BGS83, Eri92, Han80b, HM85, HSYJ83, Ike83, Kis97, SR95, SD82, BLTY81]. Master [Reh84]. Master/slave [Reh84]. Match [Ano97e]. matched [gJH81]. matching [MM79]. Material [MRH94a, FS81a, NGW92, RL92]. Materials [Col86]. Mathematics [BBD^{+}96, MRH94a]. Matrix [Cra85, DiV84, PL84, TD92, ISLW78, Jep85]. Maturity [LC96]. Maximize [Wen85]. Maximizing [VBK89]. maximum [BP76, HC75]. mbit/s [YPS76]. McCabe’s [War89]. Me [Har87, WYO^{+}87]. means [HC75, RW81, WN79b]. measured [KB71]. Measurement [Ano91f, Ano92j, Ano93-34, Ano95p, AWH^{+}87b, AWH^{+}87a, Ban84, BP82, BVAP87, BM86, Des92, ES87, EL82, HCW^{+}82, HDH96, HH95, Kar82, McK87, NK84, OH82b, PH96, Pet96, QAKP90, Ris82, Van86, Bru78, Cam82, DF86, FR87a, Foo80, FPS72, Ful83a, Gar83, IK69, KB69b, Kas72, Kla78, KW80, Lai81, Mar70, MG92, OY88, QW83, Rom91, Sch93b, Smi80, Sny75, Ste88, TST82, Wec88, WP94b]. Measurements [Ano93e, Ano93-45, Bec87, Bir96, Che86, Kar86, MiITT82, MS95, SHB84, AD75, BS90, BBT92, BHW70, Bes75, BF79, BCO^{+}93, Bot71, Bru79a, CS93, CAF78, CPR75, Cro79, Dre74, DY75, Elo93, EH81, FEG77, FR74, Fis72, FH74, Gue74b, Her73, HM71, HS80, Jac74, Joh80, Ket74, Kri81b, Lam74, Mac73, Mai91, MR91, Mar71, MVG77, MH76, MS80, NTA72, O’B93, OST75, OS75, Pee73, Pra75, Rus93, SJ80, Sch80, SFO80, Sor74, Swe89, Unt68, WKT83, Wes77, Win74]. Measures [LeV88, BGH74, BB78, Cra72, Lin69]. Measuring [AI84, Bau69, Bir91, Cas95a, Dam74, Das96, DNG84, Fol72, Gro91, Gro92, Kam72, NSM70, NTH^{+}84, PMN82, PH70, Sch70, BJM71, Cos76, HMM73, SJ75, Urq76, Wal76]. Mechanical [AGS85, Cam82, CJHE93, Cau85, Cle93, DF91, DSC^{+}90, JJ85, Ker80, Kre92, LS87, MTWG85, Pea85, PT92, RH92, Smi84b, Ste83, TCWN88, Top89, YF92, KPC^{+}82, Woo78]. mechanics [FS81b]. Mechanism [HM85, BMTW78, Goo91, JKP^{+}88, SMQ88, Ter78]. Mechanisms [Wat88, WP94c]. Media [CHW96, PSS^{+}88, WMBB94, APA92, Bri94]. Media-Rich [CHW96]. Medical [Ano91g, BPMF87, Gor91, Lar83b, Mil83, Sol81]. medium [HC75]. medium-cost [HC75]. meet [Bec79]. Meeting [CG98, UAH^{+}92]. megabyte [Smi79, Sti77]. Megasample [DWD88, Hil93a, MPR82]. megaword [WM78]. megaword-array [WM78]. members [MBH^{+}89]. Membrane [LGV^{+}98, Mat90]. memories [Ave75]. Memory [Hil93a, HMM96, LRFL83, OSN^{+}85, Row86, Bas79, BC74, CS78, EDF82, Elw74, Fin72, Fra71, Fra74, HW82, HM88, LL91, LLS88, Mar89, Mey84, TW76, US79]. Merge [ACL89, BCH^{+}89, BG89, ES89, HLM89]. mesa [Slo89]. Message [Ano92m, Cou94, NZ97, LCB90, Mai90]. Message-Based [Ano92m]. Metal [RHK^{+}83]. metallisation [Mer82]. Metallization [Ama86]. Meter [AI84, FG87, Lym87, SC83b, BHE^{+}85, Edw75, Han72, HT76, KH69, Mae74, NYS81, Sch87, SB80, Tho80, YK77]. meters [Hen91, MN79, WKT83]. Method [Wak97, BR91, Chu88, Fro77, HG79]. Methodologies [DMJCB97, BBJ^{+}95, FJ89]. Methodology [FHMP97, Gem85, KGH^{+}87, BF89, KHW89]. Methods [Ano97b, BLNP86, BRHS89, Rus92]. metric [War89]. Metrics [Ano92v, War86]. metrology [WB83b]. MHz [San72, AI84, Ano97a, BKE^{+}90, Bes75, Bot71, BFHT83, CCH^{+}91, DA75, DiP80, Fol72, GY88, GZC74, HK88, Hei88, Hil93b, HT92, IOS80, Jek75, KW90, Koh72, KE74, Lin69, LE83, LGV^{+}98, Mae74, MBW79, MT81, Rya95, RG69, SIBH88, SES90, Tut72, WD84, Wit92b, Yan92, Zel79, Zel69, SKB78]. mHz-to-50 [SKB78]. MHz/ [Jek80]. MIB [Bur93]. micro [KT81]. micro-grip [KT81]. Microassembly [BCRS88]. Microcircuits [HJG93, PAD70, BBH91]. Microcoded [Ame85, Geb78]. microcomputer [QW83, SK80, Ste80b]. microcomputer-controlled [QW83]. Microcontroller [Ano97d, Mur97]. Microelectronics [Lam74]. microinch [DG70]. micromachine [Mur82]. Micrometer [RNR^{+}98]. micropositioner [KW93]. Microprocessor [BKQW95, BB78, BS79, Lin97, Lym87, ST83, BBJ^{+}95, Ber88, BT97, Bla75, Boy80, BFHT83, Dac76, DGM80, Gor80, ISLW78, Koh72, MBW79, Moy78, Ogd78, Ogd80, Sap77, Smi77, Ste74b, SSMM81, Wag75, Wai75, Woo78]. microprocessor-based [Boy80, MBW79, Moy78, Wag75]. Microprocessor-controlled [BB78, Gor80, Ste74b, Wai75]. Microprocessor-Enhanced [Lym87]. microprocessor-scanned [Bla75]. Microprocessors [RO97, RR83, Yac80]. Microprogrammable [GJ74]. Microprogrammed [Lan77]. Microprogramming [Cou72, Lei71, Sny75]. microprograms [Dra77]. MicroScope [ABC^{+}88]. Microvolt [Ano92w]. Microwave [Ano93-28, Ano97e, Arm95, BW92, Bar82, BGR^{+}88, CS82, ES87, Fis72, FFS89, Gib86, HMPF74, IC85, Kih82, Ped86, Sch80, SFO80, AKL69, Ast77, BEFV70, BBI^{+}93, BCZ93, BB78, CM75, DF82, Det72, DW92b, Edw75, Elo93, HHR71, HPF93, Jac74, Kei71, Koe91, Koe80, LD79, Lin79b, MLS79, Mer73, MB80, Per86, PAD70, Pra75, Rob82a, SSMK92, Sch73, Str77, SS89, Tho77, Urq72, VDF72]. microwaves [KV74]. Mid [CHM76]. Mid-range [CHM76]. Middleware [Gup95]. Midrange [Bre94, BCF96, DW92a, GJM^{+}92, MBH^{+}89, RJH^{+}87]. Migrating [Joy96, TSv92]. Millimeter [Alb88, BGR^{+}88, Col88a, DG88, EJG86, Mat86, SR88, Uph88, ZFE^{+}86, Nic90, SA91]. Millimeter-Wave [Alb88, BGR^{+}88, Col88a, DG88, Mat86, SR88, Uph88, ZFE^{+}86, EJG86, Nic90, SA91]. million [Elw74]. million-word [Elw74]. MIME [Ano94-30]. Mini [Ric76]. Miniature [FEG77, Mer81]. minicomputer [CJ71, Elw74, Fra74, Mic72]. minicomputers [Ste74a]. Minimal [SJ75]. Minimize [Ano92b]. minimum [MO74, RS75]. MIPS [FMN84, MS82, WST84]. mirror [Hol82c]. mirror-motor [Hol82c]. Mixed [Ano97d, BS98, Gro91, HHT^{+}97, Mur97, SWI97, WR97, Gun94]. Mixed-Mode [SWI97]. Mixed-Signal [Ano97d, BS98, HHT^{+}97, Mur97, WR97, Gun94]. mixer [Spa85]. Mixers [Mat86, NML90, LD79]. mixing [McK83]. MLA [Cra72]. mm [Ano95-47]. MMIC [RNR^{+}98]. Mobile [Dan96, Sch84]. Mode [Ano96l, SWI97, CP85, HH95, Mor76, Tal88b]. Model [Ano94u, Ano94-35, Ano94-34, Ano95-34, BW72, Hig95, Jan96, MZ96, MRH94a, MRH94e, NSMO98, Chu79, FK84, Hal79, JY72, MMNR89, Mar89, Nea68, Osb68, Spa72a, DH76, HMLN78, LC96, MW78, RL86, Roe95, Sha78a]. Modeled [GLP^{+}97]. Modeling [Ano92-28, Ano94m, Ano94n, Ano94w, Ano97e, AWH^{+}87b, AWH^{+}87a, Bra96, DR98, ME95, MRH94e, SWI97, Bur89, FR95]. Models [Cor96, Ern95, Kru88, AM95, Kru89, RH92]. moderately [KE74]. modes [BVAP87]. MODFET [RNR^{+}98]. Modifying [Ano94s, Kar86]. Modular [JJ85, Rub86, SKK^{+}86, Sch93a, SM86, And70, BB73, BP93, BFP90, DP73, Fel86, Jes89a, JM93, KSW89, Lan75, LJ75, MP89, MP81, NPG^{+}83, PGW73, ST75, Wie90, ZHN83]. Modularity [Ano93-29, HC75, EDF82, Kas72]. modulated [BCO^{+}93, Edw79b, Hal75]. Modulation [Bir91, Edw79a, GC87, HZ83, Hoe97, Nae79, Ton85, AH89, Bru79a, HK75, Lin79a, SC89, Bru79b]. Modulator [KWR^{+}87, JM91, Koe91]. modulators [Kih82]. Module [ANSW95, BD84, Dan96, Don86, DGN98, Har82, MD96, SHB84, JBS93, Kai91, McH87, Mil84, Ogd80]. Modules [Alb88, Buf96, HGWS97, Kel92, MKW^{+}96, PH96, TD92, Car69, HMM73, RSM91, SA91, SSV90]. Molecular [Col86]. Molecular-Scale [Col86]. Monitor [Ban91, BJ82, HFKS95, TM91, CRSS77, GG77, HMW83, Hor69, Hur90]. Monitoring [Ano91a, Ano91b, DF91, Hig85a, KB69a, SW91, Tiv91, WJ91, Wes91, BN80, Bur93, Gol80, Gor91, Han80a, NSV92, Rue80, SM88, Kai91, Rei91, WH91]. Monitors [Ano91g, HPM87]. monochromator [WA93]. Monolithic [BP70, Mil75, O’N80]. Monte [NE81]. MOS [Col88b, Oh82a]. MOSFET [GBHS81, TTML81, Van97]. most [Boy80, Dac76]. Motif [DF90, KT90, ME90]. Motor [Ano94-35, FS81b, Hol82c, MM79, Mer81, TC79, WJ82]. Mount [Ano96k, BKAWBCTK98]. Move [NDS86]. MPE [BGK^{+}87, BKO87, Lal93]. MPE/iX [Lal93]. MPEG [LBLS95a, LBLS95b]. MPower [Ano94-32, BFS94, SJ94, Yod94]. MRP [Ano92r]. MS [AN90, Wol86]. MS-DOS [Wol86]. much [BJM71]. Multi [LAB97, MN79, RLB^{+}95, Rue80, AG75, Bru77, Sch76, Van77b, WKT83]. multi-color [Bru77]. Multi-frequency [MN79, WKT83]. Multi-processor [Rue80]. Multi-Quantum-Well [RLB^{+}95]. multi-terminal [AG75]. multi-user [Sch76, Van77b]. Multi-wavelength [LAB97]. Multichannel [Ano92-36, Gar92, PK75, Urq76]. Multichip [DGN98, HDH^{+}87]. Multicompany [Mil96]. multidevice [Lie88]. Multifamily [Pri76, BC76]. Multifunction [Ive88, Tho88, BC76, Hig88, SM88, WJT71, Wol75]. Multilayer [Ano92w]. Multilevel [Ano95-30]. Multilingual [BP82, Fos73]. Multimedia [Ano95-39, Fer94, GS95, LBLS95b, Man94, Mar95b, MS94, Pea95, Pet94, Ric94, ROKH94, WPS94, Yod94, LBLS95a]. Multimedia-Enhanced [LBLS95b, LBLS95a]. multimeter [CR89, Cze89, DL73, GDV73, GSVS89, Goo72, MR91, PPJ91, SFW78, SMR^{+}86, Ste89b, Tho72, TD83]. multimeters [BLMJ77]. Multimode [THHW95, SKB78]. multiple [AK79, HKC^{+}84a, NBZ^{+}97, O’B93, OGB84, Sta81a, Tal88b]. multiple-mode [Tal88b]. multiple-output [AK79]. multiple-processor [OGB84]. multiplexer [Yan93]. Multiplier [KWR^{+}87, LS83]. Multipoint [ABJ78, Bra78]. Multiprocessing [BBF^{+}97, Bre94, HVSM96]. Multiprocessor [Ano92y, Bru97, BCF96, LP92, LD93]. Multiprocessor-Based [Bru97]. Multiprogrammer [Mic72]. Multiprogramming [Rob79]. multipurpose [FG73]. Multirate [Wai93]. Multitasking [Kel92, WM85]. Multiuser [BEM95].

N [Ano93o, Ano93z]. Naming [Mar95a]. Nanometers [HMSFR87]. nanosecond [Sch70]. Narrowband [Han72, RMVT95, War72]. Native [BK85a]. natural [Rob92]. NCR [Ano92b]. Needs [CG98, GJ74]. Negrete [Neg81]. Nets [McS92]. Network [AT95, AP90, Ano90c, Ano92-29, Ano94t, Ano95-31, Ano95-37, BKD94, Bra96, Cha96a, CS88, CHNJ97, DWB^{+}97, DKW90, DG88, DGH^{+}97, FKL86, FP90, Gd86, Ks69, KHLM90, LAB97, MM97b, Moo92, NZ97, RG69, SDDC96, Uph88, Yan92, AKL69, Bak84, BWW84, Cla81a, DH76, DY75, GC90, GL88, KN93, LS76, NTA72, NK78, RW70, Rid90, REOB78, RS70, Sch78b, SKG90, Vif76, WD84, YN93, Zel79, Ano92x, Bha92, Dou92, Kre92, Pru92]. Networking [CGS^{+}86, FT86, Met86, Pea95]. Networks [BB97b, DOR97, JRM^{+}86, MM97a, WSYD97, WP94c, Bur93, Haj96, NSV92, Wil78]. Neural [McS92, Sha89]. neurons [Sha89]. NewWave [Cro89, Dys89, Ful89, LHD89, LFWEJ89, Sho89, SW89, Ste89a]. Next [Ano92y, BKO87, Ric97, HH78]. Next-Generation [Ano92y, Ric97]. nm [FRMK85, RWT97, SFHE85]. NMOS [ABH^{+}83, DL77, EM76, FSDC82, MFMS83, WSBK83, YLJ^{+}87]. NMOS-III [ABH^{+}83, MFMS83, YLJ^{+}87]. Node [Ano96r, Ano96s, Ano96x, Ano96z]. Noise [Ano92f, BS95, CLG87, NSMO98, SC83b, AP76, Ain81, AH89, Ano67, Han72, IK69, KB69a, Man69, McH87, PH70, PB82, Ras87, SCI^{+}81, Wat87a]. non [CRSS77, Joh80, MH76]. non-contact [Joh80]. non-invasive [MH76]. non-technical [CRSS77]. nonintrusive [TL89]. noninvasive [Rom91]. Nonlinear [Dre74]. Nonuniform [DR98]. Normalized [Sch96]. North [Whi93, Hoo91]. notch [Lau80]. Note [Ano95-32]. notebook [Mye96]. novel [Win81]. ns [GZC74]. NT [HD98]. nuclear [CDS71, Ups71]. numerical [GP84].

O [ABD^{+}97, Ano96-29, Bee82, DN86, EM98, FK84, GJW83, LMM84, Mic72, NMS80, NTJ^{+}91, Nel82, SLM^{+}95, SSMM81]. O/E [Ano96-29]. Object [Ano92z, Ano92-27, Ano95-34, Ano95-35, BK95, CHK^{+}95, Fie89, GLM95, Loo97, MM97b, Reh97, TW90, AC93, Bur86, Dys89, Kra89, KHW89, Mai90, Sho89, Ano95-33, Ano95-34]. object-based [Sho89]. Object-Oriented [Ano92z, Ano95-35, CHK^{+}95, GLM95, MM97b, TW90, Ano92-27, Fie89, Reh97, AC93, Bur86, Kra89, KHW89, Mai90]. Objects [Bra96, Nak96, ACL89, DL97]. observing [ABM73]. Obtain [PDW85]. ODB [AC93]. OEMF [Haj96]. off [Edw75]. offer [WKT83]. offers [BP77, BP76, BW81, Edw75, LB80, Sta81a, SMR^{+}86, Suk84]. Office [Ano94s, Ful83b, MRM83, WM94, LHD89]. Offs [DMJCB97, NIWM98]. On-Chip [NSMO98]. on-column [Kal95]. On-line [CDS71, KB69b, Sta88]. On-the-Fly [Kel94]. One [DSW83, CPK88, Emm72, Mil88, OS75, SW81]. one-gigasample-per-second [CPK88, Mil88]. Online [Can93, CHW96, HKH93, Smi94, WCH94]. only [Fin72, GSVS89, Kei71]. Open [AC93, Car95, SKWR95, Wor92, TSv92, Mar95a, RL86]. OpenCall [PA97, WA97, DOR97]. OpenGL [CL98, LCPG98]. opens [MHCH77].

OpenView [FP90, GC90, Hur90, SKG90, KHLM90, Rid90, She96, Sto96]. Operated [BLSM83]. Operating [Ano93-34, BGK^{+}87, BKO87, CHKS86, CBL96, Wat87b, WW84, AG75, Bac91, BHP85, Car79, Cha91, FRAC85, GOS84, GG77, KGTM90, KG92, KM79, LP92, LMM84, Mor93, SH77, Sta88, STE90, Van74, WM78]. Operation [EJG86, Hal85, MC84, SMGR85, ALRS89, ACK79, BS79, Edw75, HK75, KCP77, Kur81, Leo77, Reh84, SKB78]. Operational [Han94, KM97]. Operations [Ano95o, Ern95, BW76, RBB78]. Operator [Gil81, Car78]. OPNODE [Ryt77]. Opportunity [McF98]. Optic [Ano95-47, CHNJ97, RWT97, WSYD97, FR87a, FR87b, Han79, Hen91, MR91]. Optical [Ano93-30, Ano95-36, BS93, BTW83, BS95, BGG88, Bec87, BV85, BB97b, CSS^{+}91, ES85, FG87, Ger98, HGWS97, HM71, HMSFR87, LH82, MS95, MSS87, MKW^{+}96, OLR82, OB90, Vob88, Zel70, Bel93, BP93, BHE^{+}85, BCH^{+}93, CS93, DSC^{+}90, ELU81, ELN88, FRMK85, FRS88, HS80, JM91, KBHS93, KW93, Mai91, MLJP93, MS80, Rad87, RP88, SH90a, SFHE85, Sch93b, SF95, Sch87, She71, SNN88, SWP90, STE90, WA93]. Optical-Fiber [Ano95-36]. Optics [ES85, Hol82a]. Optimization [BD84, GLP^{+}97, KV94, RO97, Ryt77]. Optimizations [Ano94l, Han92]. optimized [EKW93, GRS93]. Optimizing [BMTW78, HKZ74, GMMM92]. Optimum [Per86, CJHE93]. optoelectronics [ELU81, Jon89]. ORBlite [MK97]. ORCA [Ano93v, GRS93]. Order [Ano93-42]. Ordering [Mal87]. Organization [Ano94k]. Organizing [Sha72]. Oriented [Ano92z, Ano95-35, CHK^{+}95, CGS^{+}86, Cot96, GLM95, Kjo92, MM97b, TW90, AC93, Ano92-27, Bur86, Det72, Fie89, FG75, Kra89, KHW89, Mai90, Reh97, RHM79, SW80, Sch79, Ste74a]. orifice [SBM85].

Orthopedic [Can93]. oscillator

[BW81, CF68, DF82, HEW82, MLS79, MH89, Ste81, Tut72]. Oscillators [CWFS82]. oscillographic [Bru73]. Oscilloscope [Ano92-36, Ano97f, BMD^{+}88, HT92, HHT^{+}97, Kah92, Kar82, MB93, MD96, Mon88, OH82b, REB86, Sch93a, Sch86, Toe86, Toe97, War97a, Bes75, Cam82, CJHE93, DeV70, EKW93, FEG77, FR74, HA92, HKZ74, Har76, How79, JM93, Lin69, Mil75, Mil88, Mor70, Nic88, Pax70, Pet70, RO86, Rus93, US93, Wes77, EL82, HCW^{+}82, Ris82]. Oscilloscope-Like [War97a]. Oscilloscope/Logic [Ano97f]. Oscilloscopes [Ano96h, Bru97, Gar92, Mur92, PF88, Sor88, War97b, Wit92a, YF92, Wit92b, Wya92]. OSF [DF90, KT90, ME90]. OSF/Motif [DF90, KT90, ME90]. OSI [AW90, BE90, Bor90, Joh90, Lon90, Pug90, SH90b, ST90, Tal90, Ano90c, Bra96, Chi90, Pug90, SSV90]. OTDR [Ano93c, BT98, JS93, Mai93]. Other [Ano95-37, BTW83, KB69b]. Output [BDH91, CSWR85, Hal82, KWR^{+}87, KNTV96, KM84, AK79, Aue81, CAV69, Ega72, FFS89, HW91, Hig88, HM84, JBS93, KW90, KE74, LG92, Mat79a, MS82, Nic88, Pee79b, PB81, Pon80b, SES90]. oven [DJS83]. overlay [HLM89]. Overstress [Dia94]. Overview [Ano93-31, Ano94v, Ano94-31, Ano95-39, Ano95-38, Bir96, BFS94, KD83, LCPG98, SOG98, FM84, Ful89, Joh90, LT90, Mar82, Mer82]. Own [BT98]. oxidation [Bro82, Chi82]. oximetry [KNF^{+}97]. oxygen [MH76].

P [Ano93z, DD86]. P-I-N [Ano93z].

P-Pods [DD86]. PA [BBR^{+}97, DMJCB97, JU97, MGBB97, ABG^{+}92, ADG^{+}96, Ano92s, Ano92-28, Ano95-39, Anz93, BBJ^{+}95, BKQW95, BT97, BLY92, Bre94, CHK^{+}96, DW92a, FCK97, FHDS92, GJM^{+}92, Han92, LBLS95a, LBLS95b, Pea95, Ram92, San92, SBK^{+}97, UAH^{+}92, WT97].

PA-RISC [ABG^{+}92, Ano92s, Ano92-28, Anz93, BLY92, Bre94, DW92a, FCK97, FHDS92, GJM^{+}92, Han92, Pea95, Ram92, San92, SBK^{+}97, UAH^{+}92]. PAC [MHKS94]. Package [Ano94w, BB97a, BZS83, FS86, Tan96, Ano71a, Hor79, Mac75]. Packaged [ES87]. packages [QAKP90]. Packaging [Ano94o, Bel98, DWD88, DGN98, KH82, Lil72, Lin84, MLSW83, Pie75, Ste80a, TLCB98, Cla81b, LT85, Mat92, Ste81]. Packard [Fle83, Hal79, Lan83, Lou75, MBH^{+}89, PB87, RJH^{+}87, Sta87, US79, Cag86, Chu79, CHK86, RL86, Sch95, Sch78b]. Packet [Ano95-27, AS90]. Packs [BZS83]. Pad [HPL96]. Pads [EM98]. Page [Ano92-34, Coz78]. Paging [Ano96f]. PAINTBRUSH [ADRS83]. PAINTBRUSH/ [ADRS83]. Paintjet [PSS^{+}88]. Pair [Ano95l, MCD97, DE76, Oli71]. PalmVue [SSW96]. pan [BR78]. Panel [Ano95-43, HFKS95, BL87]. Paper [HASP94, Bar88, Hol82b, KA92, MM72]. Papers [Ano94s]. Paradigm [GNR93, GTM95]. Parallel [SP92, Goo91, Mor76, WJ80]. parallel-mode [Mor76]. Parameter [MiITT82, Kai91, Rus93, Spa84b]. Parameters [AI84, Ano92j, KRDS83, WWS^{+}98, Bau69, Dam74, QAKP90]. Parametric [Ban84, ES85, NK84, WB83a]. Parasitic [Das96]. Part [MRH94c, WYO^{+}87, SBM85]. Partially [Jon89, BL87]. particle [AN90]. Partitioning [Ano92-29, Lam95]. Partnering [WP94a]. Pascal [AVB80, McD80]. Pascal/64000 [AVB80, McD80]. Pass [AGO^{+}85, WA93]. passivation [Slo89]. passive [AKL69]. Password [Ano96a]. Path [WMBB94]. Patient [Ano91g, Ban91, BN80, TM91, Gol80, Gor91, Han80a, Rue80]. Pattern [CLL81, CRSY73, RHM79, YPS76].

Patterns [HPM87, Don71, Hor69]. PC [Ano93c, Cun98, Har86, Hea85, LM86, Pea85, SR95]. PCDS [RM88]. PCI [LHIR98, Mye96]. PCI-based [Mye96]. PCIB [HL86]. PE [FR95, HWK95, SKWR95]. PE/SolidDesigner [SKWR95, FR95, HWK95]. Peak [Ano92p, Ano92-33, Ket74, SSMK92]. Peak-to-average [Ket74]. Pen [Ano92-34, BLKH77, EGH^{+}88, Has92, HM85, HPH81, PHK77, Por80]. Pen-Lift [HM85]. Pentium [Mye96, WP94b]. Performance [ALT82, Ano92b, Ano92-28, BDH91, BBG91, BEM95, BCF96, BL83, CP87, Cam94, CSS^{+}91, Col88a, CWFS82, Cun98, CT85, DWD88, Fos98, FPUB88, FSDC82, GSG^{+}87, GE86, Gib86, HASP94, Let92, LRFL83, Lym87, MY95, Pea95, PMN82, RO97, Raj95, RMVT95, Rub86, SA82, WR85, ASCC84, ABD^{+}97, Bac91, BMTW78, BLTY81, BBI^{+}93, Bel93, BNB93, Bor90, BSY89, BW81, Bur85, CJHE93, Chi81, Cou71, Cra72, DW92a, DL77, DR81, EM76, EH81, ELN88, Fis72, FS81b, FHDS92, Goe89, GBHS81, HBHC84, Ham93, HKZ74, HVSM96, HE75, HW82, HMO73, Jun78, KN93, Kei71, KPC^{+}82, KW80, KSMM97, LTC^{+}84, LM73, Lin79b, MP89, Mey90b, MG92, NBZ^{+}97, NSV92, O’B93, ORB91, PW82, PB87, Pli89, QM80, Rig77, RSM91, REOB78, REB86, SJ80, SA91, Sch93b, Sch73, SMT^{+}97, Smi73]. performance [SC89, Sny75, Sta77, Sti77, SMD95, Tal68, TD83, TRH85, TC79, Wie90, WJ80, WJ79, WB89, YPS76]. period [Ban69]. peripheral [BRHS89, Bru78, Ega72]. Peripherals [GV86, MG85, Roe95]. Permuted [Mal87]. Persistence [Kah92]. persistence/storage [Har76]. Personal [Bar84, Egb77a, Egb77b, Egb77c, Egb78, HJL^{+}83, Hig85b, Kah79, KM84, MD86, RSW^{+}86, SJL91, SPA^{+}84a, Suk84, Tun74, WBSC84, WW84, BBG91, CKP89, DE76, FRAC85, HL86, LK80, LLS88, Lyn80a, Lyn80b, MHCH77, MRH80, MG92, MWC80, Sto74, SMK84, TL89, WM85]. personal/professional [Lyn80b]. personality [HFOS81]. Perspective [Ano93-29, Ano94-34, CHK^{+}95, NH92]. perspectives [Lou75]. Perturbation [Ano95o]. Phase [Bir96, Chu88, Sch84, TOB97, AH89, Hal75, HK75, Han72, KH88a, NYS81, NTA72]. Phase-Locked [TOB97]. phase-modulated [Hal75]. Phased [Ano97h, GFK83]. Phased-Array [GFK83]. Philosophy [Ano92x, AD82, Hol79, Nel82]. Phosphide [FKO^{+}93]. Photodetector [TBB^{+}97]. Photodetectors [Ano93z, Slo89]. Photolithography [ABH^{+}83]. Photonic [ICNT93]. Physical [BBF^{+}97, CCC^{+}95, Dia81]. Physician’S [Pop83]. Physiological [WJ91]. pick [Saf96]. pick-and-place [Saf96]. picoseconds [CAF78]. picture [ADRS83]. Pin [GDKK94, TLCB98, Pri76]. Pipeline [Tha89]. pipelining [Ram92]. pixel [WYSE97]. Place [Rya95, Saf96]. Plain [HASP94, Bar88, KA92]. plain-paper [KA92]. planes [HLM89]. Planning [MRH94a, FS81a, Kru97, NGW92, WSJ87]. Plasma [Ili82, Kis97]. Plate [Ano93-28, Her73, SBM85]. platen [KT81]. Platform [Cha96a, Jan96, LMNQ96, PA97, WA97, BCK^{+}94]. Platforms [CBL96]. platinum [Fos74]. Plot [Ano92c, Ano97g]. Plotter [Ano92k, Ano92l, Ano94x, DH77, Hal85, Has92, Lo81, Lon92, MC82, PT92, SMGR85, Tre78, WJ82, APA92, ABW78, Azm82, BLKH77, BSME92, BP77, FCC79, FS81b, GAP92, HFOS81, KPC^{+}82, MEM82, MSCK92, PHK77, PL81, RC79, SSMM81, TC79]. Plotter/Printer [Tre78, ABW78]. plotters [ACK79, BAUR92, Bru77]. Plotting [HM85, ABW78, BLTY81, DF86, DKMS81].

Plug [Fro86, HRH81, Mil84, PH96, GDV73, Goo72, HEW82]. Plug-In [Fro86, HRH81, Mil84, PH96]. plug-ins [HEW82]. plug-on [GDV73]. plug-together [Goo72]. Plus [Fro86]. Pocket [EMT83, CFW75, CR73, Eve83, Lil72, NT75, Tun74]. pocket-sized [CR73]. pocketful [WRT72]. Pods [DD86]. Poincaré [Ano95-40]. Point [MC84, KS72]. point-by-point [KS72]. points [NBZ^{+}97]. Polarization [Ano93-32, Ano95u, HH95]. polarization-mode [HH95]. Pole [Adc87b]. Pole-Zero [Adc87b]. Policing [MM97a]. Polyester [Bri94]. Polynomial [Ano95-41]. Polysilicon [Kam82]. Poor [Ano92-29]. port [Bro95, SM75, Spa84b, Ano94v]. portability [Ano71a, Hay74, Res83]. Portable [BJ82, DN86, ELC^{+}86, Fro86, HSYJ83, Low69, MD86, Moo92, MT76, QH83, Row86, Yor83, Bau80, BDOT85, Jek75, KB85, Mac73, Mac75, MB80, MS80, MRM83, Smi69, SM87, TD83]. Porting [Lin97]. Portion [Ano96a]. ports [Sta81a]. Positioning [BJWW84, Kus96, RW76]. POSIX [Lal93]. possibilities [Chu79]. Potential [Ano95-42]. Power [Ano92-33, Ano97e, BDH91, Bec87, Bir96, BGR^{+}88, Che86, Dan96, DR98, DBB^{+}95, DR81, FG87, HJL^{+}83, Lym87, NK96, OL83, Pie87, SS83b, Ain81, AK79, BHW77, BS90, BBT92, Bec80, BHE^{+}85, BFMP84, Cat78, CHM76, CJ71, DiF75, DP73, Edw75, Emm72, Fel86, GBHS81, Hen91, Hey81, HM71, Jac74, Kov81, Kuh76, Lam74, MP81, Pec74, PGW73, Pra75, RSM91, Rob82b, SSMK92, Sch87, SL93, Sei81, TTML81, Van74]. Power-Line [Pie87]. power-on [Kov81]. powered [GG77]. Powerful [AD82, Ban84, Hal85, McI74, MWC80, Mur82, NT75, Pee79a, WRT72]. PPA [HJH^{+}97, Kil97, MMPT97]. practical [HW73, NR72, Pou68]. practice [BLMJ77]. Precise [Edw79b, Sch84, FD78, FR87a, SFHE85, SW90b]. Precision [BKMM87, BDFL95, BHK^{+}86, BGK^{+}87, CHKS86, DS82, FSB^{+}87, GSG^{+}87, Hey81, HMSFR87, JRC81, MBF^{+}87, MBH^{+}89, Nic88, PF88, PP69, PB87, RJH^{+}87, Sor88, Swe89, Win80, WCT87, YLJ^{+}87, BHW70, BW85, BCH^{+}93, DH76, DA75, DeV81, EHS81, EAW^{+}86, Emm72, GW77, HKA79, KGK81, KV74, Lai81, LM81, Mae74, Mar89, OST75, PHK77, Pra80, Rad87, Wal76, Wes77]. Predictive [WR86]. Preheater [Sch84]. Premonitory [Hor69]. Preparation [Gor82, BH81]. preprocessor [GMMM92]. prescaler [Jek80]. Preselected [NML90]. Presence [Van98]. Presentation [MTMP83, PS83]. Preserves [Gem85]. Pressure [WP94a, Bro82, Rom91, Sch75]. preventing [BR91, Con93]. prevention [War89]. previously [Kei71]. Price [Cou71, DW92a]. price/performance [DW92a]. priced [KE74]. Primer [Ano94z, Ano90b]. Primitive [Ano94-29]. Print [Ano92g, Ano92h, Ano92u, Ano94s, Ano94x, BJJN88, CAJ^{+}94, DMW94, BMTW78, BCK^{+}94, HSB^{+}94, MH92, MJKS92, Mon92, RL92]. Printbar [Cra85]. Printed [Dul85, Gas88, Jac88, Mar88, Pal96, Reg88, BL87, Bas79, FL80, HDH^{+}87, Rus92]. Printer [BTBW83, Bar84, Cas95a, CHP85, DiV84, GLW85, HW88, HSW88, MLPN88, PSS^{+}88, SS82, STMM88, TCWN88, Tre78, WMBB94, Woo78, ACM92, ABW78, Bar88, BD73, BTE^{+}94, BP76, Coz78, DMW94, DCN^{+}88, HSB^{+}94, HJH^{+}97, Hol82a, JKP^{+}88, KB85, Kil94, Kil97, LT85, LK80, MJKS92, MMPT97, MT76, Nad75, PPMW94, QM80, SMT^{+}97, SM87, Ter78, CH82, Hol82c, Rho88]. Printers [CHP85, Fit82, MTWG85, PL84, Spr82, Bri94, HWS87, Jep85, WM94]. Printhead [ABC^{+}94, BA85, BHH^{+}88, HW88, Nie85, Ric97]. Printheads [BCRS88, HCF^{+}88]. Printing [Bar88, Dei94, Hal82, LH82, Man94, Mar77, WCH94, ABW78, BCC^{+}94, BP76, ISLW78, KA92, SMT^{+}97, Woo78].

Priority [ASTW95]. probability

[AP69, RR69]. Probe [REB86, AH72, EZ81, Ham93, Mat90, Que76, US93]. probes [FEG77, GW77, OST75]. probing [Zel69]. Problem [Ano97c, HMPF74]. Problems [OH82b, Pet94, Chu74, Cro79, Due74, SWM78, Wag75]. Procedural [LBKY86]. Procedures [Ano92j, Fan72]. Process [Ano93-36, Ano93-43, Ano94r, Gra96, Jac88, LeV88, MFMS83, Oh82a, PT92, RNR^{+}98, RHK^{+}83, SDDC96, Chi82, Cla81a, Con93, DL77, Gil96, GHK^{+}93, GCDW91, PW82, SM88, SM82]. processes [Dia94, Far74, LD93, Tiv91]. Processing [ACK^{+}84, Ano94-27, BK85b, Gup95, Hig85a, HLH86, IC85, LS76, LMNQ96, Shr82, Slo89, Bel93, BCH81, Gro92, Gun94, HMLN78, Hue73, KB69b, Kla78, MM90, RW81, RP88, Rob92, REOB78, Ste80b]. Processor [ADG^{+}96, BBR^{+}97, DMJCB97, FCK97, GJW83, MGBB97, SM86, US79, Bee82, EM76, Fel74, Fia83, For77, GJ74, Hea85, JU97, Koh72, KSMM97, MBF^{+}87, OGB84, PB76, RH74, Rue80, WT97]. Processors [LBLS95b, SBK^{+}97, LBLS95a]. procurement [Kru97]. Produced [Ano92k]. Producers [Ano96w, Ano96p]. produces [For77]. Producing [Cor96]. Product [Ano92x, Ano94u, Ano94b, Ano94y, Ano95-43, CL98, DeV81, Hol95, Kra89, RR88, Roe95, ROKH94, SSW96, TBBM83, GCDW91, Jes89b, LF91]. Production [Ano92-32, Dul85, MY95, MRH94a, SW91, Wri96, Ano67, EAW^{+}86, HCF^{+}88, MT81, Mon92, Pat83, Ric81b, SM79]. productivity [EAW^{+}86, Gra81, O’K93, Reh84, Ric81a, SB83]. Products [Ano95x, DR95, HMM96, BB84, DP84, Res83]. Professional [KE85, Lyn80b, WM85]. professionals [Eve83, WAZH84, Wic87]. Program [ABC^{+}88, Ano96y, Bal84, Lon90, Mag94, McD84, McD80, Ped86, Cle79, CG87, DE76, HM88, Kha81, Mac93, MOT90, Ste89a]. program-compatible [DE76].

Programmable [Ada71, GBO94, Kal93, MSS87, MC82, TST82, And70, Aue81, BHE^{+}85, BKE^{+}90, CS76, Chi81, CR81, DF82, DE76, FR87b, HEW82, HKA79, Ing76, KP90, Lou72, MCCB76, Ris82, SFHE85, Sch78a, SF95, Spa72b, Tho80, Tun74, Vet91, Wag93]. Programmer [Ing70, SB83]. programmers [Tan89]. Programming [Ano92-27, Ano95-35, Atc92, Bar81, ME90, Ric81a, Sto74, BP77, Bur86, Coc68, Hal79, Joh77b, KM79, Mat84, Pir97, SKB78, Sha89, Zel70]. Programs [Ano92s, Bas79, HKC^{+}84a, Neu91]. Progress [Ano95h]. Project [Ano92r, Ano95-44, Kru88, Mar88, Mil96, MK97, Ric91]. Projects [Ano95-45, FS94]. Protecting [Ano93-33]. protection [Nie90]. Protects [Ano92w]. Protein [SD95, MB96]. Protocol [ASTW95, AW90, Ano90c, Ano95-31, BKD94, JRM^{+}86, KYKM85, Moo92, PT90, WR85, ARGA85, BDOT85, Dou92, Pru92, SSV90, SP96, Voi84, Pug90]. Protocols [Ano95t, DKW90]. Prototype [Roe96, Sch96]. Prototyping [Ano95-44]. Provides [Ban84, CHP85, Hal82, Hal85, WR85, Bol75, Lai81, Leo77, MW80, QM80, Sta88, SA80, TC79]. Providing [BLY92, BK95, Nie90, Tan89]. Proximity [RW81]. Pseudo [AR76]. Pseudo-random [AR76]. pulmonary [BF79]. pulsars [She71]. Pulse [Ano93-27, HZ83, Kal93, Aue81, BHE^{+}85, BKE^{+}90, Cat82, FL73, FD78, FRN83, GZC74, HRZE77, HLT^{+}83, HKA79, KNF^{+}97, KW90, Koe91, KE74, Rus93, SES90, SPD^{+}93, SFO80, SC89, Wag93]. Pulse/Data [Kal93]. pulse/function [FRN83, HLT^{+}83]. pulse/word [HRZE77]. pulsed [LS83]. pulser [AH72, BC76]. purity [DeV81, Has81, PB81, Ric81b, SS89, Win80]. Purpose [KM97, ADRS83, BLMJ77, CF68, GOS84, GT71, MC91, Ogd80, OGB84, Sha76]. Putting [BBD^{+}84, Kar82, RS75].

quadrature [KH88a]. Qualification

[SH90a]. Quality [Ano92a, Ano92j, Ano94s, AY98, DiV84, ES87, Hal82, Hig85b, HASP94, Mar88, MO74, MW94, SW90a, SC83a, War86, WHBP95, Azm82, BLKH77, Bau80, DCN^{+}88, FCC79, GHK^{+}93, HSB^{+}94, HMZ82, MB74, Sta81b, WM94]. Quantitative [Kin83]. Quantum [RLB^{+}95, RWT97]. quarter [Gil89, Top89]. quarter-inch [Gil89, Top89]. quartz [BW81]. quick [BCH81]. quick-turnaround [BCH81]. quiet [BD73, Cos76, Nad75].

R [EAW^{+}86, Mac93, QAKP90]. R&D [LeV88]. Radar [KH88b]. Radially [HPL96, PHL96]. Radiated [SS83b, WB89]. Radio [Duf87, HPM87, McK87, Ped86, Fis72, Rob82a, Wat87a]. radios [Hoo91]. radiosity [Bur89]. Raid [Ano95-38]. RAM [WSBK83]. random [AR76]. Range [Ano93b, CSWR85, EMT83, FR87c, Uph88, Yan92, BS93, Bol80, Cat82, CHM76, DA75, DiP80, FRS88, GDV73, HT76, Jer74, JRC81, LS76, RG69, Tho77, Tut72, Vif76, Zel79]. Ranges [Ano94-27, Wad68]. ranging [BHW77, YK77]. Rapid [Ano95-44, Gro79, Wha68]. Raster [BR78, Pra80, Roy75, SA80]. raster-scan [Pra80]. Rate [Bru97, Moa98, War97b, Wit92a, Kri81a, YWP^{+}92, PN88]. Ratio [Ano92f, Hef73, Ket74]. ray [Bre74, Cha70, GN82]. raytraced [SP92]. RCS [Goo91]. Read [Gre85, BP72, BB84, Fin72]. read-only [Fin72]. read/write [BB84, Fin72]. Reader [AT83, LB80, Tag74, Zel70]. Readers [SMH86]. Reading [CR81, GO72, Goo77, ULW81, Vif76]. readings [Ste89b]. ready [Han79]. ready-to-use [Han79]. Real [Ano93-34, Ano96u, Ave75, Can93, DH98, GWY94, Gun94, KB69b, Kel92, LBLS95a, LBLS95b, MC84, Sch76, ASCC84, AP69, AG75, Anz93, BRHS89, Bha92, BNB93, Dra77, FGH^{+}89, FMN84, HR91, LM93, Mor93, Rus93, SH77, War91, HKC^{+}84a]. Real-Time [Ano93-34, Can93, DH98, GWY94, Kel92, LBLS95b, MC84, Ave75, Gun94, KB69b, LBLS95a, Sch76, ASCC84, AP69, AG75, Anz93, BRHS89, Bha92, BNB93, Dra77, FGH^{+}89, FMN84, HR91, LM93, Mor93, Rus93, SH77, HKC^{+}84a]. realistic [MN79]. Realizing [O’K93]. reassociation [San92]. Receiver [Arm95, DS82, MKW^{+}96, YN93, Bak84, FRMK85, Mai93, San72]. Receivers [BS95, ES85, ALRS89]. recognized [Hor69]. Reconstruction [PF88]. Recorder [Ban91, BF82, FPUB88, PMN82, PBD88, San86, Bru73, GY88, HJW75, MW73, Mic71, Pat83, Tal68]. Recorders [Sor88]. Recording [HRE85, KH88b, SA82, SK69, SD82, Wad68, MM72, Pax70, Spo90]. Records [SC83b]. Recovery [Bar95, Gre85, WSYD97, Ran88]. Recrystallized [Kam82]. Red [SDC^{+}88]. redesign [Kil94]. redesigned [Jon81]. Redfield [KH88a]. Redfield-Kunz [KH88a]. Reduces [AD82]. Reducing [Byr96, Cle79, McD84, Pal96, WB89]. Reduction [Hoe97, AH89, Ano67, CDS71]. Redundancy [Ano95-41, WSBK83]. reel [BKV89]. reel-to-reel [BKV89]. Reengineering [Ano94m, GAP92]. Reference [RL86, HHS76, Ras87, RU76]. Reflection [DBB^{+}95, CS93]. reflective [Jon89]. Reflectometer [BDFL95, BGG88, Bel93, BP93, BCH^{+}93, FRS88, RP88, Vob88]. Reflectometry [MS95, MD96, SNN88, Smi69, War72]. Refractory [RHK^{+}83]. Register [Ano92-31, Ano93-39, San92, TD92]. Register-Based [Ano92-31, TD92]. registration [gJH81]. regulators [DP73].

Related [Ano95-45, MVG77]. Relational [HKH93]. Relationships [Sha98]. Relay [Dub94, TD92]. Release [Han94]. Reliability [BKAWBCTK98, DGHP85, DW87, Gil89, Kru88, Reh88, BAUR92, CS80, FCC79, KPC^{+}82, Kru89, LT85, Sha76, Sti77, Wat81]. Reliable [BDM85, BFMP84, DPA^{+}88, Dou75, FR87a, MC85, Sol81]. Remote [Ano93-35, Ano95-46, BGR71, BP77, But94, Hig85a, Bur93, Nel79]. Remotely [Bra78]. Remotely-controlled [Bra78]. removable [Ric76]. Repeatability [SC83b, DJS83]. report [CP79]. Reporting [Pet94, Ste74b]. Reproducibility [Jeg95]. Requirements [BF82, CS82, Bec79, NGW92, SAMS73]. requiring [Thr68]. Resampling [Ano93-36]. Research [Ano94n, TZ93]. Residential [Pit97]. Resist [OLR82]. resistance [Fos74]. resistances [KB71]. Resolution [Ano92-36, Gar92, Mal87, Bol75, BHH^{+}88, Car78, CS93, CK76, Coz78, Cze89, Goo77, HMW83, Mac73, PF78, Roy75, Sch70, Smi69, SMR^{+}86, WK93]. resolves [CAF78]. resonators [CE81]. Resource [DeV80]. resources [BCH^{+}89]. respiration-rate [Kri81a]. Response [Adc87a, Ano92-32, Kjo92, PH96, Gro92, OL83, Zam73]. responses [MM79]. responsibilities [BG89]. Responsibility [Ano94y]. restrictions [HJW75]. Resuscitation [BJ82]. Retargetable [IGHG^{+}94]. Retention [EGH^{+}88]. Retrofitting [RB82]. Return [Gra96, Sch93b]. Reusable [Cor96]. Reuse [Ano96p, Ano96w, Buf96]. Review [Ano92-34]. rewritable [SWP90]. RF [AH89, Ano95f, Bra78, Bur85, But95, Cra72, CGH^{+}93, DH76, DGN98, GHUW68, HH78, Inh75, KV74, LK85, SCI^{+}81, Ton85, YN93]. RF/ID [Ano95f, But95]. Rich [CHW96]. Ridge [RLB^{+}95]. Rigorous [BR91]. Ring [BKD94]. RISC [ABG^{+}92, Ano92s, Ano92-28, Anz93, BW85, BLY92, Bre94, DW92a, FCK97, FHDS92, GJM^{+}92, Han92, Let92, Pea95, Ram92, San92, SBK^{+}97, UAH^{+}92]. RL [Ros88]. RMON [Bur93]. rms [Wad68, Fol72]. robot [GRS93]. Robotic [Gas88]. Robust [Ano95o]. role [O’C86]. Roles [Ano96w, Ano96p]. ROM [CKP89, GP84, KGTM90, Mey90b, Nie90, SS90, Wol83]. Roms [Fro86, Lei71]. room [GG77]. roots [Egb77a]. rotation [WK93]. rotational [HHR71]. Route [Rya95]. routed [BL87]. routine [HHR71]. Routing [Ano97b, Ano97c, FCK97]. RPN [Wic87]. RT [Ano93i, Mor93]. RTE [WM78]. RTE-IV [WM78]. rubidium [Thr68]. rubidium-vapor [Thr68]. rugged [Bru73, Low69, US93]. Rule [GV86, DD97, WRT72]. Rule-Based [GV86, DD97]. Run [Ano93-37, LTC^{+}84]. run-time [LTC^{+}84]. Run-Time-Configurable [Ano93-37]. Running [BGK^{+}87]. Runway [Fid96].

s [GLP^{+}97, WSYD97, YPS76, Spa84b, CDLN96]. S-parameter [Spa84b]. S1010A [Ano95-43]. SABER [SWI97]. Safe [Ano92u]. Safeguarding [Har84]. safety [Nik97, Pir97]. safety-critical [Nik97, Pir97]. Salicide [Ama86]. SAM [Ano93-38]. Sample [Bru97, Lar83a, PN88, Sch95, War97b, Wit92a, Goe89, KP90]. Sample-and-hold [Lar83a]. sampler [Mer73]. Sampling [Ano96h, KH84, MB93, Swe89]. Sapphire [For77, Roy81]. Satellite [YN98, Wil78]. satellite-linked [Wil78]. SAW [Dia81]. SAWR [BC81a, MT81]. SC-cut [BW81]. Scalar [Uph88, EAW^{+}86]. Scale [Col86, Ano71b, FG73, Joh80, Van77b]. Scaling [Ano94-28, Van97]. scan [BR78, LHL83, Pra80, Roy75]. scanned [Bla75]. Scanner [GFK83, Kar86, GS93, SMR^{+}86, ULW81, WYSE97, Wol75]. Scanning [KV94, RK94a, WP94a, WHBP95].

Scheduling [LeV88]. Schematic [Bur95]. Scheme [Moa98, Cat82, McK83]. Schottky [LM73]. Schottky-barrier [LM73]. Science [Ano92h, BBD^{+}96, Eve83]. Scientific [BPA^{+}94a, EMT83, Geb78, MT76, Smi80, SMB^{+}91, WP91]. screen [Sta81a, Sta88]. SCSI [Ano91d, Ano92-35, TBTM92]. Second [BB84, DWD88, HGWS97, MPR82, ACS88, CPK88, Goe89, KT73, Mil88, Ste89b, Hil93a, MB93, Sch93a]. Second-generation [BB84, ACS88, KT73]. secondary [SWP90]. Section [CSWR85, Ton85, CCH^{+}91, KW90, PB81, SES90]. Secure [Kon95]. Security [GH95]. selectable [McK75]. selected [Fra74]. Selection [Tan96, RL92, WA93]. selective [Cos76, Tho80, Urq76, Wal76]. Self [Gol80, RK94a, DL73, Foo80, Kov81, LaF80, Sei91]. self-contained [DL73, Foo80]. Self-Test [RK94a, Gol80, Kov81, Sei91]. semiautomatic [CRSS77]. Semiconductor [AI84, Col86, Dia94, Hig85a, MiITT82, Mai91, Fin72, Fra74, Jac74, TW76, WSJ87]. Sensing [Ano93u, Hol82b]. Sensitivity [Ano93e, Ano93-32, BDFL95, BS93, CS93, GTM95, HJW75, Kal95]. Sensor [FST92, Mer81, PS81, RSM91]. Sensors [Col88a, Dia81, KNF^{+}97]. separation [BBW95]. Sequence [MRH94b, Don71, MB96]. Sequencer [MBC95]. sequences [Kov81]. sequencing [MB96]. Serial [KYKM85, MBC95, Erd79, Kil94, Mor76]. Series [AFG^{+}86, Ano92e, Ano92q, Ano95v, BWW84, BHW86, CP87, Cat78, Geb78, Har87, Hol79, KM84, KH82, MK79, Pal86, Sha76, WLHC84, YF92, ASCC84, Fra74, GMMM92, HJG93, HEW82, Joh77a, KG92, Lan77, Rig77, RSM91, Spa72b, Sta77, WS84b, Wya92, ABG^{+}92, CL92, CLL^{+}92, EDF82, Har82, LG92]. Server [Ano93h, Ano94-33, BEM95, HKH93, Kon95, LH95, Pet96, Swe96, NBZ^{+}97]. Servers [ARL^{+}94, Ano96b, BBF^{+}97, Bre94, BCF96, SR95, HVSM96]. Serves [MG85]. Service [Ano96j, Ano96k, GH95, Wen85, ARGA85, BDOT85, CDE^{+}87, Fro77, MK79, PS81, Pie75, Wri96]. serviceability [Ano71a, Gol80, Sha78a, Swe81]. Services [Bar96, BK95, Cas95b, FKL86, FT86, KT95, MS94, She96, Chi90, CFK^{+}90]. Servo [MW80, Lo81, TRH85]. Servo-controlled [MW80]. Servomechanism [Edw84, OB90]. Set [CHP85, DS82, Ern95, KGH^{+}87, MC84, AD75, Cos76, DVl91, Fia83, Geb78, GP84, Rob82a, Spa84b, Urq76, Wal76, Wat87a, DNG84, NTH^{+}84]. Sets [SC83b, Mar77, MWC80]. Setting [RR88]. settling [Pon80a]. Setup [Hol84, Kru94a, Pal96]. setups [TST82]. Seven [SMR^{+}86]. Seven-function [SMR^{+}86]. severe [HHS76, Van74]. SFC [Ano94-37]. Shadow [Ano93-39]. shaft [ELU81]. Shapes [Ano97g]. Shared [Ano93-33, CR92, Mar89, Mun77, Pou68, Wal70]. SharedX [Ano94-38, GWY94]. Sharing [BCH^{+}89, BG89, ES89, HLM89, ACL89, DeV80, Pet68]. sharper [Har76]. Shielding [Ano92w]. shifted [HT86]. Shmoo [Ano97g]. Short [Ano94z, Hen84, NMY95, Wes77]. Short-Circuit [Hen84]. Short-Wavelength [NMY95]. Shortening [MY95]. Shortens [Duf87]. shows [Pri76]. Shuttle [LT85]. Si [Bro82]. Sidebar [Ano94-27, Ano94-28, Ano94-29, Ano94-30, Ano94-31, Ano94-32, Ano94-33, Ano95-47, Ano96q, Ano96r, Ano96s, Ano96u, Ano96t, Ano96v, Ano96w, Ano96y, Ano96x, Ano96z, Ano96-27, BPA^{+}94b, Dei94, Fid96, Kru94b, Mag94, NK96, Pet94, RK94b]. Sigma [Hea94]. Signal [Ano91f, Ano92f, Ano95v, Ano97d, BW92, BMD^{+}88, BS98, CS82, CSWR85, Edw87, EH84, EEH^{+}84, Eps84, FHM^{+}85, GY88, GC87, Gro91, HZ83, HHT^{+}97, HMLN78, Hue73, IC85, JM87, JJ85, KWR^{+}87, LK85, Lau87, Mur97, NVBH87, PK84, Shr82, Ste98, Str77, Ton85, Wen85, WR97, da 94, Ain81, AH89, Ast77, BBDK93, BBI^{+}93, Bel93, BCO^{+}93, BKS93, Bru79b, Bur85, Chi81, CK76, CGH^{+}93, DF79, DO77, ES77, FFS89, GW77, Gro92, Gun94, Hal75, HK75, HHS^{+}73, Has81, Hay74, HS68, HPF93, Jer74, Kih82, Lar83a, LS83, MP89, Mer73, Mil90, PCW77, PB81, RU76, RR69, REOB78, SKB78, SCI^{+}81, SFHE85, Sha71, SAMS73, SS89, Tho77, Tut72, WB89]. signal-level [RU76]. Signal-To-Noise [Ano92f]. Signaling [Ano95-30, NZ97, CCC^{+}95]. Signals [Ano92w, ANSW95, Arm95, Bir91, HKM96, ABM73, AR76, BCO^{+}93, Chu88, Edw79b, GW77, LS76, RD75]. Signature [Fro77, HR82, Nad77, RB82, Cha77]. significantly [QM80]. Silica [LRS83]. Silicon [Col88b, For77, Kam82, WSYD97, Cow70, Roy81]. Silicon-On-Insulator [Col88b]. Silicon-on-Sapphire [For77, Roy81]. Simple [Ano94-34, Hal85, MRH94a, MRH94e, RC79, WR85, BP83, BP77]. Simplicity [Ame85, Woo78]. Simplified [Gue74b, RB82, Gor69]. Simplifies [Eps84, Hol84, SHB84, Ing76, NTH^{+}84]. simplify [TST82]. Simulating [TOB97]. Simulation [AGO^{+}85, Ano92e, Ano92-28, Ano92-30, Ano94m, Ano94n, Ano94-35, Bur95, KYKM85, MRH94e, SWI97, Cho90, MMNR89, OB88, SDA81]. Simulations [Oh82a, NE81]. simulator [Par79]. simultaneously [Pri76]. sine [Lee74]. Single [Mur97, YLJ^{+}87, AH89, BS90, Fia83, For77, Lo81]. single- [BS90]. Single-Chip [Mur97, YLJ^{+}87, Fia83, For77]. single-loop [AH89]. Site [Har87]. Six [Hea94]. Six-Sigma [Hea94]. sized [CR73]. slave [Reh84]. Slice [Loo97]. slide [WRT72]. slow [HW73]. Small [Ano92-31, BZS83, Per88, Van77b, WMBB94, Fra71, Ham79, HB72, Joh77a, Mat79b, McI74, MD79, Rom91, Smi79, UAH77, WN79a]. small-business [Ham79, Mat79b, MD79, WN79a]. Smaller [NT75]. Smalltalk [Ano92z, KF95]. SmartClock [Kus96]. smoothing [MM79]. smoothness [PHK77]. SnapLED [Ste98]. SNR [Ano93y]. Sockets [AGI^{+}91]. Soft [Mey84]. SoftBench [Ano93k, Bet97, Cag90, Car95, Cou94, Lie97, Wil97a, Wil97b]. Software [AGI^{+}91, Ano91a, Ano91b, Ano91c, Ano95x, Ano96y, AY98, Bam83, Ban84, BLNP86, BNB93, BM86, Bug86, BL83, CHK^{+}95, Cor96, Cou94, Cox96, DD86, DKW90, DW87, EH84, Fit82, Fos73, FS94, FS86, Gem85, Gra96, HSK87, Ham81, Hea94, HNL87, HKC84b, Hig95, Kru88, LBLS95b, LCPG98, Lon90, LC96, Lub91, LKHK83, MB85, Mar88, MZ96, Mil96, NZ97, Neu84, Pal86, Par79, Ram92, RW70, RC98, TBTM92, TE86, War86, War89, War92, Wat87b, WS84a, BR91, BH92, Cag90, Con93, CBH91, Fel80, Fis88, Ger90, GHK^{+}93, Gra89, Gre90, HJH^{+}97, HBHC84, Kra89, Kru89, LBLS95a, Leo83, Lie97, Mac93, MP83, NSA89, Nie90, NSP97, Nik97, NPG^{+}83, Rei91, Ric81a, Sch79, Sch93c, SPA^{+}84a, Tem83, WN79a, War91, Wil97a, WB83b, YGR95]. software/firmware [MP83]. Solder [Kee89]. Solders [MHV96b, MHV96a]. Solid [Ano94w, BP69, CS82, CWFS82, Ern95, BEFV70, Det72, Kei71, Kih82, Kuh76, Mic71, SAMS73, UAH77]. Solid-State [CS82, CWFS82, BP69, BEFV70, Det72, Kih82, Kuh76, Mic71, UAH77]. SolidDesigner [FR95, HWK95, SKWR95]. Solution [Adl86, Ano95x, Man94, YGR95, Per86, TSv92, WSJ87]. Solutions [LEV96, MB85, Pet96, Chu74, Cro79, Sel97, SKG90]. solve [HMPF74, Kah79, Vog91]. solvent [GV84, SW90b]. solver [McC87]. Solves [OH82b]. Solving [FCK97]. SONET [WSYD97]. SOS [Edw79c]. Source [Alb88, Ano93-46, Ano97e, BGR^{+}88, GLP^{+}97, Bru79b, CKP89, DH76, DF79, JBS93, Pon80a, Pon80b, Ras87, RD75, SA91, SKB78, SFHE85, Ste89b, Tut72]. Sources [Bar82, FR87c, Man69, RLB^{+}95, SR88, ALRS89, BEFV70, Mai91, MLJP93, Pee73, PP69]. space [Fra71]. spatial [Rus92]. speaks [JY72]. Special [Bru79b, GJ74]. Specialized [Gor82]. Specific [Gro91]. Specification [Bur95, CBH91, KHW89, LF91, LCB90]. specification/800 [LCB90]. Specifications [RR88]. Specifying [GR91, HR91]. speckle [Mag83]. spectra [KB69b, Wha68]. spectral [BS90, NML90, Ric81b, SS89, Win80]. SpectraLAN [LAB97]. Spectrometer [Kis97, Wak97, HHR71, KT73]. spectrophotometer [KW80, MW80, Ste80b, WJ80]. spectrophotometric [LNH^{+}84]. spectrophotometry [SW80]. spectroscopy [FK79]. Spectrum [Ano93-30, Mat86, McF98, BS93, BHW70, BEFV70, CCH^{+}91, Chu88, CM72, CW72, DA75, Elo93, GHUW68, HW73, Haw71, HH78, HMLN78, HS80, KN93, Kei71, LD79, Lin79b, MW78, MB80, PF78, REOB78, Sha72, Sha78a, Ups71, WA93, YN93]. Speech [HM84, KM84, Ano67]. Speed [Ano93z, Ano94-27, Ano95p, BMD^{+}88, Gor82, HM85, KYKM85, KRDS83, LB98, MTWG85, MCD97, MBD^{+}98, Mil94, MM95, Mon88, PHK77, PBD88, RNR^{+}98, TBB^{+}97, Yan92, AI84, BD73, CRSY73, EHS81, EH78, FRS88, For77, Geb78, HBNB83, JRC81, JM91, KGK81, KBHS93, Koe91, LNH^{+}84, MVG77, Mil90, Pax70, Pli89, Slo89, SA80, TTML81, WKT83, WJT71, WB83a, WHHC89, WHH91]. Speeds [MC84, BW76, Par79]. Sphere [Ano95-40]. SPICE [DR98]. Splitter [Ano97e]. Sponsor [Ano94-34]. spool [AG75]. spooler [Lie88]. spots [EZ81]. spread [Chu88]. spread-spectrum [Chu88]. SPU [Rub86]. Spurious [BDFL95].

Spurious-Free [BDFL95]. Sputtering [DLJ85]. Square [Egb77a]. SS7 [PA97, WA97]. stability [CK76, CF68, Mai91, Mar71, Ras87, Thr68]. stabilization [SL93]. Stable [FR87c]. Stack [AW90, BE90, SSV90]. Stage [KM97, NK84, Aue81, LM81]. Staggered [HPL96, PHL96]. stand [DF86, YK77]. stand-alone [DF86, YK77]. Standard [AT95, Ano92s, Ano92-35, CHP85, Des92, MY95, Rya95, CL92, DP84, HB88, Hal93, Jes89a, KGTM90, Thr68]. Standardization [Ano97h]. Standards [BV85, CHNJ97, LHIR98, MBD^{+}98, GSVS89, HMO73, Lin79b, Mar77, MBL95, MWC80]. Starbase [ACL89, BCH^{+}89, BG89, ES89, HLM89]. Starbase/X11 [ACL89, BCH^{+}89, BG89, ES89, HLM89]. State [CS82, CWFS82, Far74, Rip87, Sim94a, Sim94b, BEFV70, BP69, Det72, Fel74, Haa78, Kei71, Kih82, Kuh76, Mic71, Mor76, MH80, Ogd78, Ogd80, SWM78, SAMS73, SM75, SD78, Smi77, TRH85, UAH77]. State-Of-The-Art [Rip87, TRH85]. states [Pri76]. static [Bet97, DD97]. station [BN80, BW76, Gor80, Gra81, Han80a, Ker80, Sta81a, SW81]. Stations [Ano94-32]. Statistical [Cho90, KP88, Kaf88, RR88]. step [Ada71, KV74, MM79, TC79, Vet91]. Stepper [Ano94-35]. Stick [Ano93u]. still [Ups71]. Stimulus [Ano92-32, ES85, Kjo92, OL83, TST82]. stimulus/measurement [TST82]. Stimulus/Response [Ano92-32, Kjo92, OL83]. Storage [BGS83, DSB94, HSYJ83, LEV96, Sha85, SR95, SD82, ES77, EL82, Han80b, Har76, Ike83, Mun77, NSW76, Ric76, SWP90]. Store [Ano96t, Cou72]. straightness [BGH74]. Strained [RWT97]. strategies [HG81]. Strategy [CGS^{+}86, Wri96]. Streaming [CDG^{+}88, DVE85, RRC^{+}85, DPA^{+}88]. Streamlines [MiITT82]. Streams [MBC95]. Strengthening [AY98]. Stress [Har84]. stresses [Hay74]. strip [MW73, Mic71]. strip-chart [MW73, Mic71]. Stripchart [Ban91]. stripline [Kir71]. Structure [Ano93-40, NMY95, ACS88, BA85, Ega72, Sta88]. Structured [Bur95, FHMP97, Kil94, BRHS89, BM81, CBH91, FJ89, Joh77b, LM73, LM93, NSA89, Neg81]. Structured-Custom [Bur95]. Structures [Ano93-33, Sha89]. Structuring [WYO^{+}87]. Stuck [Ano95g]. Stuck-At [Ano95g]. Study [BKQW95, BB91, LHIR98, BL87, BF89, Cho90, LM86, Rus93]. stylus [Car81]. sub [PS81, Sol81]. subassembly [CS80]. Subharmonic [Mat86]. Subject [Ano94c]. Submicrometer [Ama86, NSMO98]. subnanosecond [HRZE77]. substrate [LM81]. Subsystem [Ano91d, DDT83, HLH86, ST84, Kas72, LMM84, LWBB84, LG92, McH87]. Suite [OR96, HD98]. summary [AK81]. super [VBK89]. super-blocking [VBK89]. Supercritical [NH94]. superior [Mai91, Thr68]. Superscalar [SBK^{+}97]. supersystem [MS74]. supplies [DP73, GBHS81, MP81, OL83, PGW73, Sei81]. supply [Ain81, AK79, BHW77, Bec80, BFMP84, CJ71, DiF75, Emm72, Fel86, Kru97, Pec74]. supply/amplifier [Pec74]. Support [Ano95n, Ano95p, BK85a, BFS94, DSW83, KGTM90, Ric94, SH90b, Tan96, WR86, AS90, Goo91, KG92, MS83, Ric91, SW81]. Supporters [Ano96w, Ano96p]. supports [HKC^{+}84a, Kin81, Van77b]. Surface [Ano96k, BKAWBCTK98, CE81, IKS81, ME95, Shr82, SD95, THHW95]. Surface-acoustic-wave [CE81, IKS81]. Surfaces [TLCB98]. Surgical [Can93]. survey [BW76, LD93]. Sustained [War97b]. sustaining [Mac93]. SWAMI [Chi82]. sweep [DF82, HEW82, KS72, MLS79].

Sweeper [BDH91, BBH91, Det72, HE75, PAD70, Sei91]. sweepers [BCZ93, HJG93, Kuh76, ORB91]. Sweeping [BEFV70]. sweeps [HW73]. swept [Bot71, DY75, KS72]. swept-frequency [KS72]. switch [Bra78, FR87b, Rob82b]. switched [Zel91]. Switching [Ano93i, Ain81, AK79, DP73, Has81, Sei81, Yan93]. Symbolic [Pat87, BH81]. Symmetric [BBF^{+}97, Bre94, HVSM96]. synchronous [Hol82c]. Synergy [Ano97f]. Synlib [Cha96b]. Synthesis [Lam95, AH89, Ast77, COO^{+}93, HBB88, KMT88, Sha71, SGC93, SS89, Tal88a]. Synthesizable [Lin97]. Synthesized [HK75, HHS^{+}73, HZ83, BBI^{+}93, BBH91, DF79, Has81, MLS79, ORB91, Sei91, Tho80, Tho77, Tut72]. Synthesizer [Ake85, AS85, Ano93o, Bar82, BP85, CLG87, FHM^{+}85, KH88b, Cos76, FD78, HK88, Hig88, Ive88, McH87, SKB78, SM88, Spa85, Str77, Tho88, Win80]. synthesizer/level [Win80]. Synthesizes [Adc87a]. System [ACK^{+}84, AGO^{+}85, Ano91a, Ano91b, Ano92-27, Ano93v, Ano93-34, Ano94k, Ano94-33, BDH91, Ban84, Bec80, BZS83, BLNP86, BJPH95, BEM95, BSY89, BLSM83, BGK^{+}87, BBD^{+}84, BK85b, BKO87, Byr96, Cam94, CHK^{+}95, CHKS86, CHW96, CBL96, DF91, DSW83, DR95, Des92, DLJ85, GE86, GV86, GNR93, GJW83, Hal82, HSYJ83, Hoe97, HKH93, HL85, HR82, HMM96, HSMA^{+}86, JAD83, Kar82, KSV78, KE85, Kus96, LSM^{+}95, LRFL83, LW82, LKHK83, MM90, Mar88, May88, MB93, MFW87, MPR82, NG87, NZ97, NTJ^{+}91, NK84, NK96, Not93, OMD83, PDW85, RM88, Reg88, REB86, Sch93a, SW91, ST84, SLM^{+}95, SHB^{+}82, Tan96, Van98, Wat87b, WJ91, Wes91, WCH94, WW84, da 94, AC93, ACL89, AG75, Ave75, BB80, BLKH77, BJM71, Ban83, BT88, Bel93, Bia90, Bla86]. system [BT97, BCH^{+}89, BM81, BC74, BHP85, BG89, BOOS81, BFP90, Cam82, CLL81, Car69, Car79, CKP89, CR89, CPR75, CH82, COO^{+}93, DF86, DeV80, DGM80, EW78, EHS81, Edw79c, EAW^{+}86, ES89, EKW93, Ewy69, ES69, FRAC85, Fan72, FM84, FS81a, Fin72, Fle83, Foo80, FG73, FFS89, FH80, Ful83a, Gar83, GV84, GOS84, GR91, Gor91, GB87, GT71, Gun94, HG79, HB88, HA92, Ham81, HVSM96, HW91, HPH81, HRZE77, Her73, HLM89, Hig88, HW82, HS80, HMM73, gJH81, JM93, KGTM90, KG92, Koe80, KSMM97, Kur81, Lai81, Lan75, LT85, LDNT72, Lar83a, LS83, LP92, LMM84, LH82, LM81, LE83, LD93, Lou72, MM79, Mar71, MW78, Mat79a, MS82, MVG77, MB96, MS80, Mor93, MWC80, NR72, NSV92, NS88, NPG^{+}83, OY88, OGB84, OL83, PHK77, Pee79a]. system [Pee79b, PSA83, Per88, Pli89, Pou68, QW83, Ran88, Rob92, RO86, RS70, SIBH88, SK80, SW90a, Sch78b, Sha72, Smi73, SC89, SH77, SWP90, Ste80b, STE90, Van77b, Van74, Wag88, WK93, WMZ84, WDDH84, WB83a, WM85, Win81, WM79, WM78, Won79, WCT87, YPS76, ZHN83, BSY89, EL82, HCW^{+}82, JF80, Kai91, Kar82, LH82, Rei91, Ris82, Tiv91, WH91]. Systems [Adc87a, Ano95f, BK85a, BP82, BT98, BKO87, CSS^{+}91, Che86, DH98, DU82, Eri92, GLP^{+}97, GLM95, Gre85, Jen98, KHI82, Lan94, MM95, Reh88, RL86, Ros88, ST83, WR97, Wol86, ABJ78, Anz93, Bac91, Ber88, BNB93, Bra78, Cha91, CM75, CS78, CRSY73, DiF75, Dic74, ES77, Edw75, Erd79, FG75, FPS72, Gue79, HB72, HM84, Ing70, Joh77a, JRC81, Jun78, KBHS93, KHW89, LM93, Mar70, McI74, MR93, MK97, NK78, PK75, Pec74, Pir97, PH70, RS75, RHM79, RW76, Sak78, SJ75, SWM78, ST75, Sha76, Sha78b, Smi77, SS90, SMR^{+}86, TSv92, Thr68, Urq76, Wag75, Wal70, WLHC84, War72, WP94b, Wol75, YK77]. systems/1000 [Sha78b]. systems/3000 [Sak78].

TAB [Mat92]. Table [Ano96x, Sim94a, Sta88, Ano96-28]. Table-driven [Sta88]. Tablet [Kin81, Bar81, Sta81b]. Tablet/display [Kin81]. Tachyon [SP96]. Talk [Ano95k, Ano95l, NSMO98]. Tape [BDM85, CDG^{+}88, CT85, DSB94, DVE85, Gen85, Gil89, Gre85, MG85, RRC^{+}85, Sha85, Top89, Ujv94, AMSW85, BKV89, CS80, DPA^{+}88, NSW76, VBK89]. Tape/Disc [MG85]. Task [Ano93w, Ano93x, GAL^{+}93, PW89, SW80, Ano93-41]. task-oriented [SW80]. Tasks [GAL^{+}93, Wag88]. TDR [Das96]. Teach [Ped86]. teaches [BS79]. Team [Pet96, AH72]. teams [Hey81]. TECAP [Kha81]. Technical [BMS87, CG98, HMM96, CRSS77, FRAC85, Lea89, Lie88, WAZH84, Wic87, WM85, YGR95]. technique [Fel80, SA80, Tan89]. Techniques [AM95, Ano93q, Fos98, Hea94, PF88, Cze89, GHK^{+}93, Gro92, HSB^{+}94, HH91, Hue73, Slo89, Swe89]. Technological [Hol95]. Technologies [Ano93g, Ano95-37, SAS98, Mil75]. Technology [Ano95w, Ano95-29, Ano95-38, Ano96f, Bar95, BPMF87, BZS83, DOR97, FST92, FSDC82, HJL^{+}83, LH95, MFMS83, NIWM98, Nis87, PHL96, SD95, Toe86, BLTY81, BBH91, Cag86, DKMS81, Edw79c, FL80, For77, GBHS81, ICNT93, JM80, Kra89, MBL95, MP81, SM82, Spo90, Sti77, Wec88]. Techwriter [RA85]. Telecommunication [Cha96a, GH73]. Telecommunications [Cox96, LMNQ96, SDDC96, Haj96, NSV92, Win80]. Telemetry [BJPH95, LDNT72]. Telephone [Dan96, Tuc95, AD75, CM75, PH70, SS83a]. TeleShare [Tuc95]. Temperature [MHV96b, Sch84, VCMH96, EZ81, MHV96a, SL93]. Ten [MPR82]. Ten-Bit [MPR82]. Terminal [AP90, Bam83, BGK^{+}87, Cau85, Dul85, SBB85, AS90, AG75, Bee82, BP77, BDF84, CS76, CP85, Dic78, Dou75, Gau81, HG79, Kil94, Lan75, Leo77, Mai90, MC85, MB96, Moy78, NSW76, RBB78, Spa72a, TBL^{+}83, Wai75, Woo78]. terminals [ABJ78, SW81]. Terminology [Ano92-37]. Terms [Ano94p]. Test [AW90, Ano92s, Bal84, Ban84, Des92, DS82, Duf87, FS86, GDKK94, Han94, HR82, Hol84, Kar94, Kjo92, KGH^{+}87, KV94, Lan94, LSM^{+}95, McD84, Moo92, NZ97, NK84, OR96, RK94a, SW91, ST84, War92, Whi85, AD75, AR76, BH92, CW92, DVl91, Ewy69, ES69, FM84, FG75, GUW92, Gol80, Gor91, GH73, GB87, GT71, Gun94, HA92, ICNT93, Koe80, Kov81, LaF80, MN79, Mat84, Neu91, Par79, PSA83, Pli89, Rad87, RD75, Rob82a, Sch97, Sei91, Spa84b, SM79, TST82, Tho72, Tut72, TL89, Urq72, Wag75, Wat87a, Wri96, WCT87, Ano94v]. Testability [Ano93l, DR95, Mei82, Swe81]. Testable [Lam95]. TESTAID [Bas79]. Tester [BS98, DNC96, Sto96, Van98, War92, CWF^{+}79, Ing76, Sch79]. Testing [Ano92-33, Bal84, BS98, Bug86, CHK^{+}95, CAJ^{+}94, DNG84, DW87, Dub94, Duf87, ES85, FS94, Har86, Har84, Hen84, HRE85, Jeg95, Kar94, MM97a, MBC95, Mur97, Nik97, Pet74, RH74, RC98, SM88, ST90, Sti97, TOB97, WR97, Ano67, Bas79, Bob73, BG89, Bur85, CDE^{+}87, Car69, CRSY73, Don69, EAW^{+}86, Fie89, Gri69, Ham93, HRZE77, Hoo91, HBNB83, Hue73, Mai90, MLJP93, Mat90, Mey90a, Nea68, NSP97, NTH^{+}84, PK75, Pir97, Pli89, RHM79, SM79, Win80]. Tests [Cha96b, KRDS83, Bra78, Bru79b, Sch93c]. text [ABW78, BCC^{+}94]. texture [HH90]. thanks [Mer81]. Their [BT98, DR81]. Theoretical [Ano94-36, Sha98]. Theory [DW87, EM98, LY94, Mag86, Chu74, RW81]. Thermal [ABC^{+}94, AMK85, Ano92-34, BJJN88, BCRS88, DR98, EGH^{+}88, HSW88, NH94, WMBB94, ACS88, Bar88, BSME92, Bro82, BHH^{+}88, Coz78, DCN^{+}88, HCF^{+}88, LK80, MM72, O’N80]. Thermocouple [Col88a, Jac74, Lam74]. Thermodynamics [AMK85]. thermometer [Fos74]. Thick [Toe86, Zam73, Mil75]. Thick-film [Toe86, Mil75]. Thin [AGS85, DGHP85, DLJ85, HRE85, Ili82, MSB85, OSN^{+}85, Ada71, BA85, Jac74, Mer73]. Thin-Film [AGS85, DGHP85, DLJ85, HRE85, MSB85, OSN^{+}85, BA85, Jac74, Mer73]. thin-film/semiconductor [Jac74]. ThinkJet [BA85, Nie85, SBM85]. Third [ABC^{+}94, Ano93-42, DiP80, MCCB76, Mut74, GHUW68, DNC96]. Third-Generation [ABC^{+}94, GHUW68, DNC96]. Third-Order [Ano93-42]. Threads [Ano93-31]. Three [Ano92h, NT75, FR95]. Three-Color [Ano92h]. three-dimensional [FR95]. Threshold [KRDS83]. Throughput [Cas95a, HT92, SMGR85, TD92, CR89, EKW93, Rie94, Sha76]. THz [RMVT95]. tidal/respiration [Kri81a]. Time [Ano93-37, Ano93-34, Ano96-30, BLNP86, BF82, Byr96, Can93, Chu74, DH98, Duf87, FD78, GWY94, Kel92, LBLS95b, Loo97, MS95, MC84, MY95, MD96, Pal96, War72, ASCC84, AP69, Ano71b, AG75, Anz93, Ave75, BRHS89, BGG88, Bel93, BP93, Bha92, BNB93, BCO^{+}93, CAF78, Dra77, EKW93, FGH^{+}89, FRS88, FMN84, Gas88, Gun94, HR91, Hei88, KB69b, KH88a, KBHS93, LTC^{+}84, LBLS95a, LM93, MVG77, Mor70, Mor93, NS88, OST75, PN88, Pet68, Pou68, RP88, Rus93, Saf96, SES90, Sch70, Sch76, SNN88, Smi69, SH77, Sor74, Ste88, Vob88, Wal70, Wec88, Wes77, HKC^{+}84a]. Time-Domain [MS95, MD96, BGG88, Bel93, BP93, FRS88, KBHS93, RP88, SNN88, Smi69, Vob88]. time-related [MVG77]. time-shared [Pou68, Wal70]. Time-Slice [Loo97]. time-varying [BCO^{+}93]. timer [Mac75]. timer/counter [Mac75]. times [GZC74]. timesharing [MS74]. Timing [Ano97f, Bir96, KRDS83, Neu84, War97a, FD78, FR74, SWM78, She71, ZHN83]. TIMS [DNG84, Reh84]. tiny [Tag74]. TMN [Spe96]. Today [Ano91g, FL73, SAMS73]. Today’S [CS82]. together [Goo72, RS75]. Tolerance [SR95, WP94c]. Tone [Ano95h]. Tones [Ano95h, Urq72]. Tool [Ano93-34, BTW83, Bug86, Cou94, DD86, GWY94, GSFF90, GAL^{+}93, Har86, KF95, Lub91, McK87, Moo92, Reh88, Sim94b, AGI^{+}91, Fro90, Gau81, HBHC84, Neu91, SSMK92, Sch97, TL89]. Toolkit [Cox96, CG98, Spe96, Sto96, TW90, WHBP95]. Tools [Ano92i, Ano93-43, AWH^{+}87b, AWH^{+}87a, BPA^{+}94b, DR98, FHMP97, FS86, PDW85, Rya95, Bac91, Cag90, Ger90, HG81, Jes89b, Mur82, SB83, Tiv91]. Toolset [ABC^{+}88, Ano92-30, Bra96, Bur86]. Topology [Bha94]. total [BW76, GHK^{+}93, Gor80, Ker80]. Touchscreen [SPA^{+}84a, SMK84, Suk84]. tougher [HB72]. Trace [Mal87]. Traceability [Ger98]. tracer [Bec76]. Traces [Ano93c, Har76]. Tracking [Ano93-46, Pag71, Bla86, Due74, NYS81, PSA83, PB82, SWM78]. Trade [DMJCB97, NIWM98]. Trade-Offs [DMJCB97, NIWM98]. traditional [LM86]. training [LFWEJ89]. Transaction [Gup95, Raf93]. transceiver [Bob73, Hue73]. Transconductance [KM97]. Transducer [GFK83, EH81, Lar83b, McK83, Mil83, OS76, RW76]. transducer-based [EH81]. Transducers [MW94, SS83b]. Transfer [Ano95w, BB97a, Man90]. Transient [Win74, FK79]. Transimpedance [Ano96-29]. Transistor [Kha81, Bau69, BP70, IK69]. transistors [Man69]. Transition [BW92, DW92b, GZC74, SES90]. Translation [Ano92o, Ano93-43, Buf96]. Transmission [MCD97, NTH^{+}84, Pie87, Pli89, CDE^{+}87, Gue74b, HMPF74, DNG84]. Transmitter [Mil94, Lee74]. Transmitters [Bir91]. Transport [FKL86, Pug90, CS80]. Transportable [KE85]. transversal [IKS81]. travel [BGH74]. treatment [KH88a]. Trigger [Ano97a, GE86, EKW93, Hei88]. Triggered [How79]. Triggers [Bug86]. Trigonometric [Egb77b, Egb77c]. Trilayer [OLR82]. Troubleshooting [BYA86, Mur92, AH72, BS79, Nel79, SM79, WS84b]. troubleshoots [Cha77, Que76]. True [Ano95-48, Fol72, Wad68]. true-rms [Wad68]. Tsutsuji [COO^{+}93]. Tube [Gro83, Cha70, HMO73]. Tunable [MLJP93, NML90, RLB^{+}95, JBS93]. Tuned [CWFS82, Lau80, LS83]. tuning [Tho80]. Turnaround [FHMP97, BCH81]. TV [HMPF74]. Twelve [KH69]. Twenty [MPR82]. Twenty-Megasample-Per-Second [MPR82]. Twisted [Ano95l, MCD97]. Twisted-Pair [Ano95l, MCD97]. Two [Ake85, AS85, Ban91, Buf96, Has92, RHK^{+}83, Wil84, DG70, GSVS89, KW93, Sch87, Spa84b, Ste89b]. two-axis [KW93]. Two-Channel [Ake85, AS85, Ban91, Sch87]. two-hundred-foot [DG70]. Two-Language [Buf96]. Two-Layer [RHK^{+}83]. Two-Pen [Has92]. two-port [Spa84b]. two-source [Ste89b]. type [Ril79]. Types [Ano96r]. Typing [Bar96].

ULSI [NSMO98]. Ultra [TWCD93]. Ultrasonic [Che86, MW94]. Ultrasound [GFK83, HLH86, JAD83, Kar86, KD83, LKHK83, OMD83, SC83a, SS83b, Ban83, HT86, Kin83, LHL83, Mag83, O’C86]. Unattended [SMGR85, ACK79]. Unbiased [Mat86]. understandable [AD75, Far74]. unified [Han80b]. unifying [Wat81]. unit [Fie89, FG75, Hea85, Ike83, MM90, MFPE72, STE90]. unit-under-test [FG75]. Uniting [Sha71]. Units [ACK^{+}84, BK85b, Ger98, Kel94, TST82]. Universal [Ano96-30, BB97b, CAF78, HC75, LJ75, MBW79, McO80, SJ80]. University [BC81b]. UNIX [Ano96a, Bla86, CBL96, FRAC85, HD98, WLHC84, Wat87b, Wat88, BHP85]. Unless [Ano96z]. Unravelling [Wag75]. Unshielded [Ano95l]. unstructured [FJ89]. Unsynthesized [Bar82]. Update [Ano92-35]. Updating [HD98]. Upper [Chi90, BE90]. Uptime [Wen85]. Usability [Ano93-44, Har86, Tig96]. Usable [Tig96]. Use [BRHS89, Bur95, DH77, Ada71, Bac91, Car78, Cha77, CP79, CRSS77, DiF75, Dic78, GHK^{+}93, Han79, Har76, Pec74, RBB78, SH90a, SJ75, SDS83, Sol81, Suk84]. used [LD93]. Useful [Cor96]. User [Ano92c, Ano92l, Ano93c, Bar84, BP85, BF89, BPA^{+}94b, DF90, Fer94, HSK87, HWK95, JS93, LK85, NZ97, Not93, RWW95, Sch79, AC90, Bac91, BHP85, Cha91, Dou92, GAP92, GJ74, Gra81, Hun92, JF80, Kas72, NGW92, OB88, Por80, Sch76, Sho89, SKG90, Ste74a, SMK84, Van77b, Vob88, WB90]. User-centered [BF89]. User-Network [NZ97]. User-oriented [Sch79, Ste74a]. users [HKC^{+}84a]. uses [Toe86]. Using [Ano92-36, Ano94w, Ano95o, BS95, CS76, Das96, FHMP97, Gal84, Hea94, HNL87, KP88, Kam82, Kru88, LF91, LJ75, MCD97, Mil94, NMY95, OLR82, Oh82a, Pie87, Rya95, Shr82, Sim94b, Wak97, War86, Wil97a, Bur89, CKP89, CS93, Dia81, DP73, DD97, FK79, GC90, Gro92, GBHS81, HT86, Kee89, Kra89, Kri81b, Leo83, Mar89, MB96, Sol81, Swe89, War89, Wec88, WJ79]. UTC [Ano96-30, Ano71b]. Utility [Bir97, DL73, JS74, OB88]. Utilizers [Ano96w, Ano96p]. UX [AFG^{+}86, Ano92y, BT88, BLY92, Cha91, CHKS86, CR92, KGTM90, KG92, LP92, Mar89, Not93, OY88, Ran88, SSV90, SMQ88, SAS98, STE90, WLHC84, Wan88].

v.h.f [Pet70]. Validation [Kru89].

Valuable [Har86]. Value [Ano93-44, FS94, Kam72]. Vapor [Ili82, Thr68]. variable [GZC74, Har76, SES90]. variable-persistence [Har76]. variable-persistence/storage [Har76]. variable-transition-time [SES90]. variety [Ega72, HRH81, RD75]. varying [BCO^{+}93, Wec88]. Vascular [MW94]. VCO [RMVT95]. Vector [Ano93-45, Ano95v, BCO^{+}93, DWB^{+}97, DG88, Edw87, GC87, IOS80, JM87, Kaf88, Kar94, KWR^{+}87, Lau87, MFW87, NG87, ABW78, BBDK93, CGH^{+}93, KN93]. Vectors [Han94]. Vectra [LLS88, Ano91e, Ano91d, LL91, NTJ^{+}91, SJL91]. VEE [Bee92, GJ98, Hun92]. Verification [ADG^{+}96, BBR^{+}97, Fos98, GLP^{+}97, HNL87, KGH^{+}87, MGBB97, ABG^{+}92, Neu91, TH81]. Verifying [Han94, Ric81b, WT97]. Verilog [SWI97]. Vernier [GBO94]. Versatile [Ano93-46, Dic78, Ega72, EH81, Eps84, GLW85, Hay74, MFPE72, Ter78, BHE^{+}85, CRSS77, FR87a, Fos74, MH80, NYS81, Pec74]. versatility [Goo72, MH80, Mut74, Woo78]. Version [Ano93x]. Versions [BPA^{+}94b]. versus [Ano92p, Ano93-38, Hig95, LM86]. Vertical [NMY95, Mil75, TTML81]. Vertical-Cavity [NMY95]. vertical-channel [Mil75]. Very [KB71, Pra75, Rom91, Wes77]. Very-low-level [Pra75]. VGA [TWCD93]. VHF [Hay74, Low69, SAMS73]. via [Goo91, HKH93, Lin97, Was80]. vibrometry [Kee89]. Video [LBLS95b, Ric94, LBLS95a, MC85, Roy81]. Videoscope [TL89]. View [Ano94r, Pop83]. viewing [SM75]. Virtual [Bas79, But94, CS88, HM88, Mun77, OB88, SW81]. vision [GS93, HCF^{+}88, Mon92]. visual [BH92, Cha91, LT90]. VISUALIZE [SOG98]. Visualizing [Die75]. VLSI [AGO^{+}85, Ama86, Ano90a, BZS83, BHW86, BC81b, BFHT83, Gal84, GSG^{+}87, GJM^{+}92, GJW83, HG81, HKC84b, HL85, KGH^{+}87, LRFL83, MLSW83, MBF^{+}87, Mer82, Neg81, Nis87, QAKP90, Raj95, RJH^{+}87, SDA81, SM82, SBB85, Tha89, TH81]. VLSI-Based [GSG^{+}87]. Voice [CS88, CDE^{+}87, NTH^{+}84]. voice- [CDE^{+}87]. voiceband [CPR75, Dam74]. volt [DiF75, KE74]. Voltage [Ben95, Emm72, FR74, IK69, MH89, Swe89]. voltage-controlled [MH89]. voltages [Fol72, Wad68]. voltmeter [Goo77, MVG77, McO80, VDF72]. volts [Han70]. Volume [BCRS88, MY95]. VRAMs [Ano95-32]. VUE [Joy96]. VXI [KV94]. VXI-Based [KV94]. VXIbus [Ano92m, Ano92t, Ano92-37, Atc92, CW92, Des92, Eri92, Jes89a, Jes89b, Kel92, NH92, War92].

Wafer [NK84, Ham93]. Wafer-Stage [NK84]. was [Sta77]. Wave [Alb88, BGR^{+}88, Col88a, DG88, IC85, Mat86, SR88, Uph88, ZFE^{+}86, CE81, EJG86, IKS81, Lee74, Nic90, SA91, Tho80]. Waveform [BF82, FPUB88, HSK87, HKM96, KH88b, PF88, PMN82, PBD88, San86, Sch86, SA82, Sor88, SD82, EL82, GY88, HK88, HBB88, KMT88, Nic88, Tal88a]. Waveforms [SHB84, Ive88]. Waveguide [RLB^{+}95]. Wavelength [NMY95, LAB97, MRRS93, WA93]. Wavelet [LY94]. Waves [Shr82]. Way [Ano92l, SBK^{+}97, Bec76]. Ways [PDW85, Pee73]. Weekly [MRH94b]. Weighted [Cas95a]. welcomes [Ega72]. Well [RLB^{+}95, RWT97]. White [Ben95, CS93, PB82]. white-light [CS93]. white-noise [PB82]. Whiteboard [Ano94-38]. Wide [Ano95p, ANSW95, Arm95, CSWR85, FR87c, Met86, BHW77, Cat82, HT76, JRC81, LS76, RD75, Wad68, YK77]. Wide-Bandwidth [Ano95p, ANSW95, Arm95]. wide-dynamic-range [Cat82]. Wide-Frequency-Range [CSWR85]. wide-range [HT76, LS76]. wide-ranging [BHW77, YK77]. Wideband [CWFS82, OH82b, Cho90, DeV70, HK75, LS83, McH87, Ras87]. widen [Zam73]. widgets [ME90]. widths [FD78]. Winchester [BDM85]. Window [Ano94-33, BSY89, KT90, HB88]. windows [Sta81a, Ano93c, GC90, HD98, SKG90]. Wireless [McF98, NIWM98, RNR^{+}98, SM87]. Wiring [Pie87]. word [Elw74, HRZE77, PK75]. words [Don69]. work [Ogd78]. Working [AGO^{+}85]. workspaces [Mun77]. Workstation [Ano92q, BBD^{+}84, Cle93, Cun98, DR95, Hig85b, Mar95b, MTMP83, Pea95, Rub86, SGD83, Tuc95, ABD^{+}97, BDF84, Cag86, Jon84, LWBB84, Roe95, SB84, TBL^{+}83, TBBM83]. Workstations [AFG^{+}86, BCF96, Let92, LHIR98, ROKH94, SKK^{+}86, SAS98, ABG^{+}92, BBG91, CL92, CLL^{+}92, DW92a, HVSM96, KG92, RH92, UAH^{+}92, Wan88]. World [Ano92u, Ano96u, Lyn80b, HD98]. Worlds [NVBH87]. Worldwide [HSMA^{+}86]. Wrist [MHCH77]. writable [Cou72]. Write [Gre85, BB84, Fin72]. Writing [OH82b, Bru73, Cha70, GO72]. wt [HH90]. WYSIWYG [Dei94, WCH94].

X [Ano94-32, Ano94-33, Bre74, BSY89, Bru77, Dei94, FCC79, GN82, HB88, How79, HJW75, KT81, LM81, Mar95a, Met86, RC79, Tal68, TRH85, WCH94]. X-axis [KT81]. X-Ray [GN82, Bre74]. X-Y [Bru77, FCC79, How79, HJW75, LM81, RC79, Tal68, TRH85]. X.25 [AS90]. X/Open [Mar95a]. X11 [ACL89, BCH^{+}89, BG89, ES89, HLM89]. XL [BGK^{+}87, BKO87].

Y-axis [HPH81]. yardstick [DG70]. years [Lea89]. yields [Edw79c]. YIG [LS83].

YIG-tuned [LS83].

Z [MM72]. Z-fold [MM72]. Zero [Adc87b, But95, Chi82]. zero-encroachment [Chi82]. zoom [BR78].

References

Ambras:1988:MIP

[ABC^{+}88] James P. Ambras, Lucy M. Berlin, Mark L. Chiarelli, Alan L. Foster, Vicki O’Day, and Raldolph N. Splitter. MicroScope: an integrated program analysis toolset. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):71–82, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Aden:1994:THT

[ABC^{+}94] J. Stephen Aden, Jaime H. Bohorquez, Douglas M. Collins, M. Douglas Crook, Andre Garcia, and U. E. Hess. The third-generation HP thermal InkJet printhead. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):41–45, February 1994. CODEN HPJOAX. ISSN 0018-1153.

                                                        URL http://www.hp.com/
                                                        hpj/94feb/41to45.pdf.

Allan:1997:LCW

[ABD^{+}97] S. P. Allan, B. P. Bergmann, R. P. Dean, D. Jiang, Floyd, and D. L. A low cost workstation with enhanced performance and I/O capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):82– 88, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97jun/ju97a11.htm.

Ahi:1992:DVH

[ABG^{+}92] Ali M. Ahi, Gregory D.

Burroughs, Audrey B. Gore, Steve W. LaMar, Chi-Yen R. Lin, and A. L. Wiemann. Design verification of the HP 9000 Series 700 PA-RISC workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):34–42, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Abraham:1983:NP

[ABH^{+}83] Howard E. Abraham, Keith G.

Bartlett, Gary L. Hillis, Mark Stolz, and Martin S. Wilson. NMOS-III photolithography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(8):34–37, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Anderson:1978:MTH

[ABJ78] D. B. Anderson, M. B.

Bain, and G. Johnson. Multipoint terminals for HP 1000 systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(14):28– 32, October 1978. CODEN HPJOAX. ISSN 0018-1153.

Adler:1973:LAN

[ABM73] R. Adler, M. Baker, and H. D. Marshall. The logic analyzer: a new instrument for observing logic signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(2):2–16, October 1973. CODEN HPJOAX. ISSN 0018-1153.

Azmoon:1978:DPP

[ABW78] M. Azmoon, J. H. Bohorquez, and R. A. Warp. Desktop plotter/printer does both vector graphic plotting and fast text printing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(13):24–25, 27–32, September 1978. CODEN HPJOAX. ISSN 0018-1153.

Anderson:1990:CEU

[AC90] Steven R. Anderson and Jennifer Chaffee. Creating an effective user interface for HP IVIBuild. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(5):39–48 (or 39–44??), October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Ahad:1993:HOO

[AC93] Rafiul Ahad and Tu-Ting Cheng. HP Open ODB: an object-oriented database management system for commercial applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 20–30, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Augenblick:1983:EFC

[ACCM83] K. B. Augenblick, M. A.

Casale, J. E. Cusack, and A. J. Murphy. Electronic flow control: a new level of automation for gas chromatography. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(1):35–40, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Azmoon:1979:FPE

[ACK79] M. Azmoon, R. A. Coverstone, and R. M. Kemplin. Four-color plotters enhanced for unattended operation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(12):29– 32, December 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Allen:1984:CSP

[ACK^{+}84] Kevin W. Allen, Paul C. Christofanelli, Robert E. Kuseski, Ronald D. Larson, David Maitland, and Larry J. Thayer. Compact 32-bit system processing units. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(5):31– 38, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Andreas:1989:MSD

[ACL89] James R. Andreas, Robert C. Cline, and Courtney Loomis. Managing and sharing display objects in the Starbase/X11 Merge system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (6):12–19, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Allen:1992:HDC

[ACM92] William J. Allen, Toni D. Courville, and Steven O. Miller. HP DeskWriter C printer driver development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):93–102, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Askeland:1988:STI

[ACS88] Ronald A. Askeland, Winthrop D.

Childers, and William R. Sperry. The secondgeneration thermal inkjet structure. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):28–31, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Aken:1975:UTS

[AD75] M. B. Aken and D. K.

Deaver. An understandable test set for making basic measurements on telephone lines. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(9):2–9, May 1975. CODEN HPJOAX. ISSN 0018-1153.

Ashkenas:1982:PDP

[AD82] David J. Ashkenas and Richard F. De Gabriele. Powerful diagnostic philosophy reduces downtime. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3): 11–14, March 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Adam:1971:PSA

[Ada71]  S. F. Adam.  Pro-

grammable step attenuators use distributed-thin-film attenuator cards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(8): 14–16, April 1971. CODEN HPJOAX. ISSN 0018-1153.

Adcock:1987:ASF

[Adc87a] James L. Adcock. Analyzer synthesizes frequency response of linear systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (1):25–32, January 1987. CODEN HPJOAX. ISSN 0018-1153.

Adcock:1987:CFP

[Adc87b] James L. Adcock. Curve fitter for pole-zero analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(1):33–37 (or 33–36??), January 1987. CODEN HPJOAX. ISSN 0018-1153.

Alexander:1996:VCD

[ADG^{+}96] Thomas B. Alexander, Kent A. Dickey, David N. Goldberg, Ross V. La Fetra, James R. McGee, Nazeem Noordeen, and Akshya Prakash. Verification, characterization, and debugging of the HP PA 7200 processor. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):34–43, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/pa7200_
verification.pdf;   http:
//www.hp.com/hpj/toc-
02-96.html.

Adler:1986:HSL

[Adl86] Glenn J. Adler. Hybrid solution for a 25-line LCD controller. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(7):25– 27, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Alburger:1983:PGP

[ADRS83] J. R. Alburger, J. L. Davis, D. A. Rodriguez, and B. A. Stanley. PAINTBRUSH/ 2700: a general-purpose picture creator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(9):34–37, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Ansley:1971:HC

[AE71] W. G. Ansley and S. D. Edwards. HP/CAI. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(6):2–3, February 1971. CODEN HPJOAX. ISSN 0018-1153.

Anderson:1986:NHF

[AFG^{+}86] Andrew G. Anderson, David L. Frydendall, Robert D. Gardner, Robert M. Lenk, Robert J. Schneider, Bonnie Dykes Stahlin, and Ronald G. Tolley. New HP-UX features for HP 9000 series 300 workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7): 34–41, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Anzinger:1975:ROS

[AG75] G. A. Anzinger and A. M. Gadol. A real-time operating system with multiterminal and batch/spool capabilities. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(4):12–20, December 1975. CODEN HPJOAX. ISSN 0018-1153.

Amino:1991:HSI

[AGI^{+}91] M. J. Amino, C. Givens, M. Ikemoto, A. C. Miranda, S. A. Gulland, K. A. Fulton, and I. S. Smith. HP Software Integration Sockets: a tool for linking islands of automation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (5):6–23, December 1991.

CODEN HPJOAX. ISSN 0018-1153.

Alvarez:1985:SEW

[AGO^{+}85] Patria G. Alvarez, Greg L.

Gilliom, John R. Obermeyer, Paul L. Rogers, and Malcolm E. Woodward. Simulation ensures working first-pass VLSI computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(9):13–16, September 1985. CODEN HPJOAX. ISSN 0018-1153.

Allyn:1985:TDM

[AGS85] Michael C. Allyn, Peter R.

Goglia, and Scott R. Smay. Thin-film discs: Magnetic, electrical, and mechanical design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(11):36– 40, November 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Adler:1972:LPP

[AH72] R. Adler and J. R. Hofland.

Logic pulser and probe: a new digital troubleshooting team. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(1):2–7, September 1972. CODEN HPJOAX. ISSN 0018-1153.

Andersen:1989:RSG

[AH89] Brad E. Andersen and Earl C. Herleikson. RF signal generator single-loop frequency synthesis, phase noise reduction, and frequency modulation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5): 27–33, October 1989. CO-DEN HPJOAX. ISSN 0018-1153.

Akiyama:1984:HCC

[AI84] Tomoyuki Akiyama and Kenzo Ishiguro. Highspeed 1-MHz capacitance/ conductance meter for measuring semiconductor parameters. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(6):12– 24, June 1984. CODEN HPJOAX. ISSN 0018-1153.

Ainsworth:1981:SPS

[Ain81] G. L. Ainsworth. A switching power supply for a low-noise signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (2):20–??, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Amin:1979:MSP

[AK79] D. A. Amin and T. Kriegel.

A multiple-output switching power supply for computer applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(7): 25–28, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Adams:1981:SCB

[AK81] C. A. Adams and J. A.

Kusters. The SC cut, a brief summary. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(3):22–23, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Aken:1985:ATS

[Ake85] Michael B. Aken. Applications of a two-channel synthesizer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(8):19–21, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Adam:1969:BPC

[AKL69] S. F. Adam, G. R. Kirkpatrick, and R. A. Lyon. Broadband passive components for microwave network analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(5):2–10, January 1969. CODEN HPJOAX. ISSN 0018-1153.

Albin:1988:MSM

[Alb88] Robert D. Albin. Millimeterwave source modules. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 18–19, 21, 23–25, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Albin:1989:DOH

[ALRS89] Robert D. Albin, Kent W. Leyde, Rollin F. Rawson, and Kenneth W. Shaughnessy. Design and operation of high-frequency lightwave sources and receivers (for lightwave component analyser). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(3):52– 57, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Amerson:1982:HCD

[ALT82] Frederic C. Amerson, Mark S. Linsky, and Elio A. Toschi. High-performance computing with dual ALU architecture and ECL logic. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (3):3–4, 7–11, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Aikin:1973:ACL

[AM73] E. R. Aikin and J. L.

Minck. Automating the calibration laboratory. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(8): 14–17, 20–23, April 1973. CODEN HPJOAX. ISSN 0018-1153.

Aitken:1995:BMB

[AM95] Robert C. Aitken and Peter C. Maxwell. Better models or better algorithms? techniques to improve fault diagnosis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 110–116, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95110.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Amano:1986:SAM

[Ama86] Jun Amano. Salicide: Advanced metallization for submicrometer VLSI circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(5):33–39, May 1986. CODEN HPJOAX. ISSN 0018-1153.

Amerson:1985:SMC

[Ame85] Frederic C. Amerson. Simplicity in a microcoded computer architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(9): 7–12, September 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Allen:1985:THT

[AMK85] Ross R. Allen, John D. Meyer, and William R. Knight. Thermodynamics and hydrodynamics of thermal ink jets. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(5):21– 27, May 1985. CODEN HPJOAX. ISSN 0018-1153.

Auyer:1985:CHT

[AMSW85] Walter L. Auyer, Charles H. McConica, David J. Schmeling, and Mark E. Wanger. Controlling the head/tape interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3):44–47, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Apffel:1990:IPB

[AN90] James A. Apffel, Jr. and Robert G. Nordman. Introduction to particle beam LC/MS. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(3):69–76, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Anderson:1970:PMB

[And70] G. F. Anderson. A programmable, modular, bidirectional data coupler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22 (1):2–6, September 1970. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1967:LAA

[Ano67] Anonymous. Loudness analyzer aids noise reduction, production testing, speech analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(3):15–19, November 1967. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1971:PPS

[Ano71a] Anonymous. A package for portability and serviceability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(12):9–??, August 1971. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1971:UTS

[Ano71b] Anonymous. UTC time scale to change in 1972. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (2):16–??, October 1971. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1990:CVC

[Ano90a] Anonymous. Custom VLSI chips for DMA. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):15–??, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1990:EDC

[Ano90b] Anonymous. Error detection and correction primer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (6):46–47, December 1990. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1990:OCN

[Ano90c] Anonymous. The OSI connectionless network protocol. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(1):49–??, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:CMSa

[Ano91a] Anonymous. Component monitoring system software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):14, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:CMSb

[Ano91b] Anonymous. Component monitoring system software development environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (4):15–18, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:ECS

[Ano91c] Anonymous. EISA configuration software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(4):75–77, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:HVE

[Ano91d] Anonymous. The HP Vectra 486 EISA SCSI subsystem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):70–71, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:HV

[Ano91e] Anonymous. The HP Vectra 486/33T. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(4):72, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:MCH

[Ano91f]  Anonymous.  Measure-

ment capabilities of the HP 8703A lightwave component analyzer and the HP 71400C lightwave signal analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):17–??, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1991:MET

[Ano91g] Anonymous. Medical expectations of today’s patient monitors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):9, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:AQL

[Ano92a] Anonymous. Acceptable quality level index. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6): 37–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:ANM

[Ano92b] Anonymous. Adapting the NCR 53C710 to minimize interrupt impact on performance. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):105– ??, December 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:AUP

[Ano92c] Anonymous. Average user plot. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):36–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:BLA

[Ano92d] Anonymous. Bang-bang loop analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):110– ??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:BSS

[Ano92e] Anonymous. Board-level simulation of the series 700 EISA cards. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):94–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:CEB

[Ano92f] Anonymous. Calculating effective bits from signalto-noise ratio. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):34–35, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:CIP

[Ano92g] Anonymous. Color inkjet print cartridge ink manifold design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(4):82–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:CST

[Ano92h] Anonymous. Color science in three-color inkjet print cartridge development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):71–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:DT

[Ano92i] Anonymous. Debugging tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):39–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:DMP

[Ano92j] Anonymous. Definitions and measurement procedures for cut quality parameters. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):46–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:DPE

[Ano92k]  Anonymous.  DesignJet

plotter end covers produced by coinjection. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):31–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:DPU

[Ano92l]  Anonymous.  DesignJet

plotter user interface design: Learning the hard way about human interaction. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):12–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:EMV

[Ano92m] Anonymous. Examples of message-based VXIbus instruments. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):20– 21, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:FI

[Ano92n] Anonymous. The Forth interpreter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):25–??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:FTC

[Ano92o] Anonymous. Frequency translation as convolution. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (5):61–??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:HEA

[Ano92p] Anonymous. Harmonic errors and average versus peak detection. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):94, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:HSW

[Ano92q] Anonymous. HP 9000 series 700 workstation firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):9–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:HMA

[Ano92r] Anonymous. HP MRP action manager project management. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):108–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:HSP

[Ano92s] Anonymous. HP standard PA-RISC test programs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):35–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:HVM

[Ano92t] Anonymous. The HP VX-Ibus mainframes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 9–12, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:MHP

[Ano92u]  Anonymous.  Making

HP print cartridges safe for consumers around the world. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):76–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:M

[Ano92v] Anonymous. Metrics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4): 41–??, August 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:MSP

[Ano92w] Anonymous. Multilayer shielding protects microvolt signals in highinterference environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 84–89, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:NAP

[Ano92x] Anonymous. Network Advisor product enhancement philosophy. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):9–??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:NMH

[Ano92y] Anonymous. Next-generation multiprocessor HP-UX. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):58–??, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:ODS

[Ano92z]  Anonymous.  Object-

oriented design and smalltalk. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):24–??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:OPL

[Ano92-27]  Anonymous.  Object-

oriented programming in a large system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):76–??, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:PPM

[Ano92-28]  Anonymous.  PA-RISC

performance modeling and simulation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(4):21–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:PNP

[Ano92-29] Anonymous. Poor network partitioning. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):75–??, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:ST

[Ano92-30] Anonymous. Simulation toolset. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):36–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:SLM

[Ano92-31] Anonymous. Small, lowcost mainframe with a register-based interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 22–23, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:SRD

[Ano92-32]  Anonymous.  Stimulus/

response defect diagnosis in production. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):27–28, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:TPP

[Ano92-33] Anonymous. Testing the peak power analyzer firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):99–100, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:TIR

[Ano92-34]  Anonymous.  Thermal

inkjet review, or how do dots get from the pen to the page? Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):67–??, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:USS

[Ano92-35] Anonymous. Update on the SCSI standard. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):103– ??, December 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:UHM

[Ano92-36] Anonymous. Using the high-resolution, multichannel DAC in the HP 54601A oscilloscope. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):54–56, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1992:VT

[Ano92-37] Anonymous. VXIbus terminology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):13–14, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:ADE

[Ano93a] Anonymous. Absolute digital encoder. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(3):14– ??, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:ABD

[Ano93b] Anonymous. ADC bits, distortion, and dynamic range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):38–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:AOT

[Ano93c]  Anonymous.  Analyzing

OTDR traces on a PC with a Windows user interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (1):77–??, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:AD

[Ano93d] Anonymous. Applications for demodulation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):12–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:AMI

[Ano93e]  Anonymous.  Averaging

measurements to improve sensitivity. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):44–??, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:CE

[Ano93f] Anonymous. C environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):35–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:CT

[Ano93g] Anonymous. Cellular technologies. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):92–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:CSD

[Ano93h]  Anonymous.  Client/

server database architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):78–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:CSH

[Ano93i]  Anonymous.  Context

switching in HP-RT. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4): 32–??, August 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:CFD

[Ano93j] Anonymous. Cooling of the frequency divider IC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 61–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DEC

[Ano93k] Anonymous. The debug environment connection to HP SoftBench. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(2):93– ??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DMD

[Ano93l]  Anonymous.  A design

for manufacturability, design for testability checklist. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):33–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DEF

[Ano93m] Anonymous. Developing extensible firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5): 64–??, October 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DG

[Ano93n] Anonymous. Diffraction grating. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):70–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DCF

[Ano93o] Anonymous. A digitally corrected fractional-N synthesizer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):44–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DB

[Ano93p] Anonymous. Dither and bits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):42–??, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:DTD

[Ano93q] Anonymous. DSP techniques for digital IF. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5): 90–??, October 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:ED

[Ano93r] Anonymous. Error definition. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):72–??, June 1993. CODEN HPJOAX. ISSN 0018-1153. Anonymous:1993:FDI

[Ano93s] Anonymous. Filter design for interpolation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(5):45–??, October 1993. CODEN HPJOAX. ISSN 0018-1153. Anonymous:1993:G

[Ano93t]  Anonymous.  Glossary.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4): 86–??, August 1993. CO-DEN HPJOAX. ISSN 0018-1153. Anonymous:1993:GJS

[Ano93u]  Anonymous.  Gravity-

sensing joy stick. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(3):12– ??, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:HOS

[Ano93v]  Anonymous.  The HP

ORCA system outside the analytical laboratory. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 9–??, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:HTBa

[Ano93w]  Anonymous.  HP Task

Broker and computational clusters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):16–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:HTBb

[Ano93x] Anonymous. HP Task Broker version 1.1. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):21–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:ISA

[Ano93y] Anonymous. Improving SNR by averaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):65–??, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:III

[Ano93z] Anonymous. InP/InGaAs/ InP P-I-N photodetectors for high-speed lightwave detectors. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):85–??, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:IPG

[Ano93-27] Anonymous. Internal pulse generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):27–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:MPA

[Ano93-28] Anonymous. Microwave plate assembly. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):50–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:NPE

[Ano93-29] Anonymous. A new perspective on emulation hardware modularity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(2):102– ??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:OSA

[Ano93-30] Anonymous. Optical spectrum analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):62–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:OT

[Ano93-31] Anonymous. An overview of threads. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):27–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:PS

[Ano93-32] Anonymous. Polarization sensitivity. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):71–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:PSD

[Ano93-33] Anonymous. Protecting shared data structures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):

33–??, August 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:ROS

[Ano93-34] Anonymous. A real-time operating system measurement tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):97–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:RD

[Ano93-35] Anonymous. Remote debugging. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):29–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:RP

[Ano93-36] Anonymous. The resampling process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):10–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:RHD

[Ano93-37] Anonymous. Run-time-configurable hardware drivers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):20–??, December 1993.

CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:SVM

[Ano93-38] Anonymous. SAM versus manual administration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 81–??, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:SRE

[Ano93-39] Anonymous. The shadow register environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4): 34–??, August 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:SLH

[Ano93-40] Anonymous. The structure of LEDs: Homojunctions and heterojunctions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):8–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:TBD

[Ano93-41] Anonymous. Task Broker and DCE interoperability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):19–??, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:TCA

[Ano93-42] Anonymous. Third-order curve-fit algorithm. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 41–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:TLT

[Ano93-43]  Anonymous.  Tools for

the language translation process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):68–??, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:VU

[Ano93-44] Anonymous. The value of usability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):91–??, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:VMB

[Ano93-45] Anonymous. Vector measurements beyond 1.8 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):58–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:VTA

[Ano93-46]  Anonymous.  A versa-

tile tracking and arbitrary source. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):54–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1993:WD

[Ano93-47]  Anonymous.  What is

dithering? Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):44–??, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:CI

[Ano94a] Anonymous. 1994 chronological index. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):??, December 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/toc-12-94.html.

Anonymous:1994:PAI

[Ano94b] Anonymous. 1994 product and author index. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):??, December 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/toc-12-94.html.

Anonymous:1994:SI

[Ano94c] Anonymous. 1994 subject index. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):??, December 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/toc-12-94.html.

Anonymous:1994:Aa

[Ano94d]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(1): ??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/94feb/authorj.pdf.

Anonymous:1994:Ab

[Ano94e]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 96–100, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        96to100.pdf.

Anonymous:1994:Ac

[Ano94f]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (3):??, June 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94jun/
jun94_au.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Anonymous:1994:Ad

[Ano94g]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (4):66–68, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_66.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Anonymous:1994:Ae

[Ano94h]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (5):112–115, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_11.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Anonymous:1994:ACE

[Ano94i] Anonymous. Autoloader control electronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6): 13, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_13.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Anonymous:1994:AFD

[Ano94j] Anonymous. Autoloader firmware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):15– 16, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_15.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Anonymous:1994:CSO

[Ano94k] Anonymous. CAD system organization. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):56–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:COD

[Ano94l] Anonymous. Compiler optimizations and debugging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6): 37, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_37.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:EMSa

[Ano94m] Anonymous. Enterprise modeling and simulation applications in reengineering. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):86– 87, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_86.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:EMSb

[Ano94n] Anonymous. Enterprise modeling and simulation research at HP laboratories. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):90– 91, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_90.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:EFP

[Ano94o] Anonymous. Environmentally friendly packaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (1):53–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:GTA

[Ano94p] Anonymous. Glossary of terms and abbreviations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6): 85, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_85.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Anonymous:1994:GG

[Ano94q]  Anonymous.  Graphics

glossary. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):26, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:IVD

[Ano94r] Anonymous. An inside view of the drop generation process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):11–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:MOP

[Ano94s] Anonymous. Modifying office papers to improve inkjet print quality. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(1): 16–??, February 1994. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:NBH

[Ano94t]  Anonymous.  Network

backup with the HP C1553A DDS autoloader. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6): 18, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_18.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:NPD

[Ano94u] Anonymous. A new product development model. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (1):65–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:OTA

[Ano94v]  Anonymous.  Overview

of the Test Access Port. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (6):56–67, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_56.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:PDU

[Ano94w] Anonymous. Package design using 3D solid modeling. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3):29, June 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/toc-6-
94.html.

Anonymous:1994:PCL

[Ano94x] Anonymous. Print cartridges for a large-format color inkjet drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):50–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:PDE

[Ano94y] Anonymous. Product design effect on environmental responsibility and distribution costs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):59–??, February 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SPD

[Ano94z]  Anonymous.  A short

primer on debugger internals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6):39, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94dec/dec94_39.pdf;
http://www.hp.com/hpj/
toc-12-94.html.

Anonymous:1994:BRS

[Ano94-27] Anonymous. Sidebar: Binary ranges speed processing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):18, October 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SHI

[Ano94-28] Anonymous. Sidebar: HP image library scaling functions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):41, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SLP

[Ano94-29] Anonymous. Sidebar: Legal primitive evaluation.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3): 47, June 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SMH

[Ano94-30]  Anonymous.  Sidebar:

MIME header fields. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 76, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SOA

[Ano94-31]  Anonymous.  Sidebar:

Overview of A-law and μlaw data formats. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):65, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SXS

[Ano94-32] Anonymous. Sidebar: X stations in HP MPower. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 16, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:XWS

[Ano94-33]  Anonymous.  Sidebar:

X window system client/ server architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 25, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:SMSb

[Ano94-34]  Anonymous.  The sim-

ple model: Sponsor’s perspective. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):94– 105, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_10.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Anonymous:1994:SMSa

[Ano94-35] Anonymous. Stepper motor simulation model. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(1): 75–??, February 1994. CO-DEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:TAC

[Ano94-36] Anonymous. Theoretical approach to CMOS inverter jitter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):54, October 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1994:WIS

[Ano94-37] Anonymous. What is SFC? Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (4):39, August 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94aug/
                                        aug94_39.pdf;   http:/
                                        /www.hp.com/hpj/toc-8-
                                        94.html.

Anonymous:1994:WNC

[Ano94-38] Anonymous. Whiteboard: A new component of HP SharedX. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):28–29, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Anonymous:1995:Aa

[Ano95a]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 117–123, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        febauth.pdf;   http://
                                        www.hp.com/hpj/toc-02-
                                        95.html.

Anonymous:1995:Ab

[Ano95b]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2): 96–100, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_11.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Anonymous:1995:Ac

[Ano95c]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 96–98, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_96.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Anonymous:1995:Ad

[Ano95d]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):71–74, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_71.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Anonymous:1995:Ae

[Ano95e]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (5):107–110, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95oct/
                                        authoc95.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-95.html.

Anonymous:1995:BRI

[Ano95f] Anonymous. Backscatter RF/ID systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 95, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_95.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Anonymous:1995:BSF

[Ano95g] Anonymous. Bridging and stuck-at faults. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 110, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95111.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        02-95.html.

Anonymous:1995:CPD

[Ano95h] Anonymous. Call progress, DTMF tones, and tone detection. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):73, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_73.pdf;
http://www.hp.com/hpj/
toc-04-95.html.

Anonymous:1995:CI

[Ano95i]  Anonymous.  Caller-ID.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (2):72, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_72.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Anonymous:1995:CEA

[Ano95j]  Anonymous.  Capillary

electrophoresis applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):12–13, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_12.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Anonymous:1995:CTA

[Ano95k] Anonymous. Cross talk analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):22, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_22.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Anonymous:1995:CTU

[Ano95l] Anonymous. Cross talk in unshielded twisted-pair cables. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):19, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_19.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Anonymous:1995:DCC

[Ano95m] Anonymous. Development of a common ChemStation architecture. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):46, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95jun/jun95_46.pdf;
                                        http://www.hp.com/hpj/
                                        toc-06-95.html.

Anonymous:1995:EHD

[Ano95n]  Anonymous.  Exception

handling and development support. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):55, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_55.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Anonymous:1995:FIU

[Ano95o] Anonymous. Fighting inaccuracies: Using perturbation to make Boolean operations robust. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5): 78, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_78.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Anonymous:1995:FDW

[Ano95p] Anonymous. Firmware design for wide-bandwidth IF support and improved measurement speed. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(5):84– 85, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_84.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Anonymous:1995:Ga

[Ano95q]  Anonymous.  Glossary.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):45, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95aug/
                                        aug95_45.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        08-95.html.

Anonymous:1995:Gb

[Ano95r]  Anonymous.  Glossary.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 42, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_42.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Anonymous:1995:Gc

[Ano95s]  Anonymous.  Glossary.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 60, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_60.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Anonymous:1995:HLP

[Ano95t] Anonymous. Higher-level protocols. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):36, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95aug/aug95_36.pdf;
http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:HBL

[Ano95u]  Anonymous.  The HP

8509A/B lightwave polarization analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 32, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_32.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Anonymous:1995:HSV

[Ano95v] Anonymous. The HP 89400 series vector signal analyzers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):87, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_87.pdf;

http://www.hp.com/hpj/
toc-10-95.html.

Anonymous:1995:HCT

[Ano95w] Anonymous. HP CE technology transfer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):16, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_16.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Anonymous:1995:HSS

[Ano95x] Anonymous. HP software solution broker accessible products. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):98, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_98.pdf;
http://www.hp.com/hpj/
toc-04-95.html.

Anonymous:1995:IFF

[Ano95y] Anonymous. IEEE 802.3 and 802.5 frame formats. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):30, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_30.pdf;   http:

//www.hp.com/hpj/toc-
08-95.html.

Anonymous:1995:IDH

[Ano95z] Anonymous. Industrial design of the HP CE instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):18–19, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_18.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Anonymous:1995:IPM

[Ano95-27] Anonymous. Invalid packet marker. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):41, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95aug/aug95_41.pdf;
http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:JC

[Ano95-28] Anonymous. Jones calculus. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1):28, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95feb/feb95_28.pdf;

                                        http://www.hp.com/hpj/
                                        toc-02-95.html.

Anonymous:1995:LCD

[Ano95-29] Anonymous. Liquid crystal display technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4): 53, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95aug/
                                        aug95_53.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        08-95.html.

Anonymous:1995:MS

[Ano95-30]  Anonymous.  Multilevel

signaling. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):21, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_21.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Anonymous:1995:NPL

[Ano95-31] Anonymous. Network protocol layers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):15, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_15.pdf;

http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:NAV

[Ano95-32] Anonymous. A note about VRAMs. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):59, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95aug/aug95_59.pdf;
http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:OMGa

[Ano95-33] Anonymous. Object Management Group. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):86, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_86.pdf;
http://www.hp.com/hpj/
toc-04-95.html.

Anonymous:1995:OMGb

[Ano95-34]  Anonymous.  The Ob-

ject Management Group’s distributed object model. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 76, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_76.pdf;   http:

                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Anonymous:1995:OOP

[Ano95-35]  Anonymous.  Object-

oriented programming. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 79, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_79.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Anonymous:1995:OFL

[Ano95-36] Anonymous. Optical-fiber links for 100VG-AnyLAN. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):26, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95aug/
                                        aug95_26.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        08-95.html.

Anonymous:1995:ONT

[Ano95-37] Anonymous. Other network technologies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):10, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_10.pdf;

http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:ORT

[Ano95-38] Anonymous. An overview of Raid technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 74, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_74.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Anonymous:1995:OIP

[Ano95-39] Anonymous. Overview of the implementation of the PA 7100LC multimedia enhancements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):66, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_66.pdf;
http://www.hp.com/hpj/
toc-04-95.html.

Anonymous:1995:PS

[Ano95-40] Anonymous. The Poincaré sphere. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1):29, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/

                                        95feb/feb95_29.pdf;
                                        http://www.hp.com/hpj/
                                        toc-02-95.html.

Anonymous:1995:PAC

[Ano95-41] Anonymous. Polynomial arithmetic and cyclic redundancy checks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):31, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_31.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Anonymous:1995:PD

[Ano95-42] Anonymous. Potential detection. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1):115, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95feb/feb95115.pdf;
                                        http://www.hp.com/hpj/
                                        toc-02-95.html.

Anonymous:1995:PDH

[Ano95-43] Anonymous. Product design of the HP S1010A flat panel display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(4):57– 58, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_57.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Anonymous:1995:RPH

[Ano95-44] Anonymous. Rapid prototyping for the HP CE project. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):28–29, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_28.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Anonymous:1995:RP

[Ano95-45]  Anonymous.  Related

projects. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):38, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95aug/aug95_38.pdf;
http://www.hp.com/hpj/
toc-08-95.html.

Anonymous:1995:RBE

[Ano95-46]  Anonymous.  Remote

bridge example. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):35, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/

                                        95aug/aug95_35.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Anonymous:1995:FA

[Ano95-47] Anonymous. Sidebar: 1.55mm fiber-optic amplifier. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 9, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95_9.pdf;   http://
                                        www.hp.com/hpj/toc-02-
                                        95.html.

Anonymous:1995:TC

[Ano95-48] Anonymous. True color. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (2):52, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95apr/
                                        apr95_52.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        04-95.html.

Anonymous:1996:APG

[Ano96a] Anonymous. Appendix A: A portion of a GDMO definition for a UNIX password file. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):??, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/oct96/tc-10-
96.htm.

Anonymous:1996:ASC

[Ano96b] Anonymous. Application servers and clients in CDE. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(2): 22–??, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Anonymous:1996:A

[Ano96c]  Anonymous.  Authors.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/oct96/tc-10-
96.htm.

Anonymous:1996:Ca

[Ano96d] Anonymous. Cover. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/aug96/cov0896.
htm.

Anonymous:1996:Cb

[Ano96e] Anonymous. Cover. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:DTP

[Ano96f] Anonymous. Data through paging technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 68–??, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/jun96/jun96.
                                        htm.

Anonymous:1996:D

[Ano96g] Anonymous. Departments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:EDS

[Ano96h] Anonymous. Eye diagrams and sampling oscilloscopes.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Anonymous:1996:G

[Ano96i]  Anonymous.  Glossary.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Anonymous:1996:HEBa

[Ano96j] Anonymous. HP E5200A broadband service analyzer EMC design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Anonymous:1996:HEBb

[Ano96k] Anonymous. HP E5200A broadband service analyzer surface mount assembly. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Anonymous:1996:HED

[Ano96l] Anonymous. HP eyeline display mode. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Anonymous:1996:Ia

[Ano96m] Anonymous. In this issue. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com:
                                        80/hpj/aug96/iti-aug.
                                        htm.

Anonymous:1996:Ib

[Ano96n] Anonymous. In this issue. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/oct96/tc-10-
96.htm.

Anonymous:1996:M

[Ano96o] Anonymous. Macros. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Anonymous:1996:RRP

[Ano96p] Anonymous. Reuse roles:

producers, supporters, and utilizers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):50– 55, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm.

Anonymous:1996:SA

[Ano96q] Anonymous. Sidebar: Annotation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.

                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:SCNa

[Ano96r] Anonymous. Sidebar: Correlation node types. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:SCNb

[Ano96s]  Anonymous.  Sidebar:

Count node. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:SFS

[Ano96t] Anonymous. Sidebar: Fact store and data store. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1996:SFR

[Ano96u] Anonymous. Sidebar: Fusion in the real world. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/aug96/augart3.
htm.

Anonymous:1996:SM

[Ano96v]  Anonymous.  Sidebar:

Management. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4):??, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/aug96/augart5.htm.

Anonymous:1996:SRR

[Ano96w]  Anonymous.  Sidebar:

Reuse roles: Producers, supporters, and utilizers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/aug96/augart5.
htm.

Anonymous:1996:STN

[Ano96x]  Anonymous.  Sidebar:

Table node. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/oct96/tc-10-
96.htm.

Anonymous:1996:SSI

[Ano96y] Anonymous. Sidebar: The software initiative program. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4):??, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/aug96/augart4.htm.

Anonymous:1996:SUN

[Ano96z]  Anonymous.  Sidebar:

Unless node. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5): ??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/oct96/tc-10-
96.htm.

Anonymous:1996:SWI

[Ano96-27]  Anonymous.  Sidebar:

What is fusion? Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4):??, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/aug96/augart3.htm.

Anonymous:1996:TCA

[Ano96-28] Anonymous. Table of Contents August 1996. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/aug96/tc-08-96.
htm.

Anonymous:1996:TAE

[Ano96-29] Anonymous. Transimpedance amplifier O/E converter design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Anonymous:1996:UTC

[Ano96-30]  Anonymous.  Universal

time coordinated (UTC). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.

                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Anonymous:1996:WAa

[Ano96-31]  Anonymous.  What’s

ahead. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4):??, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com:80/
                                        hpj/aug96/aheadaug.htm.

Anonymous:1996:WAb

[Ano96-32]  Anonymous.  What’s

ahead. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):??, October 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/oct96/tc-10-
                                        96.htm.

Anonymous:1997:CAC

[Ano97a]  Anonymous.  A cost-

effective architecture for a 500-MHz counter for glitch trigger. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):22–??, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        97apr/tc-04-97.htm.

Anonymous:1997:DRM

[Ano97b] Anonymous. Detailed routing methods. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):43–??, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/tc-08-97.htm.

Anonymous:1997:GRB

[Ano97c] Anonymous. Global routing — A block-level problem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):42–??, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/global_
routing.pdf;   http://
www.hp.com/hpj/97aug/
tc-08-97.htm.

Anonymous:1997:MM

[Ano97d] Anonymous. Mixed-signal microcontroller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):8–??, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97apr/tc-04-97.htm.

Anonymous:1997:MSM

[Ano97e]  Anonymous.  Modeling

source match effects for microwave power splitter design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):72–??, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97apr/tc-04-97.htm.

Anonymous:1997:OLT

[Ano97f] Anonymous. Oscilloscope/ logic timing analyzer synergy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):31–??, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97apr/tc-04-97.htm.

Anonymous:1997:SPS

[Ano97g] Anonymous. Shmoo plot shapes. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):35–??, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/shmoo.
pdf;   http://www.hp.
com/hpj/97aug/tc-08-97.
htm.

Anonymous:1997:SPA

[Ano97h] Anonymous. Standardization — A phased approach. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(4): 49–??, August 1997. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97aug/
tc-08-97.htm.

Armantrout:1995:IMW

[ANSW95] Robert J. Armantrout, Terrence R. Noe, Christopher E. Stewart, and Leonard M. Weber. An IF module for wide-bandwidth signals. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(5):89– 103, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_89.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Anzinger:1993:MPM

[Anz93] George A. Anzinger. Managing PA-RISC machines for real-time systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4): 31–37, August 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Anderson:1969:CRC

[AP69] G. C. Anderson and M. A.

Perry. A calibrated realtime correlator/averager/ probability analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(3): 9–16, November 1969. CO-DEN HPJOAX. ISSN 0018-1153.

Ainsworth:1976:ALD

[AP76] G. Ainsworth and R. Pope. An accurate, low-noise discriminator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(2):10–??, October 1976. CODEN HPJOAX. ISSN 0018-1153.

Amar:1990:NMH

[AP90] Serge Y. Amar and Michele A. Prieur. Network management for the HP 3000 Datacom and terminal controller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):76–84, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Agrafojo:1992:AMC

[APA92] Ventura Caamano Agrafojo, David Perez, and Josep Abella. An automatic media cutter for a drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):42–48, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Anderson:1976:PRT

[AR76] G. C. Anderson and G. T.

Roberts. Pseudo-random and random test signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(??): ??, ???? 1976. CODEN HPJOAX. ISSN 0018-1153.

Appleyard:1985:PAE

[ARGA85] Aileen C. Appleyard, Roger W.

Ruhnow, William Grant Grovenburg, and Wayne M. Angevine. A protocol analyzer for EDP centers and field service. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(7):4–11, July 1985. CODEN HPJOAX. ISSN 0018-1153.

Alexander:1994:CBS

[ARL^{+}94] Thomas B. Alexander, Kenneth G. Robertson, Dean T. Lindsay, Donald L. Rogers, John R. Obermeyer, John R. Keller, Keith Y. Oka, and Marlin M. Jones, II. Corporate business servers: An alternative to mainframes for business computing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (3):8–30, June 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94jun/
                                        jun94_8t.pdf;   http:/
                                        /www.hp.com/hpj/toc-6-
                                        94.html.

Armantrout:1995:MRW

[Arm95] Robert J. Armantrout. A microwave receiver for wide-bandwidth signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (5):80–88, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95oct/
                                        oct95_80.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-95.html.

Aken:1985:DTF

[AS85] Michael B. Aken and William M. Spaulding. Development of a twochannel frequency synthesizer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(8):11–18, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Allegre:1990:XPA

[AS90] Jean-Pierre Allegre and Marie-Therese Sarrasin. X.25 packet assembler/ disassembler support in the HP 3000 data communications and terminal controller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):63–73, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Allan:1984:NSH

[ASCC84] Marlu E. Allan, Nancy Schoendorf, Craig B. Chatterton, and Don M. Cross. A new series of highperformance real-time computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(2):3–6, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Astrof:1977:FSM

[Ast77] K. L. Astrof. Frequency synthesis in a microwave signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(3):8–15, November 1977. CODEN HPJOAX. ISSN 0018-1153.

Albrecht:1995:DPP

[ASTW95] Alan R. Albrecht, Michael P.

Spratt, Patricia A. Thaler, and Gregory C. A. Watson. Demand priority protocol. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):13–17, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_13.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Arnold:1983:IHH

[AT83] Thomas J. Arnold and Billy E. Thayer. Integration of the HP-75’s handpulled card reader electronics in CMOS. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(6):24– 26, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Albrecht:1995:IAI

[AT95] Alan R. Albrecht and Patricia A. Thaler. Introduction to 100VG-AnyLAN and the IEEE 802.12 local area network standard. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4): 6–12, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95aug/
                                        aug95_6t.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        08-95.html.

Atchison:1992:VPC

[Atc92] Lee Atchison. VXIbus programming in C. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):35– 40, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Aue:1981:FPP

[Aue81] P. Aue. A fast, programmable pulse generator output stage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(9):29–32, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Alonso-Velez:1980:PC

[AVB80] I. I. Alonso-Velez and J. G.

Bourque. The Pascal/ 64000 compiler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(10):20– 21, 23–24, 26–28, October 1980. CODEN HPJOAX. ISSN 0018-1153.

Averett:1975:RES

[Ave75] L. W. Averett. Real-time executive system manages large memories. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(4):21–24, December 1975. CODEN HPJOAX. ISSN 0018-1153.

Alexander:1990:ITH

[AW90] Neil M. Alexander and Randy J. Westra. Integration and test for the HP OSI Express Card’s protocol stack. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(1):72–79 (or 72–77??), February 1990. CODEN HPJOAX. ISSN 0018-1153.

Atkinson:1987:LFA

[AWH^{+}87a] Edward S. Atkinson, Gaylord L. Wahl, Jr., Michael L. Hall, Eric J. Wicklund, and Steven K. Peterson. Lowfrequency analyzer combines measurement capability with modeling and analysis tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(1):4–16, January 1987. CODEN HPJOAX. ISSN 0018-1153.

Atkinson:1987:LAC

[AWH^{+}87b] Edward S. Atkinson, Gaylord L. Wahl, Jr., Michael L. Hall, Eric J. Wicklund, and Steven K. Peterson. Lowfrequency analyzer combines measurement capability with modeling and analysis tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(5):4–15, May 1987. CODEN HPJOAX. ISSN 0018-1153.

Asada:1998:SSQ

[AY98] M. Asada and Pong Mang Yan. Strengthening software quality assurance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2): 89–??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Azmoon:1982:DLH

[Azm82] M. Azmoon. Development of a low-cost, high-quality graphics plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(12):12– 15, December 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Bhaskar:1985:DTS

[BA85] Eldurkar V. Bhaskar and J. Stephen Aden. Development of the thin-film structure for the ThinkJet printhead. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(5):27–31, 33, May 1985. CODEN HPJOAX. ISSN 0018-1153.

Backman:1991:EPT

[Bac91] Rex A. Backman. Easyto-use performance tools with a consistent user interface across HP operating systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3):65–70, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Baker:1981:ILD

[Bak81] T. H. Baker. IC layout on a desktop computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(6):20– 21, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Baker:1984:ANA

[Bak84] Alan J. Baker. An ADC for a network analyzer receiver. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(11):21–23, November 1984. CODEN HPJOAX. ISSN 0018-1153.

Balliew:1984:ATP

[Bal84] Robert E. Balliew. Automatic test program generation for digital board testing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):11– 14, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Bame:1983:HTS

[Bam83] Paul D. Bame. HP 64000 terminal software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(3):6–??, 9–10, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Band:1969:ACI

[Ban69] I. T. Band. Automatic counter inverts period to get frequency. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(9):17– 19, May 1969. CODEN HPJOAX. ISSN 0018-1153.

Banks:1983:UIS

[Ban83] Lawrence W. Banks. An ultrasound imaging system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (10):6, 9–11, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Banno:1984:PTS

[Ban84] Takuo Banno. Powerful test system software provides extensive parametric measurement capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(6):9–11, June 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Bank:1991:PMT

[Ban91]  Leslie Bank.  A pa-

tient monitor two-channel stripchart recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4): 26–28, October 1991. CO-DEN HPJOAX. ISSN 0018-1153.

Bartlett:1981:PGT

[Bar81] D. S. Bartlett. Programming the graphics tablet. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (1):20–21, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Barber:1982:SAU

[Bar82] V. Alan Barber. Synthesizer accuracy for unsynthesized microwave sources. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (9):30–36, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Barbera:1984:PCP

[Bar84] Joseph D. Barbera. Personal computer printer is user installable. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(8):30–31, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Bares:1988:PPP

[Bar88] Steven J. Bares. Printing on plain paper with a thermal inkjet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(6):39–44, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Barkans:1995:HCR

[Bar95] Anthony C. Barkans. HP color recovery technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2): 51–59, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_51.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Barstow:1996:CAD

[Bar96] Arthur F. Barstow. The CDE action and data typing services. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):24– 28, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Baskins:1979:VMT

[Bas79] D. L. Baskins. Virtual memory for TESTAID and FASTRACE (printed circuit board testing programs). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3):17–18, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Bauhaus:1969:MHT

[Bau69] R. H. Bauhaus. Measuring high-frequency transistor parameters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(5):11–14, January 1969. CODEN HPJOAX. ISSN 0018-1153.

Bausch:1980:HCD

[Bau80] J. F. Bausch. A highquality CRT display for a portable computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(7): 19–22, July 1980. CODEN HPJOAX. ISSN 0018-1153.

Beauchamp:1992:IDR

[BAUR92] Robert W. Beauchamp, Josep Giralt Adroher, Joan Uroz, and Isidre Rosello. Improved drawing reliability for drafting plotters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):35–41, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Basiji:1973:CBL

[BB73] J. Basiji and A. B. Bergh.

Central bus links modular HP 3000 hardware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(5): 9–14, January 1973. CO-DEN HPJOAX. ISSN 0018-1153.

Bologlu:1978:MHH

[BB78] A. Bologlu and V. A.

Barber. Microprocessorcontrolled harmonic heterodyne microwave counter also measures amplitudes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(9): 2–4, 6–7, 9–12, 14–16, May 1978. CODEN HPJOAX. ISSN 0018-1153.

Baeverstad:1980:DSD

[BB80] H. L. Baeverstad, Jr. and C. C. Bruderer. Display system designed for color graphics. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(12):25– 27, December 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Batey:1984:SDR

[BB84] Robert M. Batey and James D. Becker. Secondgeneration disc read/write electronics (HP 7933 and 7935 disc products). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1): 7–12, January 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Blakely:1991:CSC

[BB91] Frank W. Blakely and Mark E. Boles. A case study of code inspections. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (4):58–63, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Bash:1997:IHT

[BB97a] C. E. Bash and R. L. Blanco. Improving heat transfer from a flip-chip package. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):121– 125, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a16.htm.

Bray:1997:ONB

[BB97b] Robert C. Bray and Douglas M. Baney. Optical networks: Backbones for universal connectivity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5): ??, December 1997. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97dec/
de97a3.htm.

Burkman:1984:PCS

[BBD^{+}84] Jack L. Burkman, Robert L.

Brooks, Ronald P. Dean, Paul F. Febvre, Michael K. Bowen, S. R. Anderson, and J. R. Milner. Putting a 32-bit computer system in a desktop workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(5): 3–5, 7–9, 11, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Beers:1996:GCM

[BBD^{+}96] Ted W. Beers, Diana K.

Byrne, James A. Donnelly, Robert W. Jones, and Feng Yuan. A graphing calculator for mathematics and science classes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 45–58, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Bartz:1993:BVS

[BBDK93] Manfred Bartz, Keith A.

Bayern, Joseph R. Diederichs, and David F. Kelley. Baseband vector signal analyzer hardware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):31–46, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Beers:1991:HIA

[BBE^{+}91] Ted W. Beers, Diana K. Byrne, Gabe L. Eisenstein, Robert W. Jones, and Patrick J. Megowan. HP 48SX interfaces and applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3):13–21, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Bening:1997:PDG

[BBF^{+}97] L. C. Bening, T. M. Brewer, H. D. Foster, J. S. Quigley, R. A. Sussman, P. F. Vogel, and A. W. Wells. Physical design of 0.35-μ m gate arrays for symmetric multiprocessing servers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):95–103, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97apr/ap97a16.htm.

Blevins:1991:PAP

[BBG91] David W. Blevins, Christopher A. Bartholomew, and John D. Graf. Performance analysis of personal computer workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):92–96, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Bischof:1991:NTS

[BBH91] Richard S. Bischof, Ronald C.

Blanc, and Patrick B. Harper. New technology in synthesized sweeper microcircuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(2):36–46, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Baumgartner:1993:NHS

[BBI^{+}93] William D. Baumgartner, John S. Brenneman, John L. Imperato, Douglas A. Larson, Ricardo de Mello Peregrino, and Gregory A. Taylor. A new high-performance 0.01-to-20-GHz synthesized signal generator microwave chain. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 17–29, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Bass:1995:DMP

[BBJ^{+}95] Mick Bass, Terry W. Blanchard, D. Douglas Josephson, Duncan Weir, and Daniel L. Halperin. Design methodologies for the PA 7100LC microprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):23–35, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95apr/apr95_23.pdf;
                                        http://www.hp.com/hpj/
                                        toc-04-95.html.

Bockhaus:1997:EVH

[BBR^{+}97] J. W. Bockhaus, R. Bhatia, C. M. Ramsey, J. R. Butler, and D. J. Ljung. Electrical verification of the HP PA 8000 processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):32–39, August 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://cpus.hp.
                                        com/techreports/electrical_
                                        verification.pdf;   http:
                                        //www.hp.com/hpj/97aug/
                                        au97a4.htm.

Barnard:1992:ACE

[BBT92] David L. Barnard, Henry Black, and James A. Thalmann. Automatic calibration for easy and accurate power measurements.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 95–100 (or 95–98??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Bek:1995:HCS

[BBW95] Fritz Bek, Franz Bertsch, and Klaus Witt. HP CE separation control electronics and firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 36–43, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_36.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Bradley:1974:MMS

[BC74] H. E. Bradley and C. J.

Christopher. Mass memory system broadens calculator applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(8): 2–12, April 1974. CODEN HPJOAX. ISSN 0018-1153.

Bronson:1976:MML

[BC76] B. Bronson and A. Y.

Chan. A multifunction, multifamily logic pulser. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (4):12–17, December 1976. CODEN HPJOAX. ISSN 0018-1153.

Bray:1981:SF

[BC81a] R. C. Bray and Y. C. Chu. SAWR fabrication. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12): 11–13, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Brooksby:1981:UIC

[BC81b] M. W. Brooksby and P. L. Castro. University and industrial cooperation for VLSI. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):29–??, 32– 33, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Bohorquez:1994:LIT

[BCC^{+}94] Jaime H. Bohorquez, Brian P. Canfield, Kenneth J. Courian, Frank Drogo, Corrina A. E. Hall, C. L. Holstun, A. R. Scandalis, and M. E. Shepard. Laser-comparable inkjet text printing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(1): 9–17, February 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94feb/
9to17.pdf.

Bryg:1996:HLM

[BCF96] William R. Bryg, Kenneth K. Chan, and Nicholas S. Fiduccia. A high-performance, low-cost multiprocessor bus for workstations and midrange servers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):18–24, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/toc-02-96.html.

Brooksby:1981:BQI

[BCH81] M. W. Brooksby, P. L.

Castro, and F. L. Hanson. Benefits of quickturnaround integrated circuit processing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(6):33– 35, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Boyton:1989:SAD

[BCH^{+}89] Jeff R. Boyton, Sankar L.

Chakrabarti, Steven P. Hiebert, John J. Lang, Jens R. Owen, Keith A. Marchington, Peter R. Robinson, Michael H. Stroyan, and John A. Waitz. Sharing access to display resources in the Starbase/X11 Merge system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):20–32, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Booster:1993:DPO

[BCH^{+}93] D. Howard Booster, Harry Chou, Michael G. Hart, Steven J. Mifsud, and Rollin F. Rawson. Design of a precision optical low-coherence reflectometer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):39–48, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Beamer:1994:DHD

[BCK^{+}94] C. Beamer, T. Carlin, G. Kaplinsky, S. Bauer, D. Blair, H. Brower, E. Coiner, M. Hamlin, D. Hunt, R. Little, T. Panah, B. Reid, J. Scheffelin, J. Thoman, D. Timm, and A. Van Liew. Development of the HP DeskJet 1200C print cartridge platform. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):46–54, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/46to54.pdf.

Blue:1993:VSA

[BCO^{+}93] Kenneth J. Blue, Robert T.

Cutler, Dennis P. O’Brien, Douglas R. Wagner, and Benjamin R. Zarlingo. Vector signal analyzers for difficult measurements on time-varying and complex modulated signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6): 6–16, December 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Boeller:1988:HMC

[BCRS88] Cheryl A. Boeller, Timothy J. Carlin, Peter M. Roessler, and Steven W. Steinfield. High-volume microassembly of color thermal inkjet printheads and cartridges. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(4):32–40, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Bloom:1993:NGM

[BCZ93] Alan R. Bloom, Jason A.

Chodora, and James R. Zellers. A new generation of microwave sweepers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):38–45, April 1993.

CODEN HPJOAX. ISSN 0018-1153.

Barney:1973:QLH

[BD73] D. B. Barney and J. R. Drehle. A quiet, low-cost, high-speed line printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(9): 18–22, May 1973. CODEN HPJOAX. ISSN 0018-1153.

Blascow:1984:MAC

[BD84] Stanley M. Blascow, Jr. and James A. Donnelly. Module adds curve-fitting and optimization capabilities to the HP-71B. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7): 22–24, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Bounaix:1984:IWT

[BDF84] Jean Bounaix, Jean-Claude Dureau, and Jacques Firdmann. An industrial workstation terminal for harsh environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(11):25– 29, November 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Braun:1995:PRS

[BDFL95] David M. Braun, Dennis J. Derickson, Luis M. Fernandez, and Greg D. Le-Cheminant. Precision reflectometer with spuriousfree enhanced sensitivity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (1):39–42, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_39.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Baker:1991:HSO

[BDH91] Glen M. Baker, Mark N.

Davidson, and Lance E. Haag. A high-performance sweeper output power leveling system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):24– 30, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Becker:1985:LHR

[BDM85] John C. Becker, Donald A.

DiTommaso, and Sterling J. Mortensen. Lowcost, highly reliable tape backup for Winchester disc drives. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3):34–35, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Black:1985:LPF

[BDOT85] Vonn L. Black, Alan Delwiche, Chris L. Odell, and Stephen B. Tursich. A lowcost, portable field service protocol analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(7):24– 29, July 1985. CODEN HPJOAX. ISSN 0018-1153.

Banta:1985:NCF

[BDS85] Robert H. Banta, Jr., Peter H. Dorward, and Steven A. Scampini. New cardiograph family with ECG analysis capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36 (9):23–28, September 1985. CODEN HPJOAX. ISSN 0018-1153.

Banker:1990:ULH

[BE90] Kimball K. Banker and Michael A. Ellis. The upper layers of the HP OSI Express card stack. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):28–36, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Beckwith:1976:CTN

[Bec76] J. F. Beckwith. Current tracer: A new way to find low-impedance logic-circuit faults. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(4):2–8, December 1976. CODEN HPJOAX. ISSN 0018-1153.

Becker:1979:DME

[Bec79] J. C. Becker. Designing to meet electromagnetic interference requirements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(5): 16–17, May 1979. CODEN HPJOAX. ISSN 0018-1153.

Becker:1980:SPS

[Bec80]  J. S. Becker.  System

45C power supply considerations. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(12):28– ??, December 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Becker:1987:DOP

[Bec87] Josef Becker. Detectors for optical power measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(2):16–21, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Beck:1993:FDD

[Bec93] Patricia A. Beck. Fabrication of diffused diodes for HP lightwave applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (1):49–51, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Beetem:1982:ATP

[Bee82] J. E. Beetem. The advanced terminal processor: a new terminal I/O controller for the HP 3000. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3): 22–25, March 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Beethe:1992:HVD

[Bee92] Douglas C. Beethe. HP VEE: A dataflow architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):84–88, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Bathiany:1970:SMS

[BEFV70] R. H. Bathiany, C. J. Enlow, P. G. Foster, and S. Vitkovits. Sweeping the microwave spectrum with solid-state sources. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22 (3):2–8, November 1970. CODEN HPJOAX. ISSN 0018-1153.

Beller:1993:HSP

[Bel93]  Josef Beller.  A high-

performance signal processing system for the HP 8146A optical time-domain reflectometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):63–68, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Bellis:1998:AP

[Bel98] Matthew W. Bellis. Appendix — packaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1): ??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a7.htm.

Bowers:1995:DLC

[BEM95] Dennis A. Bowers, Gerard M. Enkerlin, and Karen L. Murillo. Development of a low-cost, high-performance, multiuser business server system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):79–84, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95apr/apr95_79.pdf;
                                        http://www.hp.com/hpj/
                                        toc-04-95.html.

Benzel:1995:BBW

[Ben95] Jack D. Benzel. Bugs in black and white: Imaging IC logic levels with voltage contrast. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):102–106, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95apr/apr95_1b.pdf;
                                        http://www.hp.com/hpj/
                                        toc-04-95.html.

Berger:1988:HIM

[Ber88] Arnold S. Berger. Host independent microprocessor development systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (6):45–51, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Best:1975:AOD

[Bes75] A. I. Best. A 100-MHz analog oscilloscope for digital measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(4):2–7, December 1975. CODEN HPJOAX. ISSN 0018-1153.

Bethke:1997:SSA

[Bet97]  R. C. Bethke.  The

SoftBench static analysis database. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):16–18, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a3.htm.

Blasser:1967:ALA

[BF67] H. Blasser and H. Finckh.

Automatic loudness analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(3):12–15, November 1967. CODEN HPJOAX. ISSN 0018-1153.

Blais:1979:APF

[BF79] M. R. Blais and J. L.

Fanton. Automated pulmonary function measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(9):20–24, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Bird:1982:TBR

[BF82] Steven C. Bird and Jack A.

Folchi. Time base requirements for a waveform recorder. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(11):29– 34, November 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Berlin:1989:UAD

[BF89] Lucy M. Berlin and C. Fletcher. User-centered application definition: a methodology and case study. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):90–97, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Burkhart:1983:VM

[BFHT83] Keven P. Burkhart, Mark A. Forsyth, Mark E. Hammer, and Darius F. Tanksalvala. An 18-MHz, 32-bit VLSI microprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(8):7–8, 10– 11, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Burkman:1984:CRP

[BFMP84] Jack L. Burkman, Howell R. Felsenthal, Thomas O. Meyer, and Warren C. Pratt. A compact, reliable power supply for an advanced desktop computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(5): 24–31, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Buttner:1990:FDM

[BFP90] Christian Büttner, Fromut Fritze, and Gerhard Ple. Firmware development for a modular liquid chromatography system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (2):44–50, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Brandt:1994:ASH

[BFS94] Ellen N. Brandt, Thomas G.

Fincher, and Monish S. Shah. Audio support in HP MPower: Overview of A-law and m-law data formats. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):62– 67, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94apr/
62to67.pdf;   http://
www.hp.com/hpj/toc-4-
94.html.

Brown:1989:STR

[BG89] John M. Brown and Thomas J. Gilg. Sharing testing responsibilities in the Starbase/X11 Merge system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):42–46, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Beck:1988:OCD

[BGG88] Juergen Beck, Siegfried Gross, and Robin Giffard. Optical component design for a correlation-based optical time-domain reflectometer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(6):22–28, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Baldwin:1974:LIM

[BGH74] R. R. Baldwin, B. E. Grote, and D. A. Harland. A laser interferometer that measures straightness of travel. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(5):10–20, January 1974. CODEN HPJOAX. ISSN 0018-1153.

Buchanan:1987:DTC

[BGK^{+}87] Gregory F. Buchanan, Francois Gaullier, Olivier Krumeich, Eric Lecesne, Jean-Pierre Picq, and Heng V. Te. Distributed Terminal Controller for HP Precision Architecture computers running the MPE XL operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(3): 21–28, March 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Baldwin:1971:RLI

[BGR71] R. R. Baldwin, G. B. Gordon, and A. F. Rude. Remote laser interferometry. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (4):14–20, December 1971. CODEN HPJOAX. ISSN 0018-1153.

Bloom:1988:HMS

[BGR^{+}88] Alan R. Bloom, Roger R.

Graeber, Kenneth A. Richter, Andrew N. Smith, and Ronald T. Yamada. High-power microwave source for millimeter-wave generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2):26–30, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Buskirk:1983:CDC

[BGS83] William A. Buskirk, Charles W.

Gilson, and David J. Shelly. Compact digital cassette drive for low-cost mass storage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(5):17–24, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Black:1981:ASA

[BH81] K. M. Black and P. K. Hardage. Advanced symbolic artwork preparation (ASAP). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):8–9, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Beethe:1992:VEE

[BH92] Douglas C. Beethe and William L. Hunt. A visual engineering environment for test software development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):72–77, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Bhat:1992:NAA

[Bha92] Sunil Bhat. The Network Advisor analysis and realtime environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):29–33, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Bhat:1994:FTM

[Bha94] Sunil Bhat. FDDI topology mapping. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):97– 105, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_97.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Berkel:1985:VPO

[BHE^{+}85] Werner Berkel, Hans Huning, Volker Eberle, Josef Becker, Bernd Maisenbacher, Wilfried Pless, and Michael Goder. A versatile, programmable optical pulse power meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36 (1):18–27, January 1985. CODEN HPJOAX. ISSN 0018-1153.

Buskirk:1988:DHT

[BHH^{+}88] William A. Buskirk, David E.

Hackleman, Stanley T. Hall, Paula H. Kanarek, Robert N. Low, Kenneth E. Trueba, and Richard R. Van de Poll. Development of a high-resolution thermal inkjet printhead. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):55–61, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Brown:1986:DBM

[BHK^{+}86] Alan S. Brown, Thomas M. Hirata, Ann M. Koehler, Krishnan Vishwanath, Jenny Ng, Michael J. Pechulis, Mark A. Sikes, David E. Singleton, and Judson E. Veazey. Data base management for HP precision architecture computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(12):33–48, December 1986. CODEN HPJOAX. ISSN 0018-1153.

Brewster:1985:FUO

[BHP85] Jon A. Brewster, Karen S. Helt, and James N. Phillips. A friendly Unix operating system user interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36 (10):28–29, 31–33, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Barrett:1970:AMP

[BHW70] P. J. Barrett, R. R. Hay, and P. G. Winninghoff. Adding more precision to spectrum analyzer measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(11):10– 12, July 1970. CODEN HPJOAX. ISSN 0018-1153.

Bailey:1977:WPS

[BHW77] P. W. Bailey, J. W. Hyde, and W. T. Walker. A wide-ranging power supply of compact dimensions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(10): 2–9, June 1977. CODEN HPJOAX. ISSN 0018-1153.

Brokish:1986:CVC

[BHW86] James A. Brokish, David J.

Hodge, and Richard E. Warner. Custom VLSI circuits for series 300 graphics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(9):17–22, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Bianchi:1990:DCS

[Bia90] M. Bianchi. Data capture system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):29–30, December 1990. CODEN HPJOAX. ISSN 0018-1153.

Birgenheier:1991:MMA

[Bir91] Raymond A. Birgenheier.

Measuring the modulation accuracy of π/4 DQPSK signals for digital cellular transmitters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):73– 82, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Birgenheier:1996:OCP

[Bir96] Raymond A. Birgenheier. Overview of code-domain power, timing, and phase measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):73–93, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/toc-02-96.html.

Birnbaum:1997:CCD

[Bir97] Joel S. Birnbaum. Communications challenges of the digital information utility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97dec/
de97a1.htm.

Bennett:1982:PDC

[BJ82] Paul I. Bennett and Victor C. Jones. Portable defibrillator-monitor for cardiac resuscitation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(2): 22–28, February 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Baker:1988:DDC

[BJJN88] Jeffrey P. Baker, David A.

Johnson, Vyomesh Joshi, and Stephen J. Nigro. Design and development of a color thermal inkjet print cartridge. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):6–15, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Band:1971:LMS

[BJM71] I. T. Band, H. J. Jekat, and E. E. May. Lilliputian measuring system does much, costs little. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(12):2–12, August 1971. CODEN HPJOAX. ISSN 0018-1153.

Boos:1995:NLF

[BJPH95] Andreas Boos, Michelle Houghton Jagger, Gunter W. Paret, and Jurgen W. Hausmann. A new, lightweight fetal telemetry system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):82– 93, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/

                                        dec95_82.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Bell:1984:HPL

[BJWW84] R. Frank Bell, Eric W. Johnson, R. Keith Whitaker, and Roger V. Wilcox. Head positioning in a large disc drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1):14–20, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Bale:1985:NLS

[BK85a] Jonathan E. Bale and Harry E. Kellogg. Native language support for computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(6):27– 32, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Busch:1985:DCS

[BK85b] John R. Busch and Alan J. Kondoff. Disc caching in the system processing units of the HP 3000 family of computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(2):21–39, February 1985. CODEN HPJOAX. ISSN 0018-1153.

Brod:1995:PCO

[BK95] Claus Brod and Max R.

Kublin. Providing CAD object management services through a base class library. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):51–60, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_51.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Ban-Kuan:1998:RES

[BKAWBCTK98] Koay Ban-Kuan, Leong Ak-Wing, Tan Boon-Chun, and Yoong Tze-Kwan. Reliability enhancement of surface mount lightemitting diodes for automotive applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3): 71–??, August 1998. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98aug/
tc-08-98.htm.

Bhat:1994:FRM

[BKD94] Sunil Bhat, Robert H.

Kroboth, and Anne L. Driesbach. The FDDI Ring Manager for the HP Network Advisor protocol analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):88–96, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        94oct/oct94_88.pdf;
                                        http://www.hp.com/hpj/
                                        toc-10-94.html.

Berkel:1990:PPG

[BKE^{+}90] Werner Berkel, Gerd Koffmane, Frederick L. Eatock, Patrick Schmid, Heino Hopke, and Hans-Jurgen Snackers. 500-MHz and 300-MHz programmable pulse generators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(4):64–78, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Bergh:1987:HEH

[BKMM87] Arndt B. Bergh, Keith Keilman, Daniel J. Magenheimer, and James A. Miller. HP 3000 emulation on HP precision architecture computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):87– 89, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Busch:1987:MXO

[BKO87] John R. Busch, Alan J. Kondoff, and Darryl Ouye. MPE XL: The operating system for HP’s next generation of commercial computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):68– 86, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Bass:1995:PMC

[BKQW95] Mick Bass, Patrick Knebel, David W. Quint, and William L. Walker. The PA 7100LC microprocessor: A case study of IC design decisions in a competitive environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(2):12– 22, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_12.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Bostak:1993:CSG

[BKS93] Christopher J. Bostak, Camala S. Kolseth, and Kevin G. Smith. Concurrent signal generator engineering and manufacturing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 30–37, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Bianchi:1989:DCH

[BKV89] Mark J. Bianchi, Jeffery J. Kato, and David J. Van Maren. Data compression in a half-inch reel-to-reel tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(3):26– 31, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Burns:1983:DSH

[BL83] Robert R. Burns and Dale A. Luck. Designing software for highperformance graphics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(9): 9–14, September 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Barrett:1987:SPD

[BL87] George E. Barrett and John H. Lau. A study of panel deflection of partially routed printed circuit boards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(4):29–34 (or 29– 33??), April 1987. CODEN HPJOAX. ISSN 0018-1153.

Blazek:1975:MK

[Bla75] O. Blazek. A microprocessorscanned keyboard. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(10):20– 21, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Blair:1986:DTS

[Bla86] Steven R. Blair. A defect tracking system for the UNIX environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 15–18, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Balazer:1977:PIS

[BLKH77] L. P. Balazer, G. W. Lynch, R. M. Kemplin, and L. W. Hennessee. Pen and ink system helps assure fourcolor plotter line quality. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (1):20–25, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Barker:1977:CDP

[BLMJ77] R. D. Barker, V. L. Laing, J. E. Marriott, and H. M. Juneau. Contemporary design practice in generalpurpose digital multimeters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(6):20–28, February 1977. CODEN HPJOAX. ISSN 0018-1153.

Bhargava:1986:NMS

[BLNP86] Raj K. Bhargava, Teri L. Lombardi, Alvina Y. Nishimoto, and Robert A. Passell. New methods for software development: System for just-in-time manufacturing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6):11–18, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Brooks:1983:HES

[BLSM83] Elizabeth Brooks, Robert J. Livengood, Rex C. Smith, and Timothy F. Myers. High-capability electronics system for a compact, battery-operated computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6): 10–15, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Baron:1981:DHL

[BLTY81] W. D. Baron, L. LaBarre, C. E. Tyler, and R. G. Younge. Development of a high-performance, lowmass, low-inertia plotting technology. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(10):3–7, October 1981. CODEN HPJOAX. ISSN 0018-1153.

Bollinger:1992:PHK

[BLY92] D. E. Bollinger, F. P. Lemmon, and D. L. Yamine. Providing HP-UX kernel functionality on a new PA-RISC architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3): 11–14, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Bergh:1979:CHC

[BM79] A. B. Bergh and K. C. Y.

Mei. Cost-effective hardware for a compact integrated business computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(7): 3–8, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Bracken:1981:IGS

[BM81] D. F. Bracken and W. J.

McCalla. An interactive graphics system for structured design of integrated circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):18–19, 21– 25, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Bockman:1986:MGS

[BM86] Francis E. Bockman and Emil Maghakian. Measurement graphics software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37 (1):27–32, January 1986. CODEN HPJOAX. ISSN 0018-1153.

Bohley:1988:FSC

[BMD^{+}88] Thomas K. Bohley, Mark E. Mathews, Lewis R. Dove, Joe K. Millard, David W. Bigelow, and Donald D. Skarke. Front-end signal conditioning for a highspeed digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):67–69, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Brown:1987:EDA

[BMS87] Preston D. Brown, Gregory J. May, and Megha Shyam. Electronic design of an advanced technical handheld calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (8):34–39, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Baily:1978:OPE

[BMTW78] E. M. Baily, W. A. McIlvanie, W. T. Thrash, and D. B. Winterrowd. Optimizing the performance of an electromechanical print mechanism. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(15):23– 31, November 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Blancke:1980:PME

[BN80] T. B. Blancke and L. L.

Nielsen. Patient monitoring enhanced by new central station. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):3–11, November 1980. CODEN HPJOAX. ISSN 0018-1153.

Blasciak:1993:SPA

[BNB93] Andrew J. Blasciak, David L.

Neuder, and Arnold S. Berger. Software performance analysis of realtime embedded systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (2):107–115, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Bobroff:1973:ATT

[Bob73]  D. A. Bobroff.  Auto-

mated transceiver testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(12):

2–7, August 1973. CODEN HPJOAX. ISSN 0018-1153.

Bologlu:1974:CAH

[Bol74] A. Bologlu. A completely automatic 4-GHz heterodyne frequency converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(10): 19–21, June 1974. CODEN HPJOAX. ISSN 0018-1153.

Bologlu:1975:ACP

[Bol75] A. Bologlu. Automatic 4.5-GHz counter provides 1-Hz resolution. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(1):14–18, September 1975. CODEN HPJOAX. ISSN 0018-1153.

Bologlu:1980:AFC

[Bol80] A. Bologlu. A 26.5-GHz automatic frequency counter with enhanced dynamic range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4):20–22, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Bugely:1981:CHE

[BOOS81] F. L. Bugely, I. F. Osborne, G. Owen, and R. B. Schudy. Calibration of the HP electron beam lithography system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(5):27– 33, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Bortolotto:1990:HOE

[Bor90] Elizabeth P. Bortolotto.

HP OSI Express design for performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):51–58, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Botka:1971:MII

[Bot71] J. K. Botka. More informative impedance measurements, swept from 0.5 to 100 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(3):15–19, November 1971. CODEN HPJOAX. ISSN 0018-1153.

Bowman:1972:EDD

[Bow72] D. J. Bowman. An efficient disc drive/computer interface. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(9):15– 16, May 1972. CODEN HPJOAX. ISSN 0018-1153.

Boyan:1980:MMM

[Boy80] C. J. Boyan. Making the most of a microprocessorbased instrument controller (HP 8903A audio analyser). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(8):8–9, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Bordon:1969:SD

[BP69] H. C. Bordon and G. P. Pighini. Solid-state displays. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(6):2–12, February 1969. CODEN HPJOAX. ISSN 0018-1153.

Brooksby:1970:MTA

[BP70] M. Brooksby and R. D. Pering. Monolithic transistor arrays for highfrequency applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21 (5):15–16, January 1970. CODEN HPJOAX. ISSN 0018-1153.

Baker:1972:CCF

[BP72] M. Baker and J. Pipkin. Clip-and-read comparator finds IC failures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(5):8–12, January 1972. CODEN HPJOAX. ISSN 0018-1153.

Bump:1976:CIP

[BP76] R. B. Bump and G. R.

Paulson. Character impact printer offers maximum printing flexibility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(10): 19–23, June 1976. CODEN HPJOAX. ISSN 0018-1153.

Bones:1977:RTP

[BP77] D. A. Bones and M. L.

Patterson. Remote terminal plotter offers simple programming and efficient communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(1):9–13, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Bidwell:1982:AMC

[BP82] John L. Bidwell and David W. Palermo. Advanced multilingual computer systems for measurement automation and computer-aided engineering applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(5):3–7, May 1982. CODEN HPJOAX.

Balazer:1983:ASA

[BP83] S. A. Balazer and J. M.

Perry. AUTOPLOT/ 2700: a simple approach to custom chart generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(9): 31–34, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Bartle:1985:SFU

[BP85] David A. Bartle and Katherine F. Potter. Synthesizer firmware for user interface and instrument control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(8):21–24, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Beller:1993:MAO

[BP93] Josef Beller and Wilfried Pless. A modular all-haul optical time-domain reflectometer for characterizing fiber links. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):60–62, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Byrne:1994:ASG

[BPA^{+}94a] Diana K. Byrne, Charles M. Patton, David Arnett, Ted W. Beers, and Paul J. McClellan. An advanced scientific graphing calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(4):6–22, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_6t.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Byrne:1994:SUV

[BPA^{+}94b] Diana K. Byrne, Charles M.

Patton, David Arnett, Ted W. Beers, and Paul J. McClellan. Sidebar: User versions of interface tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (4):20, August 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_20.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Barnett:1987:ITM

[BPMF87] G. Octo Barnett, Judith L.

Piggins, Gordon T. Moore, and Ethan A. Foster. Information technology and medical education. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(10): 28–35, October 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Blazek:1978:RSG

[BR78] O. Blazek and M. B. Raynham. Raster scan graphics with zoom and pan. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(5): 6–12, January 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Bear:1991:RSE

[BR91] S. P. Bear and T. W. Rush. Rigorous software engineering: a method for preventing software defects. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (5):24–31, December 1991. CODEN HPJOAX. ISSN 0018-1153.

Bradford:1978:RRS

[Bra78] K. J. Bradford. Remotelycontrolled RF switch for multipoint tests in communications systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(12):20– 22, August 1978. CODEN HPJOAX. ISSN 0018-1153.

Bray:1996:MTA

[Bra96] Jacqueline A. Bray. A modeling toolset for the analysis and design of OSI network management objects. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):43–51, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/oct96/oc96a5.htm.

Brewster:1974:AEC

[Bre74] J. L. Brewster. An automatic exposure control for a lab-bench X-ray camera. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(11):20–24, July 1974. CODEN HPJOAX. ISSN 0018-1153.

Bresniker:1994:PRS

[Bre94] Kirk M. Bresniker. PA-RISC symmetric multiprocessing in midrange servers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3): 31–33, June 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94jun/
jun94_31.pdf;   http:/
/www.hp.com/hpj/toc-6-
94.html.

Bartlett:1989:USM

[BRHS89] Paul F. Bartlett, Paul F.

Robinson, Tracey A. Hains, and Mark J. Simms. Use of structured methods for real-time peripheral firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):79–86, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Briley:1994:PMD

[Bri94] Daniel L. Briley. Polyester media development for inkjet printers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):28–34, February 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/94feb/28to34.pdf.

Brown:1982:HOS

[Bro82]  W. A. Brown.  High-

pressure oxidation (si IC thermal oxidation). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(8): 34–36, August 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Brown:1995:HD

[Bro95] Lisa S. Brown. 100VG-AnyLAN 15-port hub design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):39–42, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        95aug/aug95_39.pdf;
                                        http://www.hp.com/hpj/
                                        toc-08-95.html.

Brunetti:1973:MRC

[Bru73]  L. Brunetti.  A more

rugged, cleaner writing oscillographic ink recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(9): 10–17, May 1973. CODEN HPJOAX. ISSN 0018-1153.

Brunetti:1977:NFI

[Bru77] L. G. Brunetti. A new family of intelligent multi-color X-Y plotters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(1):2–4, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Brubaker:1978:IPM

[Bru78] R. H. Brubaker, Jr. An intelligent peripheral for measurement and control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (11):2–6, 8–9, July 1978. CODEN HPJOAX. ISSN 0018-1153.

Brubaker:1979:AAM

[Bru79a] L. E. Brubaker. Assuring accuracy in modulation measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(11):22– 26, November 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Brubaker:1979:SSS

[Bru79b] L. E. Brubaker. Special signal source tests Modulation Analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(11):30– 32, November 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Brunton:1997:HSR

[Bru97] R. Scott Brunton. Highsample-rate multiprocessorbased oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2): 34–36, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a7.htm.

Bronson:1979:MLT

[BS79] B. Bronson and M. Slater. Microprocessor lab teaches operation and troubleshooting. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(10):3–8, October 1979. CODEN HPJOAX. ISSN 0018-1153.

Baney:1990:LPS

[BS90] Douglas M. Baney and Wayne V. Sorin. Linewidth and power spectral measurements of single- frequency lasers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):92–96, February 1990. CODEN HPJOAX. ISSN 0018-1153.

Bailey:1993:OSA

[BS93] David A. Bailey and James R. Stimple. Optical spectrum analyzers with high dynamic range and excellent input sensitivity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):60–67, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Baney:1995:BFC

[BS95] Douglas M. Baney and Wayne V. Sorin. Broadband frequency characterization of optical receivers using intensity noise. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 6–12, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_6t.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Borg:1998:THM

[BS98] Matthew M. Borg and Kalwant Singh. Testing with the HP 9490 mixedsignal LSI tester. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):??, August 1998. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        98aug/tc-08-98.htm.

Boeller:1992:LTI

[BSME92] Robert A. Boeller, Samuel A. Stodder, John F. Meyer, and Victor T. Escobedo. A large-format thermal inkjet drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):6–15, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Bronstein:1989:SDC

[BSY89] Kenneth H. Bronstein, David J. Sweetser, and William R. Yoder. System design for compatibility of a high-performance graphics library and the X Window System. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):6–12 (or 6–10??), December 1989. CODEN HPJOAX. ISSN 0018-1153.

Bartlett:1988:DHF

[BT88] Debra S. Bartlett and Joel D. Tesler. A discless HP-UX file system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):10–14, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Blanchard:1997:PMH

[BT97] T. W. Blanchard and P. G.

Tobin. The PA 7300LC microprocessor: a highly integrated system on a chip. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 43–47, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/highly_
integrated.pdf;   http:
//www.hp.com/hpj/97jun/
ju97a6.htm.

Born:1998:OAE

[BT98] Torsten Born and Peter Thoma. OTDR APIs enable customers to build their own systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 50(1):13–??, November 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98nov/tc-11-98.htm.

Baker:1983:GCL

[BTBW83] Tamara C. Baker, William T.

Toms, James C. Bratnober, and Gerald T. Wade. Graphics capabilities on a laser printer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(11):17– 22, November 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Bockman:1994:HDP

[BTE^{+}94] Kevin M. Bockman, Anton Tabar, Erol Erturk, Robert R. Giles, and William H. Schwiebert. HP DeskJet 1200C printer architecture. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):55–66, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/55to66.pdf.

Baldwin:1983:LOC

[BTW83] Richard R. Baldwin, Larry E. Truhe, and David C. Woodruff. Laser optical components for machine tool and other calibrations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (4):14–16, 18, 20–22, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Buffenbarger:1996:ITR

[Buf96] James R. Buffenbarger. Interface translation for reuse of assembly-language modules in a two-language environment. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(3):76– 81, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Bugarin:1986:TST

[Bug86] John R. Bugarin. Triggers:

a software testing tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 35–36, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Burns:1985:HSG

[Bur85] Robert E. Burns. A highperformance signal generator for RF communications testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(12):4–6, December 1985. CODEN HPJOAX. ISSN 0018-1153.

Burroughs:1986:TOO

[Bur86] Gregory D. Burroughs. A toolset for object-oriented programming in C. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 19–23, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Burgoon:1989:GIM

[Bur89] David A. Burgoon. Global illumination modeling using radiosity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(6):78–88, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Burdick:1993:CMR

[Bur93] Matthew J. Burdick. Continuous monitoring of remote networks: the RMON MIB. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):82–89, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Burgoon:1995:DUE

[Bur95] David A. Burgoon. Development and use of electronic schematic capture in the specification and simulation of a structuredcustom ASIC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 88–91, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95jun/
                                        jun95_88.pdf;   http:

//www.hp.com/hpj/toc-
06-95.html.

Butler:1994:VRC

[But94] Hamish Butler. Virtual remote: The centralized expert. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):75–82, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94oct/oct94_75.pdf;
http://www.hp.com/hpj/
toc-10-94.html.

Buted:1995:ZBD

[But95] Rolando R. Buted. Zero bias detector diodes for the RF/ID market. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):94– 98, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_94.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Berkel:1985:OS

[BV85] Werner Berkel and Joachim Vobis. Optical standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36 (1):29–30, January 1985. CODEN HPJOAX. ISSN 0018-1153.

Blackham:1987:MMD

[BVAP87] Raymond C. Blackham, James A. Vasil, Edward S. Atkinson, and Ronald W. Potter. Measurement modes and digital demodulation for a low-frequency analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(1):17–25, January 1987. CODEN HPJOAX. ISSN 0018-1153.

Brown:1972:MMC

[BW72] C. D. Brown and J. M. Walden. Model 10 maintains compatibility, expands capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(4):5–8, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Bullock:1976:ETS

[BW76] M. L. Bullock and R. E. Warren. Electronic total station speeds survey operations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(8):2–12, April 1976. CODEN HPJOAX. ISSN 0018-1153.

Burgoon:1981:SQO

[BW81] J. R. Burgoon and R. L. Wilson. SC-cut quartz oscillator offers improved performance. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(3):20–21, 23– 25, 28–29, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Birnbaum:1985:BRH

[BW85] Joel S. Birnbaum and William S. Worley, Jr. Beyond RISC: high-precision architecture. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(8):4–10, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Ballo:1992:MTA

[BW92] David J. Ballo and John A.

Wendler. The microwave transition analyzer: A new instrument architecture for component and signal analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):48–62, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Balza:1984:LAN

[BWW84] John J. Balza, H. Michael Wenzel, and James L. Willits. A local area network for the HP 9000 Series 500 computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(3):22–23, 25– 27, March 1984. CODEN HPJOAX. ISSN 0018-1153.

                                        URL https://www.hpl.
                                        hp.com/hpjournal/pdfs/
                                        IssuePDFs/1984-03.pdf.

Button:1986:TAA

[BYA86] Brian T. Button, R. Michael Young, and Diane M. Ahart. Troubleshooting aid for asynchronous data communications links. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):42– 47, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Byrne:1996:RTI

[Byr96] Patrick J. Byrne. Reducing time to insight in digital system integration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (3):6–14, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/jun96/jun96.
                                        htm.

Beyers:1983:VTP

[BZS83] Joseph W. Beyers, Eugene R. Zeller, and S. Dana Seccombe. VLSI technology packs 32-bit computer system into a small package. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(8):3–6, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Chu:1978:UCR

[CAF78] D. C. Chu, M. S. Allen, and A. S. Foster. Universal counter resolves picoseconds in time interval measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(12):2–11, August 1978. CODEN HPJOAX. ISSN 0018-1153.

Cagan:1986:IHA

[Cag86] Martin R. Cagan. An introduction to Hewlett– Packard’s AI workstation technology. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(3):4–14, March 1986. CODEN HPJOAX. ISSN 0018-1153.

Cagan:1990:HSE

[Cag90] Martin R. Cagan. The HP SoftBench environment: an architecture for a new generation of software tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (3):36–47, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Colburn:1994:AAT

[CAJ^{+}94] William S. Colburn, Randell A. Agadoni, Michael M. Johnson, Edward Wiesmeier, III, and Glen Oldenburg. Automated assembly and testing of HP DeskJet 1200C print cartridges. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):79–84, February 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/94feb/79to84.pdf.

Campbell:1982:MDO

[Cam82] J. W. Campbell. Mechanical design of the oscilloscope measurement system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (9):24–26, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Campbell:1994:CDD

[Cam94] Von C. Campbell. Customerdriven development of a new high-performance data acquisition system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5): 6–8, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94oct/

oct94_6t.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Canestri:1993:OCL

[Can93] Franco A. Canestri. Online CO_{2} laser beam realtime control algorithm for orthopedic surgical applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):68–72, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Cargile:1969:CST

[Car69] W. P. Cargile. A computercontrolled system for testing digital logic modules. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(7): 14–20 (or 2–20??), March 1969. CODEN HPJOAX. ISSN 0018-1153.

Carua:1978:EHD

[Car78] F. P. Carua. Easy-to-use, high-resolution digitizer increases operator efficiency. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(16):2–13, December 1978. CODEN HPJOAX. ISSN 0018-1153.

Carpenter:1979:AFO

[Car79] R. L. Carpenter. AMIGO/ 300: a friendly operating system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(7):17– 24, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Cardwell:1981:CSD

[Car81] S. M. Cardwell. Capacitive stylus design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(1):17–18, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Carmichael:1995:CSO

[Car95] Cheryl Carmichael. COBOL SoftBench: An open integrated CASE environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):82–87, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_82.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Cassidy:1995:LWA

[Cas95a] John J. Cassidy, Jr. The log weighted average for measuring printer throughput. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (5):104–106, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_10.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Caswell:1995:EDA

[Cas95b] Deborah L. Caswell. An evolution of DCE authorization services. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):49– 54, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_49.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Cates:1978:FEC

[Cat78] J. A. Cates. F-series extends computing power of HP 1000 computer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (14):12–17, October 1978. CODEN HPJOAX. ISSN 0018-1153.

Catlin:1982:WPL

[Cat82] J. F. Catlin. A widedynamic-range pulse leveling scheme. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7):26–27, 29, 32, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Cauzid:1985:MDL

[Cau85] Michel Cauzid. Mechanical design of a low-cost terminal. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4):8–9, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Colpitts:1969:GOC

[CAV69] R. W. Colpitts, D. Allen, and T. Vos. Graphical output for the computing calculator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(13):2–7, September 1969. CODEN HPJOAX. ISSN 0018-1153.

Cyrus:1991:FSS

[CBH91] J. L. Cyrus, J. D. Bledsoe, and P. D. Harry. Formal specification and structured design in software development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(5):51–58, December 1991. CODEN HPJOAX. ISSN 0018-1153.

Cripe:1996:CDE

[CBL96] Brian E. Cripe, Jon A. Brewster, and Dana E. Laursen. A common desktop environment for platforms based on the UNIX operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(2): 6–13, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm;   http://www.hp.
com/hpj/apr96/ap96a1.
pdf.

Coles:1995:PS

[CCC^{+}95] Alistair N. Coles, David G.

Cunningham, Joseph A. Curcio, Jr., Daniel J. Dove, and Steven G. Methley. Physical signaling in 100VG-AnyLAN. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(4):18– 26, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_18.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Carlson:1991:SAD

[CCH^{+}91] Kristen C. Carlson, James H.

Cauthorn, Timothy L. Hillstrom, Roy L. Mason, Joseph F. Tarantino, Jay M. Wardle, and Eric J. Wicklund. A 10-hz-to-150-MHz spectrum analyzer with a digital IF section. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3): 44–64, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Carder:1987:ITI

[CDE^{+}87] Norman Carder, William I. Dunn, James H. Elliott, David W. Grieve, and W. Gordon Rhind. Inservice transmission impairment testing of voicefrequency data circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (10):4–15, October 1987. CODEN HPJOAX. ISSN 0018-1153.

Culp:1988:STD

[CDG^{+}88] Bradfred W. Culp, Douglas R. Domel, Wayne T. Gregory, Jeffery J. Kato, Gerod C. Melton, Kraig A. Proehl, David W. Ruska, Virgil K. Russon, and Peter Way. Streaming tape drive control electronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3): 43–54, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Chang:1996:FCC

[CDLN96] Justin S. Chang, Richard Dugan, Benny W. H. Lai, and Margaret M. Nakamoto. A 1.0625-Gbit/s Fibre Channel chipset with laser driver. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (1):60–67, February 1996. CODEN HPJOAX. ISSN 0018-1153. URL http://

www.hp.com/hpj/toc-02-
96.html.

Cross:1971:ODR

[CDS71] J. R. Cross, J. A. Doub, and J. M. Stedman. Online data reduction for nuclear analyzers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(7):2–3, March 1971. CODEN HPJOAX. ISSN 0018-1153.

Cross:1981:SR

[CE81] P. S. Cross and S. S. Elliott. Surface-acoustic-wave resonators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12):9–11, 13– 14, 16–17, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Colwell:1968:LGO

[CF68] J. M. Colwell and P. F. Febvre. A low-cost, generalpurpose oscillator with low distortion and high stability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(4):12–15, December 1968. CODEN HPJOAX. ISSN 0018-1153.

Cooke:1990:DSH

[CFK^{+}90] Beth E. Cooke, Colleen S. Fettig, Paul B. Koski, Darrell O. Swope, and Roy M. Vandoorn. Directory services in the HP MAP 3.0 environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(4):15–23, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Cook:1975:INP

[CFW75] M. J. Cook, G. M. Fichter, and R. E. Whicker. Inside the new pocket calculators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(3):8–12, November 1975. CODEN HPJOAX. ISSN 0018-1153.

Coleman:1987:FPD

[CG87] Derek Coleman and Robin M. Gallimore. A framework for program development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (10):37–40, October 1987. CODEN HPJOAX. ISSN 0018-1153.

Cripe:1998:DTM

[CG98] B. E. Cripe and T. A. Gaskins. The DirectModel toolkit: Meeting the 3D graphics needs of technical applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):19– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Cutler:1993:RVS

[CGH^{+}93] Robert T. Cutler, William J.

Ginder, Timothy L. Hillstrom, Kevin L. Johnson, Roy L. Mason, and J. Pietsch. RF vector signal analyzer hardware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):47–59, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Carlson:1986:HAG

[CGS^{+}86] Robert J. Carlson, Atul Garg, Arie Scope, Craig Wassenberg, and Lyle A. Weiman. HP AdvanceNet: a growth-oriented computer networking architectural strategy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):6–10, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Crumly:1982:LPM

[CH82] James D. Crumly and Von L. Hansen. Laser Printer machine control system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7):11–15, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Chaffee:1970:FHC

[Cha70] D. Chaffee. A fast-writing, high-frequency cathode-ray tube. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(5):9–10, January 1970. CODEN HPJOAX. ISSN 0018-1153.

Chan:1977:ESA

[Cha77] A. Y. Chan. Easy-to-use signature analyzer accurately troubleshoots complex logic circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(9):9–14, May 1977. CODEN HPJOAX. ISSN 0018-1153.

Champine:1991:VUI

[Cha91] Mark A. Champine. A visual user interface for the HP-UX and Domain operating systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):88–96, 98–99, February 1, 1991.

CODEN HPJOAX. ISSN 0018-1153.

Chadayammuri:1996:PBI

[Cha96a] Prabha G. Chadayammuri.

A platform for building integrated telecommunication network management applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):6–16, October 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/oct96/oc96a1.htm.

Chakrabarti:1996:SCC

[Cha96b] Sankar L. Chakrabarti.

Synlib: The core of CDE tests. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):62– 65, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Chen:1986:PIM

[Che86] James Chen. Power and intensity measurements for ultrasonic Doppler imaging systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6):31–34, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Chisholm:1981:DCH

[Chi81] H. C. Chisholm. Digital control for a highperformance programmable signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):8–??, 10–11, February 1981. CO-DEN HPJOAX. ISSN 0018-1153.

Chiu:1982:SZE

[Chi82] Kuang Yi Chiu. SWAMI: a zero-encroachment local oxidation process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(8):31–32, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Chikarmane:1990:ULA

[Chi90] Sanjay B. Chikarmane. Upper layer architecture for HP MAP 3.0 OSI services. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4):11–15, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Coutant:1986:CNG

[CHK86] Deborah S. Coutant, Carol L. Hammond, and Jon W. Kelley. Compilers for the new generation of Hewlett– Packard computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(1): 4–18, January 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Campbell:1995:OOP

[CHK^{+}95] Mark C. Campbell, David K.

Hinds, Ana V. Kapetanakis, David S. Levin, Stephen J. McFarland, David J. Miller, and J. Scott Southworth. Object-oriented perspective on software system testing in a distributed environment. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):75– 81, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_75.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Chan:1996:DHP

[CHK^{+}96] Kenneth K. Chan, Cyrus C.

Hay, John R. Keller, Gordon P. Kurpanek, Francis X. Schumacher, and Jason Zheng. Design of the HP PA 7200 CPU. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):25–33, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/pa7200_

                                        design.pdf;   http://
                                        www.hp.com/hpj/toc-02-
                                        96.html.

Clegg:1986:HOS

[CHKS86] Frederick W. Clegg, Gary Shiu-Fan Ho, Steven R. Kusmer, and John R. Sontag. HP-UX operating system on HP precision architecture computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(12): 4–12, 15–22, December 1986. CODEN HPJOAX. ISSN 0018-1153.

Clifford:1976:MCD

[CHM76] D. M. Clifford, F. T. Hickenlooper, and A. C. Mortensen. Mid-range calculator delivers more power at lower cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(10):24– 31, June 1976. CODEN HPJOAX. ISSN 0018-1153.

Cunningham:1997:DLE

[CHNJ97] David G. Cunningham, Delon C. Hanson, Mark C. Nowell, and C. Steven Joiner. Developing leadingedge fiber-optic network link standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97dec/de97a7.htm.

Chow:1990:SCS

[Cho90] Chee K. Chow. Statistical circuit simulation of a wideband amplifier: a case study in design for manufacturability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(5):78–81, October 1990. CODEN HPJOAX. ISSN 0018-1153.

Covelli:1985:PCL

[CHP85] Ernest F. Covelli, Von L.

Hansen, and David L. Price. Printer command language provides feature set standard for HP printers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(6):23–25, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Chu:1974:TIA

[Chu74] D. C. Chu. Time interval averaging: theory, problems, and solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(10): 12–15, June 1974. CODEN HPJOAX. ISSN 0018-1153.

Chumbley:1979:EPD

[Chu79] S. L. Chumbley. Extending possibilities in desktop computing (Hewlett– Packard model 9835A/B computer). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(5):11– 13, May 1979. CODEN HPJOAX. ISSN 0018-1153.

Chu:1988:PDN

[Chu88] David C. Chu. Phase digitizing: A new method for capturing and analyzing spread-spectrum signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (1):28–32, 34–35, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Cook:1996:MOH

[CHW96] Lori A. Cook, Steven P. Hiebert, and Michael R. Wilson. A media-rich online help system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):38– 49, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Crawford:1971:BPS

[CJ71] R. D. Crawford and G. Justice. A bantam power supply for a minicomputer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (2):13–15, October 1971. CODEN HPJOAX. ISSN 0018-1153.

Crouch:1995:C

[CJ95] Simon E. C. Crouch and Jonathan Jedwab. Coding in 100VG-AnyLAN. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):27–32, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_27.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Campbell:1993:MDN

[CJHE93] John W. Campbell, Kenneth W. Johnson, Wayne F. Helgoth, and William H. Escovitz. Mechanical design of a new oscilloscope mainframe for optimum performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(5):66–72, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Collison:1976:CFC

[CK76] R. R. Collison and R. E. Kmetovicz. Calibrated FM, crystal stability, and counter resolution for a low-cost signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(7): 11–17, March 1976. CO-DEN HPJOAX. ISSN 0018-1153.

Cathell:1989:CSA

[CKP89] B. David Cathell, Michael B. Kalstein, and Stephen J. Pearce. A compiled source access system using CD-ROM and personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):50–57, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Cavanna:1992:ESH

[CL92] Vicente V. Cavanna and Christopher S. Liu. The EISA standard for the HP 9000 Series 700 workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):78–82, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Casey:1998:CEO

[CL98] R. J. Casey and L. L. Lindstone. Concurrent engineering in OpenGL’s product development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):41– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Clark:1979:BC

[Cla79] G. R. Clark. A business computer for the 1980s. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(6): 3–5, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Clare:1981:PCN

[Cla81a] C. R. Clare. A process control network. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(6):30– 31, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Clarke:1981:CIP

[Cla81b] E. L. Clarke. Cost-effective industrial packaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(7): 31–??, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Clegg:1979:RCP

[Cle79] F. W. Clegg. Reducing the cost of program development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(6):9–15, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Clements:1993:MCI

[Cle93] Brad Clements. Mechanical considerations for an industrial workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):62–67, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Carey:1987:LSD

[CLG87] Thomas J. Carey, John C. Lovell, and Thomas L. Grisell. Low-noise synthesizer design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):34– 38, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Cannon:1981:PDF

[CLL81] M. J. Cannon, H. F. Lee, and R. B. Lewis. Pattern data flow in the HP electron beam system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(5):24– 26, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Clark:1992:ECH

[CLL^{+}92] David S. Clark, Andrea C.

Lantz, Christopher S. Liu, Thomas E. Parker, and Joseph H. Steinmetz. EISA cards for the HP 9000 Series 700 workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):83–96, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Cline:1972:NCD

[CM72] S. G. Cline and N. D.

Marschke. New capabilities in digital low-frequency spectrum analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(10):14– 20, June 1972. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/72jun/ju72a4.htm.

Christensen:1975:NML

[CM75] S. Christensen and I. Matthews.

A new microwave link analyzer for communications systems carrying up to 2700 telephone channels. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27 (3):13–23, November 1975.

CODEN HPJOAX. ISSN 0018-1153.

Cochran:1968:IPC

[Coc68] D. S. Cochran. Internal programming of the 9100A calculator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(1):14–16, September 1968. CODEN HPJOAX. ISSN 0018-1153.

Cochran:1972:AAH

[Coc72] D. S. Cochran. Algorithms and accuracy in the HP-35. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(10):10– 11, June 1972. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/72jun/ju72a2.htm.

Collins:1986:MEC

[Col86] Douglas M. Collins. Molecularscale engineering of compound semiconductor materials. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):4–10, November 1986. CODEN HPJOAX. ISSN 0018-1153.

Colby:1988:DPM

[Col88a] Lee H. Colby. Design and performance of millimeterwave thermocouple sensors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 35–38, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Colinge:1988:SMD

[Col88b] Jean-Pierre P. Colinge.

Silicon-on-insulator MOS devices for integrated circuit applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):87–93, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Connolly:1993:PPS

[Con93] Brian Connolly. A process for preventing software hazards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):47–52, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Culbertson:1993:HTL

[COO^{+}93] W. Bruce Culbertson, Toshiki Osame, Yoshisuke Otsuru, J. Barry Shackleford, and Motoo Tanaka. The HP Tsutsuji logic synthesis system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):38–51, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Cornwell:1996:HDA

[Cor96] Patricia Collins Cornwell. HP domain analysis: Producing useful models for reusable software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):46– 49, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart5.
htm.

Coster:1976:DQF

[Cos76] J. H. Coster. Designing a quiet frequency synthesizer for a selective level measuring set. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(5):13–15, January 1976. CODEN HPJOAX. ISSN 0018-1153.

Cotton:1996:EFC

[Cot96] Todd Cotton. Evolutionary fusion: A customeroriented incremental life cycle for fusion. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):25– 38, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.

com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart3.
htm.

Coury:1971:PPA

[Cou71] F. F. Coury. Price, performance, architecture and the 2100A computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(2): 2–3, October 1971. CO-DEN HPJOAX. ISSN 0018-1153.

Coury:1972:MWC

[Cou72]  F. F. Coury.  Micro-

programming and writable control store. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(11):16– 20, July 1972. CODEN HPJOAX. ISSN 0018-1153.

Courant:1994:SMC

[Cou94] Joseph J. Courant. Soft-Bench message connector: Customizing software development tool interactions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3):34–39, June 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94jun/jun94_34.pdf;
http://www.hp.com/hpj/
toc-6-94.html.

Cowley:1970:DAS

[Cow70] A. M. Cowley. Design and application of silicon IM-PATT diodes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(9):2–13, May 1970. CODEN HPJOAX. ISSN 0018-1153.

Cox:1996:STD

[Cox96]  Alasdair D. Cox.  A

software toolkit for developing telecommunications application components. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):62–69, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com:80/
                                        hpj/oct96/oc96a7.htm.

Cozzens:1978:ATP

[Coz78] R. J. Cozzens. Advanced thermal page printer has high-resolution graphics capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(8):22–??, April 1978. CODEN HPJOAX. ISSN 0018-1153.

Cheng:1979:EUR

[CP79] Tu-Ting Cheng and W. Peikes. An easy-to-use report generation language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(6):20– 23, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Chapuis:1985:LCB

[CP85] Jean-Louis Chapuis and Michele Prieur. A low-cost, compact, block-mode computer terminal. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4):4–7, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Callister:1987:PAH

[CP87] James R. Callister and Craig W. Pampeyan. Performance analysis of the HP 3000 series 70 hardware cache. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(1):38–48, January 1987. CODEN HPJOAX. ISSN 0018-1153.

Corcoran:1988:OAC

[CPK88] John J. Corcoran, Ken Poulton, and Knud L. Knudsen. A one-gigasample-per-second analog-to-digital converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):59–66, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Cline:1975:CSA

[CPR75] S. G. Cline, R. H. Perdriau, and R. F. Rauskolb. A computer system for analog measurements on voiceband data channels. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(9): 10–17, May 1975. CODEN HPJOAX. ISSN 0018-1153.

Crowley:1973:PAM

[CR73] W. L. Crowley and F. Rode. A pocket-sized answer machine for business and finance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(9):2–9, May 1973. CODEN HPJOAX. ISSN 0018-1153.

Conklin:1981:RBC

[CR81] D. R. Conklin and T. L. Revere, III. Reading bar codes for the HP-41C programmable calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (1):11–14, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Ceely:1989:DHT

[CR89] Gary A. Ceely and David J. Rustici. Design for high throughput in a system digital multimeter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(2):31– 38, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Coutant:1992:SLH

[CR92] C. A. Coutant and M. A.

Ruscetta. Shared libraries for HP-UX. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(3):46– 53, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Crabtree:1972:MMR

[Cra72] M. Crabtree. MLA measures RF performance with down converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(1):17–18, September 1972. CODEN HPJOAX. ISSN 0018-1153.

Craven:1985:DMP

[Cra85] John S. Craven. Dot matrix printbar design and manufacturing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(6):6–9, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Crandall:1992:DCD

[Cra92] William W. Crandall. Design challenges for distributed LAN analysis.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (1):66–76 (or 66–74??), February 1992. CODEN HPJOAX. ISSN 0018-1153.

Crook:1979:AIC

[Cro79]  D. T. Crook.  Analog

in-circuit component measurements: problems and solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3):19–22, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Crow:1989:EAN

[Cro89] William M. Crow. Encapsulation of applications in the NewWave environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):57–66 (or 57– 64??), August 1989. CO-DEN HPJOAX. ISSN 0018-1153.

Courtin:1977:VSF

[CRSS77] E. Courtin, W. Ruchay, P. Salfeld, and H. Sommer. A versatile, semiautomatic fetal monitor for non-technical use. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(5):16–23, January 1977. CODEN HPJOAX. ISSN 0018-1153.

Crawford:1973:HPG

[CRSY73] T. Crawford, J. Robertson, J. Stinson, and I. Young. A high-speed pattern generator and an error detector for testing digital systems. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(3):16–24, November 1973. CODEN HPJOAX. ISSN 0018-1153.

Carlson:1976:UPC

[CS76] J. E. Carlson and R. L.

Stickle. Using a programmable calculator as a data communications terminal. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(6):19–24, February 1976. CODEN HPJOAX. ISSN 0018-1153.

Christensen:1978:NMS

[CS78] A. H. Christensen and D. C. Salomaki. New memory systems for HP 1000 computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(14):23– 27, October 1978. CODEN HPJOAX. ISSN 0018-1153.

Collins:1980:CTT

[CS80] D. J. Collins and B. G. Spreadbury. A compact tape transport subassembly designed for reliability and low cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(7):14– 19, July 1980. CODEN HPJOAX. ISSN 0018-1153.

Chambers:1982:SMS

[CS82] Donald R. Chambers and Steven N. Sanders. Solidstate microwave signal generators for today’S exacting requirements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(7):16–20 (or 16–19??), July 1982. CODEN HPJOAX. ISSN 0018-1153.

Coackley:1988:IVD

[CS88] Robert Coackley and Howard L. Steadman. An integrated voice and data network based on virtual circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (1):42–49, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Chou:1993:HHO

[CS93] Harry Chou and Wayne V. Sorin. High-resolution and high-sensitivity optical reflection measurements using white-light interferometry. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):52–59, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Chang:1991:HPO

[CSS^{+}91] Kok-Wai Chang, Siegmar Schmidt, Wayne V. Sorin, Jimmie L. Yarnell, Harry Chou, and Steven A. Newton. A high-performance optical isolator for lightwave systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):45–50, February 1991. CODEN HPJOAX. ISSN 0018-1153.

Collison:1985:WFR

[CSWR85] Robert R. Collison, James B.

Summers, Marvin W. Wagner, and Bryan D. Ratliff. Wide-frequency-range signal generator output section design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(12):18– 24, December 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Curtis:1985:DHH

[CT85] Hoyle L. Curtis and Richard T.

Turley. Development of a high-performance, halfinch tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3):11–12, 14– 16, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Cunniff:1998:HKP

[Cun98] R. A. Cunniff. HP Kayak: A PC workstation with advanced graphics performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2):35–??, May 1998. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        98may/tc-05-98.htm.

Cunningham:1972:IAS

[CW72] M. Cunningham and L. Wheelwright. Introducing the automatic spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(6):2–6, February 1972. CODEN HPJOAX. ISSN 0018-1153.

Carlson:1992:VMT

[CW92] Larry L. Carlson and Wayne H. Willis. The VX-Ibus in a manufacturing test environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):75–77 (or 75–76??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Crook:1979:HDA

[CWF^{+}79] D. T. Crook, B. M. Wood, F. L. Fiedler, K. Firooz, and R. H. Burger. Hardware design of an automatic circuit board tester. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3): 27–32, March 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Cristal:1982:HWC

[CWFS82] Edward G. Cristal, Arthur N.

Woo, Philip G. Foster, and Ronald F. Stiglich. High-performance wideband cavity-tuned solidstate oscillators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(7):20– 26, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Czenkusch:1989:HDT

[Cze89] David A. Czenkusch. Highresolution digitizing techniques with an integrating digital multimeter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 39–49, April 1989. CODEN HPJOAX. ISSN 0018-1153.

daCunha:1994:CFS

[da 94] John M. da Cunha. A compact and flexible signal conditioning system for data acquisition. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5): 9–15, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94oct/
                                        oct94_9t.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-94.html.

Daniels:1975:PSA

[DA75] J. W. Daniels and R. L. Atchley. A precision spectrum analyzer for the 10-Hz-to-13-MHz range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(9): 18–24, May 1975. CODEN HPJOAX. ISSN 0018-1153.

Dack:1976:MMM

[Dac76]  D. G. Dack.  Making

the most of microprocessor control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(5):16–18, January 1976. CODEN HPJOAX. ISSN 0018-1153.

Damon:1974:MAP

[Dam74] N. E. Damon. Measuring analog parameters of voiceband data channels. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(12): 2–7, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Daniels:1996:HPM

[Dan96] Melanie M. Daniels. A hybrid power module for a mobile communications telephone. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):66– 72, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Dascher:1996:MPC

[Das96] David J. Dascher. Measuring parasitic capacitance and inductance using TDR. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(2): 83–96, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Derickson:1995:HPL

[DBB^{+}95] Dennis J. Derickson, Patricia A. Beck, Tim L. Bagwell, David M. Braun, Julie E. Fouquet, Forrest G. Kellert, Michael J. Ludowise, William H. Perez, Tirumala R. Ranganath, Gary R. Trott, and Susan R. Sloan. High-power, low-internal-reflection, edge emitting light-emitting diodes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 43–48, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_43.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

DiVittorio:1988:FLT

[DCN^{+}88] Mark J. DiVittorio, Brian Cripe, Claude W. Nichols, Michael S. Ard, Kevin R. Hudson, and David J. Neff. Firmware for a laser-quality thermal inkjet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(5):81–86, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Dea:1986:PSG

[DD86] Robert W. Dea and Vincent J. D’Angelo. P-Pods: a software graphical design tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3):32–35, March 1986. CODEN HPJOAX. ISSN 0018-1153.

Duesing:1997:CRC

[DD97] T. J. Duesing and J. R.

Diamant. CodeAdvisor: rule-based C++ defect detection using a static database. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):19–21, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a4.htm.

Diserens:1983:LDG

[DDT83] Craig W. Diserens, Curtis L. Dowdy, and William R. Taylor. Logic design for a graphics subsystem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(9): 15–18, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Dickinson:1976:PPP

[DE76] P. D. Dickinson and W. E.

Egbert. A pair of programcompatible personal programmable calculators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (3):2–8, November 1976. CODEN HPJOAX. ISSN 0018-1153.

Deininger:1994:SWP

[Dei94] Axel Deininger. Sidebar: WYSIWYG printing in an X application. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):86– 87, April 1994. CODEN HPJOAX. ISSN 0018-1153.

DesJardin:1992:VST

[Des92] Lawrence A. DesJardin. VXIbus: A standard for test and measurement system architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):6–14 (or 6–8??), April 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Dethlefsen:1972:CMS

[Det72] A. E. Dethlefsen. Communicationsoriented microwave solidstate sweeper. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(1):19–20, September 1972. CODEN HPJOAX. ISSN 0018-1153.

DeVilbiss:1970:WOA

[DeV70] A. J. DeVilbiss. A wideband oscilloscope amplifier. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(5):11–14, January 1970.

CODEN HPJOAX. ISSN 0018-1153.

DeVilbiss:1980:RSL

[DeV80] A. J. DeVilbiss. Resource sharing in the logic development system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(10):7–8, October 1980. CODEN HPJOAX. ISSN 0018-1153.

DeVries:1981:PDP

[DeV81]  R. L. DeVries.  Prod-

uct design for precision and purity. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):28–30, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Danielson:1979:SSS

[DF79] D. D. Danielson and S. E.

Froseth. A synthesized signal source with function generator capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30 (1):18–26, January 1979. CODEN HPJOAX. ISSN 0018-1153.

Dalichow:1982:BFP

[DF82] Rolf Dalichow and Douglas E. Fullmer. A broadband, fully programmable microwave sweep oscillator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(2):3–10, February 1982. CODEN HPJOAX. ISSN 0018-1153.

Daniels:1986:SMP

[DF86] Thomas H. Daniels and John A. Fenoglio. A standalone measurement plotting system (HP 7090A). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37 (1):20–24, January 1986. CODEN HPJOAX. ISSN 0018-1153.

Deininger:1990:MCB

[DF90] Axel O. Deininger and Charles V. Fernandez. Making computer behavior consistent: The HP OSF/Motif graphical user interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(3):6–12, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Daumuller:1991:MIH

[DF91] Karl Daumuller and Erwin Flachslander. Mechanical implementation of the HP component monitoring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):44–48, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Dukes:1970:TYG

[DG70] J. N. Dukes and G. B. Gordon. A two-hundred-foot yardstick with graduations every microinch. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(12):2–8, August 1970. CODEN HPJOAX. ISSN 0018-1153.

Dildine:1988:MVN

[DG88] Robert G. Dildine and James D. Grace. Millimeterwave vector network analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2):12–18, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Dmitroca:1997:EAN

[DGH^{+}97] R. W. Dmitroca, S. G.

Gibson, T. R. Hill, L. M. Morales, and Chong Tean Ong. Emulating ATM network impairments in the laboratory. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):45– 50, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a9.htm.

Day:1985:TDR

[DGHP85] Clifford K. Day, C. Girvin Harkins, Stephan P. Howe, and Paul Poorman. Thinfilm disc reliability — the conservative approach. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(11): 25–31, November 1985. CODEN HPJOAX. ISSN 0018-1153.

Donnelly:1980:EMS

[DGM80] J. B. Donnelly, G. A. Greenley, and M. E. Muterspaugh. Emulators for microprocessor system development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(10):13– 20, October 1980. CODEN HPJOAX. ISSN 0018-1153.

Dove:1998:LCR

[DGN98] Lewis R. Dove, Martin L. Guth, and Dean B. Nicholson. A low-cost RF multichip module packaging family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):??, August 1998. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        98aug/tc-08-98.htm.

Dalichow:1976:PRS

[DH76] R. Dalichow and D. R.

Harkins. A precision RF source and down-converter for the Model 8505A network analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(11):12– 16, July 1976. CODEN HPJOAX. ISSN 0018-1153.

Daniels:1977:EIL

[DH77] Thomas H. Daniels and Larry W. Hennessee. Easyto-use interface language controls HP-IB plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (1):5–9, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Derynck:1998:IRT

[DH98] Ronald R. Derynck and Thomas W. Hutchinson. Integrating real-time systems with corporate information systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 50(1):26–??, November 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98nov/tc-11-98.htm.

Dias:1981:PSU

[Dia81] J. Fleming Dias. Physical sensors using SAW devices. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12):18–20, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Diaz:1994:AEO

[Dia94] Carlos H. Diaz. Automation of electrical overstress characterization for semiconductor devices (CMOS processes). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):106– 111, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_10.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Dickey:1974:DCS

[Dic74] S. Dickey. Distributed computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(3):2–11, November 1974. CODEN HPJOAX. ISSN 0018-1153.

Dickinson:1978:VLG

[Dic78] P. D. Dickinson. Versatile low-cost graphics terminal is designed for ease of use. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(5):2–6, January 1978. CODEN HPJOAX. ISSN 0018-1153.

Dickie:1984:CCA

[Dic84] James P. Dickie. Custom CMOS architecture for a handheld computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7): 14–17, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Dietrich:1975:VIB

[Die75] H. E. Dietrich. Visualizing interface bus activity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(5):19–24, January 1975. CODEN HPJOAX. ISSN 0018-1153.

DiFrancesco:1975:HPS

[DiF75] M. DiFrancesco. A highcurrent power supply for systems that use 5-volt IC logic extensively. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(8):14– 19, April 1975. CODEN HPJOAX. ISSN 0018-1153.

DiPietro:1980:TIE

[DiP80] D. M. DiPietro. Third input extends range to 1300 MHz. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(9):24–25, September 1980. CODEN HPJOAX. ISSN 0018-1153.

DiVittorio:1984:HDI

[DiV84] Mark J. DiVittorio. Highquality, dot-matrix impact printer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(11):30– 32, November 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Dryden:1983:AOA

[DJS83] P. C. Dryden, H. R. Johnson, Jr., and D. H. Smith. Advanced oven assures repeatability in new gas chromatograph. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(1):30–31, 33– 34, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Dorward:1981:NPT

[DKMS81] P. H. Dorward, S. J. Koerper, M. K. Mason, and S. A. Scampini. New plotting technology leads to a new kind of electrocardiograph. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(10):16– 24, October 1981. CODEN HPJOAX. ISSN 0018-1153.

Dean:1990:CSE

[DKW90] Steven M. Dean, David A.

Kumpf, and H. Michael Wenzel. CONE: A software environment for network protocols. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):18–28, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Dudley:1973:SHD

[DL73] R. L. Dudley and V. L.

Laing. A self-contained, hand-held digital multimeter — a new concept in instrument utility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(3):2–9, November 1973. CODEN HPJOAX. ISSN 0018-1153.

DeWeese:1977:NPH

[DL77] J. E. DeWeese and T. R.

Ligon. An NMOS process for high-performance LSI circuits. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(3):26–29, 31, November 1977. CODEN HPJOAX. ISSN 0018-1153.

Dicolen:1997:DFO

[DL97] A. A. Dicolen and J. J. Liu.

Developing fusion objects for instruments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):73–85, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a10.htm.

Delic:1998:KH

[DL98] K. A. Delic and D. Lahaix. Knowledge harvesting. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2):74–??, May 1998. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98may/
tc-05-98.htm.

Drennan:1985:LSD

[DLJ85] George A. Drennan, Robert J. Lawton, and Michael B. Jacobson. In-line sputtering deposition system for thin-film disc fabrication. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(11): 21–25, November 1985. CODEN HPJOAX. ISSN 0018-1153.

Dorweiler:1997:DMC

[DMJCB97] P. J. Dorweiler, F. E. Moore, D. D. Josephson, and G. T. Colon-Bonet. Design methodologies and circuit design trade-offs for the HP PA 8000 processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):16–21, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/design_
methodologies.pdf;   http:
//www.hp.com/hpj/97aug/
au97a2.htm.

Dangelo:1994:PCF

[DMW94] M. T. Dangelo, R. Movaghar, and A. K. Wilson. Print cartridge fixturing and maintenance in the HP DeskJet 1200C printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (1):67–71, February 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94feb/
67to71.pdf.

Davidson:1986:IDC

[DN86] Andrew W. Davidson and Harold B. Noyes. I/O and data communications in portable computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7): 14–17, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Day:1996:TGH

[DNC96] Stewart W. Day, Geoffrey H. Nelson, and Thomas F. Cappellari. The Third-Generation HP ATM tester. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):68– 73, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Dodge:1984:TTT

[DNG84] Allan W. Dodge, Scott S. Neal, and Kurt R. Goldsmith. Testing the TIMS (Transmission Impairment Measuring Set). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(9):15–18, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Donahue:1977:DSA

[DO77] T. L. Donahue and J. P. Oliverio. Digital signal analyzer applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(2):17–21, October 1977. CODEN HPJOAX. ISSN 0018-1153.

Donn:1969:GWD

[Don69]  E. Donn.  Generating

words for digital testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(8):

8–13, April 1969. CODEN HPJOAX. ISSN 0018-1153.

Donn:1971:MDP

[Don71] E. S. Donn. Manipulating digital patterns with a new binary sequence generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(8):2–8, April 1971. CODEN HPJOAX. ISSN 0018-1153.

Donnelly:1986:MAD

[Don86] James A. Donnelly. Module adds data logging capabilities to the HP-71B computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3):43–46, March 1986. CODEN HPJOAX. ISSN 0018-1153.

Donnelly:1996:CHA

[Don96] James A. Donnelly. Creating HP 38G aplets. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 59–63, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Dehni:1997:INH

[DOR97] T. Dehni, J. O’Connell, and N. Raguideau. Intelligent networks and the HP OpenCall technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48 (4):46–57, August 1997. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97aug/
au97a6.htm.

Doub:1975:CRC

[Dou75] J. A. Doub. Cost-effective, reliable CRT terminal is first of a family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(10):2–5, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Doumas:1992:UIH

[Dou92] Thomas A. Doumas. The user interface for the HP 4980 Network Advisor protocol analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):22–28, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Dudley:1973:HEM

[DP73] B. W. Dudley and R. D. Peck. High efficiency modular power supplies using switching regulators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25 (4):15–20, December 1973. CODEN HPJOAX. ISSN 0018-1153.

Dunn:1984:SKF

[DP84] L. Dunn and M. R. Perkins.

A standard keyboard family for HP computer products. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(8):34–36, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Dong:1988:RAS

[DPA^{+}88] John W. Dong, Kraig A.

Proehl, Ronald L. Abramson, Leslie G. Christie, Jr., and Douglas R. Domel. A reliable, autoloading, streaming half-inch tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):36–42, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Duell:1981:PLD

[DR81] A. W. Duell and W. V.

Roland. Power line disturbances and their effect on computer design and performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(8):25–32, August 1981. CODEN HPJOAX. ISSN 0018-1153.

Dervisoglu:1995:CSL

[DR95] Bulent I. Dervisoglu and Michael Ricchetti. Component and system level design-for-testability features implemented in a family of workstation products. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):107–113, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_1a.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Deeney:1998:CFD

[DR98] Jeffrey L. Deeney and C. Michael Ramsey. Comparison of finite-difference and SPICE tools for thermal modeling of the effects of nonuniform power generation in high-power CPUs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 50 (1):37–??, November 1998. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98nov/
tc-11-98.htm.

Drake:1977:DAM

[Dra77] H. D. Drake. Development and application of microprograms in a realtime environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(7):15–17, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Dresch:1974:NDM

[Dre74] D. A. Dresch. Nonlinear distortion measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25 (12):12–13, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Douglas:1982:DPR

[DS82] J. Guy Douglas and David Stockton. Design of a precision receiver for an integrated test set. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(4): 8–17, April 1982. CODEN HPJOAX. ISSN 0018-1153.

Dimond:1994:DTA

[DSB94] Steven A. Dimond, M. Simms, and M. G. Bertagne. DDS-2 tape autoloader: Highcapacity data storage in a 5 1/4-Inch form factor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (6):12–20, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_12.pdf;   http:

                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Dauner:1990:MDO

[DSC^{+}90] Daniel R. Dauner, Raymond C. Sherman, Michael L. Christensen, Jennifer L. Methlie, and Leslie G. Christie, Jr. Mechanical design of an optical disk autochanger. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):14–23, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Davis:1983:ELD

[DSW83] Michael W. Davis, John A. Scharrer, and Robert G. Wickliff, Jr. Extensive logic development and support capability in one convenient system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3):3–5, 7, 9– 10, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Dierschow:1982:DCC

[DU82] Carl M. Dierschow and Robert P. Uhlrich. Data communications for the 9826A and 9836A computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(5):33– 36, May 1982. CODEN HPJOAX.

Dubuc:1994:FRC

[Dub94] Martin Dubuc. Frame relay conformance testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (5):83–87, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_83.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Duerr:1974:DEA

[Due74] J. R. Duerr. A data error analyzer for tracking down problems in data communications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(6):11–20, February 1974. CODEN HPJOAX. ISSN 0018-1153.

Duff:1987:ART

[Duf87]  John A. Duff.  Auto-

mated radio testing shortens test time and enhances accuracy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(7):36–37, 39, July 1987. CODEN HPJOAX. ISSN 0018-1153.

Dulphy:1985:FAP

[Dul85] Christian-Marcel Dulphy.

Fully automated production of display terminal printed circuit assemblies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4): 16–17, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Doue:1985:CEA

[DV85] John C. Doue and Anthony G. Vallance. Computeraided ECG analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(9): 29, 31–33, September 1985. CODEN HPJOAX. ISSN 0018-1153.

Dong:1985:STD

[DVE85] John W. Dong, David J. Van Maren, and Robert D. Emmerich. Streaming tape drive hardware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3): 25–29, March 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Dunsmore:1991:LTS

[DVl91] Joel P. Dunsmore and John V. Valle-lunga. 20-GHz lightwave test set and accessories. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):23–33, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Drake:1987:RTA

[DW87] H. Dean Drake and Duane E. Wolting. Reliability theory applied to software testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(4):35–39, April 1987. CODEN HPJOAX. ISSN 0018-1153.

DeBaets:1992:MPW

[DW92a] Andrew J. DeBaets and Kathleen M. Wheeler. Midrange PA-RISC workstations with price/performance leadership. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(4):6–11, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Dethlefsen:1992:DCM

[DW92b] Michael Dethlefsen and John A. Wendler. Design considerations in the microwave transition analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):63–71, October 1992. CODEN HPJOAX. ISSN 0018-1153.

David:1997:DEV

[DWB^{+}97] F. K. David, F. W. Woodhull, II, R. R. Barg, J. P. Dunsmore, D. C. Bender, B. A. Brown, and S. E. Jaffe. Design of an enhanced vector network analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):66– 77, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a12.htm.

Deane:1988:PHA

[DWD88] Patrick D. Deane, Simcoe Walmsley, Jr., and Farid Dibachi. Packaging a high-performance 250megasample-per-second analogto-digital converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(1): 49–52, February 1988. CO-DEN HPJOAX. ISSN 0018-1153.

Dupre:1975:BSN

[DY75] J. J. Dupre and C. J. Yansouni. Broadband swept network measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(7):15–17, March 1975. CODEN HPJOAX. ISSN 0018-1153.

Dysart:1989:NOM

[Dys89]  John A. Dysart.  The

NewWave object management facility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(4):17–23, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Egbert:1986:ASA

[EAW^{+}86] J. H. Egbert, K. F. Anderson, F. W. Woodhull, II, J. Rowell, Jr., D. C. Bender, K. A. Richter, and J. C. Faick. Advanced scalar analyzer system improves precision and productivity in R and D and production testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37 (2):24–39, February 1986. CODEN HPJOAX. ISSN 0018-1153.

Eaton:1982:EMM

[EDF82] John T. Eaton, Andrew W.

Davidson, and William R. Frolik. Extended memory and modularity are added to the Series 80 computer family. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(12):3–7, December 1982. CODEN HPJOAX. ISSN 0018-1153.

Edwards:1975:DPM

[Edw75]  A. P. Edwards.  Digi-

tal power meter offers improved accuracy, hands-off operation, systems compatibility (microwave). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(2): 2–7, October 1975. CO-DEN HPJOAX. ISSN 0018-1153.

Edwards:1979:MAA

[Edw79a] A. P. Edwards. Modulation analyzer applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(11): 19–21, November 1979. CODEN HPJOAX. ISSN 0018-1153.

Edwards:1979:PCA

[Edw79b] A. P. Edwards. Precise, convenient analysis of modulated signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(11):3–9, 11– 18, November 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Edwards:1979:STY

[Edw79c]  R. C. Edwards.  SOS

technology yields low-cost HP 3000 computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(9): 3–6, 8, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Edwards:1984:HDD

[Edw84] Stephen A. Edwards. Highcapacity disc drive servomechanism design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1): 23–27, January 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Edwards:1987:VSG

[Edw87] Allen P. Edwards. Vector signal generation and analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(11):4–5, December 1987. CODEN HPJOAX. ISSN 0018-1153.

Epstein:1984:HDD

[EEH^{+}84] James S. Epstein, Glenn R.

Engel, Donald R. Hiller, Glen L. Purdy, Jr., Bryan C. Hoog, Jr., and Eric J. Wicklund. Hardware design for a dynamic signal analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(12):12–17, December 1984. CODEN HPJOAX. ISSN 0018-1153.

Egan:1972:VIO

[Ega72] G. L. Egan. Versatile input/output structure welcomes peripheral variety. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24 (4):24–26, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Egbert:1977:PCAa

[Egb77a] W. E. Egbert. Personal calculator algorithms. I. square roots. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(9):22– 23, May 1977. CODEN HPJOAX. ISSN 0018-1153.

Egbert:1977:PCAb

[Egb77b] W. E. Egbert. Personal calculator algorithms. II. trigonometric functions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(10): 17–20, June 1977. CODEN HPJOAX. ISSN 0018-1153.

Egbert:1977:PCAc

[Egb77c]  W. E. Egbert.  Per-

sonal calculator algorithms. III. inverse trigonometric functions. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(3):22–23, November 1977. CODEN HPJOAX. ISSN 0018-1153.

Egbert:1978:PCA

[Egb78] W. E. Egbert. Personal calculator algorithms. IV. logarithmic functions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(8): 29–32, April 1978. CODEN HPJOAX. ISSN 0018-1153.

Erturk:1988:IRC

[EGH^{+}88] Erol Erturk, Brian D.

Gragg, Mary E. Haviland, W. Wistar Rhoads, Jim L. Ruder, and Joseph E. Scheffelin. Ink retention in a color thermal inkjet pen. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):41–44, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Ellsworth:1978:GHC

[EH78] A. L. Ellsworth and K. Hasebe.

Generating high-speed CRT displays from digital data. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(9): 17–23, May 1978. CODEN HPJOAX. ISSN 0018-1153.

Epstein:1981:VIM

[EH81] J. S. Epstein and T. J.

Heger. Versatile instrument makes high-performance transducer-based measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(7):9–15, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Engel:1984:ISD

[EH84] Glenn R. Engel and Donald R. Hiller. Instrument software for dynamic signal analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(12):17– 19, December 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Edison:1981:PHE

[EHS81] J. C. Edison, W. C. Haase, and R. K. Scudder. A precision high-speed electron beam lithography system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(5): 3–13, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Ehlers:1986:EMD

[EJG86] Eric R. Ehlers, Sigurd W. Johnsen, and Douglas A. Gray. Extending millimeterwave diode operation to 110 GHz. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):10– 14, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Eskeldson:1993:DOT

[EKW93] David D. Eskeldson, Reginald Kellum, and Donald A. Whiteman. A digitizing oscilloscope time base and trigger system optimized for throughput and low jitter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):24–30, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Evel:1982:DWS

[EL82] Eddie A. Evel and Robert M.

Landgraf. Digital waveform storage for the Oscilloscope Measurement System. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(9):15–20, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Eaton:1986:DHP

[ELC^{+}86] John T. Eaton, Carl B.

Lantz, Clifford B. Cordy, Jr., James W. Pearson, Michael J. Barbour, Courtney Loomis, and Ella M. Duyck. Design of HP’s portable computer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7): 4–5, 7–9, 11–13, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Epstein:1988:EHO

[ELN88] Howard C. Epstein, Mark G. Leonard, and Robert Nicol. Economical, high-performance optical encoders. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(5):99– 106, October 1988. CO-DEN HPJOAX. ISSN 0018-1153.

Elo:1993:DMM

[Elo93]  Mark A. Elo.  DECT

measurements with a microwave spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6): 98–106, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Epstein:1981:IOS

[ELU81] H. C. Epstein, M. G. Leonard, and J. J. Uebbing. An incremental optical shaft encoder kit with integrated optoelectronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (10):10–15, October 1981. CODEN HPJOAX. ISSN 0018-1153.

Elward:1974:MMM

[Elw74] J. S. Elward. The millionword minicomputer main memory. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(2):19–20, October 1974. CODEN HPJOAX. ISSN 0018-1153.

Eads:1976:HNL

[EM76] W. D. Eads and D. S. Maitland. High-performance NMOS LSI processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(10): 15–18, June 1976. CODEN HPJOAX. ISSN 0018-1153.

Esch:1998:TDC

[EM98] Gerald L. Esch, Jr. and Robert B. Manley. Theory and design of CMOS HSTL I/O pads. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(3):46–??, August 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98aug/tc-08-98.htm.

Emmermann:1972:VPH

[Emm72] G. G. Emmermann. Voltage precision and high current capability-both in one power supply. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(3):16–19, November 1972. CODEN HPJOAX. ISSN 0018-1153.

Evett:1983:SPC

[EMT83] Eric A. Evett, Paul J. McClellan, and Joseph P. Tanzini. Scientific pocket calculator extends range of built-in functions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(5):25– 35, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Epstein:1984:VIS

[Eps84] James S. Epstein. Versatile instrument simplifies dynamic signal analysis at low frequencies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(12):4–11, December 1984. CODEN HPJOAX. ISSN 0018-1153.

Erdmann:1979:SDA

[Erd79] R. E. Erdmann, Jr. A serial data analyzer for locating faults in decentralized digital systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(10):23– 28, October 1979. CODEN HPJOAX. ISSN 0018-1153.

Erickson:1992:MIV

[Eri92] Calvin L. Erickson. Mass interconnect for VXIbus systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):52–58, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Ernst:1995:BSO

[Ern95] Peter H. Ernst. Boolean set operations with solid models. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):74–79, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_74.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Ewy:1969:CAT

[ES69] M. D. Ewy and S. C.

Shank. Choosing an automatic test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(12):7–10, August 1969. CODEN HPJOAX. ISSN 0018-1153.

Edgerley:1977:DSS

[ES77] W. M. Edgerley, Jr. and D. C. Snyder. Display and storage systems for a digital signal analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (2):14–17, October 1977. CODEN HPJOAX. ISSN 0018-1153.

Eckert:1985:OSR

[ES85] Achim Eckert and Wolfgang Schmid. Optical stimulus and receivers for parametric testing in fiber optics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(1):4–7, January 1985. CODEN HPJOAX. ISSN 0018-1153.

Elmore:1987:QMM

[ES87] Glenn E. Elmore and Louis J. Salz. Quality microwave measurement of packaged active devices. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (2):39–48, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Elliott:1989:SID

[ES89] Ian A. Elliott and George M. Sachs. Sharing input devices in the Starbase/X11 Merge system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(6):38–41, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Evett:1983:PCC

[Eve83] Eric A. Evett. A pocket calculator for computer science professionals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(5):36– 39, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Eads:1978:HID

[EW78] W. D. Eads and J. M.

Walden. A highly integrated desktop computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(8):2–11, April 1978. CODEN HPJOAX. ISSN 0018-1153.

Ewy:1969:BAA

[Ewy69]  M. D. Ewy.  Building

at automatic test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20 (12):11–20, August 1969. CODEN HPJOAX. ISSN 0018-1153.

Ellendman:1996:AAA

[EY96] Anna Ellendman and William R.

Yoder. Accessing and administering applications in CDE. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):15– 23, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Estes:1981:NTP

[EZ81] M. F. Estes and D. Zimmer, Jr. New temperature probe locates circuit hot spots. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(3):30–32, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Fanton:1972:CHS

[Fan72] J. L. Fanton. A computeraided hospital system for cardiac catheterization procedures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(5):2–7, January 1972. CODEN HPJOAX. ISSN 0018-1153.

Farnbach:1974:LSA

[Far74] W. A. Farnbach. The Logic State Analyzer — displaying complex digital processes in understandable form. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(5):2–9, January 1974. CODEN HPJOAX. ISSN 0018-1153.

Fenoglio:1979:HDX

[FCC79] J. A. Fenoglio, B. W. C. Chin, and T. R. Cobb. A high-quality digital X-Y plotter designed for reliability, flexibility and low cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(2):2–7, February 1979. CODEN HPJOAX. ISSN 0018-1153.

Fong:1997:SII

[FCK97] J. C. Fong, Hoi-Kuen Chan, and M. D. Kruckenberg. Solving IC interconnect routing for an advanced PA-RISC processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):40–45, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/interconnect_
routing.pdf;   http://
www.hp.com/hpj/97aug/
au97a5.htm;   http://
www.hp.com/hpj/97aug/
tc-08-97.htm.

Ferguson:1978:TSG

[FD78] K. M. Ferguson and L. R.

Dickstein. Time synthesizer generates precise pulse widths and time delays for critical timing applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(12):12–19, August 1978. CODEN HPJOAX. ISSN 0018-1153.

Finch:1977:MOP

[FEG77] C. M. Finch, M. F. Estes, and L. A. Gammill. Miniature oscilloscope probes for measurements in crowded circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(8):9–12, April 1977. CODEN HPJOAX. ISSN 0018-1153.

Felsenstein:1974:PES

[Fel74] R. E. Felsenstein. The 5345A processor: an example of state machine design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(10):9–11, June 1974. CODEN HPJOAX. ISSN 0018-1153.

Felsenstein:1980:FSD

[Fel80] R. E. Felsenstein. A flexible software development technique. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4):25–26, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Felps:1986:MPS

[Fel86] Jimmie D. Felps. A modular power supply. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(4):37– 39, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Fernandez:1994:GUI

[Fer94] Charles V. Fernandez. A graphical user interface for a multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 20–22, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94apr/
20to22.pdf;   http://
www.hp.com/hpj/toc-4-
94.html.

Fried:1989:MSG

[FFS89] Steve R. Fried, Keith L.

Fries, and John M. Sims. Microwave signal generator output system design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (5):42–50, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Forbes:1973:EFM

[FG73] B. E. Forbes and M. D.

Green. An economical fullscale multipurpose computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(5):2–8, January 1973. CODEN HPJOAX. ISSN 0018-1153.

Finch:1975:AUO

[FG75] W. R. Finch and R. B. Grady. ATLAS: a unitunder-test oriented language for automatic test systems. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(1):2–13, September 1975. CODEN HPJOAX. ISSN 0018-1153.

Flade:1987:OPM

[FG87] Bernhard Flade and Michael Goder. Optical power meter firmware development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (2):12–15, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Fatehi:1989:DBR

[FGH^{+}89] Feyzi Fatehi, Cynthia Givens, Le T. Hong, Michael R. Light, Ching-Chao Liu, and Michael J. Wright. A data base for real-time applications and environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3):6–17, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Fowles:1974:EDD

[FH74] R. G. Fowles and J. J. Heinzl. Envelope delay distortion measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25 (12):14–16, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Frost:1980:CEC

[FH80] J. B. Frost and W. L. Hale.

Color enhances computer graphics system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(12):3–5, December 1980. CODEN HPJOAX. ISSN 0018-1153.

Frink:1992:HDL

[FHDS92] Craig R. Frink, Robert J.

Hammond, John A. Dykstal, and Don C. Soltis, Jr. High-performance designs for the low-cost PA-RISC desktop. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(4):55–63, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Faulkner:1985:SGF

[FHM^{+}85] Thomas R. Faulkner, Earl C.

Herleikson, Ronald J. Mayer, Brian M. Miller, and Mark A. Niemann. Signal generator frequency synthesizer design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(12): 24–31, December 1985. CODEN HPJOAX. ISSN 0018-1153.

Fisher:1997:FTS

[FHMP97] R. L. Fisher, S. R. Herbener, J. R. Morgan, and J. R. Pessetto. Fast turnaround of a structured custom IC design using advanced design tools and methodology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):104– 107, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a17.htm.

Fiasconaro:1983:ISS

[Fia83] J. G. Fiasconaro. Instruction set for a single-chip 32-bit processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(8):9–10, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Fiduccia:1996:SRB

[Fid96] Nicholas S. Fiduccia. Sidebar: Runway bus electrical design considerations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(1): ??, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/feb96/feba2sb.
html.

Fiedler:1989:OUT

[Fie89] Steven P. Fiedler. Objectoriented unit testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 69–74, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Finn:1972:AMS

[Fin72] C. L. Finn. All-semiconductor memory system includes read-only and read/write chips. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(4):22–24, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Fisher:1972:MRC

[Fis72] J. Fisher. Microwave radio communications and performance measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24 (1):10–11, September 1972. CODEN HPJOAX. ISSN 0018-1153.

Fischer:1988:HIE

[Fis88] William A. Fischer, Jr.

Host independent emulator software architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (6):52–56, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Fitzgerald:1982:ISI

[Fit82] Kathleen A. Fitzgerald. Interactive software for intelligent printers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(6):10– 16, June 1982. CODEN HPJOAX.

Fischer:1989:CSU

[FJ89] William A. Fischer, Jr. and James W. Jost. Comparing structured and unstructured methodologies in firmware development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 80–85, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Forbes:1979:CCD

[FK79] L. Forbes and U. Kaempf. Capacitance and conductance deep-level transient spectroscopy using HP-IB instruments and a desktop computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4):29–32, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Fritz:1984:IFM

[FK84] G. D. Fritz and M. L.

Kolesar. I/O features of model 520 BASIC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(5):21– 24, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Faulkner:1986:NST

[FKL86] Kevin J. Faulkner, Charles W.

Knouse, and Brian K. Lynn. Network services and transport for the HP 3000 computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):11– 18, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Fletcher:1993:HAI

[FKO^{+}93] Robert M. Fletcher, Chihping Kuo, Timothy D. Osentowski, Jiann Gwo Yu, and Virginia M. Robbins. High-efficiency aluminium indium gallium phosphide light-emitting diodes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4): 6–14, August 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Falke:1973:PGT

[FL73] R. Falke and H. Link. A pulse generator for today’s digital circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(2):17–23, October 1973. CODEN HPJOAX. ISSN 0018-1153.

Fleming:1980:HDL

[FL80] J. H. Fleming and R. N. Low. High density and low cost with printed circuit hybrid technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(3):25–26, March 1980. CODEN HPJOAX. ISSN 0018-1153.

Fleming:1983:HIS

[Fle83] J. H. Fleming. HP-IL interconnect system (Hewlett– Packard Interface Loop). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(1): 8–??, January 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Fay:1984:HBT

[FM84] Thomas R. Fay and John E. McDermid. The HP 3065 board test family: a system overview. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(10):4–9, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Fotland:1984:ARC

[FMN84] David A. Fotland, Lee S.

Moncton, and Leslie E. Neft. An adaptable 1-MIPS real-time computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (2):7–12, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Freitag:1995:EBA

[FO95] Stefan Freitag and Karsten Opitz. Enhancements in blending algorithms. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5): 24–34, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_24.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Folsom:1972:MTR

[Fol72] J. B. Folsom. Measuring true RMS AC voltages to 100 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(7):14–19, March 1972. CODEN HPJOAX. ISSN 0018-1153.

Foote:1980:CSA

[Foo80] J. D. Foote. A complete self-contained audio measurement system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(8):3–7, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Forbes:1977:STP

[For77] B. E. Forbes. Silicon-on-sapphire technology produces high-speed singlechip processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(8):2–8, April 1977. CODEN HPJOAX. ISSN 0018-1153.

Foster:1973:SMC

[Fos73] W. E. Foster. Software for a multilingual computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(5):15–19, January 1973. CODEN HPJOAX. ISSN 0018-1153.

Foster:1974:ECV

[Fos74] T. E. Foster. An easily calibrated, versatile platinum resistance thermometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(8):13–17, April 1974. CODEN HPJOAX. ISSN 0018-1153.

Foster:1998:THP

[Fos98] Harry D. Foster. Techniques for higher-performance Boolean equivalence verification. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):30–??, August 1998. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
98aug/tc-08-98.htm.

Fraley:1990:HOB

[FP90] Andrew S. Fraley and Tamra I. Perez. HP OpenView Bridge Manager: Network management for HP LAN bridges. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (2):66–70, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Fox:1972:HIA

[FPS72] K. A. Fox, M. P. Pasturel, and P. S. Showman. A human interface for automatic measurement systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(8):10–17, April 1972. CODEN HPJOAX. ISSN 0018-1153.

Frohring:1988:WRD

[FPUB88] Brian J. Frohring, Bruce E.

Peetz, Mark A. Unkrich, and Steven C. Bird. Waveform recorder design for dynamic performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):39–48 (or 39–47??), February 1988. CODEN HPJOAX. ISSN 0018-1153.

Fischer:1974:IAC

[FR74] W. A. Fischer and W. B. Risley. Improved accuracy and convenience in oscilloscope timing and voltage measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(4):2–9, December 1974. CODEN HPJOAX. ISSN 0018-1153.

Fleischer-Reumann:1987:NFP

[FR87a] Michael Fleischer-Reumann. A new family of precise, reliable, and versatile fiber optic measurement instruments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(2):4–5, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Fleischer-Reumann:1987:PFO

[FR87b] Michael Fleischer-Reumann. A programmable fiber optic switch. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(2):36–??, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Fleischer-Reumann:1987:SLS

[FR87c] Michael Fleischer-Reumann.

Stable LED sources for a wide range of applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(2): 6–8, February 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Fahlbusch:1995:HPS

[FR95] Klaus-Peter Fahlbusch and Thomas D. Roser. HP PE/SolidDesigner: dynamic modeling for threedimensional computer-aided design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):6–13, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_6t.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Frankenberg:1971:LMS

[Fra71] R. J. Frankenberg. A lot of memory in a small space. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(2):10–12, October 1971. CODEN HPJOAX. ISSN 0018-1153.

Frankenberg:1974:ASM

[Fra74] R. J. Frankenberg. All semiconductor memory selected for new minicomputer series. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(2):15–19, 20, October 1974. CODEN HPJOAX. ISSN 0018-1153.

Fajardo:1985:UOS

[FRAC85] Ray M. Fajardo, Andrew L. Rood, James R. Andreas, and Robert C. Cline. A UNIX operating system adapted for a technical personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(10):22– 28, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Fleischer-Reumann:1985:ORN

[FRMK85] Michael Fleischer-Reumann, Emmerich Mueller, and Gerd Koffmane. An optical receiver for 550 to 950 nm. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(1):27–29, January 1985. CODEN HPJOAX. ISSN 0018-1153.

Fleischer:1983:NFP

[FRN83] Michael Fleischer, Helmut Rossner, and Uwe Newmann. A new family of pulse and pulse/function generators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(6):27– 32, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Frohwerk:1977:SAN

[Fro77] R. A. Frohwerk. Signature analysis: a new digital field service method. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(9):2–8, May 1977. CODEN HPJOAX. ISSN 0018-1153.

Frolik:1986:CPR

[Fro86] William R. Frolik. Creating plug-in roms for the portable plus computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7): 28–30, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Fromme:1990:HEB

[Fro90] Brian D. Fromme. HP Encapsulator: bridging the generation gap (tool integrator). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(3):59–68, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Fleischer-Reumann:1988:HOT

[FRS88] Michael Fleischer-Reumann and Franz Sischka. A high-speed optical timedomain reflectometer with improved dynamic range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (6):6–13, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Federman:1981:IMP

[FS81a] N. C. Federman and R. M. Steiner. An interactive material planning and control system for manufacturing companies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(4): 3–12, April 1981. CODEN HPJOAX. ISSN 0018-1153.

Flower:1981:MDM

[FS81b] T. L. Flower and M. Son. Motor drive mechanics and control electronics for a high-performance plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(11): 12–15, November 1981. CODEN HPJOAX. ISSN 0018-1153.

Fuget:1986:TAS

[FS86] Craig D. Fuget and Barbara J. Scott. Tools for automating software test package execution. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 24–28, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Franz:1994:EVI

[FS94] Louis A. Franz and Jonathan C.

Shih. Estimating the value of inspections and early testing for software projects. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):60– 67, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_60.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Fotland:1987:HDF

[FSB^{+}87] David A. Fotland, John F.

Shelton, William R. Bryg, Ross V. La Fetra, Simin I. Boschma, Allan S. Yeh, and Edward M. Jacobs. Hardware design of the first HP precision architecture computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(3):4–17, March 1987. CODEN HPJOAX. ISSN 0018-1153.

Fu:1982:HHN

[FSDC82] Horng-Sen Fu, Roger To-Hoi Szeto, Anders T. Dejenfelt, and Devereaux C. Chen. HQMOS: a highperformance NMOS technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(10):21–27, October 1982. CODEN HPJOAX. ISSN 0018-1153.

Fischer:1992:GTS

[FST92] Michael C. Fischer, Michael J. Schoessow, and Peter Tong. GaAs technology in sensor and baseband design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 90–94 (or 90–93??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Fugitt:1986:NSH

[FT86] J. Christopher Fugitt and Dean R. Thompson. Networking services for HP 9000 computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):28– 32, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Fulghum:1983:HMS

[Ful83a] Kipper K. Fulghum. The HP 64000 measurement system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3):8–9, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Fuller:1983:EMI

[Ful83b]  Ian J. Fuller.  Elec-

tronic mail for the interactive office. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(2):20–29, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Fuller:1989:OHN

[Ful89] Ian J. Fuller. An overview of the HP NewWave environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):6–8, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Fuller:1996:DDM

[Ful96] John G. Fuller. Developing a design for manufacturability focus. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):79– 84, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Galen:1984:DDE

[Gal84] Peter M. Galen. Disk drive error detection and correction using VLSI. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(1):12–13, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Graf:1993:HTB

[GAL^{+}93] Terrence P. Graf, Renato G. Assini, John M. Lewis, Edward J. Sharpe, James J. Turner, and M. C. Ward. HP Task Broker: A tool for distributing computational tasks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):15–22, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Gonzalez:1992:RUI

[GAP92] Jordi Gonzalez, Jaume Ayats Ardite, and Carles Castellsague Pique. Reengineering of a user interface for a drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):49–55, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Gardner:1983:ACL

[Gar83] D. A. Gardner. Automatic compensation (for laser beam measurement system). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(4):12–??, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Garnett:1992:HMD

[Gar92] Grosvenor H. Garnett. A high-resolution, multichannel digital-to-analog converter for digital oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):48–53, 55–56, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Gast:1988:RAH

[Gas88] P. David Gast. Robotic assembly of HP DeskJet printed circuit boards in a just-in-time environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):87–90, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Gaullier:1981:TMT

[Gau81] F. Gaullier. A terminal management tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(1):30–31, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Gravitz:1987:CBT

[GB87] Michael E. Gravitz and G. Booth. A combinational board test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):53– 64, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Gyma:1981:LAP

[GBHS81] D. W. Gyma, P. W. Bailey, J. W. Hyde, and D. R. Schwartz. Laboratoryperformance autoranging power supplies using power MOSFET technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(8): 11–17, August 1981. CO-DEN HPJOAX. ISSN 0018-1153.

Goto:1994:CPD

[GBO94] Masaharu Goto, James O. Barnes, and Ronnie E. Owens. CMOS programmable delay Vernier. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (5):51–58, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_51.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Gildea:1987:VMS

[GC87] David R. Gildea and Donald R. Chambers. Vector modulation in a signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):25– 29, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Garg:1990:DDN

[GC90] Atul R. Garg and Lisa M.

Cole. Developing a distributed network management application using HP OpenView Windows. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2): 85–91, April 1, 1990. CO-DEN HPJOAX. ISSN 0018-1153.

Graves:1991:IPD

[GCDW91] Spencer B. Graves, William P.

Carmichael, Douglas Daetz, and Edith Wilson. Improving the product development process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(3):71– 76, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Graham:1986:LAN

[Gd86] Tonia G. Graham and Charles J. de Sostoa. Local area network for HP computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):18– 22, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Grace:1994:IPE

[GDKK94] James W. Grace, David M. DiPietro, Akito Kishida, and Kenji Kinsho. Integrated pin electronics for automatic test equipment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (5):42–50, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94oct/
                                        oct94_42.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-94.html.

Gardner:1973:GRC

[GDV73] R. Gardner, A. Dumont, and S. Venzke. A greater range of capabilities for the compact, plug-on digital multimeter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(7):20–24, March 1973. CODEN HPJOAX. ISSN 0018-1153.

Genther:1986:HIH

[GE86] Scott A. Genther and Eddie A. Evel. Hardware implementation of a highperformance trigger system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4):26–33, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Geber:1978:MSI

[Geb78] C. R. Geber. Microcoded scientific instruction set enhances speed and accuracy of F-series computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (14):18–22, October 1978. CODEN HPJOAX. ISSN 0018-1153.

Gembarowski:1985:SMP

[Gem85] Mark L. Gembarowski.

Software methodology preserves consistency and creativity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3):47–48, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Gennetten:1985:CTD

[Gen85] K. Douglas Gennetten.

Cartridge tape data integrity ensured at five levels. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3):39–43, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Gerety:1990:NGS

[Ger90] Colin Gerety. A new generation of software development tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(3):48–58, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Gerster:1998:UTC

[Ger98] Andreas Gerster. Units, traceability, and calibration of optical instruments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3): 17–??, August 1998. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98aug/
tc-08-98.htm.

Gatzke:1983:ESP

[GFK83] Ronald D. Gatzke, James T. Fearnside, and Sydney M. Karp. Electronic scanner for a phased-array ultrasound transducer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(12):13– 15, 17–20, December 1983.

CODEN HPJOAX. ISSN 0018-1153.

Grobstein:1977:BEM

[GG77] S. R. Grobstein and R. D.

Gatzke. A battery-powered ECG monitor for emergency and operating room environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(1):26–31, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Graham:1973:TCF

[GH73] T. R. Graham and J. M.

Hood. Telecommunication cable fault location from the test desk. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(4):9–14, December 1973. CODEN HPJOAX. ISSN 0018-1153.

Gittler:1995:DSS

[GH95] Frederic Gittler and Anne C.

Hopkins. The DCE security service. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):41– 48, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_41.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Goodnow:1993:UTQ

[GHK^{+}93] John W. Goodnow, Cindie A. Hammond, William A. Koppes, John J. Krieger, D. Kris Rovell-Rixx, and S. J. Warner. The use of total quality control techniques to improve the software localization process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 64–70, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Grisell:1968:DTR

[GHUW68] T. L. Grisell, I. H. Hawley, Jr., B. D. Unter, and P. G. Winninghoff. Design of a third-generation RF spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(12):8–14, August 1968. CODEN HPJOAX. ISSN 0018-1153.

Gibson:1986:GAL

[Gib86] Scott R. Gibson. Gallium arsenide lowers cost and improves performance of microwave counters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(2): 4–10, February 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Gilbert:1981:OID

[Gil81] R. A. Gilbert. Operator interface design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(7):29– 30, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Gills:1989:RAQ

[Gil89] David Gills. Reliability assessment of a quarterinch cartridge tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (4):74–78, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Gilray:1996:ACI

[Gil96]  J. J. Gilray.  Apply-

ing the code inspection process to hardware descriptions. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):68–72, February 1996. CODEN HPJOAX. ISSN 0018-1153.

Gordon:1974:MCP

[GJ74] P. Gordon and J. R. Jacobs. Microprogrammable central processor adapts easily to special user needs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(2):

7–9, 11–14, October 1974. CODEN HPJOAX. ISSN 0018-1153.

Greenbaum:1998:CHV

[GJ98] S. Greenbaum and S. Jefferson. A compiler for HP VEE. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):98– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Hsu:1981:DAM

[gJH81] Tsen gong Jim Hsu. Digital adaptive matched filter for fiducial mark registration (in electron beam lithography system). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(5):34– 36, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Gleason:1992:VCL

[GJM^{+}92] Craig A. Gleason, Leith Johnson, Steven T. Mangelsdorf, Thomas O. Meyer, and Mark A. Forsyth. VLSI circuits for low-end and midrange PA-RISC computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):12–22, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Gross:1983:VIP

[GJW83] Fred J. Gross, William S.

Jaffe, and Donald R. Weiss. VLSI I/O processor for a 32-bit computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (8):11–14, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Gutierrez:1988:DNF

[GL88] David O. Gutierrez and Chyuan-Shiun Lin. The design of network functions for discless clusters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):20–26, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Gittler:1995:OOA

[GLM95] Mihaela C. Gittler, Michael Z.

Luo, and Luis M. Maldonado. An object-oriented application framework for DCE-Based systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 55–60, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_55.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Gee:1997:MOE

[GLP^{+}97] Stephen M. Gee, Herbert Lage, Chris Park, Kevin A. Williams, Richard V. Penty, Ian H. White, and Joseph A. Barnard. Modeled optimization and experimental verification of a low-dispersion source for long-haul 2.488-gbit/s systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        97dec/de97a9.htm.

Gordon:1985:CVL

[GLW85] Philip Gordon, Phillip R. Luque, and Donald K. Wadley. Cost-effective, versatile line printer electronics and firmware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(6):18– 23, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Gottlieb:1992:HSF

[GMMM92] R. A. Gottlieb, D. J. Magenheimer, S. A. Meloy, and A. C. Meyer. HP 9000 series 700 FORTRAN optimizing preprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3):

24–32, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Garrettson:1982:XL

[GN82] Garrett A. Garrettson and Armand P. Neukermans. X-Ray lithography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(8): 14–19 (or 14–18??), August 1982. CODEN HPJOAX. ISSN 0018-1153.

Gronlund:1993:HED

[GNR93] Robert D. Gronlund, Richard A.

Nygaard, Jr., and John T. Rasper. The HP 64700 embedded debug environment: A new paradigm for embedded system integration and debugging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 90–106, April 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Girdner:1972:RWF

[GO72] W. I. Girdner and W. H.

Overton. Reading and writing on the fast disc. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(9): 12–14, May 1972. CODEN HPJOAX. ISSN 0018-1153.

Goeke:1989:DIA

[Goe89] Wayne C. Goeke. An 8.5digital integration analogto-digital converter with 16-bit, 100000-sample-per-second performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 8–15, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Goldberg:1980:SSD

[Gol80] J. M. Goldberg. Selftest and serviceability for dependable central patient monitoring. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):19– 23, November 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Gookin:1972:CVN

[Goo72] A. Gookin. Compactness and versatility in a new plug-together digital multimeter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(12):2–6, August 1972. CODEN HPJOAX. ISSN 0018-1153.

Gookin:1977:FHV

[Goo77] A. Gookin. A fast-reading, high-resolution voltmeter that calibrates itself automatically. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(6):2–10, February 1977. CODEN HPJOAX. ISSN 0018-1153.

Goodnow:1991:MSP

[Goo91] John W. Goodnow. A mechanism to support parallel development via RCS. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3): 84–89, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Gordon:1969:ILC

[Gor69] G. B. Gordon. IC logic checkout simplified. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(10): 14–16, June 1969. CODEN HPJOAX. ISSN 0018-1153.

Gort:1980:FIM

[Gor80] A. F. Gort. A fully integrated, microprocessorcontrolled total station. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (9):3–11, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Gordon:1982:SHE

[Gor82] Philip Gordon. Specialized high-speed electronics for document preparation flexibility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(6):30–35, June 1982. CODEN HPJOAX.

Goring:1991:ATE

[Gor91] Dieter Goring. The automated test environment for a medical patient monitoring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):49–52 (or 49–51??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Georg:1984:GOS

[GOS84] Dennis D. Georg, Benjamin D. Osecky, and Stephan D. Scheid. A general-purpose operating system kernel for a 32-bit computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(3):28–29, 31– 34, March 1984. CODEN HPJOAX. ISSN 0018-1153.

URL https://www.hpl.
hp.com/hpjournal/pdfs/
IssuePDFs/1984-03.pdf.

Grodd:1984:REN

[GP84] Laurence W. Grodd and Charles M. Patton. ROM extends numerical function set of handheld computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7): 25–36, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Gordon:1969:ICC

[GR69] G. B. Gordon and G. A.

Reeser. Introducing the computing counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(9): 2–8, May 1969. CODEN HPJOAX. ISSN 0018-1153.

Goldsack:1991:SEM

[GR91] P. C. Goldsack and T. W.

Rush. Specifying an electronic mail system with HP-SL. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(5):32–39, December 1991. CODEN HPJOAX. ISSN 0018-1153.

Graham:1981:DSU

[Gra81] G. C. Graham. Display station’s user interface is designed for increased productivity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(3):8–12, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Grady:1989:DSF

[Gra89] Robert B. Grady. Dissecting software failures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 57–63, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Grady:1996:SFA

[Gra96] Robert B. Grady. Software failure analysis for highreturn process improvement decisions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):15– 24, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart2.
htm.

Gregory:1985:WRR

[Gre85] Wayne T. Gregory. Write and read recovery systems for a half-inch tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3): 16–18, March 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Gregory:1990:HMS

[Gre90] Douglas R. Gregory. The HP MAP 3.0 software integration lifecycle. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(4):54–60, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Grimm:1969:AT

[Gri69] R. A. Grimm. Automated testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(12):2–6, August 1969. CODEN HPJOAX. ISSN 0018-1153.

Groves:1979:RDF

[Gro79] W. A. Groves. Rapid digital fault isolation with FASTRACE. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(3):8–13, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Grote:1983:MLT

[Gro83] Richard H. Grote. Manufacturing the laser tube. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(4): 17–18, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Grossbach:1991:MES

[Gro91] Wolfgang Grossbach. Measuring the ECG signal with a mixed analogdigital application-specific IC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):21–24, October 1991.

CODEN HPJOAX. ISSN 0018-1153.

Grove:1992:MFR

[Gro92] Martin B. Grove. Measuring frequency response and effective bits using digital signal processing techniques. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):29–35 (or 29– 33??), February 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Gordon:1993:OOR

[GRS93] Gary B. Gordon, Joseph C. Roark, and Arthur Schleifer. ORCA: optimized robot for chemical analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):6–8, 10– 19, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Gennetten:1993:DSC

[GS93] K. Douglas Gennetten and Michael J. Steinle. Designing a scanner with color vision. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):52–58, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Grinham:1995:MA

[GS95] John R. Grinham and Michael P. Spratt. Multimedia applications and 100VG-AnyLAN. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(4):33– 38, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_33.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Garliepp:1990:HDD

[GSFF90] Kent L. Garliepp, Irene Skupniewicz, John U. Frolich, and Kathleen A. Fulton. HP DIS: A development tool for factory-floor device interfaces. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(5):62–72, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Gassman:1987:VHH

[GSG^{+}87] Gerald R. Gassman, Michael W.

Schrempp, Ayee Goundan, Richard Chin, Robert D. Odineal, and Marlin Jones. VLSI-Based high-performance HP precision architecture computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(9):38–48, September 1987. CODEN HPJOAX. ISSN 0018-1153.

Goeke:1989:CDM

[GSVS89] Wayne C. Goeke, Ronald L. Swerlein, Stephen B. Venzke, and Scott D. Stever. Calibration of an 8.5-digit multimeter from only two external standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2): 22–30, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Gudnitz:1971:GTS

[GT71] L. Gudnitz and H. Tsuda. General-purpose test system gets digital capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(11):2–9, July 1971. CO-DEN HPJOAX. ISSN 0018-1153.

Gordon:1995:NHC

[GTM95] Gary B. Gordon, Richard P. Tella, and Henrique A. S. Martins. A new highsensitivity capillary electrophoresis detector cell and advanced manufacturing paradigm. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 62–70, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95jun/
                                        jun95_62.pdf;   http:

//www.hp.com/hpj/toc-
06-95.html.

Guest:1974:ALD

[Gue74a] D. H. Guest. An active loop-holding device. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (4):11–??, December 1974. CODEN HPJOAX. ISSN 0018-1153.

Guest:1974:SDC

[Gue74b] D. H. Guest. Simplified data-transmission channel measurements (amplitude/ delay distortion analyser). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (3):15–24, November 1974. CODEN HPJOAX. ISSN 0018-1153.

Guest:1979:HED

[Gue79] D. H. Guest. An HP-IB extender for distributed instrument systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(8):26–32, August 1979. CODEN HPJOAX. ISSN 0018-1153.

Gunji:1994:RDS

[Gun94] Keita Gunji. Real-time digital signal processing in a mixed-signal LSI test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):59–63, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94oct/oct94_59.pdf;
http://www.hp.com/hpj/
toc-10-94.html.

Gupta:1995:HEM

[Gup95] Pankaj Gupta. HP Encina/ 9000: Middleware for constructing transaction processing applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(6): 61–74, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_61.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Godlew:1992:EAI

[GUW92] Scott Godlew, Rod Unverrich, and Stephen Witt. Embedding artificial intelligence in a LAN test instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):11–21, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Geiger:1984:NSD

[GV84] Wolfgang Geiger and Heinrich Völlmer. A new solvent delivery system (liquid chromatography). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(4):13– 21, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Gottschalk:1986:RSD

[GV86] George R. Gottschalk and Roy M. Vandoorn. Rulebased system to diagnose malfunctioning computer peripherals. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):48– 53, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Grote:1977:ADS

[GW77] R. H. Grote and H. Webber McKinney. Advanced digital signal analyzer probes low-frequency signals with ease and precision. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(2):2–9, October 1977. CODEN HPJOAX. ISSN 0018-1153.

Garfinkel:1994:HST

[GWY94] Daniel Garfinkel, Bruce C.

Welti, and Thomas W. Yip. HP SharedX: A tool for real-time collaboration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 23–36, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        23to36.pdf;   http://
                                        www.hp.com/hpj/toc-4-
                                        94.html.

Gee:1988:SCA

[GY88] Albert Gee and Ronald W. Young. Signal conditioning and analog-to-digital conversion for a 4-MHz, 12-bit waveform recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):15–22, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Globas:1974:PGT

[GZC74] G. Globas, J. Zellmer, and E. Cornish. A 250-MHz pulse generator with transition times variable to less than 1 ns. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(1):2–7, September 1974. CODEN HPJOAX. ISSN 0018-1153.

Hall:1992:FBT

[HA92] Stuart O. Hall and Jay A. Alexander. A fast, built-in test system for oscilloscope manufacturing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):21–28, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Haag:1978:LSA

[Haa78] G. A. Haag. A logic state analyzer for evaluating complex state flow. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (6):2–10, February 1978. CODEN HPJOAX. ISSN 0018-1153.

Hajela:1996:HOA

[Haj96] Sujai Hajela. HP OEMF:

alarm management in telecommunications networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):22–30, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/oct96/oc96a3.htm.

Hall:1975:APS

[Hal75] J. A. Hall. Applications of a phase-modulated signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(11):21– 24, July 1975. CODEN HPJOAX. ISSN 0018-1153.

Hallissy:1979:APC

[Hal79] R. M. Hallissy. Assembly programming capability in a desktop computer (hewlett–packard model 9835A/B computer). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(5): 18–20, May 1979. CODEN HPJOAX. ISSN 0018-1153.

Hall:1982:LPS

[Hal82] James A. Hall. Laser printing system provides flexible, high-quality, costeffective computer output. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(6): 3–5, 7–8, June 1982. CO-DEN HPJOAX.

Halpenny:1985:FPS

[Hal85] Thomas J. Halpenny. Firmware provides simple and powerful plotter operation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4): 34–36, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Hall:1993:SDF

[Hal93] Michael L. Hall. A standard data format for instrument data interchange. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):85–89, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Hamilton:1979:HSC

[Ham79] A. P. Hamilton. A humanengineered small-business computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4):3–??, 5, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Hamilton:1981:SCH

[Ham81] B. Hamilton. Software control for the HP electron beam lithography system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(5): 21–23, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Hamling:1993:PFW

[Ham93] Daniel T. Hamling. A probe fixture for wafer testing high-performance data acquisition integrated circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):73–75, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Hanson:1970:HAA

[Han70] F. L. Hanson. High accuracy a.c. calibration to 1100 volts. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(8):14– 17, April 1970. CODEN HPJOAX. ISSN 0018-1153.

Hanson:1972:NNI

[Han72] R. C. Hanson. Narrowband noise immunity in a broadband gain-phase meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(9): 17–20, May 1972. CODEN HPJOAX. ISSN 0018-1153.

Hanson:1979:RFL

[Han79] D. C. Hanson. A ready-to-use fiber-optic link for data communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(12):5–6, December 1979. CODEN HPJOAX. ISSN 0018-1153.

Hanna:1980:FPM

[Han80a]  K. L. Hanna.  Firm-

ware for a patient monitoring station. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):23– 28, November 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Hanna:1980:MSM

[Han80b] W. A. Hanna. Mass storage management — a unified approach. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(6):20–24, June 1980. CODEN HPJOAX. ISSN 0018-1153.

Hansen:1992:NOP

[Han92] R. C. Hansen. New optimizations for PA-RISC compilers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3):15–23, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Han:1994:AVO

[Han94]  Joy Xiao Han.  Ap-

proaches to verifying operational test release vectors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (6):55–59, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_55.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Harrison:1976:EVS

[Har76] V. Harrison. An easierto-use variable-persistence/ storage oscilloscope with brighter, sharper traces. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (1):2–8, September 1976. CODEN HPJOAX. ISSN 0018-1153.

Harper:1982:MBC

[Har82] T. V. Harper. Module brings CP/M to HP’s latest Series 80 computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (12):8–11, December 1982. CODEN HPJOAX. ISSN 0018-1153.

Harper:1983:CIC

[Har83] Steven L. Harper. A CMOS integrated circuit for the HP-IL interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(1):16–22, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Harwood:1984:SDA

[Har84] Vance R. Harwood. Safeguarding devices against stress caused by in-circuit testing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):20– 22, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Harrington:1986:UTV

[Har86] Daniel B. Harrington. Usability testing: a valuable tool for PC design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(1):36–40, January 1986. CODEN HPJOAX. ISSN 0018-1153.

Harmon:1987:ASE

[Har87]  Paul Harmon.  Alpha

site evaluation of me series 5/10. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(5):30– 33, May 1987. CODEN HPJOAX. ISSN 0018-1153.

Hassun:1981:HFS

[Has81] R. Hassun. A high-purity, fast-switching synthesized signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):3–7, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Haselby:1992:PAT

[Has92] Robert D. Haselby. Pen alignment in a two-pen, large-format, inkjet drafting plotter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):24–27, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Hunt:1994:HQI

[HASP94] Catherine B. Hunt, Ronald A.

Askeland, Leonard Slevin, and Keshava A. Prasad. High-quality inkjet color graphics performance on plain paper. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):18–27, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/18to27.pdf.

Hawley:1971:LFS

[Haw71] I. H. Hawley, Jr. A low frequency spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (1):12–15, September 1971. CODEN HPJOAX. ISSN 0018-1153.

Hay:1974:VVS

[Hay74] R. R. Hay. Versatile VHF signal generator stresses low cost and portability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(7): 18–23, March 1974. CO-DEN HPJOAX. ISSN 0018-1153.

Herlinger:1972:FTD

[HB72] J. E. Herlinger and J. R. Barnes. A faster, tougher disc drive for small computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(9):2–5, May 1972. CODEN HPJOAX. ISSN 0018-1153.

Hall:1988:XWS

[HB88] Frank E. Hall and James B.

Byers. X: A window system standard for distributed computing environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(5):46–50, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Heikes:1988:DDW

[HBB88] Craig A. Heikes, James O.

Barnes, and Dale R. Beucler. Development of a digital waveform synthesis integrated circuit. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):62–65, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Hamilton:1984:ETA

[HBHC84] Gail E. Hamilton, Andrew J. Blasciak, Joseph A. Hawk, and Brett K. Carver. An electronic tool for analyzing software performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(6):26–32, June 1984. CODEN HPJOAX. ISSN 0018-1153.

Hubner:1983:HSD

[HBNB83] Ulrich Hübner, Werner Berkel, Heinz Nuessle, and Josef Becker. A high-speed data generator for digital testing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(7):7–14, July 1983. CODEN HPJOAX. ISSN 0018-1153.

Horner:1975:MMM

[HC75] J. F. Horner and B. S. Corya. Modularity means maximum effectiveness in medium-cost universal counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(11): 2–8, July 1975. CODEN HPJOAX. ISSN 0018-1153.

Huth:1988:CMV

[HCF^{+}88] Mark C. Huth, Robert A. Conder, Gregg P. Ferry, Brian L. Helterline, Robert F. Aman, and Timothy S. Hubley. CIM and machine vision in the production of thermal inkjet printheads. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):91–98, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Hernday:1991:DLC

[HCHR91] Paul R. Hernday, Geraldine A. Conrad, Michael G. Hart, and Rollin F. Rawson. Design of a 20-GHz Lightwave component analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1):13–22, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Harding:1982:DOM

[HCW^{+}82] Russell J. Harding, Monte R.

Campbell, William E. Watry, John R. Wilson, and Wilhelm Taylor. Designing the Oscilloscope Measurement System. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(9):5–14, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Hutchinson:1998:UUA

[HD98] Thomas W. Hutchinson and Ronald R. Derynck. Updating a UNIX Application Suite for the Windows NT World. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 50(1):21–??, November 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98nov/tc-11-98.htm.

Hauge:1987:MHP

[HDH^{+}87] B. R. Hauge, R. E.

Dunlap, C. D. Hoekstra, Chong Num Kwee, and P. R. Van Loan. A multichip hybrid printed circuit board for advanced handheld calculators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(8):25–30, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Hart:1996:FMA

[HDH96] Michael G. Hart, Christopher P. Duff, and Stephen W. Hinch. Firmware measurement algorithms for the HP 83480 digital communications analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Hernday:1975:HS

[HE75] P. R. Hernday and C. J. Enlow. A high-performance 2-to-18-GHz sweeper. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(7): 2–14, March 1975. CODEN HPJOAX. ISSN 0018-1153.

Heath:1985:CGP

[Hea85] Dean M. Heath. Custom graphics processor unit for the Integral PC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(10):10– 12, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Head:1994:SSU

[Hea94] Grant E. Head. Six-sigma software using cleanroom software engineering techniques. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3):40–50, June 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94jun/jun94_40.pdf;
http://www.hp.com/hpj/
toc-6-94.html.

Heflin:1973:CFG

[Hef73] E. H. Heflin. Compact function generator with enhanced capability/cost ratio. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(11):15–20, July 1973. CODEN HPJOAX. ISSN 0018-1153.

Heinzl:1988:IAT

[Hei88] Johann J. Heinzl. Input amplifier and trigger circuit for a 500-MHz frequency and time interval analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):24–27, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Hendricks:1984:TSF

[Hen84] T. Michael Hendricks. Testing for short-circuit failures. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):28– 30, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Hentschel:1991:CFO

[Hen91] Christian Hentschel. Calibration of fiber optic power meters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1):70–72, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Herreman:1973:LCS

[Her73] G. O. Herreman. Laser/ calculator system improves encoder plate measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(10):16–18, June 1973. CODEN HPJOAX. ISSN 0018-1153.

Hetrick:1979:LDB

[Het79]  M. V. Hetrick.  Low-

cost data base management. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4):19–25, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Holmlund:1982:NSP

[HEW82] G. W. Holmlund, G. E. Elmore, and D. C. Wood. A new series of programmable sweep oscillator plug-ins. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (2):11–21, February 1982. CODEN HPJOAX. ISSN 0018-1153.

Heyl:1981:PDA

[Hey81] L. E. Heyl. Precision data acquisition teams up with computer power. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(7):6–??, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Hodge:1995:FPD

[HFKS95] David J. Hodge, Bradly J.

Foster, Steven J. Kommrusch, and Tom J. Searby. An 11.8-in flat panel display monitor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(4):51– 60, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_51.pdf;   http:

                                        //www.hp.com/hpj/toc-
                                        08-95.html.

Hennessee:1981:FDP

[HFOS81] L. W. Hennessee, A. K. Frankel, M. A. Overton, and R. B. Smith. Firmware determines plotter personality. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(11):16– 24, November 1981. CO-DEN HPJOAX. ISSN 0018-1153.

Ha:1979:IDS

[HG79] E. P. L. Ha and J. R. Groff. The integrated display system and terminal access method (HP 300). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(6):6–9, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Haydamack:1981:VDS

[HG81] W. J. Haydamack and D. J. Griffin. VLSI design strategies and tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):5–7, 10– 12, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Hahn:1997:GPS

[HGWS97] Kenneth H. Hahn, Kirk S. Giboney, Robert E. Wilson, and Joseph Straznicky.

Gigabyte-per-second optical interconnection modules for data communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97dec/de97a6.htm.

Holdaway:1978:NGR

[HH78] S. N. Holdaway and M. D.

Humpherys. The next generation RF spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(10):2–7, June 1978. CODEN HPJOAX. ISSN 0018-1153.

Hauser:1990:EFT

[HH90] Frank E. Hauser and Nguyen P. Hung. Effect of fiber texture on the anisotropic dimensional change of Cu 1.8 wt.% Be. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6): 88–91, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Harkins:1991:ACE

[HH91] Daniel R. Harkins and Michael A. Heinzelman. Accuracy considerations and error correction techniques for 20-GHz lightwave component analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1): 34–40, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Heffner:1995:MPD

[HH95] Brian L. Heffner and Paul R. Hernday. Measurement of polarizationmode dispersion. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):27– 33, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_27.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Harrington:1971:RRM

[HHR71] H. W. Harrington, J. R. Hearn, and R. F. Rauskolb. The routine rotational microwave spectrometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(10): 2–3, June 1971. CODEN HPJOAX. ISSN 0018-1153.

Hassun:1973:SSG

[HHS^{+}73] R. Hassun, M. Humpherys, D. Scherer, D. K. Young, B. Stribling, and C. Cook.

Synthesized signal generation to 1.3 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(7):11–19, March 1973. CODEN HPJOAX. ISSN 0018-1153.

Heger:1976:CBF

[HHS76] C. E. Heger, R. C. Hyatt, and G. A. Seavey. A cesium beam frequency reference for severe environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(7):2–10, March 1976. CODEN HPJOAX. ISSN 0018-1153.

Holcomb:1997:DMO

[HHT^{+}97] M. S. Holcomb, S. O. Hall, W. S. Tustin, P. J. Burkart, and S. D. Roach. Design of a mixed-signal oscilloscope. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):13– 22, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a3.htm.

Higaki:1985:RMC

[Hig85a] Wesley H. Higaki. Remote monitoring and control of semiconductor processing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(7): 30–34, July 1985. CODEN HPJOAX. ISSN 0018-1153.

Higgins:1985:HED

[Hig85b] Marvin L. Higgins. Highquality electroluminescent display for a personal workstation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(10):12– 17, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Higgins:1988:AOS

[Hig88] Thomas M. Higgins, Jr. Analog output system design for a multifunction synthesizer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):66–69, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Higaki:1995:AIE

[Hig95] Wesley H. Higaki. Applying an improved economic model to software buy-versus-build decisions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (4):61–65, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95aug/
                                        aug95_61.pdf;   http:

//www.hp.com/hpj/toc-
08-95.html.

Hilton:1993:ACF

[Hil93a]  Howard E. Hilton.  A

10-Megasample-per-Second analog-to-digital converter with filter and memory. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (5):100–104, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Hilton:1993:ACL

[Hil93b] Howard E. Hilton. A 10-MHz analog-to-digital converter with 110-dB linearity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):105–112, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Heyman:1993:MHS

[HJG93] Eric V. V. Heyman, Rick R.

James, and Roger R. Graeber. Microcircuits for the HP 83750 series sweepers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):46–51, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Hall:1997:PPS

[HJH^{+}97] D. M. Hall, L. W. Jackson, K. Heiles, K. E. van der Veer, and T. J. Halpenny. PPA printer software driver design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3):12–21, June 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97jun/ju97a2.htm.

Hackleman:1983:CLT

[HJL^{+}83] David E. Hackleman, Norman L. Johnson, Craig S. Lage, John J. Vietor, and Robert L. Tillman. CMOSC: Low-power technology for personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(1):23–29 (or 23–28??), January 1983. CODEN HPJOAX. ISSN 0018-1153.

Huff:1975:HXR

[HJW75] D. W. Huff, D. E. Johnson, and J. M. Wade. Highsensitivity X-Y recorder has few input restrictions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(6): 2–9, February 1975. CO-DEN HPJOAX. ISSN 0018-1153.

Hall:1975:SSG

[HK75] J. A. Hall and Young Dae Kim. Synthesized signal generator operation to 2.6 GHz with wideband phase modulation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(11):15– 20, July 1975. CODEN HPJOAX. ISSN 0018-1153.

Hassun:1988:AWS

[HK88] Roland Hassun and Albert W. Kovalick. An arbitrary waveform synthesizer for DC to 50 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2):69–71, 73– 77, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Huttemann:1979:PPP

[HKA79] W. Huttemann, L. Kristen, and P. Aue. A precision, programmable pulse generator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(5):3–10, May 1979. CODEN HPJOAX. ISSN 0018-1153.

Hartman:1984:NRT

[HKC^{+}84a] Douglas O. Hartman, Steven R. Kusmer, Elizabeth A. Clark, Douglas V. Larson, and Billy Chu. New Real-Time Executive supports large programs and multiple users. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (2):26–31, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Hetrick:1984:NVC

[HKC84b] Michael V. Hetrick, Michael L. Kolesar, and J. D. Cooley. A new 32-bit VLSI computer family. II. software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(3):3–6 (or 3– 5??), March 1984. CODEN HPJOAX. ISSN 0018-1153.

                                        URL https://www.hpl.
                                        hp.com/hpjournal/pdfs/
                                        IssuePDFs/1984-03.pdf.

Hoffmann:1993:ODM

[HKH93] Brian E. Hoffmann, David A. Keefer, and Douglas K. Howell. Online defect management via a client/server relational database management system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):73–84, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Hinch:1996:NIW

[HKM96] Stephen W. Hinch, Michael J. Karin, and Christopher M. Miller. A new instrument for waveform analysis of digital communications signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Hardage:1974:ODH

[HKZ74] P. K. Hardage, S. R. Kushnir, and T. J. Zamborelli. Optimizing the design of a high-performance oscilloscope. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(1):8–14, September 1974. CODEN HPJOAX. ISSN 0018-1153.

Herlinger:1972:IDD

[HL72] J. E. Herlinger and W. J.

Lloyd. Inside the 7900 disc drive. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(9):6–11, May 1972. CODEN HPJOAX. ISSN 0018-1153.

Holl:1985:VDL

[HL85] James H. Holl and Frank E.

La Fetra, Jr. VLSI delivers low-cost, compact HP 3000 computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(9):4–7, September 1985. CODEN HPJOAX. ISSN 0018-1153.

Hughes:1986:PLF

[HL86] William L. Hughes and Kent W. Luehman. PCIB: a low-cost, flexible instrument control interface for personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(5): 11–16, May 1986. CODEN HPJOAX. ISSN 0018-1153.

Hunt:1986:DPC

[HLH86] Barry F. Hunt, Steven C. Leavitt, and David C. Hempstead. Digital processing chain for a Doppler ultrasound subsystem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6): 45–48, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Hiebert:1989:SOI

[HLM89] Steven P. Hiebert, John J. Lang, and Keith A. Marchington. Sharing overlay and image planes in the Starbase/X11 Merge system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):33–38, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Hentschel:1983:DBI

[HLT^{+}83] Christian Hentschel, Adolf Leiter, Stephan Traub, Horst Schweikardt, and Volker Eberle. Designing bipolar integrated circuits for a pulse/function generator family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6):33–38, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Hicks:1971:OPM

[HM71] C. L. Hicks and M. R.

Mellon. Optical power measurements made easy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(11): 10–16, July 1971. CODEN HPJOAX. ISSN 0018-1153.

Hueftle:1984:SOH

[HM84] E. R. Hueftle and J. R.

Murphy. Speech output for HP 1000 and HP 3000 computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(1):34–35, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Herr:1985:LLP

[HM85] Tammy V. Herr and Hatem E. Mostafa. Lowmass, low-cost pen-lift mechanism for high-speed plotting. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4):29–30, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Hwang:1988:DPE

[HM88] Ching-Fa Hwang and William T. McMahon. Discless program execution and virtual memory management. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):15–20, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Holdaway:1978:SPM

[HMLN78] S. N. Holdaway, D. H. Molinari, S. H. Linkwitz, and M. J. Neering. Signal processing in the Model 8568 A spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(10): 9–10, 12–16, June 1978. CODEN HPJOAX. ISSN 0018-1153.

Horner:1973:DDM

[HMM73] J. F. Horner, L. W. Masters, and P. T. Mingle. DMM and DAC modules expand low-cost measuring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(10):7–15, June 1973. CODEN HPJOAX. ISSN 0018-1153.

Hotchkiss:1996:NMS

[HMM96] Thomas R. Hotchkiss, Norman D. Marschke, and Richard M. McClosky. A new memory system design for commercial and technical computing products. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (1):44–51, February 1996. CODEN HPJOAX. ISSN 0018-1153. URL http://

www.hp.com/hpj/toc-02-
96.html.

Hyatt:1973:HBT

[HMO73] R. C. Hyatt, L. F. Mueller, and T. N. Osterdock. A high-performance beam tube for cesium beam frequency standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(1):14–23, September 1973. CODEN HPJOAX. ISSN 0018-1153.

Hall:1974:MIC

[HMPF74] J. A. Hall, D. J. Mellor, R. D. Pering, and A. Fong. Microwave integrated circuits solve a transmission problem in educational TV. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(12):20– 24, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Huning:1987:POH

[HMSFR87] Hans Huning, Emmerich Mueller, Siegmar Schmidt, and Michael Fleischer-Reumann. Precision optical heads for 850 to 1700 and 450 to 1020 nanometers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (2):22–27, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Hanlon:1983:HCM

[HMW83] M. Hanlon, G. G. Moyer, and P. G. Winninghoff. A high-resolution color monitor. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(9):18–21, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Hasebe:1982:FCH

[HMZ82] Kunio Hasebe, William R. Mason, and Thomas J. Zamborelli. A fast, compact, high-quality digital display for instrumentation applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(1):20–28, January 1982. CODEN HPJOAX. ISSN 0018-1153.

Herington:1987:SVU

[HNL87] Daniel E. Herington, Paul A.

Nichols, and Roger D. Lipp. Software verification using branch analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(6): 13–22, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Hoekstra:1997:FMS

[Hoe97]  C. D. Hoekstra.  Fre-

quency modulation of system clocks for EMI reduction. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):101– 106, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a13.htm.

Holl:1979:PHS

[Hol79] J. H. Holl. Philosophy of HP 3000 Series 33 diagnostics. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(9):15–16, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Holland:1982:LPO

[Hol82a]  G. L. Holland.  Laser

printer optics control and diagnostic circuit. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7):5–??, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Holland:1982:SPJ

[Hol82b] G. L. Holland. Sensing paper jams. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(7):13– ??, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Holland:1982:SMD

[Hol82c]  G. L. Holland.  A

synchronous mirror-motor drive for the Laser Printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7): 8–9, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Holmberg:1984:ELS

[Hol84] Randy W. Holmberg. Extensive library simplifies digital board test setup. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (10):23–25, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Holloway:1995:CEP

[Hol95] Robert R. Holloway. Capillary electrophoresis: A product of technological fusion. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):6–9, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_6t.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Hoover:1991:ITN

[Hoo91] David M. Hoover. An instrument for testing North American digital cellular radios. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):65–70, 72, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Horth:1969:PHP

[Hor69] T. C. Horth. Premonitory heartbeat patterns recognized by electronic monitor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(2):12–20, October 1969. CODEN HPJOAX. ISSN 0018-1153.

Horine:1979:IPD

[Hor79] D. A. Horine. Innovative package design enhances HP 300 effectiveness. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(6):

26–30, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Howard:1979:TXO

[How79] P. G. Howard. Triggered X-Y oscilloscope displays. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30 (9):25–28, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Heinz:1993:NFM

[HPF93] William W. Heinz, Ronald E. Pratt, and Peter H. Fisher. A new family of microwave signal generators for the 1990s. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):6–11, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Haselby:1981:YPH

[HPH81] R. D. Haselby, D. J. Perach, and S. R. Haugh. Yaxis pen handling system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(11): 25–32, November 1981. CODEN HPJOAX. ISSN 0018-1153.

Horner:1996:IPC

[HPL96] Rita N. Horner, Rajendra D. Pendse, and Fan Kee Loh. Implementation of pad circuitry for radially staggered bond pad arrangements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Haworth:1987:DDM

[HPM87] David J. Haworth, John R.

Pottinger, and Murdo J. McKissock. Dedicated display monitors digital radio patterns. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(7):4–10, 12– 13, July 1987. CODEN HPJOAX. ISSN 0018-1153.

Holland:1982:GSA

[HR82] Edward R. Holland and James L. Robertson. GUEST — a signature analysis based test system for ECL logic. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3):26–29, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Harry:1991:SRB

[HR91] P. D. Harry and T. W.

Rush. Specifying realtime behavior in HP-SL. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (5):40–45, December 1991. CODEN HPJOAX. ISSN 0018-1153.

Hodges:1985:DTT

[HRE85] John Hodges, Keith S. Roskelley, and Dennis R. Edson. Dynamic testing of thin-film magnetic recording discs. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(11):11– 21, November 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Heger:1981:PAV

[HRH81] T. J. Heger, P. A. Redding, and R. L. Hester. Plugin assemblies for a variety of data acquisition/control applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(7):16– 22, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Hentschel:1977:IPW

[HRZE77] C. Hentschel, G. Riebesell, J. Zellmer, and V. Eberle. An individualized pulse/ word generator system for subnanosecond testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (12):16–22, August 1977. CODEN HPJOAX. ISSN 0018-1153.

Hearn:1968:NCS

[HS68] J. R. Hearn and D. C.

Spreng. New concepts in signal generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(12):15– 19, August 1968. CODEN HPJOAX. ISSN 0018-1153.

Hopkins:1980:OSF

[HS80] G. W. Hopkins and A. Schwartz.

An optical system for fullspectrum measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (2):17–20, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Hall:1994:IPP

[HSB^{+}94] Corinna A. E. Hall, Aneesa R.

Scandalis, Damon W. Broder, Shelley I. Moore, Reza Movaghar, W. W. Rhoads, and W. H. Schwiebert. Inkjet printer print quality enhancement techniques. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):35–40, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/35to40.pdf.

Hakenjos:1987:UIS

[HSK87] Ulrich Hakenjos, Wolfgang Srok, and Ruediger Kreiser. User interface and software architecture for a data and arbitrary waveform generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(4):12–20, April 1987. CODEN HPJOAX. ISSN 0018-1153.

Hurtado-Sanchez:1986:IWE

[HSMA^{+}86] Luis Hurtado-Sanchez, Amy Tada Mueller, Robert A. Adams, Kristy Ward, and Rebecca A. Dahlberg. Implementing a worldwide electronic mail system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(9): 30–48, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Hollis:1988:CTI

[HSW88] Jennie L. Hollis, Philip C. Schultz, and William J. Walsh. Color thermal inkjet printer electronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (4):51–56, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Hoecker:1983:HMC

[HSYJ83] Kenneth R. Hoecker, James R.

Schwartz, Francis A. Young, and Dean R. Johnson. Handpulled magnetic card, mass storage system for a portable computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6): 15–19, 21–23, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Hashimoto:1976:AWD

[HT76] S. Hashimoto and T. Tamamura. An automatic widerange digital LCR meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (1):9–15, September 1976. CODEN HPJOAX. ISSN 0018-1153.

Halberg:1986:EBF

[HT86] Leslie I. Halberg and Karl E. Thiele. Extraction of blood flow information using Doppler-shifted ultrasound. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6):35–40, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Holcomb:1992:HAA

[HT92] Matthew S. Holcomb and Daniel P. Timm. A high-throughput acquisition architecture for a 100-MHz digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):11–17, 19–20, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Huenemann:1973:SPT

[Hue73] R. G. Huenemann. Signal processing techniques for automatic transceiver testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(12):8–13, August 1973. CODEN HPJOAX. ISSN 0018-1153.

Hunt:1992:DAU

[Hun92] William L. Hunt. Developing an advanced user interface for HP VEE. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):78–83, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Hurst:1990:HOD

[Hur90] Michael S. Hurst. HP OpenView Data Line Monitor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):71–75, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Harline:1996:SMW

[HVSM96] Matt J. Harline, Brendan A. Voge, Loren P. Staley, and Badir M. Mousa. Symmetric multiprocessing workstations and servers system — designed for high performance and low cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (1):8–17, February 1996. CODEN HPJOAX. ISSN 0018-1153. URL http://

www.hp.com/hpj/toc-02-
96.html.

Hale:1973:LSA

[HW73] W. L. Hale and G. E.

Weibel. A low-frequency spectrum analyzer that makes slow sweeps practical. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(1):2–13, September 1973. CODEN HPJOAX. ISSN 0018-1153.

Hodor:1982:HMS

[HW82] Ken M. Hodor and Malcolm E. Woodward. A high-performance memory system with growth capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3):15–17, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Harmon:1988:IPH

[HW88] J. Paul Harmon and John A. Widder. Integrating the printhead into the HP DeskJet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):62–66, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Harper:1991:HCI

[HW91] Steven L. Harper and Robert S. Worsley. The HP 48SX calculator input/ output system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(3):35– 40, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Hug:1995:UIH

[HWK95] Berthold Hug, Gerhard J. Walz, and Markus Kuhl. User interaction in HP PE/ SolidDesigner. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(5):14– 23, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95oct/
                                        oct95_14.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-95.html.

Harper:1987:ILL

[HWS87] Steven L. Harper, Robert S.

Worsley, and Bruce A. Stephens. An infrared link for low-cost calculators and printers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(10):16– 21, October 1987. CODEN HPJOAX. ISSN 0018-1153.

Heinz:1983:SSG

[HZ83] William W. Heinz and Paul A. Zander. 2-to-26.5-GHz synthesized signal generator has internally leveled pulse modulation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(5):3–6, 8–9, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Ishak:1985:MDM

[IC85] Waguih S. Ishak and Kok-Wai Chang. Magnetostaticwave devices for microwave signal processing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(2):10–20, February 1985. CODEN HPJOAX. ISSN 0018-1153.

Ishak:1993:PTL

[ICNT93] Waguih S. Ishak, Kent W.

Carey, Steven A. Newton, and William R. Trutna, Jr. Photonic technology for lightwave communications test applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):6–10, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Iyengar:1994:EBR

[IGHG^{+}94] Arun K. Iyengar, Thaddeus S. Grzeski, Valerie J. Ho-Gibson, Tracy A. Hoover, and John R. Vasta. An event-based, retargetable debugger. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):33– 43, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_33.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Itoh:1969:DMT

[IK69] H. Itoh and K. L. Knudsen. Direct measurement of transistor noise voltage, noise current and noise figure. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(2):2–7, October 1969. CODEN HPJOAX. ISSN 0018-1153.

Ikemoto:1983:MSU

[Ike83] J. Ikemoto. Mass storage unit exerciser. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(4):35– 36, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Ishak:1981:SDL

[IKS81] W. S. Ishak, H. E. Karrer, and W. R. Shreve. Surface-acoustic-wave delay lines and transversal filters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12):3–8, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Ilic:1982:TFF

[Ili82]  Dragan B. Ilic.  Thin

films formed by plasma enhanced chemical vapor deposition. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(8):24–27, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Ingman:1970:PKC

[Ing70] E. M. Ingman. Programmer is key to computingcounter systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(4):7–12, December 1970. CODEN HPJOAX. ISSN 0018-1153.

Ingman:1976:CDI

[Ing76] E. M. Ingman. Cardprogrammable digital IC tester simplifies incoming inspection. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(2):11–18, October 1976. CODEN HPJOAX. ISSN 0018-1153.

Inhelder:1975:NIE

[Inh75] A. E. Inhelder. A new instrument enclosure with greater convenience, better accessibility, and higher attenuation of RF interference. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(1):19–24, September 1975. CODEN HPJOAX. ISSN 0018-1153.

Ichino:1980:VIA

[IOS80] T. Ichino, H. Ohkawara, and N. Sugihara. Vector impedance analysis to 1000 MHz. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(1):22–31, January 1980. CODEN HPJOAX. ISSN 0018-1153.

Ishida:1980:IHI

[Ish80] C. J. Ishida. An interactive HP 3000/IBM mainframe link. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):29– 32, November 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Ignoffo:1978:MDP

[ISLW78] J. J. Ignoffo, M. J.

Sproviero, P. R. Luque, and K. B. Wade. Managing dot-matrix printing with a microprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(15):8–19, November 1978. CODEN HPJOAX. ISSN 0018-1153.

Ives:1988:MSB

[Ive88] Fred H. Ives. Multifunction synthesizer for building complex waveforms. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (1):52–54, 56–57, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Jackson:1974:TST

[Jac74] W. H. Jackson. A thinfilm/semiconductor thermocouple for microwave power measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(1): 16–18, September 1974. CODEN HPJOAX. ISSN 0018-1153.

Jackoway:1988:APC

[Jac88]  Gary Jackoway.  Au-

tomating the printed circuit board design process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):68–71, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Jundanian:1983:CHU

[JAD83] Richard H. Jundanian, Janet R. Accettura, and John N. Dukes. Control hardware for an ultrasound imaging system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(12):3–5, December 1983. CODEN HPJOAX. ISSN 0018-1153.

Jandourek:1996:MPD

[Jan96] Emil Jandourek. A model for platform development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (4):56–71, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/aug96/tc-08-
                                        96.htm;   http://www.
                                        hp.com:80/hpj/aug96/augart6.
                                        htm.

Jungerman:1993:DLM

[JBS93] Roger L. Jungerman, David M.

Braun, and Kari K. Salomaa. Dual-output laser module for a tunable laser source. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):32–34, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Jegle:1995:RTH

[Jeg95] Ulrike Jegle. Reproducibility testing of the HP CE instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):50–56, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_50.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Jekat:1975:PFC

[Jek75] H. J. Jekat. A portable 1100-MHz frequency counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(8): 2–8, April 1975. CODEN HPJOAX. ISSN 0018-1153.

Jekat:1980:P

[Jek80] H. J. Jekat. A 400-to-1600-MHz/ 8 prescaler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4):5–6, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Jennyc:1998:LEB

[Jen98] K. S. Jennyc. Linking enterprise business systems to the factory floor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):62– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/98may/tc-05-98.htm.

Jeppsen:1985:NFD

[Jep85] Bryce E. Jeppsen. A new family of dot matrix line printers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(6):4–6, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Jeremiasen:1974:LAA

[Jer74] R. Jeremiasen. Logarithmic amplifier accepts 100dB signal range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(7):16–17, March 1974. CODEN HPJOAX. ISSN 0018-1153.

Jessen:1989:VNI

[Jes89a]  Kenneth Jessen.  VX-

Ibus: A new interconnection standard for modular instruments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(2):91– 95, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Jessen:1989:VPD

[Jes89b]  Kenneth Jessen.  VX-

Ibus product development tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2):96–97, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Jewett:1980:SUF

[JF80] R. A. Jewett and R. W.

Fredrickson. The System 45C user’s firmware interface. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(12):6–9, December 1980. CODEN HPJOAX. ISSN 0018-1153.

Jewell:1985:IMS

[JJ85] Michael B. Jewell and Mark W. Johnson. Internally modular signal generator mechanical design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(12): 14–18, December 1985. CODEN HPJOAX. ISSN 0018-1153.

Jackson:1988:DPC

[JKP^{+}88] Larry A. Jackson, Kieran B. Kelly, David W. Pinkernell, Steve O. Rasmussen, and John A. Widder. DeskJet printer chassis and mechanism design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(5):67–75, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Johnson:1980:BCT

[JM80] N. L. Johnson and V. V. Marathe. Bulk CMOS technology for the HP-41C. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(3): 20–22, March 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Jensen:1987:FVS

[JM87] James E. Jensen and Eric D. McHenry. Firmware for a vector signal generator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):30– 33, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Jungerman:1991:DOM

[JM91] Roger L. Jungerman and David J. McQuate. Development of an optical modulator for a high-speed lightwave component analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1):41–45, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Johnson:1993:ADM

[JM93] Dana L. Johnson and Christopher J. Magnuson. Architectural design for a modular oscilloscope system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):51–58, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Johnson:1977:NSS

[Joh77a] L. Johnson. A new series of small computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(7): 2–6, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Johnston:1977:ASP

[Joh77b] R. L. Johnston. APLGOL:

structured programming facilities for APL. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(11):11– 13, July 1977. CODEN HPJOAX. ISSN 0018-1153.

Johnson:1980:ALN

[Joh80] D. R. Johnson. An approach to large-scale noncontact coordinate measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(9):16–17, 19– 20, September 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Johnson:1990:OHO

[Joh90] William R. Johnson. An overview of the HP OSI Express card. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):6–8, February 1990. CODEN HPJOAX. ISSN 0018-1153.

Jones:1981:DCR

[Jon81] V. C. Jones. Desktop computer redesigned for instrument automation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(7):23– 31, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Jones:1984:DCC

[Jon84] V. C. Jones. Data communication for a 32-bit computer workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(3):24– 25, March 1984. CODEN HPJOAX. ISSN 0018-1153.

URL https://www.hpl.
hp.com/hpjournal/pdfs/
IssuePDFs/1984-03.pdf.

Jones:1989:PRL

[Jon89] Carolyn F. Jones. Partially reflective light guides for optoelectronics applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5):98–104, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Joy:1996:MHV

[Joy96] Molly Joy. Migrating HP VUE desktop customizations to CDE. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):29– 37, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Jonker:1984:LLC

[JR84] Robert J. Jonker and Gerard P. Rozing. Lowdispersion liquid chromatography. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(4):3–6, 8, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Jones:1981:PDW

[JRC81] L. T. Jones, J. J. Ressmeyer, and C. A. Clark. Precision DVM has wide dynamic range and high systems speed. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(4):23– 31, April 1981. CODEN HPJOAX. ISSN 0018-1153.

Jensen:1986:PAL

[JRM^{+}86] Gordon A. Jensen, Stephen P. Reames, Jerry D. Morris, Jeffrey H. Smith, Jeffrey Tomberlin, and James M. Umphrey. Protocol analyzer for local area networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7):42–47, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Jeppsen:1974:IBE

[JS74] B. E. Jeppsen and S. E. Schultz. Interface bus expands instrument utility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(10):22–24, June 1974. CODEN HPJOAX. ISSN 0018-1153.

Jahn:1993:UID

[JS93] Robert Jahn and Harald Seeger. User interface design for the HP 8146A OTDR. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):72–78, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Johnson:1997:FDH

[JU97] L. Johnson and S. R. Undy.

Functional design of the HP PA 7300LC processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 48–60, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/functional_
design.pdf;   http://
cpus.hp.com/techreports/
timing_flexibility.pdf;
http://www.hp.com/hpj/
97jun/ju97a7.htm.

Juncker:1978:HHC

[Jun78] R. K. Juncker. Higherperformance HP 1000 computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(14):2–5, October 1978. CODEN HPJOAX. ISSN 0018-1153.

James:1972:IMS

[JY72] R. L. James and F. J.

Yockey. Interactive model 20 speaks algebraic language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(4):8–13, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Kearl:1992:LPC

[KA92] Daniel A. Kearl and Michael S. Ard. Low-cost plain-paper color inkjet printing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):64–68, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Kilian:1995:CLE

[KA95] Jens Kilian and Heinz-Peter Arndt. Common Lisp as an embedded extension language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):69–73, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_69.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Kafadar:1988:SCV

[Kaf88] Karen Kafadar. Statistical calibration of a vector demodulator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(3):18– 25, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Kahan:1979:PCK

[Kah79] William M. Kahan. Personal calculator has key to solve any equation f(x) = 0. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(12):20–26, December 1979. CODEN HPJOAX. ISSN 0018-1153.

Kahan:1980:HCE

[Kah80] William M. Kahan. Handheld calculator evaluates integrals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(8):23–32, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Kahkoska:1992:DOP

[Kah92] James A. Kahkoska. Digital oscilloscope persistence. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (1):45–47, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Kaiser:1991:CMS

[Kai91] Winfried Kaiser. Component Monitoring System parameter module interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):19–21 (or 19–20??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Kalkuhl:1993:PDC

[Kal93] Christoph Kalkuhl. Pulse/ data channel extends programmable pulse generator applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(2):56– 59, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Kaltenbach:1995:HDA

[Kal95] P. Kaltenbach. A highsensitivity diode array detector for on-column detection in capillary electrophoresis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(3):20– 24, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_20.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Kameoka:1972:MHC

[Kam72]  Y. Kameoka.  Measur-

ing high-value capacitors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(7):

8–13, March 1972. CODEN HPJOAX. ISSN 0018-1153.

Kamins:1982:SIC

[Kam82] Theodore I. Kamins. Silicon integrated circuits using beam-recrystallized polysilicon. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(8):10–13, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Karin:1982:PMS

[Kar82] M. J. Karin. Putting the measurement system on the bus (HP 1980A/B Oscilloscope Measurement System on HP-IB). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(9):21–24, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Karp:1986:MUI

[Kar86] Sydney M. Karp. Modifying an ultrasound imaging scanner for Doppler measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(6):41– 44, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Karube:1994:VET

[Kar94] Koji Karube. Vector error testing by automatic test equipment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):64– 66, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_64.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Kasson:1972:FMH

[Kas72]  J. M. Kasson.  Func-

tional modularity helps designer and user of new measurement and control subsystem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(12):13–19, August 1972. CODEN HPJOAX. ISSN 0018-1153.

Kato:1990:EI

[Kat90] R. Kato. Error injection. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (6):33–34, December 1990. CODEN HPJOAX. ISSN 0018-1153.

Kapuskar:1969:MAN

[KB69a] W. T. Kapuskar and C. J. Balmforth. Monitoring airport noise. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(11):11– 15, July 1969. CODEN HPJOAX. ISSN 0018-1153.

Kapuskar:1969:RMO

[KB69b] W. T. Kapuskar and C. J.

Balmforth. Real-time measurement and on-line processing of acoustical and other audio-frequency spectra. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(11):2–10, July 1969. CODEN HPJOAX. ISSN 0018-1153.

Kameoka:1971:VHV

[KB71] Y. Kameoka and J. E.

Bonhomme. Very high and very low resistances — why and how they are measured. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(7):11–16, March 1971. CODEN HPJOAX. ISSN 0018-1153.

Katen:1985:IPI

[KB85] Cheryl V. Katen and Thomas R. Braun. An inexpensive, portable ink-jet printer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(5):11–17, 19–20, May 1985. CODEN HPJOAX. ISSN 0018-1153.

King:1993:HTL

[KBHS93] Randall King, David M. Braun, Stephen W. Hinch, and Karl Shubert. Highspeed time-domain lightwave detectors (for optical fibre systems). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):83–86, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Kirkpatrick:1977:CCA

[KCP77] G. R. Kirkpatrick, D. R. Chambers, and R. E. Pratt. Coaxial components and accessories for broadband operation to 26.5 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(10): 10–16, June 1977. CODEN HPJOAX. ISSN 0018-1153.

Karrer:1983:UIO

[KD83] H. Edward Karrer and Arthur M. Dickey. Ultrasound imaging: an overview. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(10):3–6, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Krauss:1974:MPP

[KE74] G. Krauss and R. Eggert. A moderately priced 20-MHz pulse generator with 16-volt output. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(7):10–15, March 1974. CODEN HPJOAX. ISSN 0018-1153.

Kepler:1985:EST

[KE85] David L. Kepler and James A. Espeland. Electronics system for a transportable professional computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(10):6–9, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Keely:1989:SJI

[Kee89] Catherine A. Keely. Solder joint inspection using laser Doppler vibrometry. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5):81–85, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Keiter:1971:FCS

[Kei71] R. C. Keiter. A fully calibrated, solid state microwave spectrum analyzer. microwave spectrum analysis with performance advantages previously associated only with lower frequency instruments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(1): 4–9, September 1971. CO-DEN HPJOAX. ISSN 0018-1153.

Kelly:1992:RMI

[Kel92] Christopher P. Kelly. Realtime multitasking of instruments in the VXIbus command modules. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):29– 34, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Kelly:1994:FEU

[Kel94] Christopher P. J. Kelly. On-the-fly engineering units conversion. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):21– 24, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94oct/
                                        oct94_21.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-94.html.

Kerschner:1980:MDC

[Ker80] R. K. Kerschner. Mechanical design constraints for a total station. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(9):12–14, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Ketelsen:1974:PRM

[Ket74]  E. Ketelsen.  Peak-to-

average ratio measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(12):17–18, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Keremitsis:1995:HDS

[KF95] Eileen Keremitsis and Ian J. Fuller. HP Distributed Smalltalk: A tool for developing distributed applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(2):85– 92, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_85.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Kerschen:1992:HOS

[KG92] K. Kerschen and J. R.

Glasson. HP-UX operating system kernel support for the HP 9000 series 700 workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3):6–10, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Kohlhardt:1987:DVT

[KGH^{+}87] Charles Kohlhardt, Tony W. Gaddis, Daniel L. Halperin, Stephen R. Undy, and Robert A. Schuchard. Design, verification, and test methodology for a VLSI chip set. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(9):18–26, September 1987. CODEN HPJOAX. ISSN 0018-1153.

Kelly:1981:PHH

[KGK81] J. Kelly, T. R. Groves, and Huei Pei Kuo. A precision, high-current, highspeed electron beam lithography column. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(5):14–15, 18–20, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Kao:1990:SIH

[KGTM90] Ping-Hui Kao, William A. Gates, Bruce A. Thompson, and Dale K. Mc-Cluskey. Support for the ISO 9660/HSG CD-ROM file system standard in the HP-UX operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6): 54–59, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Kay:1969:TFN

[KH69] B. Kay and J. L. Harmon. Twelve functions in a new digital meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(7):2–13, March 1969. CODEN HPJOAX. ISSN 0018-1153.

Kohli:1982:PHS

[KH82] Manmohan Kohli and Bennie E. Helmso. Packaging the HP 3000 series 64. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3):30–32, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Kretz:1984:ALC

[KH84] Wolfgang Kretz and Hans-Georg Härtl. Automatic liquid chromatograph injection and sampling. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(4): 21–24, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Keller:1988:GRT

[KH88a] Alexander Keller and Ulrich H. Haeberlen. Generalization of the Redfield-Kunz treatment of quadrature phase time domain. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(6): 74–76 (or 74–75??), December 1988. CODEN HPJOAX. ISSN 0018-1153.

Kovalick:1988:AWS

[KH88b] Albert W. Kovalick and Roland Hassun. Arbitrary waveform synthesizer applications in magnetic recording and radar. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 86–93, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Khalily:1981:TEC

[Kha81]  E. Khalily.  Transis-

tor electrical characterization and analysis program (TECAP). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(6):16– 17, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Kwinn:1982:LS

[KHI82] Kathryn Y. Kwinn, Robert M. Hallissy, and Roger E. Ison. 9826A/9836A language systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(5):24– 32, May 1982. CODEN HPJOAX.

Klemba:1990:HON

[KHLM90] Keith S. Klemba, Mark L.

Hoerth, Hui-Lin Lim, and Maureen C. Mellon. HP OpenView network management architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2): 54–59, April 1, 1990. CO-DEN HPJOAX. ISSN 0018-1153.

Kurtz:1989:OMS

[KHW89] Barry D. Kurtz, Donna Ho, and Teresa A. Wall. An object-oriented methodology for systems analysis and specification. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(2):86– 90, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Kihlstrom:1982:MSA

[Kih82]  K. P. Kihlstrom.  Mi-

crowave solid-state amplifiers and modulators for broadband signal generators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7):30–31, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Kilcrease:1994:SAD

[Kil94] Catherine L. Kilcrease.

Structured analysis and design in the redesign of a terminal and serial printer driver. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(4):52–61, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94aug/aug94_52.pdf;
http://www.hp.com/hpj/
toc-8-94.html.

Kilk:1997:PPF

[Kil97] E. Kilk. PPA printer firmware design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):22– 30, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97jun/ju97a3.htm.

Kinsell:1981:TDC

[Kin81] D. A. Kinsell. Tablet/ display combination supports interactive graphics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (1):22–23, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Kinicki:1983:QAU

[Kin83] R. M. Kinicki. Quantitative analysis for ultrasound imaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(10):8–9, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Kirkpatrick:1971:ESD

[Kir71] G. R. Kirkpatrick. Effective stripline device characterization. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(9):12– 16, May 1971. CODEN HPJOAX. ISSN 0018-1153.

Kiss:1970:CCF

[Kis70] A. Z. Kiss. A calibrated computer-based Fourier analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(10):10–20, June 1970. CODEN HPJOAX. ISSN 0018-1153.

Kishi:1997:BIC

[Kis97] Y. Kishi. A benchtop inductively coupled plasma mass spectrometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(4): 72–79, August 1997. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97aug/
au97a9.htm.

Kjosness:1992:MDS

[Kjo92] David P. Kjosness. A manufacturing-oriented digital stimulus/response test instrument. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):59– 68, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Klaiss:1978:FIM

[Kla78] D. E. Klaiss. Firmware intelligence for measurement and control processing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(11):10–12, 14–15, 17– 18, July 1978. CODEN HPJOAX. ISSN 0018-1153.

Knoll:1979:IPO

[KM79] A. F. Knoll and N. D. Marschke. An innovative programming and operating console. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(7):13– 17, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Koehler:1984:SOH

[KM84] Loren M. Koehler and Timothy C. Mackey. Speech output for HP series 80 personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1):29–33, 35– 36, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Knee:1997:GCO

[KM97] D. L. Knee and C. E.

Moore. General-purpose 3V CMOS operational amplifier with a new constanttransconductance input stage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(4):114–120, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a15.htm.

Kikuchi:1988:WGL

[KMT88] Derrick T. Kikuchi, Rafael F.

Miranda, and Peter A. Thysell. A waveform generation language for arbitrary waveform synthesis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 94–96, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Kawabata:1993:HVN

[KN93] Shigeru Kawabata and Akira Nukiyama. A highperformance 1.8-GHz vector network and spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):76–84, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Kastle:1997:NFS

[KNF^{+}97] S. Kastle, F. Noller, S. Falk, A. Bukta, E. Mayer, and D. Miller. A new family of sensors for pulse oximetry. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):39–53, February 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97feb/fe97a7.htm.

Kjos:1996:HCC

[KNTV96] Todd J. Kjos, Helen Nusbaum, Michael K. Traynor, and Brendan A. Voge. Hardware cache coherent input/output. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(1):52–59, February 1996. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/toc-02-96.html.

Koepke:1980:AMF

[Koe80] L. L. Koepke. An automatic microwave frequency counter test system to 40 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4):11–12, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Koenig:1991:HMP

[Koe91]  Mary K. Koenig.  A

high-speed microwave pulse modulator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):34–36 (or 34–35??), April 1991. CODEN HPJOAX. ISSN 0018-1153.

Kohoutek:1972:PIM

[Koh72] H. J. Kohoutek. 9800 processor incorporates 8-MHz microprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(4):19–22, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Kong:1995:DES

[Kon95] Michael M. Kong. DCE:

An environment for secure client/server computing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (6):6–15, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_6t.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Kovalick:1979:HBB

[Kov79] M. Y. Kovalick. HP 300 business BASIC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(6):23– 26, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Kovalick:1981:PSS

[Kov81] A. W. Kovalick. 8662A power-on and self-test sequences. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(2):9–10, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Kafadar:1988:CDU

[KP88] Karen Kafadar and Lynn M. Plouse. Circuit design using statistical data analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):12–17, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Kretz:1990:CPS

[KP90] Wolfgang Kretz and Gerhard Ple. A compact, programmable sample injector and autosampler for liquid chromatography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(2):17–23, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Kemplin:1982:GPM

[KPC^{+}82] R. M. Kemplin, D. M.

Petersen, C. Ta. Chuong, D. C. Tribolet, and R. J. Porcelli. Graphics plotter mechanical design for performance and reliability at low cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(12):28– 33, December 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Kraemer:1989:PDU

[Kra89] Thomas F. Kraemer. Product development using object-oriented software technology. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):87–97, 99– 100, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Kible:1983:HDA

[KRDS83] Dieter Kible, Bernhard Roth, Martin Dietze, and Ulrich Schöttmer. Highspeed data analyzer tests threshold and timing parameters. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(7):14– 25, July 1983. CODEN HPJOAX. ISSN 0018-1153.

Krebs:1992:MDH

[Kre92] Kenneth R. Krebs. Mechanical design of the HP 4980 Network Advisor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (5):41–47, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Krieger:1981:ERA

[Kri81a] J. J. Krieger. An endtidal/respiration-rate algorithm. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(9):12–15, 18, 20, 21, September 1981. CO-DEN HPJOAX. ISSN 0018-1153.

Krieger:1981:MAC

[Kri81b] J. J. Krieger. Making accurate CO_{2} measurements (using IR). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(9):19–20, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Kruger:1988:PMU

[Kru88] Gregory A. Kruger. Project management using software reliability growth models. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):

30–35, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Kruger:1989:VFA

[Kru89] Gregory A. Kruger. Validation and further application of software reliability growth models. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(2):75– 79, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Krucky:1994:FFS

[Kru94a] Jan Krucky. Fuzzy family setup assignment and machine balancing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3): 51–64, June 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94jun/
jun94_51.pdf;   http:/
/www.hp.com/hpj/toc-6-
94.html.

Krucky:1994:SGB

[Kru94b] Jan Krucky. Sidebar: The greedy board family assignment heuristic. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(3):54, June 1994. CODEN HPJOAX. ISSN 0018-1153.

Kruger:1997:SCA

[Kru97] G. A. Kruger. The supply chain approach to planning and procurement management. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):28–38, February 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97feb/fe97a6.htm.

Kingsford-smith:1969:NAA

[Ks69] C. A. Kingsford-smith. Network analysis at low frequencies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(11):16–20, July 1969. CODEN HPJOAX. ISSN 0018-1153.

Kingsford-Smith:1972:ISG

[KS72] C. A. Kingsford-Smith. The incremental sweep generator — point-by-point accuracy with sweptfrequency convenience. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(11): 9–15, July 1972. CODEN HPJOAX. ISSN 0018-1153.

Kubicek:1997:HPD

[KSMM97] D. C. Kubicek, T. J. Sullivan, A. Mehra, and J. G. McBride. Highperformance processor design guided by system costs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):61– 68, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/high_
performance.pdf;   http:
//www.hp.com/hpj/97jun/
ju97a8.htm.

Keith:1978:SHD

[KSV78] J. C. Keith, L. T. Schulte, and A. K. Vogen. System 45 hardware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(8): 11–22, April 1978. CODEN HPJOAX. ISSN 0018-1153.

Kanago:1989:FDM

[KSW89] Kerwin D. Kanago, Mark A.

Stambaugh, and Brian D. Watkins. Firmware development for modular instrumentation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):20–26, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Kelly:1973:SES

[KT73] M. A. Kelly and C. E.

Tyler. A second-generation ESCA spectrometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(11): 2–14, July 1973. CODEN HPJOAX. ISSN 0018-1153.

Kaplan:1981:XMD

[KT81] R. J. Kaplan and R. S. Townsend. X-axis microgrip drive and platen design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(11):33–36, November 1981. CODEN HPJOAX. ISSN 0018-1153.

Krizan:1990:HOM

[KT90] Brock C. Krizan and Keith M. Taylor. The HP OSF/Motif window manager. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(3):12–22, 24– 26, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Kong:1995:DDS

[KT95] Michael M. Kong and David Truong. DCE directory services. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(6):23– 27, December 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_23.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Kuhlman:1976:HSS

[Kuh76] L. J. Kuhlman, Jr. Highpower solid-state 5.9-12.4-GHz sweepers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(3):19–24, November 1976. CODEN HPJOAX. ISSN 0018-1153.

Kurtz:1981:AAS

[Kur81] B. D. Kurtz. Automating application system operation and control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(4):19– 22, April 1981. CODEN HPJOAX. ISSN 0018-1153.

Kusters:1996:GPS

[Kus96] John A. Kusters. The global positioning system and HP SmartClock. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): 60–67, December 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Kirkpatrick:1974:EPS

[KV74] G. R. Kirkpatrick and D. R. Veteran. Economical precision step attenuators for RF and microwaves. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(9): 21–24, May 1974. CODEN HPJOAX. ISSN 0018-1153.

Kolts:1994:MTO

[KV94] Bertram S. Kolts and Rodney K. Village. Manufacturing test optimization for VXI-Based scanning analog-to-digital converters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):30–34, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
94oct/oct94_30.pdf;
http://www.hp.com/hpj/
toc-10-94.html.

Knudsen:1980:LDM

[KW80] K. L. Knudsen and R. W. Widmayer. Light detection and measurement in a high-performance spectrophotometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(2):20–24, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Klein:1990:PGO

[KW90] Stefan G. Klein and Hans-Jurgen Wagner. A 500-MHz pulse generator output section. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4):79–84, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Knight:1993:TMO

[KW93] J. Douglas Knight and Joseph N. West. A two-axis micropositioner for optical fiber alignment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(6):80–84, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Kelly:1987:VMO

[KWR^{+}87] Wayne M. Kelly, Mark J.

Woodward, Eric B. Rodal, Pedro A. Szente, and James D. McVey. Vector modulator, output amplifier, and multiplier chain assemblies for a vector signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):48– 52, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Keisling:1985:SDA

[KYKM85] Mark D. Keisling, Dorothy J.

Yackle, David B. Karlin, and Elizabeth Gates Moore. Serial data acquisition and simulation for a high-speed protocol analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(7):18–24, July 1985. CODEN HPJOAX. ISSN 0018-1153.

Lemoff:1997:SLC

[LAB97] Brian E. Lemoff, Lewis B. Aronson, and Lisa A. Buckman. SpectraLAN: A lowcost multi-wavelength local area network. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97dec/de97a5.htm.

LaFollette:1980:STD

[LaF80] R. J. LaFollette. 5335A self test and diagnostics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (9):30–??, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Laing:1981:ISP

[Lai81] V. L. Laing. Instrument system provides precision measurement and control capabilities. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(7):3–8, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Lalwani:1993:PIM

[Lal93] Rajesh Lalwani. POSIX interface for MPE/iX. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 41–46, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Lamy:1974:MET

[Lam74] J. C. Lamy. Microelectronics enhances thermocouple power measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (1):19–23, September 1974. CODEN HPJOAX. ISSN 0018-1153.

Lam:1995:SDF

[Lam95] William K. Lam. Synthesis of 100% delay fault testable combinational circuits by cube partitioning. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1):105–109, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95feb/feb95105.pdf;
http://www.hp.com/hpj/
toc-02-95.html.

Lane:1975:FML

[Lan75] A. B. Lane. A functionally modular logic system for a CRT terminal. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(10):6– 10, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Lane:1977:MFE

[Lan77] T. A. Lane. Microprogrammed features of the 21MX E-series. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(7):24–27, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Landsness:1983:EIH

[Lan83] C. J. Landsness. The electronics interface for the Hewlett–Packard Interface Loop. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(1):11–16, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Landis:1994:DDT

[Lan94] Adele S. Landis. Datadriven test systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(4): 62–66, August 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_62.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Larson:1983:SLS

[Lar83a] R. K. Larson. Sample-and-hold leveling system (in signal generator). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(5):7–??, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Larson:1983:ATA

[Lar83b] John D. Larson, III. An acoustic transducer array for medical imaging. I. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (10):17–22, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Lau:1980:DTN

[Lau80] C. Y. Lau. A digitally tuned notch filter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(8):14–15, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Lau:1987:DVB

[Lau87] Chung Y. Lau. Digital and vector baseband circuits for a vector signal generator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):39– 44, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Lowe:1980:CRO

[LB80] D. J. Lowe and P. V. Boyd. Card reader offers compatibility and expanded capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(3):12–15, March 1980. CODEN HPJOAX. ISSN 0018-1153.

Lefebvre:1998:AII

[LB98] K. T. Lefebvre and J. M. Brown. An API for interfacing interactive 3D applications to high-speed graphics hardware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2): 6–??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Lewis:1986:ICP

[LBKY86] Jeffrey A. Lewis, Andrew A. Berlin, Allan J. Kuchinsky, and Paul K. Yip. Integrated circuit procedural language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(6):4–7, 9– 10, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Lee:1995:RSM

[LBLS95a] Ruby B. Lee, John P.

Beck, Joel Lamb, and Kenneth E. Severson. Realtime software MPEG video decoder on multimediaenhanced PA 7100LC processors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):60–68, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_60.pdf.

Lee:1995:RTS

[LBLS95b] Ruby B. Lee, John P.

Beck, Joel Lamb, and Kenneth E. Severson. Realtime software MPEG video decoder on multimediaenhanced PA 7100LC processors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):63–68, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
toc-04-95.html.

Lowe:1996:ICM

[LC96] Douglas E. Lowe and Guy M. Cox. Implementing the Capability Maturity Model for software development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):6– 14, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/aug96/tc-08-
                                        96.htm;   http://www.
                                        hp.com:80/hpj/aug96/augart1.
                                        htm.

Lagoni:1990:HMM

[LCB90] Peter A. Lagoni, Christopher Crall, and Thomas G. Bartz. HP MAP 3.0 manufacturing message specification/800. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4):31–39, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Lefebvre:1998:OHO

[LCPG98] K. T. Lefebvre, R. J. Casey, M. J. Phelps, and C. D. Goeltzenleuchter. An overview of the HP OpenGL software architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2):9–??, May 1998. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        98may/tc-05-98.htm.

Lamy:1979:BIM

[LD79] J. C. Lamy and F. K. David. Broadband input mixers for a microwave spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(8): 8–13, August 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Long:1993:SPU

[LD93] David W. Long and Christopher P. Duff. A survey of processes used in the development of firmware for a multiprocessor embedded system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):59–65, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Larsen:1972:EET

[LDNT72] J. L. Larsen, R. F. Dillman, A. M. Nardizzi, and R. N. Tverdoch. An effective ECG telemetry system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(8): 2–8, April 1972. CODEN HPJOAX. ISSN 0018-1153.

Lob:1983:CDS

[LE83] C. G. Lob and A. O. Elkins.

18-MHz clock distribution system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(8):17–??, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Leath:1989:YCT

[Lea89]  Charles L. Leath.  40

years of chronicling technical achievement. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):6–13, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Lee:1974:SWT

[Lee74] R. T. Lee. The 4940A sine wave transmitter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(12):9–11, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Leis:1971:MRF

[Lei71] C. T. Leis. Microprogramming, ROMs, firmware and all that. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(2):4–9, October 1971. CODEN HPJOAX. ISSN 0018-1153.

Leong:1977:CTP

[Leo77] W. W. Leong. CRT terminal provides both APL and ASCII operation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(11):25– 28, July 1977. CODEN HPJOAX. ISSN 0018-1153.

Leonard:1983:DAC

[Leo83] D. F. Leonard. Data acquisition and control software for the 3421A using the HP-85 computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(2):13–14, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Lettang:1992:ECH

[Let92] Frank J. Lettang. ECL clocks for high-performance RISC workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4): 23–25, August 1992. CO-DEN HPJOAX. ISSN 0018-1153.

LeVitt:1988:PMI

[LeV88] Richard M. LeVitt. Process measures to improve R&D scheduling accuracy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 61–65, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Lomb:1996:SMS

[LEV96] Reiner Lomb, Kelly A.

Emo, and Roy M. Vandoorn. Storage management solutions for distributed computing environments. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):81–89, October 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/oct96/oc96a10.
htm.

Ladeau:1991:UFS

[LF91] B. R. Ladeau and C. W. Freeman. Using formal specification for product development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(5):46–50, December 1991. CODEN HPJOAX. ISSN 0018-1153.

Lynch-Freshner:1989:NCT

[LFWEJ89] Lawrence A. Lynch-Freshner, R. Thomas Watson, Brian B. Egan, and John J. Jencek. NewWave computer-based training development facility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):48–56, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Li:1992:HSI

[LG92] Daniel Li and Audrey B. Gore. HP 9000 Series 700 input/output subsystem. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (4):26–33, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Lum:1998:MBM

[LGV^{+}98] Paul Lum, Michael Greenstein, Edward D. Verdonk, Charles Grossman, Jr., and Thomas L. Szabo. A 150-MHz-bandwidth membrane hydrophone for acoustic field characterization. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):6–??, August 1998. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
98aug/tc-08-98.htm.

Lewis:1982:OSD

[LH82] John R. Lewis and Laurence M. Hubby, Jr. Optical system design for the Laser Printing System. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(7): 3–4, 6–7, 9–10, July 1982. CODEN HPJOAX. ISSN 0018-1153.

Lloyd:1995:ADT

[LH95] Paul Lloyd and Samuel D.

Horowitz. Adopting DCE technology for developing client/server applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (6):16–22, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_16.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-95.html.

Lam:1989:NO

[LHD89] Beatrice Lam, Scott A. Hanson, and Anthony J. Day. The NewWave Office. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):23–31, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Lewis:1998:DPH

[LHIR98] R. L. Lewis, E. A. Handgen, N. J. Ingegneri, and G. T. Robinson. Delivering PCI in HP B-class and C-class workstations: A case study in the challenges of interfacing with industry standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):51– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/98may/tc-05-98.htm.

Leavitt:1983:SCA

[LHL83] Steven C. Leavitt, Barry F. Hunt, and Hugh G. Larsen.

A scan conversion algorithm for displaying ultrasound images. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(10):30– 34, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Lienhart:1988:MST

[Lie88] Deborah H. Lienhart.

A multidevice spooler for technical applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):77–80, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Lienhart:1997:SEI

[Lie97] D. A. Lienhart. Soft-Bench 5.0: the evolution of an integrated software development environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48 (1):6–11, February 1997. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97feb/
fe97a1.htm.

Liljenwall:1972:PPC

[Lil72] E. T. Liljenwall. Packaging the pocket calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(10):

12–13, June 1972. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/72jun/ju72a3.htm.

Linkwitz:1969:FON

[Lin69] S. Linkwitz. Frequencydomain oscilloscope now measures to 1250 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(8): 14–20, April 1969. CODEN HPJOAX. ISSN 0018-1153.

Lingane:1979:IMA

[Lin79a] P. J. Lingane. Interactive modulation analyzer control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(11):26–30, November 1979. CODEN HPJOAX. ISSN 0018-1153.

Linkwitz:1979:NPS

[Lin79b] S. H. Linkwitz. New performance standards in microwave spectrum analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(8):3–7, August 1979. CODEN HPJOAX. ISSN 0018-1153.

Lindberg:1984:PHH

[Lin84] Thomas B. Lindberg. Packaging the HP-71B handheld computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(7):17– 20, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Lin:1997:FSM

[Lin97] J. J. Lin. Fully synthesizable microprocessor core via HDL porting. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):107– 113, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a14.htm.

Langguth:1975:UMU

[LJ75] A. Langguth and W. D.

Jackson. Using a modular universal counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(11):9– 14, July 1975. CODEN HPJOAX. ISSN 0018-1153.

Lo:1980:CTP

[LK80] C. C. Lo and R. W.

Keil. A compact thermal printer designed for integration into a personal computer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(7):22– 26, July 1980. CODEN HPJOAX. ISSN 0018-1153.

Lassiter:1985:UII

[LK85] Albert Einstein Lassiter and Charles R. Kogler. User interface and internal controller for an RF signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(12):6–10, December 1985. CODEN HPJOAX. ISSN 0018-1153.

Luszcz:1983:USS

[LKHK83] Joseph M. Luszcz, William A. Koppes, David C. Hempstead, and Robert J. Kunz. Ultrasound system software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(12):6–13, December 1983. CODEN HPJOAX. ISSN 0018-1153.

Lang:1991:HVM

[LL91] Marilyn J. Lang and Gary W. Lum. The HP Vectra 486 memory controller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):78–83 (or 78–82??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Lum:1988:EMH

[LLS88] Gary W. Lum, Milton J. Lau, and Wesley H. Stelter. Expanded memory for the HP vectra ES personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(6):57–63, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Lepoff:1973:SDS

[LM73] J. H. Lepoff and R. A.

Morris. Schottky-barrier diodes structured for better high-frequency performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(10):2–6, June 1973. CODEN HPJOAX. ISSN 0018-1153.

Lindberg:1981:PXS

[LM81] E. E. Lindberg and C. L.

Merja. A precision X-Y stage and substrate handling system for electron beam lithography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(5):16– 18, May 1981. CODEN HPJOAX. ISSN 0018-1153.

Laczynski:1986:CSP

[LM86] Edward Laczynski and Robert V. Miller. Case study: PC instruments counter versus traditional counters. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(5):29– 32, May 1986. CODEN HPJOAX. ISSN 0018-1153.

Luszcz:1993:BGB

[LM93] Joseph M. Luszcz and Daniel G. Maier. Bridging the gap between structured analysis and structured design for real-time systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):90–99, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Lenk:1984:ISC

[LMM84] R. M. Lenk, C. E. Mear, Jr., and M. E. Meier. An I/O subsystem for a 32-bit computer operating system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(3):38– 41, March 1984. CODEN HPJOAX. ISSN 0018-1153.

                                        URL https://www.hpl.
                                        hp.com/hpjournal/pdfs/
                                        IssuePDFs/1984-03.pdf.

Leong:1996:DPE

[LMNQ96] Frank Leong, Satya P. Mylavarabhata, Trong Nguyen, and Frank Quemada. Distributed processing environment: A platform for distributed telecommunications applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):17–21, October 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/oct96/oc96a2.htm.

Leyrer:1984:HSS

[LNH^{+}84] Joachim Leyrer, Günter E.

Nill, Detlev Hadbawnik, Günter Hoeschele, and Joachim Dieckmann. A high-speed spectrophotometric LC detector. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(4): 31–41, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Lo:1981:PSE

[Lo81] C. C. Lo. Plotter servo electronics contained on a single IC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(10):8–9, October 1981. CODEN HPJOAX. ISSN 0018-1153.

Longo:1990:HOE

[Lon90] Joseph R. Longo, Jr. The HP OSI Express Card software diagnostic program. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(1): 59–67, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Longust:1992:DPC

[Lon92] Timothy A. Longust. DesignJet plotter chassis design: A concurrent engineering challenge. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):28–31, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Loomis:1997:AHT

[Loo97] T. P. Loomis. Audit history and time-slice archiving in an object DBMS for laboratory databases. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):80–89, August 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97aug/au97a10.htm.

Loughry:1972:CDI

[Lou72] D. C. Loughry. A common digital interface for programmable instruments: The evolution of a system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(2): 8–11, October 1972. CO-DEN HPJOAX. ISSN 0018-1153.

Loughry:1975:HIB

[Lou75]  D. C. Loughry.  The

Hewlett–Packard Interface Bus: current perspectives.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(5): 2–4, January 1975. CO-DEN HPJOAX. ISSN 0018-1153.

Lowe:1969:PRC

[Low69] R. D. Lowe. Portable, rugged cable fault locator for VHF communications and CATV. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(10):2–8, June 1969. CODEN HPJOAX. ISSN 0018-1153.

Larson:1992:MHO

[LP92] Douglas V. Larson and Kyle A. Polychronis. A multiprocessor HP-UX operating system for HP 9000 computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):56–61, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Lob:1983:HVM

[LRFL83] Clifford G. Lob, Mark J.

Reed, Joseph P. Fucetola, and Mark A. Ludwig. Highperformance VLSI memory system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(8):14–20, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Larson:1983:FSC

[LRS83] Paul A. Larson, Bruce L. Ryder, and Thomas J. Stark. Fused silica capillary columns for gas chromatography. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(12):35– 40, December 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Lawson:1976:PWN

[LS76] W. S. Lawson and D. D. Sharrit. Processing widerange network analyzer signals for analog and digital display. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(11):8– 11, July 1976. CODEN HPJOAX. ISSN 0018-1153.

Larson:1983:WYM

[LS83] Ronald K. Larson and Lawrence A. Stark. A wideband YIG-tuned multiplier and pulsed signal generation system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(5):10–11, 13– 16, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Layman:1987:MDH

[LS87] Judith A. Layman and Mark A. Smith. Mechanical design of the HP-18C and HP-28C handheld calculators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(8):17–20, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Leckel:1995:EDF

[LSM^{+}95] Edgar Leckel, Jurgen Sang, Rolf Muller, Clemens Ruck, and Christian Hentschel. Erbium-doped fiber amplifier test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):13– 19, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_13.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Lantz:1985:SSP

[LT85] Jeffrey M. Lantz and Ben B. Tyson. Shuttle system and packaging of a lowcost, high- reliability, 300lpm line printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(6):9–12, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Lau:1990:OHI

[LT90] Roger K. Lau and Mark E. Thompson. An overview of the HP interactive visual interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(5):6–10, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Landers:1984:IRC

[LTC^{+}84] David M. Landers, Timothy W. Tillson, Jack D. Cooley, Richard R. Rupp, and G. L. Shults. An interactive run-time compiler for enhanced BA-SIC language performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (3):15–19, 21, March 1984. CODEN HPJOAX. ISSN 0018-1153. URL https://

                                        www.hpl.hp.com/hpjournal/
                                        pdfs/IssuePDFs/1984-03.
                                        pdf.

Lubkin:1991:DSC

[Lub91] David C. Lubkin. DSEE: A software configuration management tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(3):77– 83, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Lowe:1982:CDS

[LW82] Lawrence Lowe and Brian W.

Woodroffe. Control and display system for a baseband analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(4):18–20, 22, April 1982. CODEN HPJOAX. ISSN 0018-1153.

Lewis:1984:BLG

[LWBB84] Kenneth W. Lewis, Alan D.

Ward, Xuan Bui, and R. J. Bury. BASIC language graphics subsystem for a 32-bit workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(5): 16–21, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Lee:1994:WAT

[LY94] Daniel T. L. Lee and Akio Yamamoto. Wavelet analysis: Theory and applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):44– 54, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_44.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Lymer:1987:MPA

[Lym87] Anthony Lymer. Microprocessorenhanced performance in an analog power meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(7): 26–30, July 1987. CODEN HPJOAX. ISSN 0018-1153.

Lynch:1980:CLA

[Lyn80a] T. R. Lynch. A custom LSI approach to a personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(8):18–22, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Lynch:1980:NWP

[Lyn80b] T. R. Lynch. A new world of personal/professional computation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(7):3–7, July 1980. CODEN HPJOAX. ISSN 0018-1153.

MacLeod:1973:PHC

[Mac73]  K. J. MacLeod.  A

portable high-resolution counter for low-frequency measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(3):10–15, November 1973. CODEN HPJOAX. ISSN 0018-1153.

MacLeod:1975:BTC

[Mac75] K. J. MacLeod. Big timer/ counter capability in a portable package. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(8):9–10, 11– 13, April 1975. CODEN HPJOAX. ISSN 0018-1153.

MacLeod:1993:ISS

[Mac93] Jean M. MacLeod. Implementing and sustaining a software inspection program in an R and D environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):60–63, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Maeda:1974:APD

[Mae74] K. Maeda. An automatic, precision 1-MHz digital LCR meter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(7):2–9, March 1974. CODEN HPJOAX. ISSN 0018-1153.

Magnin:1983:CSU

[Mag83] P. A. Magnin. Coherent speckle in ultrasound images. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(10):39– 40, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Magnin:1986:DEH

[Mag86] Paul A. Magnin. Doppler effect: History and theory. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6): 26–31, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Magenis:1994:SHI

[Mag94]  Sue Magenis.  Sidebar:

The HP instant ignition program. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):17, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Maioli:1990:OMI

[Mai90] Frederic Maioli. An objectoriented message interface for testing the HP 3000 data communications and terminal controller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6): 74–80, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Maier:1991:SLS

[Mai91] Frank A. Maier. Semiconductor laser sources with superior stability for optical loss measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1): 73–76, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Maier:1993:DCH

[Mai93] Frank Maier. Design considerations for the HP 8146A OTDR receiver. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (1):69–71, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Malzbender:1987:PTO

[Mal87] Thomas Malzbender. Permuted trace ordering allows low-cost, high-resolution graphics input. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(6):4–7, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Mantena:1969:SNT

[Man69] N. R. Mantena. Sources of noise in transistors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(2): 8–11, October 1969. CO-DEN HPJOAX. ISSN 0018-1153.

Manweiler:1990:HMF

[Man90] Steven W. Manweiler. HP MAP 3.0 file transfer, access, and management/800. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (4):24–30, August 1990. CODEN HPJOAX. ISSN 0018-1153.

Mandler:1994:PSM

[Man94] John Mandler. A printing solution for a multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):44– 52, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        44to52.pdf;   http://
                                        www.hp.com/hpj/toc-4-
                                        94.html.

Martin:1970:CMS

[Mar70] D. Martin. Computingcounter measurement systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(4):2–6, December 1970. CODEN HPJOAX. ISSN 0018-1153.

Martin:1971:FSM

[Mar71]  D. Martin.  Frequency

stability measurements by computing counter system.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (3):9–14, November 1971. CODEN HPJOAX. ISSN 0018-1153.

Martin:1977:PFC

[Mar77] R. E. Martin. Printing financial calculator sets new standards for accuracy and capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(2):22–28, October 1977. CODEN HPJOAX. ISSN 0018-1153.

Marcoux:1982:DEO

[Mar82] P. J. Marcoux. Dry etching: an overview. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(8):19–23, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Martin:1988:SQA

[Mar88] David E. Martin. Software quality assurance on the HP printed circuit design system project. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):84–86, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Martelli:1989:MHS

[Mar89] Anastasia M. Martelli. A model for HP-UX shared libraries using shared memory on HP precision architecture computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5): 86–90, October 1989. CO-DEN HPJOAX. ISSN 0018-1153.

Martin:1995:XOF

[Mar95a] Elizabeth A. Martin. X/ Open federated naming. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (6):28–33, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95dec/
dec95_28.pdf;   http:
//www.hp.com/hpj/toc-
12-95.html.

Martin:1995:IGA

[Mar95b] Paul Martin. An integrated graphics accelerator for a low-cost multimedia workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):43–50, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_43.pdf;

http://www.hp.com/hpj/
toc-04-95.html.

Matheson:1979:CIO

[Mat79a] W. G. Matheson. A computer input/output system based on the HP Interface Bus. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(7):9–10, 12– 13, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Mathis:1979:HSC

[Mat79b]  B. Mathis.  Human-

engineering the smallbusiness computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4): 4–5, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Mathieu:1984:IBT

[Mat84] Mark A. Mathieu. An interpreter-based board test programming environment. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):25– 28, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Matreci:1986:USM

[Mat86] Robert J. Matreci. Unbiased subharmonic mixers for millimeter-wave spectrum analyzers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):22– 26, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Matta:1990:AIT

[Mat90] Farid Matta. Advances in IC testing: the membrane probe card. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(3):77–85, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Matta:1992:AIC

[Mat92] Farid Matta. Advances in integrated circuit packaging: demountable TAB. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):62–77, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Mayotte:1988:IAD

[May88] Mark E. Mayotte. Integrating applications in a design management system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):80–83, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Maze:1980:FHL

[Maz80] C. Maze. The first HP liquid crystal display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(3): 22–24, March 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Marrocco:1974:QCD

[MB74] J. A. Marrocco and B. Bronson. A quality course in digital electronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(3):12–14, November 1974. CODEN HPJOAX. ISSN 0018-1153.

Molinari:1980:EPM

[MB80] D. H. Molinari and R. L.

Belding. An economical, portable microwave spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(3):27–32, March 1980. CODEN HPJOAX. ISSN 0018-1153.

Malin:1985:HMM

[MB85] Joseph L. Malin and Irving Bunton. HP maintenance management: a new approach to software customer solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(3):4–10, March 1985. CODEN HPJOAX. ISSN 0018-1153.

McTigue:1993:DAS

[MB93] Michael T. McTigue and Patrick J. Byrne. An 8-Gigasample-per-Second, 8-Bit data acquisition system for a sampling digital oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):11–23, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Miller:1996:ACP

[MB96] Chad G. Miller and Jerome M. Bailey. Automated Cterminal protein sequence analysis using the HP G1009A C-terminal protein sequencing system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(2): 73–82, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

McAuliffe:1995:NFS

[MBC95] Robert E. McAuliffe, James L. Benson, and Christopher B. Cain. A new, flexible sequencer architecture for testing complex serial bit streams. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):76– 90, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_76.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Millar:1998:ISH

[MBD^{+}98] Iain Millar, Martin Beale, Bryan J. Donoghue, Kirk W. Lindstrom, and Stuart Williams. The IrDA standards for high-speed infared communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1): ??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a2.htm.

Mangelsdorf:1987:VPH

[MBF^{+}87] Steven T. Mangelsdorf, Darrell M. Burns, Paul K. French, Charles R. Headrick, and Darius F. Tanksalvala. A VLSI processor for HP Precision Architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (9):4–11, September 1987. CODEN HPJOAX. ISSN 0018-1153.

Meyer:1989:NMM

[MBH^{+}89] Thomas O. Meyer, Russell C. Brockmann, Jeffrey G. Hargis, John Keller, and Floyd E. Moore. New midrange members of the Hewlett–Packard Precision Architecture computer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3):18–25, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Martinez:1995:BSA

[MBL95] Antonio A. Martinez, Aloke S. Bhandia, and Henry H. W. Lie. Benchmark standards for ASIC technology evaluation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4): 66–70, August 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95aug/
aug95_66.pdf;   http:
//www.hp.com/hpj/toc-
08-95.html.

Masters:1979:LMM

[MBW79] L. W. Masters, K. M. Blankenship, and M. J. Ward. A low-cost, microprocessorbased, 100 MHz universal counter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(1):2–7, 9–11, January 1979. CODEN HPJOAX. ISSN 0018-1153.

Miller:1982:CGP

[MC82] Robert M. Miller and Randy A. Coverstone. Controlling a graphics plotter with a handheld programmable calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(12): 16–18, December 1982. CODEN HPJOAX. ISSN 0018-1153.

McAllister:1984:FCS

[MC84] William H. McAllister and John R. Carlson. Floatingpoint chip set speeds realtime computer operation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (2):17–23, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Martinelli:1985:LRA

[MC85] Rene Martinelli and Jean Yves Chatron. A low-cost, reliable analog video display terminal design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(4):18– 22, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Miller:1991:BGI

[MC91] Christopher M. Miller and Roberto A. Collins. A broadband, generalpurpose instrumentation lightwave converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1): 51–57, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

McClellan:1987:ESH

[McC87] Paul J. McClellan. An equation solver for a handheld calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(8):30–34, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Morris:1976:TGP

[MCCB76] D. E. Morris, C. J. Christopher, G. W. Chance, and D. B. Barney. Third generation programmable calculator has computer-like capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(10):2–14, June 1976. CODEN HPJOAX. ISSN 0018-1153.

McDonley:1980:PDP

[McD80] P. A. McDonley. Program debugging with Pascal/64000. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(10):22– 23, October 1980. CODEN HPJOAX. ISSN 0018-1153.

McDermid:1984:REA

[McD84] John E. McDermid. Reducing errors in automated analog in-circuit test program generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):31– 36, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Methley:1997:DTH

[MCD97] Steven G. Methley, Alistair N. Coles, and Eric Deliot. Data transmission for higher-speed IEEE 802 LANs using twisted-pair copper cabling. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97dec/de97a4.htm.

McFarland:1998:WCS

[McF98] William J. McFarland.

Wireless communications: A spectrum of opportunity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1): ??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a1.htm.

McHenry:1987:WFS

[McH87]  Eric D. McHenry.  A

wideband FM subsystem for a low-noise synthesizer module. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(11):45–47, December 1987. CODEN HPJOAX. ISSN 0018-1153.

McIntire:1974:PDB

[McI74] R. E. McIntire. Powerful data base management systems for small computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(11):2– 10, July 1974. CODEN HPJOAX. ISSN 0018-1153.

McKinney:1975:BFA

[McK75] H. W. McKinney. Bandselectable Fourier analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(8):20–24, April 1975. CODEN HPJOAX. ISSN 0018-1153.

McKnight:1983:MSF

[McK83] R. N. McKnight. A mixing scheme to focus a transducer array dynamically. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(12):16–17, December 1983. CODEN HPJOAX. ISSN 0018-1153.

McKissock:1987:CMT

[McK87] Murdo J. McKissock. Constellation measurement: a tool for evaluating digital radio. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(7):13– 17, July 1987. CODEN HPJOAX. ISSN 0018-1153.

McQuate:1993:CLD

[MCM93] David J. McQuate, Kok Wai Chang, and Christopher J. Madden. Calibration of lightwave detectors to 50 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):87–92, February 1993. CODEN HPJOAX. ISSN 0018-1153.

McOmber:1980:VUC

[McO80]  V. D. McOmber.  A

voltmeter for a universal counter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(9):28–29, September 1980. CODEN HPJOAX. ISSN 0018-1153.

McShane:1992:INN

[McS92] John McShane. An introduction to neural nets. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (1):62–65, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Meyer:1979:CES

[MD79] G. L. Meyer and V. Delloy Forbes. Cost-effective electronics for the smallbusiness computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4): 6–10, 12–14, April 1979. CODEN HPJOAX. ISSN 0018-1153.

May:1986:PAM

[MD86] Robert B. May and Alesia Duncombe. Personal applications manager for HP portable computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(7): 18–21, July 1986. CODEN HPJOAX. ISSN 0018-1153.

McTigue:1996:DTD

[MD96] Michael M. McTigue and Christopher P. Duff. Differential time-domain reflectometry module for a digital oscilloscope and communications analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

McMinds:1990:PHO

[ME90] Donald L. McMinds and Benjamin J. Ellsworth. Programming with HP OSF/Motif widgets. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(3): 26, 29–35, June 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Metzger:1995:FSM

[ME95] Michael Metzger and Sabine Eismann. Freeform surface modeling. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):61–68, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95oct/oct95_61.pdf;
http://www.hp.com/hpj/
toc-10-95.html.

Meinert:1982:DTG

[Mei82] K. L. Meinert. Designing for testability with GUEST. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3):28–??, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Martini:1982:LPE

[MEM82] N. J. Martini, D. M. Ellement, and P. L. Ma. Low-cost plotter electronics design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(12):23– 26, December 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Merkelo:1973:DTS

[Mer73] J. Merkelo. A dc-to-20-GHz thin-film signal sampler for microwave instrumentation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(8):10– 13, April 1973. CODEN HPJOAX. ISSN 0018-1153.

Merrick:1981:MMC

[Mer81] E. B. Merrick. A miniature motor for the CO_{2} sensor (with thanks to Kettering). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(9):8–9, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Merchant:1982:EOV

[Mer82] P. P. Merchant. Electromigration: an overview (VLSI metallisation). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(8):28–31, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Mettetal:1986:XWA

[Met86] Pierry Mettetal. X. 25 wide area networking for HP computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):36– 40, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Meyers:1984:SCE

[Mey84] Nathan Meyers. Soft configuration enhances flexibility of handheld computer memory (HP-71B). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7): 10–13, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Meyer:1990:ITH

[Mey90a] Jeffrey D. Meyer. Interoperability testing for HP MAP 3.0. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4):50–53, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Meyer:1990:ECI

[Mey90b] John C. Meyer. Error correction implementation and performance in a CD-ROM drive. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):42–48, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Mikkelson:1983:NPT

[MFMS83] James M. Mikkelson, Fung-Sun Fei, Arun K. Malhotra, and S. Dana Seccombe. NMOS-III process technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(8):27–30, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Morrison:1972:VDU

[MFPE72] D. J. Morrison, B. W. Finnie, R. S. Patel, and K. H. Edwards. Versatile display unit extends correlator capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(3):8–15, November 1972. CODEN HPJOAX. ISSN 0018-1153.

Messenger:1987:FSD

[MFW87] Brian S. Messenger, Peter H. Fisher, and Stanley P. Woods. Firmware system design for a vector analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(11):17– 24, December 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Miller:1985:TDC

[MG85] Craig L. Miller and Mark L.

Gembarowski. Tape/disc controller serves integrated peripherals. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(3):37–39, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Moore:1992:HMC

[MG92] Mike P. Moore and Eric N.

Gullerud. A high-performance measurement coprocessor for personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 110–116 (or 110–111??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Mangelsdorf:1997:FVH

[MGBB97] S. T. Mangelsdorf, R. P.

Gratias, R. M. Blumberg, and R. Bhatia. Functional verification of the HP PA 8000 processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):22–31, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/functional_
verification.pdf;   http:

                                        //www.hp.com/hpj/97aug/
                                        au97a3.htm.

Merrick:1976:CNM

[MH76] E. B. Merrick and T. J. Hayes. Continuous, noninvasive measurements of arterial blood oxygen levels. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(2):2–9, October 1976. CODEN HPJOAX. ISSN 0018-1153.

Morrill:1980:IVV

[MH80] J. S. Morrill, Jr. and J. D. Hansen. Increased versatility for a versatile logic state analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(5):14– 18, May 1980. CODEN HPJOAX. ISSN 0018-1153.

McJunkin:1989:DCF

[MH89] Barton L. McJunkin and David M. Hoover. Design considerations in a fast hopping voltage-controlled oscillator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5):34–36, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Mason:1992:AAH

[MH92] Lee S. Mason and Mark C. Huth. Automated assembly of the HP DeskJet 500C/ DeskWriter C color print cartridge. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):77–83, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Marion:1977:WIO

[MHCH77] A. F. Marion, E. A. Heinsen, R. Chin, and B. E. Helmso. Wrist instrument opens new dimension in personal information. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (4):2–10, December 1977. CODEN HPJOAX. ISSN 0018-1153.

Mahn:1994:HPN

[MHKS94] Johannes Mahn, Jurgen Haberle, Siegfried Kopp, and Tim Schwegler. HP-PAC: A new chassis and housing concept for electronic equipment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(4):23– 28, August 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_23.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Mei:1996:LTS

[MHV96a] Zequn Mei, H. A. Holder, and H. A. Vander Plas. Low-temperature solders. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (4):91–98, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/aug96/tc-08-
                                        96.htm.

Mei:1996:LS

[MHV96b] Zequn Mei, Helen A. Holder, and Hubert A. Vander Plas. Lowtemperature solders. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): ??, August 1996. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com:
                                        80/hpj/aug96/augart10.
                                        htm.

Michener:1971:AAS

[Mic71] C. K. Michener. An almost all solid-state stripchart recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(12):13– 16, August 1971. CODEN HPJOAX. ISSN 0018-1153.

Mickowski:1972:MMM

[Mic72] J. Mickowski. Multiprogrammer magnifies minicomputer I/O capacity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (12):20–23, August 1972. CODEN HPJOAX. ISSN 0018-1153.

Maeda:1982:IIS

[MiITT82] Kohichi Maeda, Jin ichi Ikemoto, Fumiro Tsurada, and Teruo Takeda. Intelligent instrument streamlines DC semiconductor parameter measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (10):3–15, October 1982. CODEN HPJOAX. ISSN 0018-1153.

Millard:1975:OVA

[Mil75] J. K. Millard. An oscilloscope vertical-channel amplifier that combines monolithic, thick-film hybrid, and discrete technologies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27 (4):8–11, December 1975. CODEN HPJOAX. ISSN 0018-1153.

Miller:1983:ATA

[Mil83]  David G. Miller.  An

acoustic transducer array for medical imaging. II. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (10):22–26, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Miller:1984:PMA

[Mil84] Robert M. Miller. Plugin module adds Forth language and assembler to a handheld computer (HP-71B). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7):37–40, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Millard:1988:ODO

[Mil88] Joe K. Millard. A onegigasample-per-second digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3): 58–59, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Miller:1990:HLS

[Mil90] Christopher M. Miller. High-speed lightwave signal analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):80–91, February 1990. CODEN HPJOAX. ISSN 0018-1153.

Miller:1994:HDT

[Mil94] Christopher M. Miller.

High-speed digital transmitter characterization using eye diagram analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (4):29–37, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_29.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Miller:1996:MMS

[Mil96] Robert M. Miller. Managing a multicompany software development project. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(2): 50–53, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Maze:1992:IPC

[MJKS92] Craig Maze, Loren E. Johnson, Daniel A. Kearl, and James P. Shields. Ink and print cartridge development for the HP DeskJet 500C/DeskWriter C printer family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):69–76, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Matsui:1979:FEC

[MK79] Y. Matsui and M. Kohli. A friendly, easy-to-service computer (HP 3000 Series 33). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(9):9–12, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Moore:1997:BES

[MK97] K. E. Moore and E. R. Kirshenbaum. Building evolvable systems: the OR-Blite project. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):62–72, February 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97feb/fe97a9.htm.

Marcus:1995:HIL

[MKR95] Jane B. Marcus, Navaneet Kumar, and Lawrence J. Rose. HP integrated login. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (6):34–40, December 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95dec/
                                        dec95_34.pdf;   http:

//www.hp.com/hpj/toc-
12-95.html.

Miller:1996:DOR

[MKW^{+}96] Christopher M. Miller, Randall King, Mark J. Woodward, Tim L. Bagwell, Donald L. Faller, Jr., Joseph Straznicky, and Naily L. Whang. Design of optical receiver modules for digital communications analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Maisenbacher:1993:TLS

[MLJP93] Bernd Maisenbacher, Edgar Leckel, Robert Jahn, and Michael Pott. Tunable laser sources for optical amplifier testing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):11–19, February 1993. CODEN HPJOAX. ISSN 0018-1153.

May:1988:DDH

[MLPN88] Donna J. May, Mark D.

Lund, Thomas B. Pritchard, and Claude W. Nichols. Data to dots in the HP DeskJet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(5):76–80, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Martin:1979:SML

[MLS79] L. R. Martin, K. L. Lange, and S. T. Sparks. A synthesized microwave local oscillator with continuoussweep capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(8):13–19, August 1979. CODEN HPJOAX. ISSN 0018-1153.

Malhotra:1983:FNC

[MLSW83] Arun K. Malhotra, Glen E. Leinbach, Jeffrey J. Straw, and Guy R. Wagner. Finstrate: a new concept in VLSI packaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(8):24–26, August 1983. CODEN HPJOAX. ISSN 0018-1153.

McGrath:1972:FTR

[MM72] W. R. McGrath and A. Miller. Fine-line thermal recording on Z-fold paper. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23 (6):17–19, February 1972. CODEN HPJOAX. ISSN 0018-1153.

Maiorca:1979:CSS

[MM79] P. P. Maiorca and N. H.

MacNeil. A closed-loop system for smoothing and matching step motor responses. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(2):18–23, February 1979. CODEN HPJOAX. ISSN 0018-1153.

Mansingh:1990:SLA

[MM90] Vivek Mansingh and Kent P.

Misegades. System level air flow analysis for a computer system processing unit. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(5):82–87, October 1990. CODEN HPJOAX. ISSN 0018-1153.

Miller:1995:JAH

[MM95] Christopher M. Miller and David J. McQuate. Jitter analysis of high-speed digital systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):49– 56, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_49.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Makarechian:1997:TPA

[MM97a] M. Makarechian and N. J. Malcolm. Testing policing in ATM networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):90–95, August 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97aug/au97a11.htm.

Mellquist:1997:ONM

[MM97b] P. E. Mellquist and T. Murray. Object-oriented network management development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):59– 65, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a11.htm.

Marcoux:1989:NSM

[MMNR89] Paul J. Marcoux, Paul P. Merchant, Vladimir Naroditsky, and Wulf D. Rehder. A new 2D simulation model of electromigration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3):79–84, June 1989. CODEN HPJOAX. ISSN 0018-1153.

McWilliams:1997:PPC

[MMPT97] J. L. McWilliams, L. M.

MacMillan, B. Pathak, and H. A. Talley. PPA printer controller ASIC development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3):31–37, June 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97jun/ju97a4.htm.

Maeda:1979:MLM

[MN79] K. Maeda and Y. Narimatsu. Multi-frequency LCR meters test components under realistic conditions. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(2):24–31, February 1979. CODEN HPJOAX. ISSN 0018-1153.

Masters:1974:QFC

[MO74] L. W. Masters and W. J.

O’Buch. Quality frequency counters designed for minimum cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(11):11–14, July 1974. CODEN HPJOAX. ISSN 0018-1153.

Moayeri:1998:LCF

[Moa98] Nader Moayeri. A lowcomplexity, fixed-rate compression scheme for color images and documents. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 50 (1):46–??, November 1998. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98nov/
tc-11-98.htm.

Monnier:1968:NEC

[Mon68] R. E. Monnier. A new electronic calculator with computerlike capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20 (1):3–9, September 1968. CODEN HPJOAX. ISSN 0018-1153.

Montijo:1988:DFH

[Mon88] B. Allen Montijo. Digital filtering in a highspeed digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):70–76, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Monroe:1992:MVC

[Mon92] Michael J. Monroe. Machine vision in color print cartridge production. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):

87–92, August 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Moore:1992:HNA

[Moo92] Edmund G. Moore. The HP network advisor: A portable test tool for protocol analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):6–10, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Mordan:1970:FTB

[Mor70] W. Mordan. A fast time base for a high-frequency oscilloscope. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(5):17–20, January 1970. CODEN HPJOAX. ISSN 0018-1153.

Morrill:1976:IPL

[Mor76] J. S. Morrill, Jr. Interfacing a parallel-mode logic state analyzer to serial data. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(4):21–24, December 1976. CODEN HPJOAX. ISSN 0018-1153.

Morgan:1993:HRO

[Mor93] Kevin D. Morgan. The HP-RT real-time operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(4):23–30, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Munsch:1990:HIA

[MOT90] Pamela W. Munsch, Warren I. Otsuka, and Gary D. Thomsen. HP IVI application program interface design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(5):21–28, 30–31, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Moyer:1978:FCM

[Moy78] J. J. Moyer. Firmware control of a microprocessorbased graphics terminal. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29 (5):12–16, January 1978. CODEN HPJOAX. ISSN 0018-1153.

Myers:1981:PFT

[MP81] R. Myers and R. D. Peck. 200-kHz power FET technology in new modular power supplies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(8):3–7, 10, August 1981. CODEN HPJOAX. ISSN 0018-1153.

Mottola:1983:DAS

[MP83] Roberto Mottola and Eckhard Paul. Data analyzer software/firmware design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(7): 25–28, July 1983. CODEN HPJOAX. ISSN 0018-1153.

McNamee:1989:MFH

[MP89] Michael D. McNamee and David L. Platt. A modular family of high-performance signal generators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):14–20, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Muto:1982:DTT

[MPR82] Arthur S. Muto, Bruce E.

Peetz, and Robert C. Rehner, Jr. Designing a ten-bit, twentymegasample-per-second analogto-digital converter system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (11):9–20, November 1982. CODEN HPJOAX. ISSN 0018-1153.

Maisenbacher:1991:LMB

[MR91] Bernd Maisenbacher and Wolfgang Reichert. A lightwave multimeter for basic fiber optic measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1):58–62, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Montijo:1993:AIA

[MR93] Allen Montijo and Kenneth Rush. Accuracy in interleaved ADC systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (5):38–46, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Mills:1980:EBL

[MRH80] N. A. Mills, H. C. Russell, and K. R. Henscheid. Enhanced BASIC language for a personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(7): 26–32, July 1980. CODEN HPJOAX. ISSN 0018-1153.

Mujtaba:1994:AMP

[MRH94a] M. Shahid Mujtaba, R. Ritter, and J. Harmon. Appendix I: Mathematics of production and material planning for the simple model. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6):109–111, December 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94dec/dec94109.pdf.

Mujtaba:1994:AIW

[MRH94b] M. Shahid Mujtaba, R. Ritter, and J. Harmon. Appendix II: Weekly event sequence. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6):111, December 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94dec/dec94109.pdf.

Mujtaba:1994:AIDa

[MRH94c] M. Shahid Mujtaba, R. Ritter, and J. Harmon. Appendix III: Details of part commonality experiments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (6):112, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94109.pdf.

Mujtaba:1994:AIDb

[MRH94d] M. Shahid Mujtaba, R. Ritter, and J. Harmon. Appendix IV: Details of explanations for experiments 0 and 1a. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(6):113, December 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94dec/dec94109.pdf.

Mujtaba:1994:EMS

[MRH94e] M. Shahid Mujtaba, R. Ritter, and J. Harmon. Enterprise modeling and simulation: Complex dynamic behavior of a simple model of manufacturing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):80– 113, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_80.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Morris:1983:PCF

[MRM83] Donald E. Morris, Anthony S. Ridolfo, and Donald L. Morris. A portable computer for field, office, or bench applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6): 3–4, 6–7, 9, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Muller:1993:ELD

[MRRS93] Emmerich Muller, Wolfgang Reichert, Clemens Ruck, and Rolf Steiner. External-cavity laser design and wavelength calibration. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):20–27, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Mack:1974:SBT

[MS74] N. Mack and L. E. Shar.

A supersystem for BASIC timesharing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(4):12–17, December 1974. CODEN HPJOAX. ISSN 0018-1153.

Moore:1980:COS

[MS80] C. E. Moore and D. J.

Sims. A compact optical system for portable distance and angle measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(9):14–16, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Matheson:1982:IOS

[MS82] W. Gordon Matheson and J. Marcus Stewart. An input/output system for a 1-MIPS computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(3):18–21, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Morris:1983:DSC

[MS83] Janet Elich Morris and Richard J. Simms, Jr. A decision support chartmaker. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (11):5–9, November 1983. CODEN HPJOAX. ISSN 0018-1153.

Munro:1994:ISM

[MS94] Andrew Munro and Ahmad H. Shekarabi. Imaging services in a multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):37– 43, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94apr/
37to43.pdf;   http://
www.hp.com/hpj/toc-4-
94.html.

Maier:1995:AOT

[MS95] Frank A. Maier and Harald Seeger. Automation of optical time-domain reflectometry measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (1):57–62, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_57.pdf;   http:

//www.hp.com/hpj/toc-
02-95.html.

Moore:1985:MTD

[MSB85] Glenn E. Moore, Jr., Richard S. Seymour, and Darrel R. Bloomquist. Manufacturing thin-film discs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(11):34–35, November 1985. CODEN HPJOAX. ISSN 0018-1153.

Mebane:1992:EFD

[MSCK92] Alfred Holt Mebane, IV, James R. Schmedake, Iue-Shuenn Chen, and Anne P. Kadonaga. Electronic and firmware design of the HP DesignJet drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):16–23, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Maisenbacher:1987:DAP

[MSS87] Bernd Maisenbacher, Siegmar Schmidt, and Michael Schlicker. Design approach for a programmable optical attenuator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(2):31–35, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Musch:1976:PSC

[MT76] B. E. Musch and R. B. Taggart. Portable scientific calculator has builtin printer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(3):9–18, November 1976. CODEN HPJOAX. ISSN 0018-1153.

Mierzwinski:1981:PS

[MT81] M. E. Mierzwinski and M. E. Terrien. 280-MHz production SAWR. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12): 15–16, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Mead:1983:CPG

[MTMP83] Sharon O. Mead, William R. Taylor, Kenneth A. Mintz, and Catherine M. Potter. A color presentation graphics workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(9):3–9, September 1983. CODEN HPJOAX. ISSN 0018-1153.

McIlvaine:1985:MDF

[MTWG85] George V. McIlvaine, Stephen L. Testardi, Daniel D. Wheeler, and Peter Gysling. Mechanical design of a family of high-speed impact line printers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(6):13–14, 17– 18, June 1985. CODEN HPJOAX. ISSN 0018-1153.

Munsey:1977:ADV

[Mun77] G. J. Munsey. APL data:

virtual workspaces and shared storage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(11):6– 10, July 1977. CODEN HPJOAX. ISSN 0018-1153.

Murillo:1982:DAM

[Mur82] R. D. Murillo. Dual ALU micromachine has powerful development tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(3): 5–6, March 1982. CODEN HPJOAX. ISSN 0018-1153.

Murphy:1992:CAD

[Mur92] Jerald B. Murphy. Comparing analog and digital oscilloscopes for troubleshooting. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):57–61 (or 57–59??), February 1992. CODEN HPJOAX. ISSN 0018-1153.

Murphy:1997:TMD

[Mur97] J. B. Murphy. Testing a mixed-signal design based on a single-chip microcontroller. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):10– 12, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a2.htm.

Muto:1974:TIC

[Mut74] A. S. Muto. Third input channel increases counter versatility. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 25(10):16– 18, June 1974. CODEN HPJOAX. ISSN 0018-1153.

McDermid:1977:HSV

[MVG77] J. E. McDermid, J. B. Vyduna, and J. M. Gorin. A high-speed system voltmeter for time-related measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(6):11–19, February 1977. CODEN HPJOAX. ISSN 0018-1153.

Merrill:1973:GSR

[MW73] H. L. Merrill and R. A. Warp. A go-anywhere strip-chart recorder that has laboratory accuracy.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25 (4):2–8, December 1973. CODEN HPJOAX. ISSN 0018-1153.

Marzalek:1978:DDC

[MW78] M. S. Marzalek and L. M.

Wheelwright. Developing the digital control system for the Model 8568 A spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(10):16– 17, 19–20, June 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Morgenthaler:1980:SBD

[MW80] M. P. Morgenthaler and L. Weber. Servo-controlled beam director provides major benefits (spectrophotometer). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(2):24–28, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Mooney:1994:LAT

[MW94] Matthew G. Mooney and Martha Grewe Wilson. Linear array transducers with improved image quality for vascular ultrasonic imaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(4):43–51, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94aug/
                                        aug94_43.pdf;   http:/
                                        /www.hp.com/hpj/toc-8-
                                        94.html.

Musch:1980:PPC

[MWC80] B. E. Musch, J. J. Wong, and D. R. Conklin. Powerful personal calculator system sets new standards. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (3):3–6, 8–12, March 1980. CODEN HPJOAX. ISSN 0018-1153.

McDougal:1995:STV

[MY95] Jay D. McDougal and William E. Young. Shortening the time to volume production of highperformance standard cell ASICs. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):91– 96, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95_91.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        02-95.html.

Myers:1996:FPP

[Mye96] Timothy F. Myers. A full-featured Pentium PCIbased notebook computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 38–44, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

May:1996:EDM

[MZ96] Elaine L. May and Barbara A. Zimmer. The evolutionary development model for software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): 39–45, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart4.
htm.

Nadig:1975:QHC

[Nad75] H. J. Nadig. A quiet, HP-IB compatible printer that listens to both ASCII and BCD. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(5):14–15, January 1975. CODEN HPJOAX. ISSN 0018-1153.

Nadig:1977:SAE

[Nad77] H. J. Nadig. Signature analysis-concepts, examples, and guidelines. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(9): 15–21, May 1977. CODEN HPJOAX. ISSN 0018-1153.

Naegeli:1979:IFM

[Nae79] A. H. Naegeli. IF filters for the 8901A Modulation Analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(11):10– 11, November 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Nakulski:1996:MOI

[Nak96] John P. Nakulski. Managed objects for internal application control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):74– 78, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Nease:1997:ESM

[NBZ^{+}97] L. A. Nease, K. M. Bresniher, C. J. Zacky, M. J. Greenside, and A. Sandoval. An entry-level server with multiple performance points. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3):73–81, June 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97jun/ju97a10.htm.

Navarro:1986:DMT

[NDS86] Nancy L. Navarro, Deepak V.

Desai, and Timothy C. Shafer. Dmi/3000: a move toward integrated communication. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):41– 48, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Neukermans:1981:MCS

[NE81] A. P. Neukermans and S. G. Eaton. Monte Carlo simulations for electron beam exposures. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(12):24– 25, December 1981. CO-DEN HPJOAX. ISSN 0018-1153.

Near:1968:CHM

[Nea68] C. W. Near. Computertesting the HP model 9100A calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(1):17–19, September 1968. CODEN HPJOAX. ISSN 0018-1153.

Negrete:1981:MNS

[Neg81] M. Negrete. Marco Negrete on structured VLSI design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):3–4, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Nelson:1979:RCT

[Nel79] D. L. Nelson. A remote computer troubleshooting facility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(9):13–14, 16, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Nelson:1982:IPA

[Nel82]  Loyd F. Nelson.  I/O

philosophy and architecture for instrument control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(5): 17–21, May 1982. CODEN HPJOAX.

Neuder:1984:NSI

[Neu84] David L. Neuder. New software increases capabilities of logic timing analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (2):32–38, February 1984.

CODEN HPJOAX. ISSN 0018-1153.

Neuder:1991:TVT

[Neu91] David L. Neuder. A test verification tool for C and C++ programs. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):83– 92, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Naegeli:1987:HSD

[NG87] Andrew H. Naegeli and Juan Grau. Hardware system design for a vector analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(11):6–17, December 1987. CODEN HPJOAX. ISSN 0018-1153.

Nishimoto:1992:IUI

[NGW92] Alvina Y. Nishimoto, William J.

Gray, and Barbara J. Williams. An interactive user interface for material requirements planning. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (4):103–110, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Narciso:1992:VID

[NH92] Steven J. Narciso and Gregory A. Hill. The VXIbus from an instrument designer’s perspective. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 15–23 (or 15–19??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Nathan:1994:TMS

[NH94] Connie Nathan and Barbara A. Hackbarth. Thermal management in supercritical fluid chromatography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(4):38–42, August 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94aug/
aug94_38.pdf;   http:/
/www.hp.com/hpj/toc-8-
94.html.

Nichols:1988:PDO

[Nic88] Douglas C. Nichols. Precision digitizing oscilloscope waveform analysis, display, and input/output. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):53–64, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Nicholson:1990:HFM

[Nic90] Dean B. Nicholson. Hexagonal ferrites for millimeterwave applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(5):59–61, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Nielsen:1985:HTP

[Nie85] Niels J. Nielsen. History of ThinkJet printhead development. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(5):4–10, May 1985. CODEN HPJOAX. ISSN 0018-1153.

Nielsen:1990:PSP

[Nie90] Kenneth R. Nielsen. Providing software protection capability for a CD-ROM drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):49–53, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Nikolaropoulos:1997:TSS

[Nik97] E. Nikolaropoulos. Testing safety-critical software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 89–94, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97jun/ju97a12.htm.

Nishi:1987:DVC

[Nis87] Yoshio Nishi. Direction of VLSI CMOS technology.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(6): 24–25, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Negus:1998:TTO

[NIWM98] Kevin J. Negus, Bryan T. Ingram, John D. Waters, and William J. McFarland. Technology trade-offs for wireless data applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1): ??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a3.htm.

Nielsen:1978:DEC

[NK78] J. R. Nielsen and D. S. Kaplan. Data entry and communications systems have network capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(7): 21–26, March 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Narimatsu:1984:PTS

[NK84] Yoh Narimatsu and Keiki Kanafuji. A parametric test system for accurate measurement of waferstage ICs. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(6):3–8, June 1984. CODEN HPJOAX. ISSN 0018-1153.

Nelson:1996:SKC

[NK96] Gerald J. Nelson and James K. Koch. Sidebar: K class power system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(1): ??, February 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/feb96/feba1sb.
html.

Nicholson:1990:PMB

[NML90] Dean B. Nicholson, Robert J.

Matreci, and Michael J. Levernier. 26.5-to-75-GHz preselected mixers based on magnetically tunable barium ferrite filters (for spectral analyzers). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(5):49–58, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Nairn:1980:AIC

[NMS80] J. H. Nairn, T. I. Mikkelsen, and D. J. Sweetser. Adding I/O capability to the HP-85. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(7):7–14, July 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Nakagawa:1995:GSW

[NMY95] Shigeru Nakagawa, Danny E. Mars, and Norihide Yamada. Generating shortwavelength light using a vertical-cavity laser structure. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):72– 75, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95_72.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        02-95.html.

Notess:1993:UIM

[Not93] Mark H. Notess. A user interface management system for HP-UX system administration applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3): 80–84, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Nygaard:1983:MAS

[NPG^{+}83] Richard A. Nygaard, Jr., Fredrick J. Palmer, Bryce S. Goodwin, Jr., Stan W. Bowlin, and Steven R. Williams. A modular analyzer for software analysis in the 64000 system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3):

16–23, March 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Nelson:1972:PIS

[NR72] G. E. Nelson and D. W.

Ricci. A practical interface system for electronic instruments. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(2):2–7, October 1972. CODEN HPJOAX. ISSN 0018-1153.

Nimori:1988:FSD

[NS88] Terrance K. Nimori and Lisa B. Stambaugh. Firmware system design for a frequency and time interval analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):13–21, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Nakajo:1989:SAS

[NSA89] Takeshi Nakajo, Katsuhiko Sasabuchi, and T. Akiyama. A structured approach to software defect analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2):50–56, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Noguchi:1970:MCA

[NSM70] H. Noguchi, T. Shimizu, and K. Maeda. Measuring capacitance automatically. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(9):14– 20, May 1970. CODEN HPJOAX. ISSN 0018-1153.

Nakagawa:1998:CCT

[NSMO98] Samuel O. Nakagawa, Dennis M. Sylvester, John G. McBride, and Soo-Young Oh. On-chip cross talk noise model for deepsubmicrometer ULSI interconnect. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):39–??, August 1998. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
98aug/tc-08-98.htm.

Nikolaropoulos:1997:ETL

[NSP97] E. Nikolaropoulos, J. Schwering, and A. Pirrung. Effective testing of localized software. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):109– 111, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97jun/ju97a15.htm.

Nieddu:1992:PMS

[NSV92] Giovanni Nieddu, Fernando M. Secco, and Alberto Vallerini. A performance monitoring system for digital telecommunications networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(5):89–99, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Nordman:1976:NCT

[NSW76] R. G. Nordman, R. L.

Smith, and L. A. Witkin. New CRT terminal has magnetic tape storage for expanded capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(9): 2–5, 7–11, 13–14, May 1976. CODEN HPJOAX. ISSN 0018-1153.

Neff:1975:TNP

[NT75] R. B. Neff and L. Tillman.

Three new pocket calculators: Smaller, less costly, more powerful. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(3):2–7, November 1975. CODEN HPJOAX. ISSN 0018-1153.

Nelson:1972:FGM

[NTA72] G. E. Nelson, P. L.

Thomas, and R. L. Atchley. Faster gain-phase measurements with new automatic 50Hz-to-13MHz network analyzers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(2):12–19, October 1972. CODEN HPJOAX. ISSN 0018-1153.

Novotny:1984:TIM

[NTH^{+}84] David R. Novotny, Jeffrey Tomberlin, Charles P. Hill, James P. Quan, Gordon A. Jensen, Jerry D. Morris, and K. R. Goldsmith. Transmission Impairment Measuring Set simplifies testing of complex voice and data circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(9):4–12, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Narayanan:1991:HVB

[NTJ^{+}91] Viswanathan S. Narayanan, Thomas Tom, Irvin R. Jones, Philip Garcia, and Christophe Grosther. The HP Vectra 486 basic I/O system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):83–92 (or 83–91??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Neumann:1987:DSG

[NVBH87] Uwe Neumann, Michael Vogt, Friedhelm Brilhaus, and Frank Husfeld. Digital signal generator combines digital and analog worlds. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(4): 4–12, April 1987. CODEN HPJOAX. ISSN 0018-1153.

Narimatsu:1981:VLI

[NYS81] Y. Narimatsu, K. Yagi, and T. Shimizu. A versatile low-frequency impedance analyzer with an integral tracking gain-phase meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (9):22–28, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Naganawa:1997:MHS

[NZ97] S. Naganawa and R. Z.

Zuo. A message handling system for B-ISDN usernetwork interface signaling test software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):51– 58, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a10.htm.

Olsson:1988:VUS

[OB88] Kjell A. Olsson and Mark Bergman. A virtual user simulation utility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(2):48– 53, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Oliver:1990:ODA

[OB90] Thomas C. Oliver and Mark J. Bianchi. Optical disk autochanger servomechanism design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6): 24–28, 30–34, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

OBrien:1993:FAM

[O’B93] Dennis P. O’Brien. A firmware architecture for multiple high-performance measurements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):17–30, December 1993. CODEN HPJOAX. ISSN 0018-1153.

OConnell:1986:RDU

[O’C86] Raymond G. O’Connell, Jr. The role of Doppler ultrasound in cardiac diagnosis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(6):20–25, June 1986. CODEN HPJOAX. ISSN 0018-1153.

Osecky:1984:DGM

[OGB84] Benjamin D. Osecky, Dennis D. Georg, and Robert J. Bury. The design of a general-purpose multipleprocessor system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(3):34– 38, March 1984. CODEN HPJOAX. ISSN 0018-1153.

URL https://www.hpl.
hp.com/hpjournal/pdfs/
IssuePDFs/1984-03.pdf.

Ogden:1978:ALS

[Ogd78] D. J. Ogden. Adapting the 1611A logic state analyzer to work with the F8 microprocessor family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(6):28–32, February 1978. CODEN HPJOAX. ISSN 0018-1153.

Ogden:1980:GMA

[Ogd80] D. J. Ogden. Generalpurpose module adapts dedicated logic state analyzer to almost any microprocessor. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(5):19– 24, May 1980. CODEN HPJOAX. ISSN 0018-1153.

Oh:1982:MDP

[Oh82a] Soo-Young Oh. MOS device and process design using computer simulations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (10):28–32, October 1982. CODEN HPJOAX. ISSN 0018-1153.

Oldfield:1982:WFO

[OH82b] Danny J. Oldfield and James F. Haley. Wideband, fast-writing oscilloscope solves difficult measurement problems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(4): 26–32 (or 26–31??), April 1982. CODEN HPJOAX. ISSN 0018-1153.

Ohme:1967:LE

[Ohm67] W. E. Ohme. Loudness evaluation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(3):2–14, November 1967. CODEN HPJOAX. ISSN 0018-1153.

OKonski:1993:RPG

[O’K93] Timothy C. O’Konski. Realizing productivity gains with C++. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(4):85–89, August 1993. CODEN HPJOAX. ISSN 0018-1153.

Otto:1983:PSS

[OL83] Ulrich Otto and Horst Link. Power supplies for the stimulus/response system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(7):28–31, July 1983. CODEN HPJOAX. ISSN 0018-1153.

Oliver:1971:DCC

[Oli71] B. M. Oliver. Distortion in complementary-pair class-B amplifiers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(6):11–16, February 1971. CODEN HPJOAX. ISSN 0018-1153.

OToole:1982:OIL

[OLR82] Michael M. O’Toole, E. David Liu, and Gary W. Ray. Optical IC lithography using trilayer resist. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(8):5–9, August 1982. CODEN HPJOAX. ISSN 0018-1153.

OConnell:1983:DSU

[OMD83] Raymond G. O’Connell, James R. Mniece, and Alwyn P. D’Sa. Display system for ultrasound images. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(12): 20–28, December 1983. CODEN HPJOAX. ISSN 0018-1153.

ONeill:1980:MTC

[O’N80] P. M. O’Neill. A monolithic thermal converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(5):12– ??, May 1980. CODEN HPJOAX. ISSN 0018-1153.

Orton:1996:DDC

[OR96] Kristann L. Orton and Paul R. Ritter. Design and development of the CDE 1.0 test suite. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(2):54– 61, April 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/apr96/apr96.
htm.

Oblad:1991:FHS

[ORB91] Roger P. Oblad, John R. Regazzi, and James E. Bossaller. A family of high-performance synthesized sweepers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):6–16, April 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Osada:1975:DHM

[OS75] K. Osada and J. Suehiro.

Digital high-capacitance measurements to one farad. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (6):10–16, February 1975. CODEN HPJOAX. ISSN 0018-1153.

Olson:1976:ELT

[OS76] W. E. Olson and R. B.

Smith. Electronics for the laser transducer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(6):7–18, February 1976. CODEN HPJOAX. ISSN 0018-1153.

Osborne:1968:HDM

[Osb68] T. E. Osborne. Hardware design of the model 9100A calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(1):10–13, September 1968. CODEN HPJOAX. ISSN 0018-1153.

Opfer:1985:TMD

[OSN^{+}85] James E. Opfer, Bruce F.

Spenner, Bangalore R. Natarajan, Richard A.

Baugh, Edward S. Murdock, Charles C. Morehouse, and David J. Bromley. Thin-film memory disc development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(11):4–10, November 1985. CODEN HPJOAX. ISSN 0018-1153.

Offermann:1975:API

[OST75] R. W. Offermann, S. E. Schultz, and C. R. Trimble. Active probes improve precision of time interval measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(2):11–16, October 1975. CODEN HPJOAX. ISSN 0018-1153.

Olsson:1988:HKL

[OY88] Kjell A. Olsson and Grace T. Yee. An HP-UX kernel load and measurement system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 54–60, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Pierrot:1997:HOS

[PA97] D. Pierrot and J.-P. Allegre. The HP OpenCall SS7 platform. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):58–64, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a7.htm.

Pratt:1970:MMS

[PAD70] R. E. Pratt, R. W. Austin, and A. Dethiefsen. Microcircuits for the microwave sweeper. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(3):9–16, November 1970. CODEN HPJOAX. ISSN 0018-1153.

Page:1971:TG

[Pag71] J. Page. Tracking generators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(1):10–11, September 1971. CODEN HPJOAX. ISSN 0018-1153.

Palombo:1986:SCS

[Pal86] Rosemarie Palombo. Software compatibility for series 200 and series 300 computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(9):22–27, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Palm:1996:RST

[Pal96] Richard C. Palm, Jr. Reducing setup time for printed circuit assembly. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47 (4):84–90, August 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart9.
htm.

Parker:1979:SSS

[Par79] K. P. Parker. Software simulator speeds digital board test generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3):13–16, 18– 19, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Patton:1983:HPC

[Pat83] D. B. Patton. The HP-75 production card recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6): 20–21, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Patton:1987:SCH

[Pat87] C. M. Patton. Symbolic computation for handheld calculators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(8):21–25, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Paxton:1970:NCH

[Pax70] D. Paxton. A new camera for high-speed oscilloscope recording. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(8):18–20, April 1970. CODEN HPJOAX. ISSN 0018-1153.

Poole:1976:DED

[PB76] J. S. Poole and L. Bilen.

Designing efficiency into a digital processor for an analytical instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(8):13– 19, April 1976. CODEN HPJOAX. ISSN 0018-1153.

Platt:1981:HSG

[PB81] D. L. Platt and D. T.

Borowski. A high-purity signal generator output section. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):22–26, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Pottinger:1982:CTW

[PB82] J. R. Pottinger and S. A.

Biddle. A combined tracking and white-noise generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(4):22–25, April 1982. CODEN HPJOAX. ISSN 0018-1153.

Pettis:1987:HPA

[PB87] Karl W. Pettis and William B. Buzbee. Hewlett–Packard Precision Architecture compiler performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(3):29– 35, March 1987. CODEN HPJOAX. ISSN 0018-1153.

Pon:1988:DDH

[PBD88] Rayman W. Pon, Steven C. Bird, and Patrick D. Deane. Digital design of a high-speed waveform recorder. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):32–38, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Patkay:1977:FDD

[PCW77] J.-P. D. Patkay, F. R. F. Chu, and H. A. M. Wiggers. Front-end design for digital signal analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(2):9–14, October 1977. CODEN HPJOAX. ISSN 0018-1153.

Parrish:1985:CWO

[PDW85] William M. Parrish, Eric B.

Decker, and Edwin G. Wong. Creative ways to obtain computer system debug tools. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(9):17–22, September 1985. CODEN HPJOAX. ISSN 0018-1153.

Pearo:1985:MDI

[Pea85] Thomas A. Pearo. Mechanical design of the integral PC: not just a desktop computer with a handle. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36 (10):18–22, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Pearson:1995:LCH

[Pea95]  Roger A. Pearson.  A

low-cost, high-performance PA-RISC workstation with built-in graphics, multimedia, and networking capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2):6–11, April 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95apr/apr95_6t.pdf;
http://www.hp.com/hpj/
toc-04-95.html.

Pecchio:1974:VBP

[Pec74]  S. Pecchio.  A versa-

tile bipolar power supply/ amplifier for lab and systems use. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(11):15–19, July 1974. CODEN HPJOAX. ISSN 0018-1153.

Pedersen:1986:PHT

[Ped86] Christen K. Pedersen. Program helps teach digital microwave radio fundamentals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4):40–46, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Peerboom:1973:DCS

[Pee73] R. Peerboom. Digitallycontrolled current sources for new ways of making automatic measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24 (12):14–19, August 1973. CODEN HPJOAX. ISSN 0018-1153.

Peery:1979:HBF

[Pee79a] D. L. Peery. HP 250 BASIC: a friendly, interactive, powerful system language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4):14–19, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Peery:1979:HIO

[Pee79b] D. L. Peery. HP 250 input/ output system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(4):11– 12, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Peregrino:1986:OSI

[Per86] L. Peregrino. Optimum solution for IF bandwidth and LO frequencies in a microwave counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(2):11–14, February 1986. CODEN HPJOAX. ISSN 0018-1153.

Perlmetter:1988:SCS

[Per88] Paul Q. Perlmetter. Small computer system interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):39–45, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Peterson:1968:LTS

[Pet68]  G. L. Peterson.  The

language of time sharing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20 (3):2–8, November 1968. CODEN HPJOAX. ISSN 0018-1153.

Pettit:1970:DO

[Pet70] J. Pettit. A d.c.-to-v.h.f. oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(5):2–8, January 1970. CODEN HPJOAX. ISSN 0018-1153.

Peterson:1974:THL

[Pet74] K. W. Peterson. Testing the HP-65 logic board. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(9): 18–20, May 1974. CODEN HPJOAX. ISSN 0018-1153.

Peterson:1994:SDR

[Pet94] John V. Peterson. Sidebar: Diagnosing and reporting problems in the multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2):18, April 1994. CODEN HPJOAX. ISSN 0018-1153.

Peters:1996:DBM

[Pet96] Gregory J. Peters. Determining a best-fit measurement server implementation for digital design team solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 22–29, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Pendergrass:1978:HLS

[PF78] N. A. Pendergrass and J. S. Farnbach. A highresolution, low-frequency spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(13): 2–9, 11–13, September 1978. CODEN HPJOAX. ISSN 0018-1153.

Page:1988:WRT

[PF88] Richard W. Page and Allen S. Foster. Waveform reconstruction techniques for precision digitizing oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):26–31, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Pierce:1973:CDM

[PGW73] W. C. Pierce, Jr., J. S.

Gallo, and W. T. Walker. Computer-aided design of modular power supplies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24 (6):4–19, February 1973. CODEN HPJOAX. ISSN 0018-1153.

Plumb:1970:MNL

[PH70] J. Plumb and J. Holtzinger. Measuring noise and level on international telephone systems. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(4):13–16, December 1970. CODEN HPJOAX. ISSN 0018-1153.

Park:1996:FRM

[PH96] Rin Park and Paul D. Hale. Frequency response measurement of digital communications analyzer plugin modules. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Patterson:1977:SPS

[PHK77] M. L. Patterson, R. D. Haselby, and R. M. Kemplin. Speed, precision, and smoothness characterize four-color plotter pen drive system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(1):13–19, September 1977. CODEN HPJOAX. ISSN 0018-1153.

Pendse:1996:RSB

[PHL96] Rajendra D. Pendse, Rita N.

Horner, and Fan Kee Loh. Radially staggered bonding technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6): ??, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/dec96/tc-12-
96.htm.

Pierce:1975:PFM

[Pie75] R. B. Pierce. Packaging for function, manufacturability, and service. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(10):22– 24, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Piety:1987:IDT

[Pie87] Robert A. Piety. Intrabuilding data transmission using power-line wiring. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(5): 35–40, May 1987. CODEN HPJOAX. ISSN 0018-1153.

Pirrung:1997:HPL

[Pir97] A. Pirrung. A high-level programming language for testing complex safetycritical systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):95– 102, June 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97jun/ju97a13.htm.

Pitt:1997:RC

[Pit97] Daniel A. Pitt. Residential communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97dec/de97a2.htm.

Pannach:1975:MWG

[PK75] A. Pannach and W. Kappler. A multichannel word generator for testing digital components and systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (12):17–24, August 1975. CODEN HPJOAX. ISSN 0018-1153.

Panek:1984:CDF

[PK84] Charles R. Panek and Steven F. Kator. Custom digital filters for dynamic signal analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(12):28– 36, December 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Patterson:1981:DLD

[PL81] M. L. Patterson and G. W.

Lynch. Development of a large drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (11):3–7, November 1981. CODEN HPJOAX. ISSN 0018-1153.

Pritchard:1984:CIC

[PL84] Thomas B. Pritchard and David S. Lee. Custom IC controls dot-matrix impact printers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(11):33– 36, November 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Plitschka:1989:TLE

[Pli89] Rainer Plitschka. Transmission line effects in testing high-speed devices with a high-performance test system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):58–68 (or 58–67??), December 1989. CODEN HPJOAX. ISSN 0018-1153.

Peetz:1982:MWR

[PMN82] Bruce E. Peetz, Arthur S. Muto, and J. Martin Neil. Measuring waveform recorder performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(11): 21–29, November 1982. CODEN HPJOAX. ISSN 0018-1153.

Page:1988:ASR

[PN88] Richard W. Page and Nancy W. Nelson. Adaptive Sample Rate: a firstgeneration automatic time base. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):23–25, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Pontis:1980:DLF

[Pon80a] G. D. Pontis. Design for a low-distortion, fast-settling source. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(8):10–11, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Pontis:1980:FSO

[Pon80b] G. D. Pontis. Floating a source output. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(8):12–17, August 1980. CODEN HPJOAX. ISSN 0018-1153.

Popp:1983:PVE

[Pop83] Richard L. Popp. Physician’S view of echocardiographic imaging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(10):13– 16, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Porter:1980:LPA

[Por80] F. J. Porter. Light pen aids user interaction with display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(12):10–19, December 1980. CODEN HPJOAX. ISSN 0018-1153.

Poulter:1968:PTC

[Pou68] T. C. Poulter, Jr. A practical time-shared computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 19(11):2–7, July 1968. CODEN HPJOAX. ISSN 0018-1153.

Perkinson:1969:PDC

[PP69] J. C. Perkinson and W. C.

Pierce, Jr. Precision dc current sources. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(13):15– 20, September 1969. CO-DEN HPJOAX. ISSN 0018-1153.

Pless:1991:LMF

[PPJ91] Wilfried Pless, Michael Pott, and Robert Jahn. Lightwave multimeter firmware design. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(1):77–87 (or 77–83??), February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Parkhurst:1994:CHD

[PPMW94] Anthony D. Parkhurst, Padmanabhan, Steven D. Mueller, and Kirt A. Winter. Connectivity of the HP DeskJet 1200C printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (1):85–97, February 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94feb/
85to97.pdf.

Pratt:1975:VMP

[Pra75] R. E. Pratt. Very-low-level microwave power measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(2):8–10, October 1975. CODEN HPJOAX. ISSN 0018-1153.

Pratt:1980:PCR

[Pra80] W. C. Pratt. A precision color raster-scan display for graphics applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(12):19–24, December 1980. CODEN HPJOAX. ISSN 0018-1153.

Priebe:1976:MLC

[Pri76]  D. Priebe.  Multifam-

ily logic clip shows all pin states simultaneously. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28 (4):18–20, December 1976. CODEN HPJOAX. ISSN 0018-1153.

Primmer:1996:IFC

[Pri96] Meryem Primmer. An introduction to Fibre Channel. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):94–98, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/oct96/oc96a11.htm.

Prufer:1992:NAP

[Pru92] Rona J. Prufer. Network Advisor protocol analysis: decodes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):34–40, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Parker:1981:ICS

[PS81] R. A. Parker and R. J. Solomon. In-service CO/ sub 2/ sensor calibration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (9):16–18, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Purnaveja:1983:CCM

[PS83] Chayaboon Purnaveja and Janet Swift. Convenient creation and manipulation of presentation aids. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(11): 13–17, November 1983. CODEN HPJOAX. ISSN 0018-1153.

Perkins:1983:CTT

[PSA83] Michael R. Perkins, Susan Snitzer, and Charles W. Andrews. A computeraided test and tracking system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(9):25–28, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Palmer:1988:IMD

[PSS^{+}88] Donald J. Palmer, John Stoffel, Ronald J. Selensky, Peter C. Morris, M. Beth Heffernan, and Mark S. Hickman. Ink and media development for the HP Paintjet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(4):45–50, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Park:1990:HMA

[PT90] Collin Y. W. Park and Bruce J. Talley. HP Manufacturing Automation Protocol 3.0. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4):6–10, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Petersen:1992:DPM

[PT92] David M. Petersen and Chuong Ta. DesignJet plotter mechanical architecture development process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (6):32–34, December 1992.

CODEN HPJOAX. ISSN 0018-1153.

Pugh:1990:IOC

[Pug90] Rex A. Pugh. Implementation of the OSI Class 4 Transport Layer Protocol in the HP OSI Express card. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):36–45, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Pecenco:1982:HBI

[PW82] Irene V. Pecenco and Albert S. Wang. A highperformance bipolar integrated circuit process. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (9):27–29, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Packard:1989:EAT

[PW89] Barbara B. Packard and Charles H. Whelan. An extensible agent task language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):38–42, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Quint:1990:MRC

[QAKP90] David W. Quint, Asad Aziz, Ravi Kaw, and Frank J. Perezalonso. Measurement of R, L, and C parameters in VLSI packages. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (5):73–77, October 1990. CODEN HPJOAX. ISSN 0018-1153.

Quick:1983:HLD

[QH83] Roger D. Quick and Steven L. Harper. HP-IL: a low-cost digital interface for portable applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(1): 3–7, 9–10, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Quick:1980:EPP

[QM80] R. D. Quick and D. L. Morris. Evolutionary printer provides significantly better performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(3):15–20, March 1980. CODEN HPJOAX. ISSN 0018-1153.

Quenelle:1976:NLP

[Que76] R. C. Quenelle. New logic probe troubleshoots many logic families. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(4):9–11, December 1976. CODEN HPJOAX. ISSN 0018-1153.

Quenelle:1983:NML

[QW83] Robert C. Quenelle and Lawrence J. Wuerz. A new microcomputer-controlled laser dimensional measurement and analysis system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (4):3–4, 6–13, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Regelson:1985:HTI

[RA85] Elaine C. Regelson and Roy E. Anderson. HP techwriter: Illustrated documents for engineers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(2): 4–10 (or 4–9??), February 1985. CODEN HPJOAX. ISSN 0018-1153.

Radermacher:1987:HOC

[Rad87] Wilhelm Radermacher. A high-precision optical connector for optical test and instrumentation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(2):28–30, February 1987. CODEN HPJOAX. ISSN 0018-1153.

Rafnel:1993:TAE

[Raf93] Bruce A. Rafnel. A transaction approach to error handling. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):71–77, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Raje:1995:FII

[Raj95] Prasad Raje. A framework for insight into the impact of interconnect on 0.35μ m VLSI performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (1):97–104, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_97.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Ramakrishnan:1992:SPP

[Ram92] S. Ramakrishnan. Software pipelining in PA-RISC compilers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(3):39– 45, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Randel:1988:CDR

[Ran88] Annette Randel. Crash detection and recovery in a discless HP-UX system.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (5):27–32, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Rasaratnam:1987:AWN

[Ras87] Dayananda K. Rasaratnam. An accurate wideband noise generator and a high-stability reference source. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(7):30–32, 34– 36, July 1987. CODEN HPJOAX. ISSN 0018-1153.

Rhodes-Burke:1982:RSA

[RB82] Robert Rhodes-Burke. Retrofitting for signature analysis simplified. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(1):9–16, January 1982. CODEN HPJOAX. ISSN 0018-1153.

Ripert:1978:EDC

[RBB78] J. A. Ripert, D. C. Berthier, and M. E. Bernard. An easy-to-use data capture terminal for industrial operations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(11): 19–23, July 1978. CODEN HPJOAX. ISSN 0018-1153.

Royce:1979:SEE

[RC79] W. G. Royce and P. Chu.

Simple, efficient electronics for a low cost X-Y plotter. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(2):14–18, February 1979. CODEN HPJOAX. ISSN 0018-1153.

Robinson:1998:NAC

[RC98] Harry J. Robinson and Sankar L. Chakrabarti. New approaches to creating and testing internationalized software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 50(1):29–??, November 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98nov/tc-11-98.htm.

Riedel:1975:DFG

[RD75] R. J. Riedel and D. D.

Danielson. The dual function generator: A source of a wide variety of test signals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(7):18–24, March 1975. CODEN HPJOAX. ISSN 0018-1153.

Rush:1986:HPS

[REB86] Kenneth Rush, William H.

Escovitz, and Arnold S.

Berger. High-performance probe system for a 1-GHz digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4): 11–19, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Regelson:1988:DPC

[Reg88] Elaine C. Regelson. Developing a printed circuit board design system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (1):65–67, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Reh:1984:MST

[Reh84] Teresa L. Reh. Master/ slave TIMS operation increases productivity. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(9): 13–15, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Rehder:1988:ARC

[Reh88] Wulf D. Rehder. An availability and reliability calculation tool for computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):26–29, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Rehder:1997:OCE

[Reh97]  W. Rehder.  Object-

oriented customer education. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(1):96–102, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a12.htm.

Reiche:1991:CMS

[Rei91] Martin Reiche. Component Monitoring System software architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(4):13–18, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Roos:1978:ADS

[REOB78] M. D. Roos, J. H. Egbert, R. P. Oblad, and J. T. Barr. Add-on digital signal processing enhances the performance of network and spectrum analyzers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(5):17–24, January 1978. CODEN HPJOAX. ISSN 0018-1153.

Ressmeyer:1983:LPC

[Res83] James J. Ressmeyer. Lowcost and portability come to data acquisition/control products. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(2):10–12, 15–16, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Riggen:1972:AGD

[RF72] J. Riggen and D. Fogg. An agile graphic display device. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(8):18–23, April 1972. CODEN HPJOAX. ISSN 0018-1153.

Rytand:1969:NAR

[RG69] W. A. Rytand and D. R. Gildea. Network analysis in the range 100 kHz to 110 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(4):2–11, December 1969. CODEN HPJOAX. ISSN 0018-1153.

Riggins:1974:TP

[RH74] C. C. Riggins and R. L. Hammons. Testing the 21MX processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(2):10–11, October 1974. CODEN HPJOAX. ISSN 0018-1153.

Roesner:1992:MDH

[RH92] Arlen L. Roesner and John P. Hoppal. Mechanical design of the HP 9000 models 720 and 730 workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(4):43–48, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Roland:1983:TRM

[RHK^{+}83] James P. Roland, Norman E. Hendrickson, Daniel D. Kessler, Donald E. Novy, Jr., and David W. Quint. Two-layer refractory metal IC process. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(8):30–32, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Riebesell:1979:DPG

[RHM79] G. Riebesell, U. Hubner, and B. Moravek. A digital pattern generator for functional testing of busoriented digital systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30 (8):20–25, August 1979. CODEN HPJOAX. ISSN 0018-1153.

Rhodes:1988:MDH

[Rho88] John D. Rhodes. Managing the development of the HP DeskJet Printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(5):51–54, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Richards:1976:MDC

[Ric76] A. J. Richards. Mini data cartridge: a convincing alternative for low-cost, removable storage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(9):6–7, May 1976. CODEN HPJOAX. ISSN 0018-1153.

Richard:1981:PPF

[Ric81a] S. H. Richard. Programming for productivity: factory data collection software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(1):25–29, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Richardson:1981:VHS

[Ric81b] J. W. Richardson. Verifying high spectral purity and level accuracy in production. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):30–32, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Richardson:1991:BMI

[Ric91] Ronald F. Richardson.

Building and managing an integrated project support environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(3):90– 96, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Richard:1994:VSM

[Ric94] Craig S. Richard. Video support in a multimedia environment. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):68– 70, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94apr/
68to70.pdf;   http://
www.hp.com/hpj/toc-4-
94.html.

Rice:1997:NIP

[Ric97]  H. W. Rice.  Next-

generation inkjet printhead drive electronics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(3):38– 42, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97jun/ju97a5.htm.

Ridolfo:1990:HON

[Rid90] Anthony S. Ridolfo. HP OpenView network management. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):51–53, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Riedel:1994:HAA

[Rie94] Ronald J. Riedel. Highthroughput amplifier and analog-to-digital converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (5):16–20, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94oct/
                                        oct94_16.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        10-94.html.

Riggins:1977:EDC

[Rig77] C. C. Riggins. E-series doubles 21MX computer performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(7):18–19, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Riley:1979:NTF

[Ril79] R. B. Riley. A new type of FM demodulator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(11):13– ??, November 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Riper:1987:MSH

[Rip87] Richard W. Riper. Manufacturing state-of-the-art handheld calculators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(10): 24–27, October 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Riper:1991:MHC

[Rip91] Richard W. Riper. Manufacturing the HP 48SX calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3):40–43, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Risley:1982:OMS

[Ris82] William B. Risley. Oscilloscope Measurement System is programmable and autoranging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(9):3–4, September 1982. CODEN HPJOAX. ISSN 0018-1153.

Robinson:1987:MVH

[RJH^{+}87] Craig S. Robinson, Leith Johnson, Robert J. Horning, Russell W. Mason, Mark A. Ludwig, Howell R. Felsenthal, Thomas O. Meyer, and Thomas V. Spencer. A midrange VLSI Hewlett–Packard Precision Architecture computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (9):26–34, September 1987. CODEN HPJOAX. ISSN 0018-1153.

Raak:1994:BSC

[RK94a] Gerald I. Raak and Christopher P. J. Kelly. Builtin self-test and calibration for a scanning analog-to-digital converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):25– 29, October 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94oct/
oct94_25.pdf;   http:
//www.hp.com/hpj/toc-
10-94.html.

Raak:1994:SHC

[RK94b] Gerald I. Raak and Christopher P. J. Kelly. Sidebar: A hierarchy of calibration commands. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(5):28, October 1994. CODEN HPJOAX. ISSN 0018-1153.

Reusser:1986:HOS

[RL86] Gertrude G. Reusser and Donald C. Loughry. Hewlett– Packard and the Open Systems Interconnection Reference Model. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(10):4–5, October 1986. CODEN HPJOAX. ISSN 0018-1153.

Reed:1992:AME

[RL92] Douglas J. Reed and Terry M. Lambright. Adhesive material and equipment selection for the HP DeskJet 500C/DeskWriter C color print cartridge. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (4):84–86, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Ranganath:1995:MQW

[RLB^{+}95] Tirumala R. Ranganath, Michael J. Ludowise, Patricia A. Beck, Dennis J. Derickson, William H. Perez, Tim L. Bagwell, and David M. Braun. Multiquantum-well ridge waveguide lasers for tunable external-cavity sources. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (1):20–26, February 1995.

CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95_20.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        02-95.html.

Reese:1988:MHP

[RM88] Paul S. Reese and Mark E. Mayotte. Managing HP PCDS with the Design System Manager. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(1):71–76, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Robrish:1995:DPN

[RMVT95] Peter R. Robrish, Christopher J. Madden, Rory L. VanTuyl, and William R. Trutna, Jr. Design and performance of a narrowband VCO at 282 THz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (1):63–66, February 1995. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95feb/
                                        feb95_63.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        02-95.html.

Rohdin:1998:MGL

[RNR^{+}98] Hans Rohdin, Avelina Nagy, Virginia Robbins, Chung-Yi Su, Arlene S. Wakita, Judith Seeger, Tony Hwang, Patrick Chye, Paul E. Gregory, Sandeep R. Bahl, Forrest G. Kellert, Lawrence G. Studebaker, Donald C. D’Avanso, and Sigurd Johnsen. 0.1-micrometer gate-length AlInAs/GaInAs/ GaAs MODFET MMIC process for applications in high-speed wireless communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(1):??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a4.htm.

Rush:1986:DAS

[RO86] Kenneth Rush and Danny J.

Oldfield. A data acquisition system for a 1-GHz digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4): 4–11, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Rahmat:1997:OIG

[RO97] K. Rahmat and Soo-Young Oh. Optimization of interconnect geometry for high-performance microprocessors. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):78– 83, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a13.htm.

Robinson:1979:AME

[Rob79] C. Robinson, Jr. Adapting the Multiprogramming Executive to a new hardware environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(9):7–8, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Roberts:1982:ITS

[Rob82a] Richard J. Roberts. An integrated test set for microwave radio link baseband analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(4):3–7, April 1982. CODEN HPJOAX. ISSN 0018-1153.

Robertson:1982:IAP

[Rob82b] R. A. Robertson. An improved AC power switch. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(12): 34–40, December 1982. CODEN HPJOAX. ISSN 0018-1153.

Roberts:1992:IED

[Rob92] D. C. Roberts. Integrating an electronic dictionary into a natural language processing system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(3):54– 65, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Roesner:1995:PDM

[Roe95] Arlen L. Roesner. Product design of the Model 712 workstation and external peripherals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(2):75– 78, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_75.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Roeca:1996:PAA

[Roe96] Jeffrey E. Roeca. Prototype analyzer architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 15–21, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Rose:1994:DMP

[ROKH94] Gary P. Rose, Jeffery T.

Oesterle, Joseph E. Kasper, and Robert J. Hammond. Development of a multimedia product for HP workstations. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(2):6– 9, April 1994. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/94apr/6to9.pdf;
                                        http://www.hp.com/hpj/
                                        toc-4-94.html.

Rometsch:1991:VSN

[Rom91] Rainer Rometsch. A very small noninvasive blood pressure measurement device. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):25–26, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Rosenberg:1988:HES

[Ros88] Steven T. Rosenberg. HP-RL: an expert systems language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):57–65, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Roth:1970:DFA

[Rot70] P. R. Roth. Digital Fourier analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(10):2–9, June 1970. CODEN HPJOAX. ISSN 0018-1153.

Rowland:1983:WGC

[Row83] F. W. Rowland. What is gas chromatography? Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (1):32–33, January 1983. CODEN HPJOAX. ISSN 0018-1153.

Rowe:1986:MMP

[Row86] Mark S. Rowe. Memory management for portable computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(7):21– 25, July 1986. CODEN HPJOAX. ISSN 0018-1153.

Roy:1975:HRS

[Roy75]  J.-C. Roy.  A high-

resolution raster scan display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(10):11–15, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Roy:1981:SIV

[Roy81] J.-C. Roy. A silicon-on-sapphire integrated video controller. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(3):16–19, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Rivoir:1988:DPC

[RP88] Jochen Rivoir and Wilfried Pless. Data processing in the correlating optical time-domain reflectometer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(6):29–34, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Rex:1969:CSA

[RR69] R. L. Rex and G. T. Roberts. Correlation, signal averaging, and probability analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(3):2–8, November 1969. CODEN HPJOAX. ISSN 0018-1153.

Richey:1983:EM

[RR83] David B. Richey and John P. Romano. Emulators for 16-bit microprocessors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3):31–38, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Read:1988:SIS

[RR88] Sherry L. Read and Timothy R. C. Read. Statistical issues in setting product specifications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(3):6–11, June 1988. CODEN HPJOAX. ISSN 0018-1153.

Ruska:1985:FST

[RRC^{+}85] David W. Ruska, Virgil K. Russon, Bradfred Culp, Alan J. Richards, and John A. Ruf. Firmware for a streaming tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(3): 29–31, March 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Rytting:1970:SAN

[RS70] D. K. Rytting and S. N.

Sanders. A system for automatic network analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(6):2–10, February 1970. CODEN HPJOAX. ISSN 0018-1153.

Ricci:1975:PTI

[RS75] D. W. Ricci and P. S.

Stone. Putting together instrumentation systems at minimum cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(5):5–11, January 1975. CODEN HPJOAX. ISSN 0018-1153.

Rivoir:1991:DSH

[RSM91] Jochen Rivoir, Horst Schweikardt, and Emmerich Müller. Design of a series of high-performance lightwave power sensor modules. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):63–69, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Rothschild:1986:LAI

[RSW^{+}86] Charles J. Rothschild, III, Robert C. Sismilich, William T. Walker, A. Levine, D. J. Weller, G. Kononenko, and D. Schlesinger. Lowcost automated instruments for personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(5):4–11 (or 4– 10??), May 1986. CODEN HPJOAX. ISSN 0018-1153.

Raynham:1991:EC

[RTD91] Michael B. Raynham, Douglas M. Thom, and T. Dowden. The EISA connector. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):73–77 (or 73–74??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Reid-Urquhart:1976:SR

[RU76] J. Reid-Urquhart. A signallevel reference. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(2):10–??, October 1976. CODEN HPJOAX. ISSN 0018-1153.

Rubinstein:1986:HSM

[Rub86] Jonathan J. Rubinstein.

High-performance SPU for a modular workstation family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(9):12–16, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Rueter:1980:MAC

[Rue80]  J. M. Rueter.  Multi-

processor architecture and communications for patient monitoring. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):15– 18, November 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Russell:1992:ASF

[Rus92] D. D. Russell. Application of spatial frequency methods to evaluation of printed images. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3):68–75, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Rush:1993:SPP

[Rus93] Kenneth Rush. A study of pulse parameter accuracy in real-time digitizing oscilloscope measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):47–50, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Ray:1970:SAN

[RW70] W. A. Ray and W. W. Williams. Software for the automatic network analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 21(6):11–15, February 1970. CODEN HPJOAX. ISSN 0018-1153.

Rude:1976:LTS

[RW76] A. F. Rude and M. J. Ward. Laser transducer systems for high-accuracy machine positioning. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(6):2–6, February 1976. CODEN HPJOAX. ISSN 0018-1153.

Rissman:1981:PEC

[RW81] P. Rissman and M. P. C. Watts. Proximity effect corrections by means of processing: theory and applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(12):21–23, 25–27, December 1981. CODEN HPJOAX. ISSN 0018-1153.

Ring:1997:NSQ

[RWT97] William S. Ring, Simon J. Wrathall, and Adrian J. Taylor. 1300nm strained quantum well lasers for fiber-optic communications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97dec/de97a8.htm.

Ritzmann:1995:UIC

[RWW95] Alwin Ritzmann, Klaus Witt, and H. Wiederoder. A user interface for capillary electrophoresis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 44–49, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_44.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Ryan:1995:DDB

[Rya95] R. E. Ryan. Design and development of a 120-MHz bus interface block using standard cells and automatic place and route tools. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 92–95, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95jun/
                                        jun95_92.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        06-95.html.

Rytand:1977:OIL

[Ryt77] W. A. Rytand. OPNODE: interactive linear circuit design and optimization. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(7): 28–31, March 1977. CO-DEN HPJOAX. ISSN 0018-1153.

Stettiner:1980:HRT

[SA80] R. L. Stettiner and G. L. Adleman. High-speed raster technique provides flexible display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(11):11– 15, November 1980. CO-DEN HPJOAX. ISSN 0018-1153.

Sorden:1982:WRH

[SA82] James L. Sorden and Mark S. Allen. Waveform recording with a highdynamic-performance instrument. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(11):3–9, November 1982. CODEN HPJOAX. ISSN 0018-1153.

Sayed:1991:HMS

[SA91] Mohamed M. Sayed and Giovonnae F. Anderson. 50-to-110-GHz high-performance millimeter-wave source modules. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(2):50–64, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Safai:1996:CTI

[Saf96]  Fereydoon Safai.  Cy-

cle time improvement for Fuji IP2 pick-and-place machines. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):80– 83, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart8.
htm.

Sakakihara:1978:DS

[Sak78] P. M. Sakakihara. Distributed systems/3000. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(7): 7–14, March 1978. CODEN HPJOAX. ISSN 0018-1153.

Shannon:1973:SSV

[SAMS73] R. M. Shannon, K. L. Astrof, M. S. Marzalek, and L. C. Sanders. A solid state VHF signal generator for today’s exacting requirements. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(6):2–13, February 1973. CODEN HPJOAX. ISSN 0018-1153.

Sanders:1972:AR

[San72] S. N. Sanders. Automating the 10-MHz-to-18-GHz receiver. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(6):10–13, February 1972. CODEN HPJOAX. ISSN 0018-1153.

Sanchez:1986:ACL

[San86] Jorge Sanchez. Analog channel for a low-frequency waveform recorder. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(1):

32–35, January 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Santhanam:1992:RRP

[San92] V. Santhanam. Register reassociation in PA-RISC compilers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3):33–38, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Saponas:1977:FMA

[Sap77] T. A. Saponas. Firmware of a microprocessor analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(5):12–15, January 1977. CODEN HPJOAX. ISSN 0018-1153.

Spencer:1998:ADT

[SAS98] T. M. Spencer, P. M. Anderson, and D. Sweetser. Advanced display technologies on HP-UX workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(2):46–??, May 1998. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
98may/tc-05-98.htm.

Sayed:1980:FCH

[Say80] M. M. Sayed. 40-GHz frequency converter heads.

Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4): 14–19, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Smith:1980:IDM

[SB80] D. E. Smith and T. L. Brown. Industrial distance meter applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(6): 11–19, June 1980. CODEN HPJOAX. ISSN 0018-1153.

Shenoy:1983:ITI

[SB83] A. K. Shenoy and C. M. Bircher. Integrated tools improve programmer productivity. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(2):30–36, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Schmidt:1984:CGD

[SB84] Daniel G. Schmidt and M. K. Bowen. Color graphics display for an engineering workstation (and detaching the keyboard). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (5):12, 14–15, May 1984. CODEN HPJOAX. ISSN 0018-1153.

Simon:1985:VDH

[SBB85] Jean-Jacques Simon, F. Barbut, and R. Brabant. VLSI design in the HP 2392A terminal. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4):9–16, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Scott:1997:FSP

[SBK^{+}97] A. P. Scott, K. P. Burkhart, A. Kumar, R. M. Blumberg, and G. L. Ranson. Four-way superscalar PA-RISC processors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):8–15, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/four_
way_superscalar.pdf;
http://www.hp.com/hpj/
97aug/au97a1.htm.

Siewell:1985:TOP

[SBM85] Gary L. Siewell, William R.

Boucher, and Paul H. Mc-Clelland. The ThinkJet orifice plate: a part with many functions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(5):33– 37, May 1985. CODEN HPJOAX. ISSN 0018-1153.

Snyder:1983:UIQ

[SC83a] Richard A. Snyder and Richard J. Conrad. Ultrasound image quality. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (10):34–38, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Swain:1983:NFM

[SC83b] Howard L. Swain and Rick M. Cox. Noise figure meter sets records for accuracy, repeatability, and convenience. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(4):23–25, 27, 29–34, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Snook:1989:DHP

[SC89] Douglas R. Snook and G. Stephen Curtis. Design of a high-performance pulse modulation system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (5):51–59, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Schmidhauser:1970:MNT

[Sch70] R. Schmidhauser. Measuring nanosecond time intervals by averaging. what kind of resolution can you get and how do you get it? what about accuracy? Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(8):11–13, April 1970. CODEN HPJOAX. ISSN 0018-1153.

Schneider:1973:HAM

[Sch73] R. F. Schneider. A highperformance automatic microwave counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(8):2–9, April 1973. CODEN HPJOAX. ISSN 0018-1153.

Schrenker:1975:FCH

[Sch75] H. Schrenker. Flow control in high-pressure liquid chromatography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 27(2):17–24, October 1975. CODEN HPJOAX. ISSN 0018-1153.

Schultz:1976:RMB

[Sch76] J. T. Schultz. Real-time multi-user BASIC. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(5): 19–24, January 1976. CO-DEN HPJOAX. ISSN 0018-1153.

Schmidt:1978:DPD

[Sch78a] L. A. Schmidt. Designing programmable digital filters for LSI implementation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(13):15–23, September 1978. CODEN HPJOAX. ISSN 0018-1153.

Schwager:1978:HDS

[Sch78b]  A. O. Schwager.  The

Hewlett–Packard distributed system network. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(7):2–6, March 1978. CODEN HPJOAX. ISSN 0018-1153.

Schlotzhauer:1979:USA

[Sch79] E. O. Schlotzhauer. Useroriented software for an automatic circuit-board tester. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3):22–27, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Schneider:1980:MCM

[Sch80]  R. F. Schneider.  Mi-

crowave counter measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(4):23–25, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Schrenker:1984:MPP

[Sch84] Helge Schrenker. Mobile phase preheater ensures precise control of LC column temperature. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(4):24– 26, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Schlater:1986:WGD

[Sch86] Rodney T. Schlater. Waveform graphics for a 1-GHz digitizing oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4): 20–25, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Schweikardt:1987:ATO

[Sch87] Horst Schweikardt. An accurate two-channel optical average power meter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(2): 8–12 (or 8–11??), February 1987. CODEN HPJOAX. ISSN 0018-1153.

Scharrer:1993:MDO

[Sch93a] John A. Scharrer. An 8-Gigasample-per-Second modular digitizing oscilloscope system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(5):6–10, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Schmidt:1993:HOR

[Sch93b] Siegmar Schmidt. Highperformance optical return loss measurement. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):79–82, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Schroath:1993:CMS

[Sch93c] Leonard T. Schroath. Configuration management for software tests. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(3):53– 59, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Schneider:1995:SIH

[Sch95] Werner Schneider. Sample injection in HP CE (Hewlett–Packard capillary electrophoresis). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 32–35, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_32.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Schnaible:1996:NDL

[Sch96] Mark P. Schnaible. A normalized data library for prototype analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 30–37, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Schwering:1997:ATE

[Sch97] Jörg Schwering. An automated test evaluation tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 103–108, June 1997. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97jun/
ju97a14.htm.

Scherer:1981:LRS

[SCI^{+}81] D. Scherer, B. S. Chan, F. H. Ives, W. J. Crilly, Jr., and D. W. Mathiesen. Low-noise RF signal generator design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(2):12–22, February 1981. CODEN HPJOAX. ISSN 0018-1153.

Small:1978:ELL

[SD78] C. T. Small and A. J. DeVilbiss. Entry level logic state analyzer has highlevel capability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(6):21–27, February 1978. CODEN HPJOAX. ISSN 0018-1153.

Szeto:1982:DMS

[SD82] Christina M. Szeto and Michael C. Detro. Display and mass storage for waveform recording. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(11):34– 36, November 1982. CO-DEN HPJOAX. ISSN 0018-1153.

Swedberg:1995:ICT

[SD95] Sally A. Swedberg and Monika Dittmann. The impact of column technology on protein analysis by capillary electrophoresis: Surface coatings and analytical approaches for assessment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3):57–61, June 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95jun/jun95_57.pdf;
http://www.hp.com/hpj/
toc-06-95.html.

Scheffer:1981:DSV

[SDA81] L. K. Scheffer, R. I. Dowell, and R. M. Apte. Design and simulation of VLSI circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):12–15, 17– 18, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Steranka:1988:RAL

[SDC^{+}88] Frank M. Steranka, Dennis C. DeFevere, Michael D. Camras, Chin-Wang Tu, David K. McElfresh, Serge L. Rudaz, Louis W. Cook, and Wayne L. Snyder. Red AlGaAs light-emitting diodes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):84–88, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Shan:1996:BPF

[SDDC96] Ming-Chien Shan, James W.

Davis, Weimin Du, and Qiming Chen. Business process flow management and its application in the telecommunications management network. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):70–76, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com:80/
                                        hpj/oct96/oc96a8.htm.

Seaver:1983:EC

[SDS83] M. Seaver, R. W. Dea, and R. J. Simms, Jr. An easy-to-use chartmaker. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(11): 10–12, November 1983. CODEN HPJOAX. ISSN 0018-1153.

Seipel:1981:MCH

[Sei81] W. Seipel. Magnetic components for high-frequency switching power supplies. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(8): 8–9, August 1981. CODEN HPJOAX. ISSN 0018-1153.

Seibel:1991:BSS

[Sei91] Michael J. Seibel. Builtin synthesized sweeper selftest and adjustments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(2): 17–23, April 1, 1991. CO-DEN HPJOAX. ISSN 0018-1153.

Seliger:1997:AAE

[Sel97] R. A. Seliger. An approach to architecting enterprise solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):86–95, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a11.htm.

Schinzel:1990:VPG

[SES90] Peter Schinzel, Volker Eberle, and Gunter Steinbach. A 300-MHz, variabletransition-time pulse generator output section. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(4): 85–92 (or 85–91??), August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Schmidt:1995:NDA

[SF95] Siegmar Schmidt and Halmo Fischer. A new design approach for a programmable optical attenuator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(1): 34–38, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_34.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Schmid:1985:PPO

[SFHE85] Wolfgang Schmid, Bernhard Flade, Klaus Hoeing, and Rainer Eggert. A precise, programmable 850nm optical signal source. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(1): 7–18, January 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Schneider:1980:MCP

[SFO80] R. F. Schneider, R. E. Felsenstein, and R. W. Offermann. Microwave CW and pulse frequency measurements to 40 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (4):3–5, 7–12, 13–14, April 1980. CODEN HPJOAX. ISSN 0018-1153.

Scruggs:1978:CLM

[SFW78] J. E. Scruggs, M. L. Faber, and D. L. Wolpert. A compact logging multimeter that can manipulate data. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(16): 28–32, December 1978. CODEN HPJOAX. ISSN 0018-1153.

Short:1993:BFF

[SGC93] Brian R. Short, Thomas L. Grisell, and Edward G. Cristal. Broadband fundamental frequency synthesis from 2 to 20 GHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(2):12– 16, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Smith:1983:MDI

[SGD83] Jeffrey H. Smith, Carlton E. Glitzke, and Alan J. DeVilbiss. Mainframe design for an integrated engineering workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3): 11–15, March 1983. CO-DEN HPJOAX. ISSN 0018-1153.

Snow:1977:HOS

[SH77] D. L. Snow and K. F. Hahn.

HP 1000 operating system is enhanced real-time executive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(7):7–14, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Saldanha:1990:QOD

[SH90a] Kevin S. Saldanha and Colette T. Howe. Qualification of an optical disk drive for autochanger use. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6): 35–37, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Shah:1990:SFH

[SH90b] Jayesh K. Shah and Charles L. Hamer. Support features of the HP OSI Express card. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (1):67–72, February 1990. CODEN HPJOAX. ISSN 0018-1153.

Shanahan:1971:USG

[Sha71] J. C. Shanahan. Uniting signal generation and signal synthesis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(4):2–13, December 1971. CODEN HPJOAX. ISSN 0018-1153.

Shaffer:1972:OAS

[Sha72] W. H. Shaffer. Organizing the automatic spectrum analyzer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(6):7–9, February 1972. CODEN HPJOAX. ISSN 0018-1153.

Shar:1976:SIG

[Sha76] L. E. Shar. Series II general-purpose computer systems: designed for improved throughput and reliability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(12):2–7, August 1976. CODEN HPJOAX. ISSN 0018-1153.

Sharrit:1978:DSM

[Sha78a] D. D. Sharrit. Designing serviceability into the Model 8568 A spectrum analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(10):20– 21, 23–24, June 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Shatzer:1978:DS

[Sha78b] R. R. Shatzer. Distributed systems/1000. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(7):15–20, March 1978. CODEN HPJOAX. ISSN 0018-1153.

Shafer:1985:DFC

[Sha85] Jimmy L. Shafer. Digital formatting and control electronics for half-inch tape data storage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(3):19–24, March 1985. CODEN HPJOAX. ISSN 0018-1153.

Shackleford:1989:NDS

[Sha89] J. Barry Shackleford. Neural data structures: programming with neurons. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3): 69–78, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Shan:1998:TDR

[Sha98]  J. Z. Shan.  A the-

oretical derivation of relationships between forecast errors. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):82– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Stewart:1982:HDI

[SHB^{+}82] Don D. Stewart, Robert J. Horning, Ken L. Burgess, Ronald G. Rogers, and James W. McLucas. Hardware design for an integrated instrumentation computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(5):7–17, May 1982. CODEN HPJOAX.

Smith:1984:CMS

[SHB84] Donald J. Smith, Johnnie L. Hancock, and Thomas K. Bohley. Counter module simplifies measurements on complex waveforms. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(6):33–40, June 1984. CODEN HPJOAX. ISSN 0018-1153.

Shergalis:1971:AFO

[She71]  L. D. Shergalis.  As-

tronomers find optical timing of pulsars more accurate. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(8):12–13, April 1971. CODEN HPJOAX. ISSN 0018-1153.

Sheers:1996:HOE

[She96] Kenneth R. Sheers. HP OpenView event correlation services. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(5):31–42, October 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/oct96/oc96a4.htm.

Showman:1989:OUI

[Sho89] Peter S. Showman. An object-based user interface for the HP NewWave environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):9–17, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Shreve:1982:SPU

[Shr82] William R. Shreve. Signal processing using surface acoustic waves. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(1):3–8, January 1982. CODEN HPJOAX. ISSN 0018-1153.

Sagun:1988:DCS

[SIBH88] Wilfredo T. Sagun, Fred H. Ives, Gary L. Baldwin, and Thomas Hornak. A 125-MHz 12-bit digital-to-analog converter system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2): 78–85, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Simms:1994:AST

[Sim94a]  Mark J. Simms.  Au-

tomatic state table generation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):21– 26, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_21.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Simms:1994:USM

[Sim94b] Mark J. Simms. Using state machines as a design and coding tool. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):27– 32, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_27.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Sasaki:1975:MCM

[SJ75] G. D. Sasaki and L. P.

Johnson. Minimal cost measuring instruments for systems use. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(5):18–??, January 1975. CODEN HPJOAX. ISSN 0018-1153.

Sasaki:1980:AMH

[SJ80] G. D. Sasaki and R. C.

Jensen. Automatic measurements with a highperformance universal counter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (9):21–23, 25–27, 29–31, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Sung:1994:FDH

[SJ94] Francis P. Sung and Mark A. Johnson. Faxing documents in HP MPower. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 53–61, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        53to61.pdf;   http://
                                        www.hp.com/hpj/toc-4-
                                        94.html.

Shintaku:1991:HVP

[SJL91] Larry Shintaku, M. Jerbic, and M. Linsley. The HP Vectra 486 personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):69–72, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Steinmetz:1969:RDC

[SK69] W. J. Steinmetz and R. L. Knapp. Recording data for computer analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(5): 15–20, January 1969. CO-DEN HPJOAX. ISSN 0018-1153.

Saponas:1980:LDS

[SK80] T. A. Saponas and B. W. Kerr. Logic development system accelerates microcomputer system design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(10): 3–10, 12–13, October 1980.

CODEN HPJOAX. ISSN 0018-1153.

Schad:1978:MMS

[SKB78] T. Schad, D. Kible, and P. Brunner. 1 mHzto-50 MHz signal source combines synthesizer accuracy, multimode operation, and easy programming. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(16):19–27, December 1978. CODEN HPJOAX. ISSN 0018-1153.

Smith:1990:HOW

[SKG90] Catherine J. Smith, Arthur J.

Kulakow, and Kathleen L. Gannon. HP OpenView Windows: a user interface for network management solutions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):60–65, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Sandberg:1986:AME

[SKK^{+}86] Gilbert I. Sandberg, Daryl E.

Knoblock, John C. Keith, Michael K. Bowen, and Ronald P. Dean. Advanced modular engineering workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(9):4–9, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Schild:1995:ODE

[SKWR95] Peter J. Schild, Wolfgang Klemm, Gerhard J. Walz, and Hermann J. Ruess. Open data exchange with HP PE/ SolidDesigner. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(5):35– 50, October 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95oct/
oct95_35.pdf;   http:
//www.hp.com/hpj/toc-
10-95.html.

Schweikardt:1993:ELT

[SL93] Horst Schweikardt and Edgar Leckel. Externalcavity laser temperature stabilization and power control. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 44(1):28–31, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Spencer:1995:SC

[SLM^{+}95] Thomas V. Spencer, Frank J. Lettang, Curtis R. McAllister, Anthony L. Riccio, Joseph F. Orth, and Brian K. Arnold. An I/O system on a chip. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2): 36–42, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_36.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Sloan:1989:PPT

[Slo89] Susan R. Sloan. Processing and passivation techniques for fabrication of high-speed InP/InGaAs/ InP mesa photodetectors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40 (5):69–75, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Small:1975:LSA

[SM75] C. T. Small and J. S. Morrill, Jr. The logic state analyzer, a viewing port for the data domain. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(12):2–11, August 1975. CODEN HPJOAX. ISSN 0018-1153.

Stone:1979:CBT

[SM79] P. S. Stone and J. F. Mc-Dermid. Circuit board testing: cost-effective production test and troubleshooting. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(3):2–8, March 1979. CODEN HPJOAX. ISSN 0018-1153.

Schwettmann:1982:IPT

[SM82] F. N. Schwettmann and J. L. Moll. IC process technology: VLSI and beyond. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(8):3–4, August 1982. CODEN HPJOAX. ISSN 0018-1153.

Speer:1986:MCL

[SM86] Martin L. Speer and Nicholas P. Mati. Modular computer low-end processor board design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(9):9–12, September 1986. CODEN HPJOAX. ISSN 0018-1153.

Smith:1987:LWP

[SM87] David L. Smith and Masahiko Muranami. A low-cost wireless portable printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38 (10):21–23, October 1987. CODEN HPJOAX. ISSN 0018-1153.

Schwartz:1988:TPM

[SM88] David J. Schwartz and Alan L. McCormick. Testing and process monitoring for a multifunction synthesizer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(1):77–80 (or 77– 79??), February 1988. CO-DEN HPJOAX. ISSN 0018-1153.

Smith:1991:HDH

[SMB^{+}91] Mark A. Smith, Lester S.

Moore, Preston D. Brown, James P. Dickie, David L. Smith, Thomas B. Lindberg, and M. Jack Muranami. Hardware design of the HP 48SX scientific expandable calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3): 25–34, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Strohmeier:1995:NHC

[SMD95] Fred Strohmeier, A. Maute, and R. Dinter. A new high-performance capillary electrophoresis instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 10–19, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_10.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Stone:1985:IPH

[SMGR85] Martin L. Stone, Peter L. Ma, Jeffery W. Groenke, and Todd L. Russell. An intelligent plotter for highthroughput, unattended operation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4):25–29, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Slater:1986:AEA

[SMH86] Lynn R. Slater, Jr., Craig M. Myles, and Keith A. Harrison. Aida: an expert assistant for dump readers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):34– 41, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Smith:1969:HTR

[Smi69]  J. H. Smith.  High-

resolution time-domain reflectometry with a portable 30-lb instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(13):8–14, September 1969. CODEN HPJOAX. ISSN 0018-1153.

Smith:1973:HPF

[Smi73] D. H. Smith. High performance flame-ionization detector system for gas chromatography. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 24(7):2–10, March 1973. CODEN HPJOAX. ISSN 0018-1153.

Smith:1977:LSA

[Smi77] J. H. Smith. A logic state analyzer for microprocessor systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(5):2–11, January 1977. CODEN HPJOAX. ISSN 0018-1153.

Smith:1979:SLF

[Smi79] R. L. Smith. A small, low-cost 12-megabyte fixed disc drive. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(7):11– 12, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Smith:1980:EDM

[Smi80] D. E. Smith. Electronic distance measurement for industrial and scientific applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(6):3–11, June 1980. CODEN HPJOAX. ISSN 0018-1153.

Smit:1984:HAH

[Smi84a] Paul R. Smit. A hyphenation algorithm for HPWord. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(9):26–30, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Smith:1984:MDL

[Smi84b] James H. Smith. Mechanical design of a large disc drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1):20–22, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Smith:1994:DOA

[Smi94] Dex Smith. Developing online application help. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 90–95, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        90to95.pdf;   http://
                                        www.hp.com/hpj/toc-4-
                                        94.html.

Straton:1984:HTI

[SMK84] Peter R. Straton, Scott R. McClelland, and Thomas E. Kilbourn. The HP 150 Touchscreen: an interactive user input device for a personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (8):11–15, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Scott:1988:BMD

[SMQ88] Perry E. Scott, John S.

Marvin, and Robert D. Quist. Boot mechanism for discless HP-UX. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(5):33–36, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Stever:1986:SSM

[SMR^{+}86] S. D. Stever, J. E. Mueller, T. G. Rodine, D. W. Olsen, and R. K. Tuttle. Seven-function systems multimeter offers extended resolution and scanner capabilities. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(2):15–23, February 1986. CODEN HPJOAX. ISSN 0018-1153.

Shelley:1997:LIP

[SMT^{+}97] D. J. Shelley, J. T. Majewski, M. R. Thackray, McWilliams, and J. L. A lower-cost inkjet printer based on a new printing performance architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 6–11, June 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97jun/ju97a1.htm.

Sischka:1988:CCO

[SNN88] Franz Sischka, Steven A. Newton, and Moshe Nazarathy. Complementary correlation optical time-domain reflectometry. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(6):14–21, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Snyder:1975:CPI

[Sny75] D. C. Snyder. Computer performance improvement by measurement and microprogramming. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(6):17–24, February 1975. CODEN HPJOAX. ISSN 0018-1153.

Scott:1998:OVF

[SOG98] N. D. Scott, D. M. Olsen, and E. W. Gannett. An overview of the VISUAL-IZE fx graphics accelerator hardware. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 49(2):28– ??, May 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98may/tc-05-98.htm.

Solomon:1981:RAC

[Sol81] R. J. Solomon. A reliable, accurate CO/ sub 2/ analyzer for medical use (using IR). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(9):3–7, September 1981. CODEN HPJOAX. ISSN 0018-1153.

Sorden:1974:NGF

[Sor74] J. L. Sorden. A new generation in frequency and time measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(10):2–8, June 1974. CODEN HPJOAX. ISSN 0018-1153.

Sorden:1988:PDO

[Sor88] James L. Sorden. Precision digital oscilloscopes and waveform recorders. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(1): 6–11, 14, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Spach:1992:PRI

[SP92] S. S. Spach and R. W.

Pulleyblank. Parallel raytraced image generation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3): 76–83, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Smith:1996:TGF

[SP96] Judith A. Smith and Meryem Primmer. Tachyon: a gigabit Fibre Channel protocol chip. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):99–112, October 1996. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com:
80/hpj/oct96/oc96a12.
htm.

Spangler:1972:BMC

[Spa72a] R. M. Spangler. BASIClanguage model 30 can be calculator, computer, or terminal. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(4):14–18, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Spangler:1972:NSP

[Spa72b] R. M. Spangler. A new series of programmable calculators. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(4):2–4, December 1972. CODEN HPJOAX. ISSN 0018-1153.

Showman:1984:AST

[SPA^{+}84a] Peter S. Showman, Karl W.

Pettis, Karlie J. Arkin, Jeffrey A. Spoelstra, John Price, W. Bruce Culbertson, and Robert D. Shurtleff, Jr. Applications software for the Touchscreen Personal Computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(8): 15–24, August 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Spaulding:1984:BTS

[Spa84b] William M. Spaulding. A broadband two-port Sparameter test set. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(11): 17–20, November 1984. CODEN HPJOAX. ISSN 0018-1153.

Spaulding:1985:HAM

[Spa85] William M. Spaulding. A high-level active mixer (for HP 3326A frequency synthesizer). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(8):25–29, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Schinzel:1993:DPG

[SPD^{+}93] Peter Schinzel, Andreas Pfaff, Thomas Dippon, Thomas Fischer, and Allan R. Armstrong. Design of a 3-GHz pulse generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2): 60–72, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Speaker:1996:TDT

[Spe96]  Lisa A. Speaker.  A

toolkit for developing TMN manager/agent applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):52–61, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/oct96/oc96a6.htm.

Sponheimer:1990:MRT

[Spo90] E. Sponheimer. Magnetooptical recording technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):8–9, December 1990. CODEN HPJOAX. ISSN 0018-1153.

Sproviero:1982:LPC

[Spr82] Michael J. Sproviero. Lowcost printers for the 9826A and 9836A computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(5):

22–24, May 1982. CODEN HPJOAX.

Sayed:1988:MSI

[SR88] Mohamed M. Sayed and John R. Regazzi. Millimeterwave sources and instrumentation. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(2):6–7, 9– 11, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Skeie:1995:HDA

[SR95] Tom A. Skeie and Michael R.

Rusnack. HP disk array: Mass storage fault tolerance for PC servers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 71–81, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95jun/
jun95_71.pdf;   http:
//www.hp.com/hpj/toc-
06-95.html.

Schwiebert:1982:EIF

[SS82] Erwin H. Schwiebert and Paul R. Spencer. Electrostatic image formation in a laser printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 33(6):16– 19, June 1982. CODEN HPJOAX.

Snell:1983:TIH

[SS83a] Sidnee Snell and Brian G. Spreadbury. A telephone interface for HP-IL controllers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6):5–7 (or 5– 6??), June 1983. CODEN HPJOAX. ISSN 0018-1153.

Szabo:1983:RPC

[SS83b] Thomas L. Szabo and Gary A. Seavey. Radiated power characteristics of diagnostic ultrasound transducers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(10):26– 29, October 1983. CODEN HPJOAX. ISSN 0018-1153.

Summers:1989:HFS

[SS89] James B. Summers and Douglas R. Snook. Highspectral-purity frequency synthesis in a microwave signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):37–41, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Sponheimer:1990:CDH

[SS90] Edward W. Sponheimer and John C. Santon. A CD-ROM drive for HP 3000 and HP 9000 computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):38–41, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Scherer:1992:PPA

[SSMK92] Dieter Scherer, William E.

Strasser, James D. McVey, and Wayne M. Kelly. The peak power analyzer, a new microwave tool. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):81–89 (or 81–83??), April 1992. CODEN HPJOAX. ISSN 0018-1153.

Stewart:1981:AMI

[SSMM81] L. J. Stewart, D. W.

Schaper, N. J. Martini, and H. E. Mostafa. Aspect microprocessor and I/O design for a drafting plotter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32 (11):7–11, November 1981. CODEN HPJOAX. ISSN 0018-1153.

Scoredos:1990:HKC

[SSV90] Eric C. Scoredos, Kimberly K. Scott, and Richard H. Van Gaasbeck. HP-UX kernel communications modules for a card-based OSI protocol stack. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(4):40–49, August 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Schmuhl:1996:HPN

[SSW96] Edward H. Schmuhl, Allan P. Sherman, and Jon D. Waisnor. HP PalmVue: A new healthcare information product. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3): 64–69, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/jun96/jun96.
                                        htm.

Schultz:1975:FGA

[ST75] S. E. Schultz and C. R. Trimble. Filling in the gaps-modular accessories for instrument systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (5):12–13, 15–18, January 1975. CODEN HPJOAX. ISSN 0018-1153.

Smith:1983:HLC

[ST83] Martin W. Smith and Joel D. Tesler. Highlevel language compilers for developing microprocessor systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(3):38–40, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Snook:1984:DSB

[ST84] Matthew L. Snook and Michael A. Teska. Digital subsystem for a board test system. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(10):14– 20, October 1984. CODEN HPJOAX. ISSN 0018-1153.

Smith:1990:DLL

[ST90] Judith A. Smith and Bill Thomas. Data link layer design and testing for the HP OSI Express card. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(1): 45–51, February 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Stallard:1977:HEP

[Sta77] S. J. Stallard. How the E-series performance was attained. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(7):20–23, March 1977. CODEN HPJOAX. ISSN 0018-1153.

Staas:1981:NDS

[Sta81a] G. C. Staas. New display station offers multiple screen windows and dual data communications ports. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(3):3–8, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Stavely:1981:HLG

[Sta81b] D. J. Stavely. A highquality, low-cost graphics tablet. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(1):15–24, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Starr:1987:HHI

[Sta87] Robert R. Starr. The Hewlett–Packard Human Interface Link. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(6):8–12, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Stambaugh:1988:TDH

[Sta88] Lisa B. Stambaugh. Tabledriven help screen structure provides on-line operating manual. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):21–24, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Stedman:1974:UFM

[Ste74a] J. M. Stedman. A useroriented family of minicomputers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(2):2–6, October 1974. CODEN HPJOAX. ISSN 0018-1153.

Stefanski:1974:DRC

[Ste74b]  A. Stefanski.  Deriv-

ing and reporting chromatograph data with a microprocessor-controlled integrator. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(4):18–24, December 1974. CODEN HPJOAX. ISSN 0018-1153.

Steiger:1980:PH

[Ste80a] G. W. Steiger. Packaging the HP-41C. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(3):7–??, March 1980. CODEN HPJOAX. ISSN 0018-1153.

Steiner:1980:MSS

[Ste80b] G. C. Steiner. A microcomputer system for spectrophotometer data processing. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(2):29–31, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Stewart:1980:L

[Ste80c] J. B. Stewart. The 64000 linker. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(10):25–26, October 1980. CODEN HPJOAX. ISSN 0018-1153.

Steinmetz:1981:FCP

[Ste81] J. H. Steinmetz. Flexible circuit packaging of a crystal oscillator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(3):26–28, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Steinmetz:1983:MDF

[Ste83] Charles R. Steinmetz. Mechanical design features of the laser head. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(4):19– 20, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Stephenson:1988:FTI

[Ste88] Paul S. Stephenson. Frequency and time interval analyzer measurement hardware. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):35–41, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Stearns:1989:AHN

[Ste89a] Glenn R. Stearns. Agents and the HP NewWave application program interface. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):32–37, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Stever:1989:DDM

[Ste89b] Scott D. Stever. An 8.5digit digital multimeter capable of 100000 readings per second and two-source calibration. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2):6–7, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Stolte:1990:IOL

[STE90] D. C. Stolte, B. A. Thompson, and D. Ellis. Integrating the optical library unit into the HP-UX operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(6):11–12, December 1990. CODEN HPJOAX. ISSN 0018-1153.

Stewart:1998:HSL

[Ste98] James W. Stewart. HP SnapLED: LED assemblies for automotive signal lighting. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 50(1):1–??, November 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98nov/tc-11-98.htm.

Stickel:1977:NDD

[Sti77] H. P. Stickel. New 50-Megabyte disc drive: high performance and reliability from high-technology design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(12):2–12, 14–15, August 1977. CODEN HPJOAX. ISSN 0018-1153.

Stimple:1997:TED

[Sti97]  Jim Stimple.  Testing

erbium-doped fiber amplifiers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97dec/de97a12.htm.

Smith:1988:DCG

[STMM88] James C. Smith, David C. Tribolet, Hatem E. Mostafa, and Emil Maghakian. Development of a color graphics printer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(4):16–21, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Stockwell:1974:PPC

[Sto74] R. K. Stockwell. Programming the personal computer (HP-65). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(9):8–14, May 1974. CODEN HPJOAX. ISSN 0018-1153.

Stoecker:1996:HOA

[Sto96] Paul A. Stoecker. HP OpenView agent tester toolkit. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(5):77–80, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com:80/
hpj/oct96/oc96a9.htm.

Stribling:1977:SGF

[Str77]  B. C. Stribling.  Sig-

nal generator features for a microwave synthesizer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(3):15–16, 18– 19, 21, November 1977. CODEN HPJOAX. ISSN 0018-1153.

Sukumar:1984:TPC

[Suk84] Srinivas Sukumar. Touchscreen Personal Computer offers ease of use and flexibility. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(8):4–6, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Schleifer:1980:TAS

[SW80] A. Schleifer and B. G. Willis. A task-oriented approach to spectrophotometry. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(2):11–17, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Sukumar:1981:HFS

[SW81] S. Sukumar and J. D. Wiese. Hardware and firmware support for four virtual terminals in one display station. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(3):13–15, March 1981. CODEN HPJOAX. ISSN 0018-1153.

Sismilich:1986:ICG

[SW86] Robert C. Sismilich and William T. Walker. Interactive computer graphics for manual instrument control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(5):17–23, 26, May 1986. CODEN HPJOAX. ISSN 0018-1153.

Spilman:1989:HNE

[SW89] Vicky Spilman and Eugene J. Wong. The HP NewWave environment help facility. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(4):43–47, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Schrenker:1990:QEL

[SW90a] Helge Schrenker and Wolfgang Wilde. Quality engineering for a liquid chromatography system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2): 11–16, April 1, 1990. CO-DEN HPJOAX. ISSN 0018-1153.

Strohmeier:1990:FPS

[SW90b] Fred Strohmeier and Klaus Witt. Flexible, precise solvent delivery for liquid chromatography. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):24–29, 31– 35, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Schuster:1991:PFT

[SW91] Otto Schuster and Joachim Weller. Production and final test of the HP component monitoring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (4):52–54, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Sweetser:1981:DTS

[Swe81] D. J. Sweetser. Designing testability and serviceability into the 9915A. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(7):27– 28, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Swerlein:1989:PAV

[Swe89] Ronald L. Swerlein. Precision AC voltage measurements using digital sampling techniques. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(2):15– 21, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Swenson:1996:CAS

[Swe96] Jill E. Swenson. Constructing an application server. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(3):

70–75, June 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/jun96/jun96.
htm.

Sheng:1997:ABM

[SWI97] B. B. Sheng, H. S. C.

Wallace, and J. S. Ignowski. Analog behavioral modeling and mixed-mode simulation with SABER and verilog. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):89– 94, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a15.htm.

Scharrer:1978:ILS

[SWM78] J. A. Scharrer, R. G. Wickliff, Jr., and W. D. Martin. Interactive logic state and timing analyses for tracking down problems in digital systems. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(6):14–20, February 1978. CODEN HPJOAX. ISSN 0018-1153.

Stavely:1990:ROD

[SWP90] Donald J. Stavely, Mark E.

Wanger, and Kraig A. Proehl. A rewritable optical disk library system for direct access secondary storage. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(6):6–7, 10, 12–13, December 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Taggart:1974:DTM

[Tag74] R. B. Taggart. Designing a tiny magnetic card reader. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(9): 15–17, May 1974. CODEN HPJOAX. ISSN 0018-1153.

Talle:1968:HDP

[Tal68] O. S. Talle, Jr. High dynamic performance X-Y recorder. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(4):8–11, December 1968. CODEN HPJOAX. ISSN 0018-1153.

Talbot:1988:DWS

[Tal88a] Mark D. Talbot. Digital waveform synthesis IC architecture. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):57–62, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Talbot:1988:FDM

[Tal88b] Mark D. Talbot. Firmware design for a multiplemode instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):70–73, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Talbott:1990:HOE

[Tal90] Glenn F. Talbott. The HP OSI Express card backplane handler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 41(1):8–18, February 1990. CODEN HPJOAX. ISSN 0018-1153.

Tanner:1989:PPD

[Tan89] Eve M. Tanner. Providing programmers with a driver debug technique. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(5):76–80, October 1989. CODEN HPJOAX. ISSN 0018-1153.

Tanner:1996:DSS

[Tan96] Craig J. Tanner. A decision support system for integrated circuit package selection. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(4):72– 79, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.

hp.com:80/hpj/aug96/augart7.
htm.

Tan:1997:FCP

[TBB^{+}97] Tun S. Tan, David M. Braun, Tim L. Bagwell, Chris Kocot, Joseph Straznicky, and Susan R. Sloan. Flip-chip photodetector for high-speed communications instrumentation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5):??, December 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97dec/de97a10.htm.

Thompson:1983:PDF

[TBBM83] D. C. Thompson, K. D. Boetzer, M. A. D. Bona, and B. M. Mousa. Product design of a friendly color graphics workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (9):28–30, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Tang:1983:GWE

[TBL^{+}83] Edward Tang, Otakar Blazek, Thomas K. Landgraf, Paula H. Ng, and Stephen P. Pacheco. The graphics workstation as an extensible computer terminal. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(9):22–25, September 1983. CODEN HPJOAX. ISSN 0018-1153.

Thomas:1992:SHE

[TBTM92] Bill Thomas, Alan C.

Berkema, Eric G. Tausheck, and Brian D. Mahaffy. Software for the HP EISA SCSI card. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(6):97– 108, December 1992. CO-DEN HPJOAX. ISSN 0018-1153.

Tsai:1979:LSM

[TC79] Lung-Wen Tsai and R. L.

Ciardella. Linear step motor design provides high plotter performance at low cost. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(2):7–9, 12– 14, February 1979. CO-DEN HPJOAX. ISSN 0018-1153.

Ta:1988:MDC

[TCWN88] Chuong Cam Ta, Lawrence W.

Chan, P. Jeffrey Wield, and Ruben Nevarez. Mechanical design of a color graphics printer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 39(4):21–27, August 1988. CODEN HPJOAX. ISSN 0018-1153.

Trautman:1983:PLH

[TD83] Jack P. Trautman and Lawrence A. DesJardin. A portable, low-cost, highperformance digital multimeter for the HP-IL. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (2):3–10, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Tsai:1992:AHT

[TD92] Sam S. Tsai and James B. Durr. Achieving high throughput with registerbased dense matrix relay modules. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):41–51, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Thompson:1986:HCL

[TE86] Bruce A. Thompson and David J. Ellis. Hierarchy chart language aids software development. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 37–42, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Temple:1983:DSB

[Tem83]  Y. Temple.  Device-

independent software for business graphics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(11):3–4, November 1983. CODEN HPJOAX. ISSN 0018-1153.

Terry:1978:VLP

[Ter78] F. D. Terry. Versatile 400-Ipm line printer with a friction-free mechanism that assures long life. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(15): 20–22, November 1978. CODEN HPJOAX. ISSN 0018-1153.

Tucker:1981:VDA

[TH81] M. G. Tucker and W. J.

Haydamack. VLSI design and artwork verification. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(6):25–29, June 1981. CODEN HPJOAX. ISSN 0018-1153.

Thayer:1989:CVG

[Tha89] Larry J. Thayer. Custom VLSI in the 3D graphics pipeline. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(6):74–77, December 1989. CODEN HPJOAX. ISSN 0018-1153.

Tan:1995:SEL

[THHW95] Michael R. T. Tan, Kenneth H. Hahn, Yu-Min D. Houng, and Shih-Yuan Wang. Surface emitting laser for multimode data link applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 46(1):67– 71, February 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95feb/
feb95_67.pdf;   http:
//www.hp.com/hpj/toc-
02-95.html.

Thompson:1972:NFM

[Tho72] L. Thompson. A new fivedigit multimeter that can test itself. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(12):7–12, August 1972. CODEN HPJOAX. ISSN 0018-1153.

Thomason:1977:ESS

[Tho77] J. L. Thomason. Expanding synthesized signal generation to the microwave range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(3):2–8, November 1977. CODEN HPJOAX. ISSN 0018-1153.

Thomas:1980:PSL

[Tho80] P. L. Thomas. A programmable selective level meter (wave analyzer) with synthesized tuning, autoranging, and automatic calibration. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(5):3–8, May 1980. CODEN HPJOAX. ISSN 0018-1153.

Thompson:1988:MSA

[Tho88] Kenneth S. Thompson.

Multifunction synthesizer applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(1):73–76, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Throne:1968:RFS

[Thr68] D. H. Throne. A rubidiumvapor frequency standard for systems requiring superior frequency stability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 19(11): 8–14, July 1968. CODEN HPJOAX. ISSN 0018-1153.

Tighe:1996:UU

[Tig96] Peter G. Tighe. Usable usability. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(6):88–93, December 1996. CODEN HPJOAX. ISSN 0018-1153.

Tivig:1991:GTP

[Tiv91] Gerhard Tivig. Globalization tools and processes in the HP Component Monitoring System. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):37–40 (or 37–39??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Tuttle:1989:VNT

[TL89] Myron R. Tuttle and Danny Low. Videoscope: a nonintrusive test tool for personal computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3): 58–64, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Tan:1998:ESC

[TLCB98] Ningxia Tan, Kenneth H. H. Lim, Bernard Chin, and Anthony J. Bourdillon. Engineering surfaces in ceramic pin grid array packaging to inhibit epoxy bleeding. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(3):81–??, August 1998. CODEN HPJOAX.

ISSN 0018-1153. URL

http://www.hp.com/hpj/
98aug/tc-08-98.htm.

Tivig:1991:PMH

[TM91] Gerhard Tivig and Wilhelm Meier. Patient monitor human interface design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (4):29–36, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Thatcher:1997:DST

[TOB97] T. J. Thatcher, M. M.

Oshima, and C. Botelho. Designing, simulating, and testing an analog phaselocked loop in a digital environment. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):84– 88, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a14.htm.

Toeppen:1986:DOU

[Toe86] Derek E. Toeppen. 1-GHz digitizing oscilloscope uses thick-film hybrid technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(4):33–36, April 1986. CODEN HPJOAX. ISSN 0018-1153.

Toeppen:1997:ACD

[Toe97] D. E. Toeppen. Acquisition clock dithering in a digital oscilloscope. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):26– 28, April 1997. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/97apr/ap97a5.htm.

Tong:1985:AMS

[Ton85] Gary L. Tong. Audio modulation section for an RF signal generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(12):31– 35, December 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Topham:1989:MDN

[Top89] Andrew D. Topham. Mechanical design of a new quarter-inch cartridge tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(4):67–73, August 1989. CODEN HPJOAX. ISSN 0018-1153.

Taylor:1979:MDH

[TPG79] P. N. Taylor, A. T. Pare, and J. R. Groff. Managing data: HP 300 files and data bases. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(6):16– 19, June 1979. CODEN HPJOAX. ISSN 0018-1153.

Trego:1978:PPI

[Tre78] M. P. Trego. Plotter/ printer interface languages: HP-GL and ASCII. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(13): 26, September 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Tribolet:1985:HXS

[TRH85] David C. Tribolet, Kenneth A. Regas, and Thomas J. Halpenny. The HP 7550A X-Y servo: state-of-the-art performance on a budget. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(4): 31–34, April 1985. CODEN HPJOAX. ISSN 0018-1153.

Takagi:1982:PSM

[TST82] S. Takagi, H. Sakayori, and T. Takeda. Programmable stimulus/measurement units simplify device test setups. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33 (10):15–20, October 1982. CODEN HPJOAX. ISSN 0018-1153.

Thompson:1992:AMO

[TSv92] Michael E. Thompson, Gregson P. Siu, and Jonathan van den Berg. An architecture for migrating to an open systems solution. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(6):109–114, December 1992. CODEN HPJOAX. ISSN 0018-1153.

Tiefert:1981:VPM

[TTML81] K. H. Tiefert, Dah Wen Tsang, R. L. Myers, and V. Li. The vertical power MOSFET for high-speed power control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(8):18–23, August 1981. CODEN HPJOAX. ISSN 0018-1153.

Tucker:1995:HTI

[Tuc95]  S. Paul Tucker.  HP

TeleShare: Integrating telephone capabilities on a computer workstation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2): 69–74, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/95apr/
apr95_69.pdf;   http:
//www.hp.com/hpj/toc-
04-95.html.

Tung:1974:CFP

[Tun74] C. C. Tung. The ‘personal computer’: A fully programmable pocket calculator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(9):2–7, May 1974. CODEN HPJOAX. ISSN 0018-1153.

Tuttle:1972:STO

[Tut72] R. K. Tuttle. The synthesized test oscillator — a new signal source for the 0.1 Hz—13 MHz range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 23(11): 2–8, July 1972. CODEN HPJOAX. ISSN 0018-1153.

Toschi:1976:AMF

[TW76] E. A. Toschi and T. Watanabe. An all-semiconductor memory with fault detection, correction, and logging. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(12):8–13, August 1976. CODEN HPJOAX. ISSN 0018-1153.

Tran:1990:HIO

[TW90] Mydung Thi Tran and David G. Wathen. The HP IVI object-oriented toolkit. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41 (5):11–20, October 1990. CODEN HPJOAX. ISSN 0018-1153.

Tuttle:1993:HUV

[TWCD93] Myron R. Tuttle, Kenneth M. Wilson, Samuel H. Chau, and Yong Deng. The HP Ultra VGA graphics board. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(3):31–41, June 1993. CODEN HPJOAX. ISSN 0018-1153.

Trutna:1993:REL

[TZ93] William R. Trutna, Jr. and Paul Zorabedian. Research on external-cavity lasers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (1):35–38, February 1993. CODEN HPJOAX. ISSN 0018-1153.

Uebbing:1977:SSA

[UAH77] J. T. Uebbing, P. B. Ashkin, and H. L. Hines. A small, solid-state alphanumeric display. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 28(8):13– 20, April 1977. CODEN HPJOAX. ISSN 0018-1153.

Ure:1992:MMC

[UAH^{+}92] Spencer M. Ure, Kevin W.

Allen, Anna M. Hargis, Samuel K. Hammel, and Paul Roeber. Meeting manufacturing challenges for PA-RISC workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (4):49–54, August 1992. CODEN HPJOAX. ISSN 0018-1153.

Ujvarosy:1994:FDD

[Ujv94] Damon R. Ujvarosy. Fast DDS-2 digital audio tape drive. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):6–11, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94dec/
dec94_6t.pdf;   http:
//www.hp.com/hpj/toc-
12-94.html.

Uebbing:1981:HSM

[ULW81] J. J. Uebbing, D. L. Lubin, and E. G. Weaver, Jr. Handheld scanner makes reading bar codes easy and inexpensive. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(1):3–10, January 1981. CODEN HPJOAX. ISSN 0018-1153.

Unter:1968:FCF

[Unt68] B. D. Unter. Fully calibrated frequency-domain measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 19(12):27– ??, August 1968. CODEN HPJOAX. ISSN 0018-1153.

Upham:1988:MDE

[Uph88] Herbert L. Upham. Millimeterwave detectors extend range of scalar network analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2):31–34, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Upshinsky:1971:HMN

[Ups71] S. Upshinsky. How to make a nuclear spectrum hold still. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22(8):9–11, April 1971. CODEN HPJOAX. ISSN 0018-1153.

Urquhart:1972:NML

[Urq72] R. Urquhart. A new microwave link analyzer with high-frequency test tones. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24 (1):8–9, September 1972.

CODEN HPJOAX. ISSN 0018-1153.

Urquhart:1976:ASL

[Urq76] J. R. Urquhart. An automatic selective level measuring set for multichannel communications systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(5): 2–7, January 1976. CO-DEN HPJOAX. ISSN 0018-1153.

Ujvarosy:1979:PEE

[US79] D. R. Ujvarosy and D. T.

Shaffer. Processor enhancements expand memory (Hewlett–Packard 9835A/B desktop computer). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(5): 13–15, May 1979. CODEN HPJOAX. ISSN 0018-1153.

Uhling:1993:RAO

[US93] Thomas F. Uhling and John R. Sterner. A rugged 2.5-GHz active oscilloscope probe. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):31–37, October 1993. CODEN HPJOAX. ISSN 0018-1153.

VanBrunt:1974:CPS

[Van74] R. C. Van Brunt. A computer power system for severe operating conditions. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26 (2):21–24, October 1974. CODEN HPJOAX. ISSN 0018-1153.

VanBree:1977:CDC

[Van77a]  K. A. Van Bree.  A

controller for the dynamic compiler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(11):21–23, July 1977. CODEN HPJOAX. ISSN 0018-1153.

VanBree:1977:SCS

[Van77b] K. A. Van Bree. Small computer system supports large-scale multi-user APL. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(11): 2–5, July 1977. CODEN HPJOAX. ISSN 0018-1153.

VanDyke:1977:DIC

[Van77c]  E. J. Van Dyke.  A

dynamic incremental compiler for an interpretive language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(11):17–20, July 1977. CODEN HPJOAX. ISSN 0018-1153.

VanVoorhis:1986:DCM

[Van86] Steven T. Van Voorhis.

Digital control of measurement graphics. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(1):24–26, January 1986. CODEN HPJOAX. ISSN 0018-1153.

VandeVoorde:1997:MSF

[Van97] P. Vande Voorde. MOS-FET scaling into the future. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(4):96–100, August 1997. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
97aug/au97a12.htm.

VanWinkle:1998:HCC

[Van98] Daniel D. Van Winkle. HP CaLan: A cable system tester that is accurate even in the presence of ingress. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1): ??, February 1998. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/98feb/fe98a8.htm.

VanMaren:1989:MTC

[VBK89] David J. Van Maren, Mark J. Bianchi, and Jeffery J. Kato. Maximizing tape capacity by superblocking. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(3):32–34, June 1989. CODEN HPJOAX. ISSN 0018-1153.

VanderPlas:1996:ALT

[VCMH96] Hubert A. Vander Plas, Russell B. Cinque, Zequn Mei, and Helen A. Holder. Assessment of lowtemperature fluxes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 47(4): 99–103, August 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

//hpcc997.external.hp.
com:80/hpj/aug96/tc-08-
96.htm;   http://www.
hp.com:80/hpj/aug96/augart11.
htm.

Vifian:1972:ME

[VDF72] H. Vifian, F. K. David, and W. L. Frederick. A ‘voltmeter’ for the microwave engineer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(3):2–7, November 1972. CODEN HPJOAX. ISSN 0018-1153.

Veteran:1991:DPS

[Vet91] David R. Veteran. DC-to-50-GHz programmable step attenuators. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(2):47– 49, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Vifian:1976:DNA

[Vif76]  H. Vifian.  A direct-

reading network analyzer for the 500-kHz-to-1.3-GHz frequency range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(11):2–6, July 1976. CODEN HPJOAX. ISSN 0018-1153.

Vobis:1988:OTR

[Vob88] Joachim Vobis. Optical time-domain reflectometer user interface design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39 (6):35–38, December 1988. CODEN HPJOAX. ISSN 0018-1153.

Vogel:1991:HSE

[Vog91] Eric L. Vogel. HP solve equation library application card. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3):22–25, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Voigt:1984:CLI

[Voi84] D. L. Voigt. A command language for improved disc protocol. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1):5–6, January 1984. CODEN HPJOAX. ISSN 0018-1153.

Wildnauer:1993:DMW

[WA93] Kenneth R. Wildnauer and Zoltan Azary. A doublepass monochromator for wavelength selection in an optical spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):68–74, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Wyld:1997:HAH

[WA97] B. C. Wyld and J.-P. Allegre. High availability in the HP OpenCall SS7 platform. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(4):65–71, August 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97aug/au97a8.htm.

Wade:1968:RTV

[Wad68] J. M. Wade. Recording true-rms voltages over wide dynamic ranges. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(2):9–14, October 1968. CODEN HPJOAX. ISSN 0018-1153.

Wagner:1975:UPD

[Wag75]  W. E. Wagner.  Un-

ravelling problems in the design of microprocessorbased systems (test instruments). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(12):12–16, August 1975. CODEN HPJOAX. ISSN 0018-1153.

Wagner:1988:DSC

[Wag88] Kimberly S. Wagner. Discless system configuration tasks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(5):37–39, October 1988. CODEN HPJOAX. ISSN 0018-1153.

Wagner:1993:PPG

[Wag93] Hans-Jürgen Wagner. A programmable 3-GHz pulse generator. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):52–55, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Waitman:1975:FMC

[Wai75] T. F. Waitman. Firmware for a microprocessorcontrolled CRT terminal. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 26(10): 16–19, June 1975. CODEN HPJOAX. ISSN 0018-1153.

Waite:1993:MBD

[Wai93] James W. Waite. A multirate bank of digital bandpass filters for acoustic applications (frequency analyzer). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(2):73–81, April 1993. CODEN HPJOAX. ISSN 0018-1153.

Wakamatsu:1997:DSL

[Wak97] H. Wakamatsu. A dielectric spectrometer for liquid using the electromagnetic induction method. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):37– 44, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a8.htm.

Walko:1970:ISC

[Wal70] N. E. Walko. Instrumentation systems controlled by time-shared computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 22 (1):7–9, September 1970. CODEN HPJOAX. ISSN 0018-1153.

Walker:1976:DPS

[Wal76] H. P. Walker. Designing precision into a selective level measuring set. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(5): 8–12, January 1976. CO-DEN HPJOAX. ISSN 0018-1153.

Wang:1988:DHW

[Wan88] Scott W. Wang. Discless HP-UX workstations. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(5): 6–9, October 1988. CO-DEN HPJOAX. ISSN 0018-1153.

Ware:1972:TDR

[War72] G. A. Ware. Time domain reflectometry in narrowband systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(7):2–7, March 1972. CODEN HPJOAX. ISSN 0018-1153.

Ward:1986:UQM

[War86] William T. Ward. Using quality metrics for critical application software. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3): 28–31, March 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Ward:1989:SDP

[War89] William T. Ward. Software defect prevention using McCabe’s complexity metric. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(2):64–69, April 1989. CODEN HPJOAX. ISSN 0018-1153.

Ward:1991:CRC

[War91] William T. Ward. Calculating the real cost of software defects. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):55–58 (or 55–57??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Ward:1992:DTD

[War92] Kenneth A. Ward. Digital test development software for a VXIbus tester. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2): 69–74, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Warntjes:1997:OLT

[War97a]  S. B. Warntjes.  An

oscilloscope-like logic timing analyzer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):29– 33, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a6.htm.

Warntjes:1997:SSR

[War97b] S. B. Warntjes. Sustained sample rate in digital oscilloscopes. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(2):23– 25, April 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97apr/ap97a4.htm.

Wasinger:1980:IH

[Was80]  G. F. Wasinger.  In-

terfacing the 3820A via the HP-IB. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(9):18–19, September 1980. CODEN HPJOAX. ISSN 0018-1153.

Watts:1981:UAD

[Wat81] K. F. Watts. A unifying approach to designing for reliability. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(7):24– 25, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Waters:1987:DRN

[Wat87a] Geoffrey Waters. A digital radio noise and interference test set. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(7):19– 26, July 1987. CODEN HPJOAX. ISSN 0018-1153.

Watkins:1987:SAU

[Wat87b] Marvin L. Watkins. Software architecture and the UNIX operating system: an introduction to interprocess communication. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(6): 26–36, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Watkins:1988:AUL

[Wat88] Marvin L. Watkins. Adapting UNIX logon mechanisms to automation applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(2):39–47, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Wechsler:1984:NHC

[WAZH84] Susan L. Wechsler, S. Abell, N. Zelle, and J. Hunt. A new handheld computer for technical professionals (HP-71B). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7):3–10, July 1984. CODEN HPJOAX. ISSN 0018-1153.

Wilbs:1983:HSA

[WB83a] Andreas Wilbs and Klaus-Peter Behrens. A highspeed system for AC parametric digital hardware analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(7):3–7, July 1983. CODEN HPJOAX. ISSN 0018-1153.

Wuerz:1983:DMS

[WB83b] L. J. Wuerz and C. Burns.

Dimensional metrology software eases calibration. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(4): 4–5, April 1983. CODEN HPJOAX. ISSN 0018-1153.

Wright:1989:RRE

[WB89] Larry R. Wright and Donald T. Borowski. Reducing radiated emissions in the performance signal generator family. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(5):59–68 (or 59– 65??), October 1989. CO-DEN HPJOAX. ISSN 0018-1153.

Witten:1990:HII

[WB90] Steven P. Witten and Hai-Wen L. Bienz. HP IVIBuild: interactive user interface builder for HP IVI. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(5):32–38, October 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Watkins:1984:HDH

[WBSC84] J. E. Watkins, P. A. Brown, G. Syeman, and S. E. Carrier. Hardware design of the HP 150 Personal Computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(8):25–30, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Wilson:1994:FIO

[WCH94] Michael R. Wilson, Lori A. Cook, and Steven P. Hiebert. A fast and intuitive online help system: WYSIWYG printing in an X application. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 79–89, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        79to89.pdf;   http://

www.hp.com/hpj/toc-4-
94.html.

Wylegala:1987:ATS

[WCT87] Thomas B. Wylegala, Long C. Chow, and Randy J. Teegarden. An automated test system for the first HP Precision Architecture computers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(3):18– 20, March 1987. CODEN HPJOAX. ISSN 0018-1153.

Witte:1984:ANA

[WD84] Robert A. Witte and Jerry W. Daniels. An advanced 5-Hz-to-200-MHz network analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(11):4–16, November 1984. CODEN HPJOAX. ISSN 0018-1153.

Wiese:1984:LLF

[WDDH84] Axel Wiese, Bernhard Dehmer, Thomas Dörr, and Guenter Höschele. A low-cost LC filterphotometric detection system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(4): 26–30, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Wechsler:1987:HBC

[Wec87] Susan L. Wechsler. A handheld Business Consultant. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(8): 4–10, August 1987. CO-DEN HPJOAX. ISSN 0018-1153.

Wechsler:1988:CTV

[Wec88] Mark Wechsler. Characterization of time varying frequency behavior using continuous measurement technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 40(1):6–12, February 1988. CODEN HPJOAX. ISSN 0018-1153.

Wende:1985:SGS

[Wen85] Michael T. Wende. Signal generator service features maximize uptime. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(12):10– 13, December 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Westlund:1977:HPO

[Wes77] R. C. Westlund. Higher precision in oscilloscope measurements of very short time intervals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(4):11–17, December 1977. CODEN HPJOAX. ISSN 0018-1153.

Westerteicher:1991:IHC

[Wes91] Christoph Westerteicher.

Introduction to the HP component monitoring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(4):6–8, 10, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Westerteicher:1991:CMS

[WH91] Christoph Westerteicher and Werner E. Heim. Component Monitoring System hardware architecture. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42 (4):10–13 (or 10–12??), October 1991. CODEN HPJOAX. ISSN 0018-1153.

Whatley:1968:RAL

[Wha68] L. A. Whatley. Rapid analysis of low frequency spectra. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 20(4):2–7, December 1968. CODEN HPJOAX. ISSN 0018-1153.

Webb:1995:HAT

[WHBP95] Steven L. Webb, Steven G.

Henry, Kevin S. Burke, and George Prokop. HP AccuPage 2.0: a toolkit for high-quality document scanning. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(4):43–50, August 1995. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
95aug/aug95_43.pdf;
http://www.hp.com/hpj/
toc-08-95.html.

Wong:1991:HLC

[WHH91] Roger W. Wong, Paul R. Hernday, and Daniel R. Harkins. High-speed lightwave component analysis to 20 GHz. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(1):6–13, February 1, 1991. CODEN HPJOAX. ISSN 0018-1153.

Wong:1989:HLC

[WHHC89] Roger W. Wong, Paul Hernday, Michael G. Hart, and Geraldine A. Conrad. High-speed lightwave component analysis. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 40(3):35– 51, June 1989. CODEN HPJOAX. ISSN 0018-1153.

White:1985:ATD

[Whi85] Reed I. White. Automated test data collection for IC manufacturing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 36(8):30–36, August 1985. CODEN HPJOAX. ISSN 0018-1153.

Whipple:1993:NAC

[Whi93] David P. Whipple. North American cellular CDMA. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44 (6):90–97, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Wickes:1987:ERC

[Wic87] William C. Wickes. An evolutionary RPN calculator for technical professionals. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(8):11–17, August 1987. CODEN HPJOAX. ISSN 0018-1153.

Wiederoder:1990:NMH

[Wie90] Herbert Wiederoder. A new modular high-performance liquid chromatograph. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2): 6–10, April 1, 1990. CO-DEN HPJOAX. ISSN 0018-1153.

Williams:1978:ESC

[Wil78]  R. W. Williams.  Ex-

perimenting with satellitelinked computer networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(7): 27–32, March 1978. CO-DEN HPJOAX. ISSN 0018-1153.

Wilken:1984:THD

[Wil84] Kent Wilken. Two highcapacity disc drives (HP 7933 and HP 7935). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(1): 3–6, January 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Williams:1997:USI

[Wil97a] S. A. Williams. Using SoftBench to integrate heterogeneous software development environments. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48 (1):22–27, February 1997. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/97feb/
                                        fe97a5.htm.

Wilson:1997:CSC

[Wil97b] J. B. Wilson. The C++ SoftBench class editor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48 (1):12–15, February 1997. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/97feb/
fe97a2.htm.

Winninghoff:1974:TM

[Win74] P. F. Winninghoff. Transient measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 25(12): 8–9, August 1974. CODEN HPJOAX. ISSN 0018-1153.

Winslow:1980:PSL

[Win80] P. D. Winslow. Precision synthesizer/level generator has high spectral purity for telecommunications testing. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31(5):9–11, 13, May 1980. CODEN HPJOAX. ISSN 0018-1153.

Winston:1981:NAC

[Win81] L. E. Winston. A novel approach to computer application system design and implementation. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 32(4):13– 18, April 1981. CODEN HPJOAX. ISSN 0018-1153.

Witte:1992:SRD

[Wit92a] R. A. Witte. Sample rate and display rate in digitizing oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):18–20 (or 18–19??), February 1992. CODEN HPJOAX. ISSN 0018-1153.

Witte:1992:LDO

[Wit92b] Robert A. Witte. Low-cost, 100-MHz digitizing oscilloscopes. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(1):6–11 (or 6–10??), February 1992. CODEN HPJOAX. ISSN 0018-1153.

Wong:1979:HBL

[WJ79] B. W. Wong and W. D. Jackson. A high-performance bipolar LSI counter chip using EFL and I^{2}L circuits. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30 (1):12–17, January 1979. CODEN HPJOAX. ISSN 0018-1153.

Willis:1980:DPH

[WJ80] B. G. Willis and G. E. James. Design and performance of a highly integrated parallel access spectrophotometer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 31(2):3–11, February 1980. CODEN HPJOAX. ISSN 0018-1153.

Wilson:1982:PDM

[WJ82] A. K. Wilson and D. E.

Johnson. Plotter drive motor encoder design. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 33(12): 26–27, December 1982. CODEN HPJOAX. ISSN 0018-1153.

Weisner:1991:PCA

[WJ91] Steven J. Weisner and Paul Johnson. The physiological calculation application in the HP component monitoring system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 42(4):40–43, October 1991. CODEN HPJOAX. ISSN 0018-1153.

Walter:1971:NHM

[WJT71] C. Walter, H. M. Juneau, and L. Thompson. A new high-speed multifunction DVM. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(5):1–15, January 1971. CODEN HPJOAX. ISSN 0018-1153.

West:1993:HDD

[WK93] Joseph N. West and J. Douglas Knight. A high-resolution direct-drive diffraction grating rotation system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(6):75–79, December 1993. CODEN HPJOAX. ISSN 0018-1153.

Wakasugi:1983:NML

[WKT83] Tomio Wakasugi, Takeshi Kyo, and Toshio Tamamura. New multi-frequency LCZ meters offer higherspeed impedance measurements. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(7):32–38, July 1983. CODEN HPJOAX. ISSN 0018-1153.

Wang:1984:HIU

[WLHC84] Scott W. Y. Wang, Jeff B. Lindberg, M. V. Hetrick, and M. L. Connor. HP-UX: implementation of UNIX on the HP 900 Series 500 computer systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(3):7–8, 10–11, 13– 15, March 1984. CODEN HPJOAX. ISSN 0018-1153.

                                        URL https://www.hpl.

hp.com/hpjournal/pdfs/
IssuePDFs/1984-03.pdf.

Wong:1978:RMO

[WM78] E. J. Wong and C. M.

Manley. RTE-IV: the megaword-array operating system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(14):6–11, October 1978. CODEN HPJOAX. ISSN 0018-1153.

Wise:1979:CLA

[WM79] D. M. Wise and J. C. Mc-Cullough. Configuring and launching the AMIGO/ 300 system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(7):20–21, July 1979. CODEN HPJOAX. ISSN 0018-1153.

Williams:1985:MPC

[WM85] Tim J. Williams and Nelson A. Mills. A multitasking personal computer system for the technical professional. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(10):4–6, October 1985. CODEN HPJOAX. ISSN 0018-1153.

Watson:1994:HCI

[WM94] Douglas R. Watson and Hatem E. Mostafa. Highquality color inkjet office printers. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):6–8, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/6to8.pdf.

Witte:1994:MPS

[WMBB94] Stephen B. Witte, Shelley I. Moore, Damon W. Broder, and David C. Burney. Media path for a small, low-cost, color thermal inkjet printer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(1):72–78, February 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94feb/72to78.pdf.

Wiederoder:1984:DHC

[WMZ84] Herbert Wiederoder, Roland Martin, and Juergen Ziegler. Design of the HP 1090 control system (liquid chromatography). Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(4): 8–13, April 1984. CODEN HPJOAX. ISSN 0018-1153.

Wang:1979:ASS

[WN79a] S. W. Y. Wang and L. F. Nelson. Applications software for the small-business computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(4):25–28, April 1979. CODEN HPJOAX. ISSN 0018-1153.

Wickliff:1979:ELA

[WN79b] R. G. Wickliff, Jr. and R. A. Nygaard, Jr. Expanding logic analyzer capabilities by means of the HP-IB. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 30(10):18–22, October 1979. CODEN HPJOAX. ISSN 0018-1153.

Wolpert:1975:MSC

[Wol75] D. L. Wolpert. A Multifunction Scanner for calculator-based data acquisition systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 26(5):17–??, January 1975. CODEN HPJOAX. ISSN 0018-1153.

Wolpert:1981:DLE

[Wol81] D. L. Wolpert. Data logging is easy with an HP-85/ 3054A combination. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 32(7): 7–8, July 1981. CODEN HPJOAX. ISSN 0018-1153.

Wolpert:1983:LIC

[Wol83] David L. Wolpert. Lowcost instrument control: a new ROM for the HP-41 handheld computers. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (2):16–19, February 1983. CODEN HPJOAX. ISSN 0018-1153.

Wolpert:1986:HCL

[Wol86] David L. Wolpert. HP-IB command library for MS-DOS systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(5):27– 29, May 1986. CODEN HPJOAX. ISSN 0018-1153.

Wong:1979:CEI

[Won79] D. T. Y. Wong. Controlling electromagnetic interference generated by a computer system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(9):17–19, September 1979. CODEN HPJOAX. ISSN 0018-1153.

Woodcock:1978:PPT

[Woo78] T. M. Woodcock. Printer and printing terminal grain versatility and mechanical simplicity with microprocessor control. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 29(15):2–7, November 1978. CODEN HPJOAX. ISSN 0018-1153.

Worrell:1992:DMF

[Wor92] Paul B. Worrell. Design of mainframe firmware in an open architecture environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(2):24–28, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Wickes:1991:HSE

[WP91] William C. Wickes and Charles M. Patton. The HP 48SX scientific expandable calculator: innovation and evolution. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(3):6–12, June 1991. CODEN HPJOAX. ISSN 0018-1153.

Warren:1994:DLP

[WP94a] Richard E. Warren and Conrad R. Proft. Design leverage and partnering in the design of a pressure scanning analog-to-digital converter. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):35–41, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        94oct/oct94_35.pdf;
                                        http://www.hp.com/hpj/
                                        toc-10-94.html.

Williams:1994:CDM

[WP94b] Michael K. Williams and Andreas M. R. Pfaff. Clock design and measurement issues in Pentium systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45 (6):68–77, December 1994. CODEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_68.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Williams:1994:TMC

[WP94c] Michael K. Williams and Andreas M. R. Pfaff. Tolerance mechanisms in clock distribution networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):70– 71, December 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94dec/
                                        dec94_70.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        12-94.html.

Williams:1994:MFM

[WPS94] Robert B. Williams, Harry K. Phinney, and Kenneth L. Steege. Mail facilities in a multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 71–78, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/94apr/
71to78.pdf;   http://
www.hp.com/hpj/toc-4-
94.html.

Witt:1985:SAP

[WR85] Stephen H. Witt and Roger W. Ruhnow. Simple architecture provides high performance for protocol analysis. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 36(7):12–13, 15– 18, July 1985. CODEN HPJOAX. ISSN 0018-1153.

Wasmuth:1986:PSA

[WR86] David B. Wasmuth and Bruce J. Richards. Predictive support: Anticipating computer hardware failures. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):30– 33, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Witte:1997:FIT

[WR97] R. A. Witte and W. Reid.

A family of instruments for testing mixed-signal circuits and systems. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(2):6–9, April 1997. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        97apr/ap97a1.htm.

Wright:1996:PTS

[Wri96] Cary J. Wright. Production test strategy for the HP E5200A broadband service analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 47(6):85– 87, December 1996. CO-DEN HPJOAX. ISSN 0018-1153. URL http:

                                        //hpcc997.external.hp.
                                        com:80/hpj/dec96/tc-12-
                                        96.htm.

Whitney:1972:PEC

[WRT72] T. M. Whitney, F. Rode, and C. C. Tung. The ‘powerful pocketful’: an electronic calculator challenges the slide rule. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 23(10):2– 9, June 1972. CODEN HPJOAX. ISSN 0018-1153.

                                        URL http://www.hp.com/
                                        hpj/72jun/ju72a1.htm.

Wilson:1984:DSI

[WS84a] Heather Wilson and Michael J. Shaw. Designing software for the international market. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(9):31–35, September 1984. CODEN HPJOAX. ISSN 0018-1153.

Winters:1984:CFD

[WS84b] Michael T. Winters and John F. Shelton. Comprehensive, friendly diagnostics aid A-series troubleshooting (HP 1000 computers). Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(2):23–26, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Wheeler:1983:NDR

[WSBK83] John K. Wheeler, John R.

Spencer, Dale R. Beucler, and Charlie G. Kohlhardt. 128K-Bit NMOS dynamic RAM with redundancy. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34 (8):20–24, August 1983. CODEN HPJOAX. ISSN 0018-1153.

Wilson:1987:PSS

[WSJ87] Edward L. Wilson, Kelly A.

Sznaider, and Clemen Jue. A planning solution for the semiconductor industry. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(4):21–27, April 1987. CODEN HPJOAX. ISSN 0018-1153.

Williamson:1984:DLC

[WST84] Donald A. Williamson, Steven C. Steps, and Bruce A. Thompson. Designing a low-cost 3-MIPS computer. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 35(2):12–17, February 1984. CODEN HPJOAX. ISSN 0018-1153.

Walker:1997:GSB

[WSYD97] Richard Walker, Cheryl Stout, Chu-Sun Yen, and Lewis R. Dove. A 2.488gbit/s silicon bipolar clock and data recovery circuit for SONET fiber-optic communications networks. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(5): ??, December 1997. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/97dec/
                                        de97a11.htm.

Weir:1997:VCP

[WT97] D. Weir and P. G. Tobin. Verifying the correctness of the PA 7300LC processor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(3): 69–72, June 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://cpus.hp.
com/techreports/random_
code.pdf;   http://cpus.
hp.com/techreports/verifying.
pdf;   http://www.hp.
com/hpj/97jun/ju97a9.
htm.

Wiese:1990:NGL

[WTB^{+}90] Axel Wiese, Konrad Teitz, Volker Brombacher, Guenter Hoeschele, and Hubert Kuderer. A new generation of LC absorbance detectors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 41(2):36–43, April 1, 1990. CODEN HPJOAX. ISSN 0018-1153.

Wood:1984:OSF

[WW84] Laurie E. Pollero Wood and Charles H. Whelan. Operating system and firmware of the HP 150 personal computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(8):6–10, August 1984. CODEN HPJOAX. ISSN 0018-1153.

Wu:1998:AFP

[WWS^{+}98] Der-Woei Wu, John S.

Wei, Chung-Yi Su, Ray M. Parkhurst, Shyh-Liang Fu, Shih-Shun Chang, and Richard B. Levitsky. Appendix A — FET parameters. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1):??, February 1998. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        98feb/fe98a5.htm.

Wyatt:1992:EDH

[Wya92] Kenneth D. Wyatt. EMC design of the HP 54600 series oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(1):41–45 (or 41–44??), February 1992. CODEN HPJOAX. ISSN 0018-1153.

Werner:1987:MCG

[WYO^{+}87] Kalr-Heinz Werner, Stephen Yie, Friedhelm M. Ottliczky, Harold B. Prince, and Heinz Diebel. ME CAD geometry construction, dimensioning, hatching, and part structuring. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 38(5): 16–29, May 1987. CODEN HPJOAX. ISSN 0018-1153.

Webb:1997:DCS

[WYSE97] S. L. Webb, K. J. Youngers, M. J. Steinle, and J. A. Eccher. Design of a 600pixel-per-inch, 30-bit color scanner. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 48(1):54–61, February 1997. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/97feb/fe97a8.htm.

Yackle:1980:AAM

[Yac80] B. E. Yackle. An assembler for all microprocessors. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 31 (10):28–30, October 1980. CODEN HPJOAX. ISSN 0018-1153.

Yanagawa:1992:ANA

[Yan92]  Koichi Yanagawa.  An

advanced 5-Hz-to-500-MHz network analyzer with high speed, accuracy, and dynamic range. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 43(2):101– 109, April 1992. CODEN HPJOAX. ISSN 0018-1153.

Yanagimoto:1993:FHM

[Yan93]  Yoshiyuki  Yanagimoto.

A fast-switching, highisolation multiplexer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5): 95–99, October 1993. CO-DEN HPJOAX. ISSN 0018-1153.

Yergenson:1992:MDH

[YF92] Robin P. Yergenson and Timothy A. Figge. Mechanical design of the HP 54600 series oscilloscopes. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43 (1):36–40, February 1992. CODEN HPJOAX. ISSN 0018-1153.

Yousefi:1995:SSB

[YGR95] Manny Yousefi, Adel Ghoneimy, and Wulf Rehder. A Software Solution Broker for technical consultants. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(2): 93–101, April 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95apr/
                                        apr95_93.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        04-95.html.

Yokokawa:1977:WAL

[YK77] M. Yokokawa and K. Kanafuji. A wide-ranging, automatic LCR meter for stand-alone or systems applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 29(4):18–23, December 1977. CODEN HPJOAX. ISSN 0018-1153.

Yetter:1987:HPA

[YLJ^{+}87] Jeffry D. Yetter, Jonathan P.

Lotz, William S. Jaffe, Mark A. Forsyth, and Eric R. DeLano. HP precision architecture NMOS-III single-chip CPU. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 38(9):12–18, September 1987. CODEN HPJOAX. ISSN 0018-1153.

Yanagimoto:1993:RDC

[YN93] Yoshiyuki Yanagimoto and A. Nukiyama. Receiver design for a combined RF network and spectrum analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 44(5):85–94, October 1993. CODEN HPJOAX. ISSN 0018-1153.

Yajima:1998:DBS

[YN98] Shunichiro Yajima and Antoni C. Niedzwiecki. Direct broadcast satellite applications. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 49(1):??, February 1998. CODEN HPJOAX. ISSN 0018-1153. URL

http://www.hp.com/hpj/
98feb/fe98a6.htm.

Yoder:1994:HMC

[Yod94] William R. Yoder. HP MPower: A collaborative multimedia environment. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(2): 10–19, April 1994. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/94apr/
                                        10to19.pdf;   http://
                                        www.hp.com/hpj/toc-4-
                                        94.html.

Yonekura:1994:HFI

[Yon94] Takanori Yonekura. Highfrequency impedance analyzer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 45(5):67–74, October 1994. CODEN HPJOAX. ISSN 0018-1153. URL

                                        http://www.hp.com/hpj/
                                        94oct/oct94_67.pdf;
                                        http://www.hp.com/hpj/
                                        toc-10-94.html.

York:1983:HHP

[Yor83] D. C. York. HP-IL and the HP-75 Portable Computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 34(6):8–9, June 1983. CODEN HPJOAX. ISSN 0018-1153.

Young:1976:PGE

[YPS76] I. R. Young, R. Pearson, and P. M. Scott. A 50-mbit/s pattern generator and error detector for evaluating digital communications system performance. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 27(7):18–24, March 1976. CODEN HPJOAX. ISSN 0018-1153.

Yen:1992:GLC

[YWP^{+}92] Chu-Sun Yen, Richard C.

Walker, Patrick T. Petruno, Cheryl Stout, Benny W. H. Lai, and W. J. McFarland. G-link: a chipset for gigabit-rate data communication. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(5):103–116, October 1992. CODEN HPJOAX. ISSN 0018-1153.

Zamborelli:1973:TFW

[Zam73] T. Zamborelli. Thick films widen attenuator response. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 24(9): 9–??, May 1973. CODEN HPJOAX. ISSN 0018-1153.

Zander:1983:A

[Zan83]  P. A. Zander.  Auto-

peaking. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(5):12– 13, May 1983. CODEN HPJOAX. ISSN 0018-1153.

Zettler:1969:HHC

[ZC69] R. A. Zettler and A. M. Cowley. Hybrid hot carrier diodes. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 20(6):13–20, February 1969. CODEN HPJOAX. ISSN 0018-1153.

Zellmer:1969:HIP

[Zel69] J. Zellmer. High impedance probing to 500 MHz. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 21(4): 12–16, December 1969. CODEN HPJOAX. ISSN 0018-1153.

Zeller:1970:OCR

[Zel70] G. Zeller. Optical card reader for fast calculator programming. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 22(2):8–12, October 1970. CODEN HPJOAX. ISSN 0018-1153.

Zellers:1979:ENA

[Zel79] J. R. Zellers. An economical network analyzer for the 4-to-1300-MHz frequency range. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 30(10):9–17, October 1979. CODEN HPJOAX. ISSN 0018-1153.

Zellers:1991:SFD

[Zel91] James R. Zellers. A 0.01-to-40-GHz switched frequency doubler. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 42(2):31–33, April 1991. CODEN HPJOAX. ISSN 0018-1153.

Zurakowski:1986:DIC

[ZFE^{+}86] Mark P. Zurakowski, Domingo A.

Figueredo, Scott S. Elliott, George A. Patterson, William J. Anklam, and Susan R. Sloan. Diode integrated circuits for millimeter-wave applications. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 37(11):14– 21, November 1986. CO-DEN HPJOAX. ISSN 0018-1153.

Zellmer:1983:MLT

[ZHN83] Joel A. Zellmer, John E.

Hanna, and David L. Neuder. A modular logic timing analyzer for the 64000 system. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 34(3):23– 30, March 1983. CODEN HPJOAX. ISSN 0018-1153.

Zimmermann:1995:CHH

[Zim95] Hans-Peter Zimmermann. Capillary handling in the HP capillary electrophoresis instrument. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(3): 25–31, June 1995. CO-DEN HPJOAX. ISSN 0018-1153. URL http:/

                                        /www.hp.com/hpj/95jun/
                                        jun95_25.pdf;   http:
                                        //www.hp.com/hpj/toc-
                                        06-95.html.